A Bibliography of Papers in Lecture Notes in Computer Science (1995), Part 2 of 2

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

29 August 2024 Version 1.26

Title word cross-reference

(1, +k(n)) [553]. (k, n) [320]. 0 − 1, 0, +1 [243]. 1 [385]. 2 [821, 688, 2180, 165, 858, 2001, 934]. 3 [2213, 361, 602, 584, 597, 833, 588, 1150, 2170, 2173, 2226, 2212, 1539, 1535, 685, 2176, 653, 45, 332, 806, 600, 858, 2202, 2001, 2219, 1529, 590, 2181, 596, 840, 2179, 664]. 3+ [1326]. 4 [624]. a [714]. c [1084]. d [1102]. f [1731, 146]. h [524]. k [150, 59, 683, 122, 1855, 1731, 616]. κ [153]. λ [375]. μ [575]. N [442, 214, 77]. NN [683, 616]. p [330]. π [375, 258, 252, 255, 251]. q [1106]. r [384]. t [320]. v [576].

-algebra [122]. -Body [442]. -Branching [553]. -Calculus [375, 258, 252, 255, 714, 251, 575, 576]. -Cheater [320]. -Colorings [1731, 146]. -Coloured [150]. -D [821, 602, 584, 597, 1150, 2226, 1539, 685, 653, 600, 590, 934, 596, 664]. -D-Motion [588]. -Dimensional [1102, 688, 45, 2001]. -Domination [384]. -generated [165]. -Gram [1106]. -nearest [1855]. -Optimal [1084]. -person [77]. -Processor [330]. -RED [2137]. -Relation [524]. -Trees [59, 1731].

0/1 [1117]. 0/1-Integer [1117]. 0p2 [132]. 0p2-Hard [132].

1 [448]. 1-Integer [1117]. 100 [358]. 17th [1027].

2 [2142, 941, 308, 1726, 144, 352]. 2-Adic [308]. 2-Monotonic [1726]. 2-Trees [2142, 144]. 20m [932].

3 [2045, 173, 397, 1712, 1993]. 3-Colourability [1993]. 3-Competitive [173]. 3-SAT [1712]. 3-Space [397]. 3/2 [173]. 3kappa [155]. 3n [1721]. 3rd [1002].

4 [72]. 4- [39]. 4-OP [72].

64 [309].

’88 [307]. 8mm [1774].

90 [338]. 93 [1961].

= [168, 2062, 155, 115].

A.I. [856]. Abduction [1437]. Abductive [102, 1486, 1857]. Abelian [165]. abilities [75]. Abnormal [1172]. Abstract [570, 377, 380, 257, 1218, 566, 1641, 1377, 805, 397, 1187, 1228, 1232, 1211, 1235, 523, 1309]. Abstracting [1681]. Abstraction [370, 277, 285, 2017, 477, 1486, 1368]. Abstractions [1215, 1646]. Abstracts [339]. AC [510, 943]. AC-Termination [510]. Acceleration [1766]. Accept [552]. Acceptance [265]. Access [1917, 1621, 1589, 426, 1628, 66]. ACCESS.bus [2286]. Accessibility [551]. according [1961]. Account [570]. Accounting [1750]. Accuracy [1245, 1770, 587]. Achievements [1765]. Achilles [2311]. Acoustic [844, 900]. Acquisition [647, 1455, 849, 939]. Acquisitions [590]. Across [527, 236]. Action [2151, 1172, 1171, 1619, 1438]. Actions [1408, 1476, 1172, 1686, 1170]. Activation [1270]. Active [1281, 1940, 1278, 292, 1271, 676, 1938, 1272, 2320, 1262, 1260, 1687, 1265, 2111, 1279, 1280, 656, 1277, 2083, 1939, 1039, 1261, 781, 764, 837, 712, 1267, 1269, 791, 1615, 619, 1941, 1282, 1866, 1802]. Active-Camera [676]. Activity [1968, 1981, 1263, 832]. Acyclic [60, 2006]. adaptable [1839]. Adaptation [1858, 1452, 1836, 1874, 1849, 1853, 87, 1860]. adaptation-guided [1860]. Adaptative [1448]. Adapted [1148]. ADAPtER [1857]. Adaptive [860, 1122, 2042, 2198, 836, 1127, 1519, 801, 2224, 706, 2194, 2080, 424, 121, 1525, 2031, 1894, 1514, 852, 2211, 625, 804, 616, 1867, 85]. ADBMS [1260]. Adder [931]. Adding [951]. Additive [1788, 850, 193]. Address [876]. Adequacy [720]. adequate [1882]. Adic [308]. Adjusting [2302]. Administrations [2103]. Administrative [1468]. Admission [1003, 2034]. Adoption [1395]. Advance [2012, 2013, 2014]. Advanced [770, 338, 903, 1608, 1618, 1978, 930]. Advances [694]. Advancing [2203]. Adversary [1121]. Advice [536]. Affine [154, 622, 633, 695]. after [2126]. Against [1121, 292, 2065, 750]. Age [1005]. Agency [454]. Agent [1281, 1160, 102, 91, 99, 1470, 1397, 100, 1441, 1450, 1508, 1472, 786, 1416, 96, 472, 1614, 1509, 2107, 1452, 1458]. Agents [1449, 98, 1280, 1507, 94, 1645, 576, 117, 93]. Agglomerative [2123]. Aggregates [1782]. Aggregating [1889]. Agreement [755, 1421, 1450, 1110, 2265, 2249]. Agricultural [1592]. AI [2215, 785]. Aid [717, 1057]. Aided [292, 2209, 2078, 293, 2244]. Aiding [1453, 1836]. aimed [1844]. Aircraft [1626, 896, 1834]. Airport [896]. Algebra [370, 1664, 672, 119, 665, 285, 2151, 122, 105]. Algebraic [458, 1050, 1322, 1669, 476, 196, 108, 1537, 1929, 1823, 120, 2132]. Algebras [1910, 205]. Algorithm [47, 968, 177, 1147, 2037, 1138, 889, 1415, 757, 127, 2272, 2298, 16, 158, 1730, 2230, 1419, 2220, 1510, 671, 1124, 2281, 1985, 56, 952, 826, 1739, 941, 915, 2260, 1733, 26, 1920, 301, 131, 1697, 173, 190, 192, 62, 628, 1737, 683, 195, 1726, 1096, 686, 1523, 2204, 2187, 300, 1712, 1448, 186, 21, 637, 106, 2177, 2148, 1445, 1993, 1511, 178, 1894, 1696, 878, 668, 1461, 261, 1526, 2211, 1113, 710, 2160, 1524, 2035, 599, 1790, 82, 1807, 1789, 73, 1581, 1882, 83, 1811, 387]. Algorithmic [1666, 1671, 1127, 139, 1695, 52]. Algorithms [812, 327, 53, 1104, 42, 1084, 2165, 2158, 1741, 2003, 665, 54, 1156, 1038, 1720, 156, 383, 408, 33, 1401, 1126, 240, 39, 2010, 157, 1150, 1095, 391, 189, 38, 1137, 853, 1893, 1895, 2241, 69, 756, 1667, 1039, 333, 121, 787, 444, 1517, 1151, 1673, 1683, 1518, 2282, 1152, 137, 2007, 199, 1111, 1731, 1937, 78, 1305, 81, 89, 1864, 79, 80]. Aliasing [823]. alibi [155]. Alignment [1531]. Alignments [68]. All-Pairs [2304, 1096]. Allocation [1769, 1222, 1131, 465, 1284, 1283]. Allow [1372]. Almost [671, 1718, 539]. along [587]. Alphabet [1696]. Alphanumeric [2193]. Alternating [153, 180, 1685]. Alternative [449, 1627]. Always [1718]. Ambiguities [1952]. Ambiguity [1032]. Amdahl [1728]. Among [94, 1513, 93]. Amortization [48]. Amount [2189]. Analogical [1581]. Analogue [2259, 2265, 1864]. Analogy [1407, 1848]. Analyse [556]. Analyses [1681, 2062]. Analysis [140, 647, 1270, 1008, 1271, 2027, 854, 1004, 798, 2253, 1140, 608, 1378, 818, 711, 833, 829, 1219, 667, 305, 1372, 1007, 831, 1221, 1222, 1009, 1223, 1003, 453, 189, 1224, 792, 706, 2212, 1225, 853, 433, 3, 879, 1190, 598, 1227, 1188, 880, 1545, 2234, 1123, 1768, 895, 2277, 443, 1929, 892, 615, 629, 1232, 1966, 1234, 1511, 790, 253, 1934, 1005, 1353, 2201, 1148, 1235, 997, 1284, 1370, 268, 898, 2245, 2246, 1328, 1655, 87, 1312, 1307, 1855]. Analytic [2058]. Analytical [882, 1380, 700, 2120]. Analyzers [1214, 1236, 1237]. Analyzing [1011, 2242, 336]. Anchored [2308]. And- [410]. And-parallelism [411]. AND/OR [545]. ANDF [441]. Angel [971]. Angles [811]. Animating [1970]. Animation [2248, 2215]. Anisotropic [625]. Anisotropy [904]. Annealing [1134, 1151, 1518, 79, 80]. Announcement [1638]. Anomalies [1154]. Anonymous [560]. Ant [1512]. Ants [90]. anytime [1882]. anytime-algorithm [1882]. Aperture [698]. API [2042]. ApOFIS [872]. Apparatus [2032]. Appearance [1529]. Appearance-Based [1529]. Applicability [1138]. Application [1626, 995, 48, 798, 1821, 608, 1519, 1741, 830, 2122, 1819, 2072, 1320, 176, 510, 983, 902, 727, 940, 1892, 154, 2075, 1971, 1043, 1123, 290, 181, 1885, 1526, 72, 1850]. Application-Specific [940]. Applications [188, 328, 1075, 1812, 57, 1938, 1758, 1026, 1145, 1050, 434, 1265, 1016, 1018, 948, 1693, 766, 1981, 867, 932, 598, 374, 302, 555, 1971, 2256, 1768, 993, 764, 1527, 1774, 335, 2024, 2036, 34, 590, 992, 991, 2086, 129, 943, 2035, 1877, 1345]. Applied [1594, 2076, 1256, 1039]. Applying [2163, 901]. Approach [1551, 812, 1239, 1990, 1048, 1482, 1394, 1769, 813, 2069, 102, 13, 592, 2238, 608, 1099, 470, 1611, 882, 100, 1449, 1421, 1450, 1483, 1488, 1587, 577, 1454, 1896, 1499, 654, 2084, 2109, 556, 1254, 1409, 1777, 1465, 809, 688, 3, 2221, 2116, 2105, 519, 1939, 8, 1119, 2190, 1070, 1958, 1412, 978, 405, 2191, 794, 1522, 1610, 1149, 2087, 713, 800, 1213, 1025, 682, 133, 2040, 1398, 1844, 1317, 1837, 1880, 85, 1856, 74, 1830]. Approaches [1938, 369, 948, 761, 71]. Appropriate [1342]. Approximability [544, 1988]. Approximate [1116, 1004, 1421, 1009, 620, 1723, 1106, 2058]. approximately [1583]. Approximating [1708, 2003, 31, 1114, 1208]. Approximation [63, 185, 2293, 1143, 125, 1137, 308, 1643, 64, 688, 1712, 846, 1993, 1112, 1696, 638, 199, 1113, 1111, 972, 1852, 1792, 2315]. Approximations [1709, 1710, 541, 2059, 1801]. Arabic [2239, 2240, 877]. Arbitrary [550, 563, 2021, 1192, 2179]. Arboricity [2282]. Arc [1394, 427, 1704, 1866]. Architectural [360, 100, 1043]. Architecture [1075, 1051, 363, 1890, 885, 942, 1441, 1594, 357, 1374, 2020, 1978, 2044, 731, 1588, 92, 1299, 1675, 906, 2188, 78, 2135]. Architecture- [2020]. Architectures [461, 355, 1620, 356, 340, 466, 950, 945, 1017, 341, 1416, 1366, 1624, 1148, 2314]. Architecturing [1385]. Arcs [649]. Area [784, 2267]. Areal [1331]. Argument [1403]. Argument-Based [1403]. Arguments [228, 312, 1403, 52]. Arising [1695]. Arithmetic [327, 113, 962, 532, 1213, 1797, 1807]. Arithmetical [2311]. Arrangement [177, 829]. Arrangements [34]. Array [810, 455, 433, 1190, 644, 1524]. Arrays [946, 430, 923, 1190, 2048]. Arrival [997]. Arrive [32]. Art [1692, 1754, 769, 1665]. Artifacts [1968, 1428]. Artificial [1396, 473, 1428, 1157, 815, 1658]. Ask [972]. Aspect [1775, 2237]. Aspects [1600, 1024, 1913]. Assertion [2081]. Assertions [987]. Assessing [1379]. Assessment [1434, 1002, 935, 1866]. Assets [2074]. Assignment [1216, 1740, 1132, 1133, 559]. Assistant [1074]. Assisted [831, 2208, 2212, 1298, 1304]. Associative [503, 333, 1885]. Associative-Commutative [503]. ASTERIX [1594]. Asteroidal [55, 1114]. ASTRAL [1363]. Astronomical [942, 834]. Asymmetric [298]. Asymptotically [542]. Asynchronous [325, 1154, 2310, 1225, 756, 2288, 269, 956, 163, 1312]. ATM [1622, 2042, 1595, 1594, 1120, 1003, 1604, 2020, 1625].

Atomic [1408, 745, 1619]. attack [1793, 1794]. Attacks [292, 305, 2258]. Attention [1446]. Attentive [1494, 707]. Attributed [798, 2192, 1207]. Attributes [699, 1365, 2110, 2186]. Atypical [1171]. Audio [2027, 2026, 2028]. Augmentation [1724, 1119, 1117]. Augmented [554]. Authenticated [755]. Authentication [289, 154, 290, 2267]. Author [324]. Authoring [1692]. Authorization [2082, 1965, 1966]. Autocorrelation [1343]. Automata [326, 287, 536, 385, 724, 1638, 1641, 567, 392, 575, 192, 2310, 573, 398, 620, 1556, 531, 401, 1815, 1649, 264, 1685, 1213, 1562, 1312, 2050, 1316]. Automata-Theoretic [1213]. Automated [670, 903, 1167, 2317, 876, 1750]. Automatic [2231, 905, 828, 987, 1632, 874, 1473, 818, 921, 679, 439, 946, 431, 246, 1530, 1176, 777, 4, 217, 479, 2190, 1468, 2242, 1429, 1448, 623, 2243, 649, 1302, 599, 2060, 1870]. Automatically [717, 110]. Automation [276, 727]. Automaton [2148]. Automaton-Driven [2148]. Automotive [791]. Autonomous [1160, 2174, 1497, 1492, 1496, 94, 472, 1513, 1490, 2171, 93, 1615]. Autotransformation [337]. Average [1722, 1090, 1701, 1245, 2295, 169, 1855]. average-case [1855]. Aversion [1032]. AVL [1818]. Avoidance [672]. Avoiding [698, 1678, 1224]. avoids [73]. Axiomatisations [263]. Axiomatising [279]. Axiomatization [2290]. Axis [1736, 39].

B [57, 2235, 1627, 2179]. B-ISDN [1627]. B-Spline [2235, 2179]. B-Trees [57]. Back [445, 573]. Back-Propagation [445]. Backprojection [643]. Backpropagation [1447, 2234]. Backtracking [1152]. Balance [2302]. Balanced [29, 1787]. Balancing [752, 28, 1146, 1526, 1108]. Band [2159]. Bandwidth [1114, 414]. Bank [2189]. Bars [404]. Base [116, 1456, 1956, 1835]. Based [1055, 860, 140, 1505, 1238, 1239, 2082, 1965, 1075, 1048, 813, 2069, 2000, 802, 318, 1968, 894, 1916, 814, 1053, 1173, 119, 1443, 1547, 1974, 1494, 2158, 2233, 1339, 675, 2235, 1536, 1449, 671, 1243, 2073, 1252, 1441, 651, 921, 1488, 778, 1781, 2072, 372, 989, 1253, 2068, 1896, 839, 962, 1351, 1373, 240, 915, 1222, 1892, 881, 2170, 2063, 777, 2017, 2208, 1254, 2212, 1255, 308, 805, 2250, 824, 809, 853, 957, 1393, 1467, 2221, 2261, 1293, 2075, 1256, 613, 1543, 1513, 1060, 2105, 2259, 1606, 2204, 849, 919, 2195, 1980, 151, 897, 1623, 2077]. Based [645, 1299, 815, 2071, 846, 913, 642, 1760, 1411, 1246, 793, 978, 1774, 615, 1211, 1165, 593, 2024, 1161, 2191, 264, 1529, 1249, 1427, 1466, 1628, 1367, 239, 1550, 199, 1403, 522, 856, 650, 1250, 710, 2247, 2160, 1025, 1654, 1259, 438, 1656, 2240, 1937, 616, 2274, 1834, 1844, 1845, 1865, 1846, 1867, 13, 1847, 311, 1869, 1872, 1873, 1875, 1876, 1618, 1851, 1852, 2078, 1878, 2051, 1879, 1889, 1838, 2128, 1841, 1807, 429, 1857, 1881, 1859, 1800, 83, 1860, 668, 1884, 120, 79, 80, 2042, 920, 922, 927, 935, 912, 451, 934, 1476, 1588]. Bases [1162, 1775, 1244, 148, 1951]. Basic [386, 2139, 1981, 1877, 1971, 842, 2136]. Basis [1674, 1764, 176, 2152, 2190, 1247, 298, 1570]. Batch [2287]. Baum [668]. Bayesian [884, 2162, 644, 687, 855]. BDDs [1670]. Be [66, 981, 266, 1301, 1728]. Beam [2212]. Beast [2065, 1279]. Beauty [2065]. Behavior [1265, 1944, 2090, 2092]. Behaviors [1160, 1490]. Behaviour [1271, 1264, 2068, 92, 2297, 1525]. Behavioural [2215, 260, 1001, 2089, 87]. Behaviours [90]. behind [1665]. Belief [1389, 1431, 1489, 2200, 1488, 530]. Belief-Goal [1489]. Beliefs [101]. Believers [1507]. Believing [1506]. Belnap [1484]. Benchmark [1279]. Benchmarking [1036]. Benefits [1593, 416]. bent [1787]. Best [300]. Better [1191, 1761, 1152, 175]. Between [1753, 745, 517, 1244, 411, 1402, 272, 765, 1598, 1619, 604]. Betweenness [1099]. Beyond [1699, 1088, 1234]. Bicoercibility [2290]. Biconnectivity [1095]. Bicriteria [1988]. Bidirectional [449]. Bigloo [1233]. Bill [1072, 10]. Binarization [865]. Binary [1703, 2302, 2298, 2002, 1101, 651, 1786, 691, 1105, 2097, 655, 2146, 2268, 1577]. Binary-Decision [1703]. Binding [1219, 1195, 1229, 1235]. Binding-Time [1219, 1229, 1235]. Binomial [1123]. Biology [198]. Bioreactors [626]. Bisection [1088, 2001]. Bisimilarity [2299, 251]. Bisimulation [272, 563, 568, 1727, 2055]. Bisimulations [284]. Bit [1773, 2044, 632, 2049, 937]. Bit-Rate [632]. BIT-Serial [937]. Bit-Sliced [1773]. Bitemporal [1917]. bitvector [2062]. Bivariate [176]. Blackboard [1416]. Bladder [885]. Blind [648]. Block [860, 2250, 1790, 1800, 1810]. Blocking [745]. Blocks [1971]. Blood [2212]. Blue [879]. Blue-Print [879]. Board [929, 1840]. Boards [939, 912]. Body [1541, 442]. Body-Centered [1541]. Bolts [1738]. Bond [903]. Boolean [1505, 525, 1482, 1310, 1787, 1999, 172, 122, 545, 1726, 241, 404]. Boosting [462]. Boot [592]. Boot-Strapping [592]. Boppana [186]. Boppana-Halldorsson [186]. Borderline [1402]. Both [1722, 275]. Bottleneck [2296]. Bottom [2285]. Bottom-Up [2285]. Bound [1156, 1154, 2280, 1088, 190, 154, 2232, 982, 1155, 1747, 553]. Boundaries [649, 236]. Boundary [364, 848, 638]. Boundary-Value [364]. Bounded [25, 1110, 294, 547, 1107, 532, 1649]. Bounds [243, 67, 552, 425, 25, 1702, 2041, 537, 544, 797, 303, 1132, 414, 551, 186, 532]. Box [394, 1802]. boxes [1791]. Boyer [1303]. Brain [827, 828, 2211, 832]. BRAM [400]. Branch [1156, 1154, 2280, 190, 982, 1155]. Branch-and-Bound [1156, 1155].

Branching [278, 553]. Branchings [2298]. Break [298]. Breast [829]. Bridging [527]. Brightness [891]. Broadband [1611, 1605]. Broadcast [747, 2250, 1513, 178]. Broadcasting [177, 754, 284, 179]. Browser [502]. Bruijn [1181]. BSP [1084]. Bucket [290]. Budget [182]. Buechi [2310]. Buffer [53, 997]. Buffers [1107]. Build [2123, 474]. Building [1746, 1397, 1530, 1015, 1971, 288, 1259, 1839, 85]. Built [507, 936, 1382]. Built-in [507]. Burst [2043]. Business [1358, 2102, 2101]. Buy [1759]. Byzantine [754, 755, 756].

C [111, 2126, 1594, 805, 479]. Cache [428, 1299, 429, 1302]. CAD [1036, 910, 782]. Calculation [347, 58, 2205]. Calculus [1408, 375, 504, 1631, 1916, 1415, 1183, 258, 286, 2167, 371, 577, 252, 372, 2149, 2289, 1552, 284, 279, 1181, 394, 122, 255, 969, 569, 714, 251, 1184, 1230, 1560, 1232, 1325, 575, 2053, 576, 83]. Calibration [1767, 698, 676, 838, 634, 633, 2243, 695]. Call [517, 1229]. Call-By-Name [1229]. Call-by-Need [517]. Camera [2217, 676, 633, 712]. Cameras [674]. Can [90, 1903, 2070, 1672, 981, 2030, 1581, 1015]. CAN-Simulator [1015]. Cancer [2208]. Canonical [513]. Can’t [1759]. cap [168]. Capacities [102, 50]. Capacity [37, 997]. CAPITL [1742]. Capturing [2068, 1031, 222]. Carcinoma [885]. Card [1258]. Cardiological [886]. Carriers [1056]. carry [1796]. Cartesian [519]. Cartographic [1331]. Cascades [299]. Case [1055, 1238, 1239, 2069, 2065, 1833, 2073, 1252, 2067, 1869, 1483, 2072, 1831, 1253, 2068, 240, 491, 2064, 1849, 1818, 1851, 2063, 1254, 1701, 1088, 1255, 2078, 1879, 1244, 2075, 1256, 1257, 1841, 2077, 645, 429, 2071, 1246, 1247, 1624, 1249, 2120, 1360, 1250, 1259, 1834, 1844, 1845, 1865, 1846, 1867, 1835, 1847, 1870, 1872, 1873, 1875, 1876, 1852, 1853, 1838, 1855, 1857, 1881, 1859, 1860, 2105]. Case-Based [1055, 1238, 1239, 2069, 2073, 1252, 1253, 2068, 240, 1254, 1255, 2075, 1256, 2077, 2071, 1246, 1249, 1250, 1259, 1869, 2078, 1879, 1841, 1844, 1845, 1865, 1846, 1867, 1847, 1872, 1873, 1875, 1838, 1857, 1881, 1859, 1860]. Cases [1240, 1241, 1406, 1722, 2074, 1856, 1248, 1866, 1847, 1839, 1854, 1882, 1861]. Cash [322, 1879]. Catastrophic [1715]. Categorial [1414, 390]. Categorical [215]. Categorically [2297, 1326]. Categories [519, 1447, 1559, 2118]. Categorisation [2123]. Category [380]. Catenation [395]. Causal [1344, 1487, 1011, 1474, 1498, 2291, 1849]. Caused [1952]. Cavity [361]. Cayley [381, 165, 561]. CBR [1251, 1868, 1837, 2076, 1850, 1853, 2075, 1880, 2070, 1842, 1883]. CCP [972]. CCS [2300]. CDL [341]. CDMA [2277]. Cell [639, 826, 1832, 626, 869]. Cells [670, 829]. Cellular [325, 328, 385, 331, 332]. Cellular-Neural [328]. Census [489]. Center [26]. Centered [1332, 1541, 1428]. Centric [2016]. Centroid [1116]. Century [1027]. CERN [1772]. Certain [1988]. Certificates [1083, 1095, 2263]. Chain [1120, 2197]. Chains [1006]. Challenge [1542]. Challenges [1765, 1692, 475]. Change [1044, 2111, 1499, 1474, 1172, 1686, 1170]. Changed [353]. Changes [2088, 1779]. Channel [272]. Channels [2266, 2259]. Chaos [1672]. Character [703, 679, 2193, 2192, 2190, 661, 2188, 877]. Characterisation [510]. Characterising [260]. Characteristic [2198, 1173, 1564]. Characteristics [1332, 1800]. Characterization [802, 1401, 196]. Characterized [1443]. Characterizing [212, 1473]. Characters [883, 2207, 878, 2240, 1937]. Charging [1598]. Chart [1426]. Cheap [562].

Cheater [320]. Check [2189, 2056]. Checker [1361]. Checkers [1558]. Checking [968, 1394, 2313, 1684, 1215, 249, 276, 1320, 273, 278, 250, 373, 574, 251, 248, 737, 1317, 2059, 1301, 2053, 2050, 2056, 2315]. Checks [1224, 2258]. Chemical [1056, 1497]. Cheques [2263]. Chess [90, 1255, 1870, 1877]. Chest [2209]. Chinese [2192, 2188]. CHIP [992]. Choice [285, 969]. Choices [481, 2249]. Cholesky [1147]. Chomsky [395]. Choosing [591]. Chord [1698]. Chordal [2009]. Chorus [1911]. Chromatic [48, 168]. Chu [1680]. chunking [1870]. Church [2271, 151, 521]. CICERO [1074]. CIM [771, 777, 766, 786, 779, 785, 789, 781, 1081, 764, 784, 790, 768]. CIM-Laboratory [790]. CIM-Structures [784]. cipher [1794, 1790]. Ciphering [915]. Ciphers [2258, 1795, 1800, 1810]. Circle [155]. Circuit [1090, 724, 2135]. Circuits [243, 82, 1312, 1303, 1311]. Circular [427, 635, 1484, 649]. Circumnavigating [2174]. City [1346]. Class [187, 19, 1028, 1306]. Classes [535, 760, 185, 537, 1179, 222, 539, 1579, 1583, 1576]. Classical [204]. Classification [639, 1890, 2220, 1401, 234, 240, 950, 1962, 809, 646, 879, 897, 816, 694, 1258, 15, 623, 1399, 603, 2246, 693, 1861]. Classifier [1469, 1245, 661, 1855]. Classifiers [801, 658, 2239]. Classifying [235]. Classroom [1031]. ClearCase [1755]. Clepsydra [470]. Clifford [672]. Climbing [2311]. Clipper [302]. Clipper-like [302]. Clique [2009]. Clock [1784]. Clock-controlled [1784]. Closed [1732, 112, 636, 2218, 519, 637, 861, 1579]. Closer [1396]. Closest [131]. Closing [627]. Closure [1097, 1227, 135]. Cloud [894]. CLP [962, 983, 964, 990]. Cluster [2204, 406]. Cluster-Based [2204]. Clustered [126, 1098]. Clustering [827, 13, 1443, 1038, 2085, 1737, 598].

Clusters [2309, 348]. Cluttered [631]. CM [1761, 1762]. CML [566]. Co [2069, 1619]. Co-Operative [2069]. Co-ordinate [1619]. Coal [2224]. Coalgebra [1557]. Coalition [94, 93]. Cocomparability [1707]. Code [813, 949, 1217, 431, 488, 354]. Codebook [889]. Coded [2021]. Codes [2266, 2251, 2278, 2272, 2255, 2260, 154, 2261, 58, 2271, 2252, 148, 2267, 2274]. Codesign [928]. Codifying [1553]. Coding [2170, 1697, 932, 2232, 890, 632, 888, 2273, 2267, 2197, 1807]. Coefficients [2006, 937]. Coercions [114]. Cognition [1350]. Cognitive [1470, 1487, 1423, 1449, 667, 1496, 1340, 1460, 1867, 1883]. Cograph [379]. Coherence [519, 429, 1302]. Coherency [428]. Coin [313, 173]. Coincidence [268]. Collection [1295, 1291, 1285, 1293, 1296, 1294, 1192, 1298, 1290, 1286]. Collections [1942]. collective [75]. Collector [1288, 1287, 753, 1297]. Collision [2225, 672]. Collisions [1492]. Collusion [323]. Collusion-Secure [323]. Colony [1512]. Color [795, 879, 2201, 713, 2091]. COLOR-X [2091]. Coloring [1995, 1115, 1667]. Colorings [1731, 146]. Colour [820]. Colourability [2283, 1993]. Coloured [150, 1654]. Combination [973, 1432]. combinations [1570]. Combinatorial [1510, 1672, 1151, 1461, 985, 336]. Combinatory [1415, 2153]. Combined [1008]. Combining [868, 719, 119, 675, 1962, 107, 2050, 1545, 1037, 761, 2285, 1857]. combustion [1842]. Come [1395]. Comer [2180]. Commercial [1761, 1768, 1527]. Commitment [745, 294, 1357, 956]. Committed [296]. Common [40, 707]. Communicating [1551, 2060]. Communication [524, 2039, 1129, 744, 2038, 1021, 2000, 1519, 457, 1090, 339, 1783, 445, 1608, 343, 2226, 2020, 394, 1600, 1225, 182, 551, 334, 887, 1623, 1607, 349, 178].

Communications [439, 773]. Commutative [503, 386, 867]. Commutativity [1960, 114]. COMOGA [1522]. Compact [2010, 182]. Compaction [440, 1289, 1313]. Comparative [1832]. Compared [657, 953]. Comparing [535, 219, 198, 2152, 1518]. Comparison [1281, 1433, 1620, 1908, 2233, 1244, 333, 1249, 790, 693]. Comparisons [1721, 1845, 2279]. Compelling [1407]. Compensation [823]. Competitive [42, 1699, 2029, 74, 173]. Compilation [989, 341, 479, 1276, 441, 1964]. Compile [1285, 1192, 1013]. Compile-Time [1285, 1013]. Compiler [925, 2138, 964, 448, 1233, 1828]. Compilers [438]. Compiling [1063, 924, 923]. Complementary [1287]. Complementation [1218]. Complements [59]. Complete [951, 1737, 1184, 259, 125, 544, 547]. Completeness [418, 2161]. Complex [835, 13, 1536, 1274, 1535, 2096, 1066, 1550, 2081, 1846]. Complexity [1666, 1708, 1703, 540, 2299, 380, 1090, 206, 31, 582, 526, 220, 529, 557, 543, 1701, 1988, 518, 278, 208, 2307, 222, 374, 1400, 551, 2259, 2152, 1525, 1576, 1715, 93, 129, 124, 1579, 2314]. Component [1383, 1277, 1076]. Components [1383, 1196, 2242, 137, 784, 1315]. Comportment [1188]. Composing [232, 1483]. Composite [1454, 1939]. Composition [570, 1064, 2188]. Compositional [249, 250, 264]. Compound [997]. Comprehension [1190]. Comprehensions [1190]. Compressed [1773, 2303]. Compression [651, 1300, 2232, 467, 822, 2056]. Computable [733, 169]. Computation [337, 984, 296, 588, 586, 445, 50, 517, 932, 432, 396, 293, 221, 632, 2154, 2288, 1100, 2258, 1662, 52, 533, 664, 76]. Computation/Communication [445]. Computational [1666, 1519, 205, 206, 1449, 408, 1428, 1141, 1668, 529, 557, 1693, 208, 2015, 1688, 1334, 1354, 92, 1174, 349, 75]. Computations [325, 326, 328, 1144, 1089, 116, 1136, 414, 123, 135]. Compute [1903, 402]. Computer [1682, 1968, 1657, 119, 847, 2209, 917, 1758, 675, 667, 1918, 831, 1150, 2215, 2170, 2208, 2212, 1255, 2078, 999, 708, 332, 773, 444, 1537, 895, 354, 978, 909, 1080, 1935, 105, 2244, 2222, 358, 2139, 2125, 2314, 2135, 83]. Computer-Aided [2209, 2078]. Computer-Assisted [2212]. Computer-Based [1968]. Computers [2158, 445, 2207, 364, 351, 1306]. Computing [1698, 35, 1912, 1730, 415, 55, 1110, 368, 2308, 989, 176, 1175, 389, 41, 529, 1669, 107, 741, 2017, 40, 2309, 742, 45, 1660, 913, 1673, 1719, 934, 1658, 1935, 133]. Concatenated [2274]. Concentration [626]. Concept [1505, 2069, 1473, 1442, 1475, 709, 3, 1165, 1391, 842, 1503, 1583]. Concepts [1338, 1819, 1045, 1981, 1614, 1607, 790, 1080]. Conceptual [10, 2123, 13, 2134, 2108, 16, 1494, 19, 2098, 11, 18, 17, 12, 2101, 2097, 6, 1456, 21, 5, 15, 1485, 103, 9, 23, 20]. Conceptualizations [1348]. Concerning [401]. Concrete [1557]. Concurrency [274, 1078, 1960, 2152, 441]. Concurrent [375, 2284, 329, 345, 1216, 783, 989, 281, 1225, 1754, 969, 2154, 282, 1680, 2153, 986, 1367, 1209, 2164, 1297, 1317]. Condensation [1630]. Condition [1278, 2283]. Conditional [1495, 1433, 506, 512, 514, 520]. Conditions [1432, 2104, 485]. Condorcet [382]. Cone [2212]. Cone-Beam [2212]. Conf [1073]. Conferencing [2036]. confident [1582]. Configurable [99, 911, 909]. Configuration [1743, 1590, 1603, 1753, 1772, 1744, 1467, 1748, 1754, 1760, 1747, 2036, 1756, 1763, 1750, 1368]. Conflict [517, 1513, 1933]. Conflictfreeness [1764]. Conflicting [1432]. Conflicts [2115].

Confluence [507, 512, 262]. Confluent [969, 975]. Confocal [2202]. Congestion [468]. Congruences [564, 260]. Conics [673]. Conjecture [2255]. Conjectures [210, 1575]. Conjunctive [132]. Connected [2158, 137]. Connecting [336]. Connection [2038, 1339, 1594, 1003]. Connectionist [1886, 1393, 1880]. Connectionist/Symbolic [1393]. Connectionless [1018]. Connections [2013, 1643]. Connectivity [1083, 1724]. Connects [1334]. coNP [125]. Conscious [1941]. Consciousness [1198]. Consensus [745, 141, 1042]. Consequences [210]. Conservative [157]. Considerations [232]. Consistency [1394, 965, 2293, 1488, 954, 1019, 2291, 2081, 225]. Consistent [639, 1633, 2122, 2090, 280]. Constancy [713]. Constant [841, 1709, 1275, 759, 2044, 221]. Constant-Space [221]. Constrained [752, 2198, 128, 1139, 1137, 961, 2034, 2184, 1522, 1149, 1103]. Constraining [699]. Constraint [973, 1482, 988, 1202, 976, 1414, 1274, 989, 962, 959, 958, 220, 979, 1640, 954, 957, 64, 1467, 96, 969, 1178, 1924, 963, 980, 982, 993, 978, 1208, 986, 1521, 953, 1209, 2164, 991, 1392, 985, 956, 974, 1656, 1566, 1876, 2051, 1858]. Constraint-Based [1467, 978, 1656]. Constraint-Programming [980]. Constraint-satisfaction [1482]. Constraints [973, 609, 194, 788, 2293, 952, 2021, 1162, 196, 967, 182, 2154, 1525, 1093, 1276, 1460, 901, 977, 2089, 118, 578, 1213, 1573]. Construct [1992, 2081]. Constructing [988, 142, 1645, 1463, 1444, 1326]. Construction [150, 1821, 1722, 1473, 2073, 1787, 147, 1125, 397, 44, 242, 1411, 1628, 1269, 1979, 1824, 1169, 2138]. Constructions [751, 2278, 1552, 567, 2274]. Constructive [971, 206]. Constructivist [102]. Constructor [1179, 520].

Constructor-Sharing [520]. Consumer [977]. Consumers [1191]. Contact [698, 712]. Containing [833, 1953]. Containment [132, 1316]. Content [1454, 2064, 670]. Contex [1923]. Contex-Sensitive [1923]. Context [554, 563, 386, 556, 8]. Context-Free [554, 563, 386]. Contexts [1551, 18, 7, 1386, 1267, 1]. Contextual [901]. Contextuality [1480]. contiguity [81]. Contiguous [393]. Continuous [1512, 1632, 2019, 2049, 1387, 1647, 2035]. Contour [656, 843, 637, 2236]. Contours [828, 875, 602, 851, 636, 587, 2179]. Control [2225, 914, 1630, 2046, 1078, 1636, 1637, 1772, 1223, 256, 2215, 1892, 772, 1589, 932, 1643, 1645, 1296, 896, 2116, 1509, 1412, 1648, 1617, 1647, 1891, 712, 1652, 468, 769, 239, 943, 2023]. Control-Flow [1223]. controllable [1791]. Controlled [1399, 1784]. Controller [1629, 1652]. Controllers [940, 1895, 1644, 1656]. Controlling [2218, 114, 1257]. Convergence [752]. Conversations [100]. Conversion [1931]. Conversions [736]. Convert [2196]. Convex [812, 31, 30, 2196]. Convexity [161, 1339]. COO [1753]. Cooperating [117]. Cooperation [99, 95, 897]. Cooperative [1160, 1280, 1753, 1018, 1778, 1080]. Coordinated [96]. Coordinates [594]. Coordination [362, 412, 1751, 1509, 746]. Cope [314, 1820]. copies [1571]. Coplanar [673]. Coprocessor [460, 465]. Copy [2020]. Coq [1560]. CORBA [1609]. Core [1505, 1920]. Cornell [1642]. Corners [610]. Corporate [2122]. Correct [1812, 1826, 1813]. Correcting [2278]. Correction [2322]. Correctness [1194, 2148]. Correlation [1803, 2258, 1793]. Correspondence [2175]. Correspondences [2199]. Corrosion [646]. Cortical [816]. COs [1598]. Cosine [933]. Cost [1767, 744, 1835, 808, 1034, 2007, 359].

Costly [51]. Costs [1021, 1593, 1783, 416]. counter [1562]. Counterfeit [173]. Counterflow [731]. Counting [535, 159, 1593, 1983]. countries [1841]. Coupled [556, 1329]. Coupled-Context-Free [556]. Course [498, 2136, 2128]. Courtesy [2189]. Covariance [847]. Covering [651]. Covers [361, 345]. Cow [2281]. Cow-Path [2281]. CPA [1307]. CPS [1229]. CPS-Translation [1229]. CPU [994, 2035]. CPUs [999]. Crack [904]. Craig [142]. Creating [2207]. Creation [917]. Credit [1258]. Crest [611]. Crew [979]. Criterion [1244]. Critical [1909, 1283]. CritiGUI [1980]. Critique [2070]. Crop [615]. Cross [1354, 160]. Cross-3-intersection [160]. Cross-Linguistic [1354]. Crossover [1893]. CRP [1651]. Cryptanalysis [321, 310, 301, 308, 307, 1809, 1804, 2276, 1802, 1795, 1801, 1800]. Cryptographic [2082, 720, 1788]. Cryptographical [210]. Cryptography [2262, 2273, 1805]. Cryptology [2267]. Cryptosystem [151]. Cryptosystems [318, 304]. CSCW [1071, 2109]. CSD [937]. CSP [955, 2056]. CSPs [966]. CSYN [925]. CT [2212, 337]. CTR [286]. Cube [423, 2176]. Cubic [127, 2007]. Cultural [1346, 1350]. Curriculum [2078, 2129]. Currying [143]. Cursive [875, 2197]. Curvature [803]. Curve [886, 1124, 582, 291, 2265]. Curved [802, 1532, 1535]. Curves [698, 851, 581, 861]. Custom [356, 913, 934]. Customised [915]. Customizable [1288]. Customization [1761, 1762]. Cut [2304, 1111]. Cutting [1149, 81]. Cutwidths [2001]. Cycle [1732]. Cycles [391, 165, 2090]. cyclotomic [1786]. Cylinder [2308].

D [821, 2213, 361, 602, 584, 597, 833, 588, 385, 1150, 2170, 2173, 2226, 2212, 1539, 1535, 2180, 685, 2176, 653, 332, 806, 600, 858, 2202, 1326, 2219, 1529, 590, 934, 2181, 596, 840, 2179, 664]. DAD [2279]. DAG [1048, 393]. Dags [1097]. Data [776, 2123, 1040, 338, 1336, 323, 1632, 418, 808, 2046, 174, 1062, 1216, 2099, 577, 1775, 1045, 439, 959, 2102, 2010, 464, 410, 430, 749, 26, 50, 1948, 1300, 805, 365, 1072, 843, 141, 806, 2116, 1192, 2117, 467, 1782, 2187, 1766, 351, 895, 436, 1770, 349, 405, 1209, 2164, 1360, 238, 939, 1847, 1852, 2128]. Data-Flow [895]. Data-Parallel [439, 464, 1209, 2164, 418]. Data-Parallelism [410, 467]. Data-Types [805]. Database [2082, 1075, 1765, 1068, 1026, 1038, 1260, 1064, 471, 2072, 1961, 1454, 1045, 1057, 1279, 1280, 1608, 725, 2121, 2095, 1034, 763, 1904, 2110, 766, 1036, 1277, 1777, 134, 1925, 1939, 350, 1766, 1456, 1906, 1261, 1329, 1037, 443, 764, 1770, 1607, 1945, 2113, 1069, 1065, 2092, 2081, 1269, 791, 2086, 1691, 1964, 1941, 1834, 2136, 1345, 995]. Databases [1976, 1281, 1356, 1965, 1278, 1048, 1271, 1021, 2088, 1917, 1473, 1272, 1265, 958, 1992, 2112, 1959, 1039, 2093, 1966, 1201, 1267, 1957, 1035, 133, 1282]. Databdrige [765]. Dataflow [1202, 344, 1237]. Datalog [111, 1275, 1944, 1958, 960]. Dataparallel [343]. Date [915]. Dating [2245]. Davis [981]. DB [1954]. DBMS [1767, 1022, 2121, 762]. DBPL [1266]. DCE [1603]. DCT [821]. Deadlines [1716]. deadlock [2056]. Dealing [1489, 1474]. Debate [406]. Debugging [1204, 1272, 963]. Decentral [784]. deception [73]. Decidability [272, 287]. Deciding [720, 2071, 1727]. Decimation [680]. Decision [1703, 540, 719, 563, 2073, 1032, 1177, 1483, 778, 1898, 658, 2260, 1157, 1523, 2074, 1866, 1578, 1836, 1577, 1837]. Decision-Making [2074]. Declarative [788, 2084]. Declustering [1035]. Decoding [2260, 2256, 2252]. Decomposable [525, 574]. Decomposing [851].

Decomposition [2005, 2004, 1142, 176, 347, 436, 2236]. Decrease [275]. Decreasing [506, 513]. Dedicated [724, 910, 1625]. Deduction [1824, 2050]. Deductive [1746, 1946, 1948, 1959, 117, 1945, 1957, 1964, 1956]. Deep [1949, 732]. Default [1481, 1169, 1439]. Defeasible [1433, 1169]. Defeat [1403]. Defects [689]. Definability [958]. Define [91, 1555]. Defined [1953]. Defining [484]. Definite [12]. Definition [1500, 1028]. Definitions [715, 1553, 727, 1484]. Defocus [617]. Deforestation [1191]. Deformable [593]. Deformations [1545]. Deformed [2176]. Degree [1708, 63, 162, 1710, 547]. Degrees [1110]. Delaunay [799, 621, 1103]. Delay [1086, 2041, 919, 2040]. Delay-Sensitive [2040]. Delayed [285]. Delays [51, 1020, 50]. Delegation [372]. Delegation-Based [372]. Deletability [2181]. Deletions [2184]. Delineation [841]. Delivering [415]. Delivery [2025, 1380, 1714]. Delta [1278, 1325, 168]. Delta-Calculus [1325]. Demand [1122, 2039, 2044, 2047, 413]. Demand-Driven [413]. Demands [1355]. Demon [971]. Demonstrator [1622]. Dempster [616]. Dempster-Shafer [616]. Denormalisation [2119]. Denotation [717]. Denotational [267, 1262]. Dense [1705, 609, 155]. Density [922]. Dependability [369]. Dependencies [961, 1874, 1854]. Dependency [1477, 2122, 1393, 1132]. Dependent [426]. Deployment [1584]. Depository [1046]. Depth [243, 674, 153, 1090, 2214, 626, 2196]. Derandomization [541]. Derivation [1395]. Derivations [514]. Derivative [2235]. Derivatives [701]. Deriving [2163]. Described [1241]. Describing [885, 870, 1303]. Description [1166, 1336, 874, 1491, 1176, 1457, 1314]. Descriptions [1397, 1320, 1208, 1503, 1309, 1307].

Descriptor [19]. Descriptors [694]. Design [1626, 1512, 1765, 1968, 345, 889, 1681, 363, 1938, 2045, 1068, 498, 470, 1449, 2019, 1252, 1186, 473, 1590, 922, 1253, 1817, 1918, 1071, 926, 2121, 2109, 2063, 2110, 1023, 1777, 999, 1816, 2080, 1256, 2070, 1376, 2049, 1980, 2242, 469, 2077, 2107, 1047, 1652, 1029, 2092, 906, 1269, 2249, 1360, 1311, 1076, 930, 1655, 1846, 2060, 1837, 1875, 1850, 2061, 1860, 1842, 1345, 1306]. Designer [1279]. Designing [1968, 780, 1384, 1972, 104, 1592, 2219, 1747]. Designs [305, 2250, 154]. Detecting [671, 226, 654, 437, 1451, 2156]. Detection [2225, 689, 1940, 670, 1053, 949, 580, 904, 608, 2209, 760, 830, 419, 899, 701, 652, 891, 853, 2180, 613, 611, 815, 2202, 913, 447, 848, 852, 761, 631]. Detectors [804]. Determinacy [282]. Determinants [1347]. Determination [2224, 809, 685, 626, 852, 1335]. Determining [1216, 2041, 876, 599]. Deterministic [554, 1089, 419, 233, 1007, 759, 259, 693, 2168, 1563, 1316]. Deterministically [1203]. Determinizing [2310]. DEVA [1823]. Developing [770, 1026, 1237, 1627, 939, 1844, 1841]. Development [2166, 1742, 1755, 1973, 903, 197, 1016, 1822, 1374, 916, 1895, 1597, 1606, 1813, 612, 1909, 1778, 1823, 2120, 1824, 79, 80, 2285]. Developments [1826, 782]. Device [546, 1806]. Devising [1999]. DEVS [1655]. DFS [359]. DFS-Super [359]. Diagnosing [978]. Diagnosis [1420, 1425, 1499, 2076, 1500, 2208, 1256, 1464, 1462, 2244, 1869, 1840, 1311]. Diagnostic [1498, 1857]. Diagnostics [1411]. Diagram [1703, 1523, 1100, 1103]. Diagrams [2312, 397, 34, 1969, 1577, 86]. Dialogue [1479, 1442]. Diameter [757, 1997]. Diametral [2008]. Dictionary [1723]. Difference [613, 2236]. Differences [859, 1350, 1569]. Differencing [803].

Different [851, 341, 1024, 1501, 1486, 2285, 1867]. Differential [1670, 609, 301, 687, 1791, 1800]. differentials [1798]. Difficult [1399]. Difficulty [65]. Diffident [1461]. Diffie [2249]. Diffusion [811]. Digital [1021, 1027, 323, 811, 1124, 2021, 1548, 4, 681, 2028, 935, 807, 918]. Digitized [697]. Digraphs [2144, 1085, 383, 1704, 165]. Dilemma [1405, 1037, 88, 77]. Dimension [327]. Dimensional [2266, 1102, 184, 31, 1749, 748, 1991, 495, 1087, 391, 1534, 688, 2232, 45, 1998, 401, 2001, 354, 624, 1149, 129]. Dimensions [1071, 34, 1982]. dining [2056]. Diophantine [951]. Direct [701, 12, 73, 702, 1969]. Directed [1739, 61, 1201, 1439]. Direction [1968, 803, 1901]. Directional [180]. Directionals [1347]. Directions [854, 769]. Directory [429]. Directory-based [429]. Disabilities [1628]. Discipline [1821]. Disciplines [1382]. Disconnected [2316]. Discourse [7]. Discourses [8]. Discovery [1453, 1775, 2118]. Discrete [1629, 866, 370, 1632, 1004, 665, 347, 933, 1538, 997, 664]. Discrete-Time [997]. Discriminant [658]. Discriminating [673]. Discriminator [864]. Discs [636]. Disease [832]. Dishonest [313]. Disjoint [1704]. Disjunction [971]. Disjunctions [159]. Disjunctive [1049, 1201]. Disk [2048, 443]. Disks [1300, 155, 1077]. Disparity [860, 700]. Displacements [587]. Display [2195]. Disposal [2148]. Dissecting [746]. Distance [630, 858, 402, 612]. Distances [1324]. Distinction [1159]. Distinctive [1532]. Distinguishing [572]. Distorted [696]. Distributed [140, 1755, 754, 355, 1769, 1771, 1027, 1829, 1144, 721, 360, 1519, 409, 2026, 1026, 739, 1011, 1613, 743, 1115, 434, 2013, 97, 1141, 1358, 1905, 331, 1669, 1017, 2155, 1920, 1693, 346, 1978, 29, 395, 1596, 463, 742, 708, 2015, 1754, 1366, 1601, 104, 1294, 1661, 1926, 1292, 1290, 429, 1014, 1683, 746, 1945, 447, 442, 349, 1778, 753, 750, 2292, 1148, 261, 1615, 1042, 2156, 956, 2182, 193, 2285, 2061, 2316]. Distributed-Memory [708]. Distributing [362, 567]. Distribution [54, 594, 2250, 433, 297, 2269]. Distributions [170, 2220, 171, 436]. Diversity [604]. Divisible [322]. Divorce [2074]. DNA [670, 1140, 1571]. DNA-Content [670]. Do [170, 1395, 29, 1744, 2279]. DOACROSS [1130]. Document [879, 880, 2246]. Documentation [1429]. Documents [873]. Does [785, 1763]. DOM [1866]. DOM-ArC [1866]. Domain [1055, 1491, 97, 654, 1618, 1382, 1444, 1259, 972, 1843, 1597, 1570]. Domain-specific [1382]. Domains [1134, 973, 1819, 983, 1854]. Dominant [2236]. Dominating [161, 183, 55]. Domination [1707, 384]. Done [981]. DOOD [1962, 1967]. Dorsch [2260]. Dot [1951]. Down [2118, 2285]. Draw [126]. Drawing [127, 1349]. Drawings [45, 2247, 2197]. Drippers [689]. Drive [943]. Driven [1072, 2148, 413, 2175]. DSP [912]. DSP-Boards [912]. DSS [319]. DTS [1144]. Dual [929, 656, 199]. Duality [411]. Dummy [388]. Duplicating [521]. Duration [2167, 371]. During [1332, 242]. Dyck [33]. Dynamic [1121, 703, 966, 886, 2037, 2022, 184, 771, 1985, 2098, 2021, 98, 434, 577, 33, 2085, 1253, 2068, 1095, 1163, 1097, 1041, 1893, 1073, 1334, 2195, 1779, 1326, 454, 335, 2036, 1133, 1267, 24, 1284, 1368, 1283, 1876]. Dynamic-Resharing [1121]. Dynamically [1633, 946, 112, 947, 1747, 353]. Dynamics [2091, 1489, 1141, 531]. dys [2140]. dys-functional [2140].

each [637]. Eagon [668]. Early [369, 1618]. Easier [388]. Easy [1092, 1718, 2070].

Eavesdroppers [750]. ECA [1272, 1266]. ECA-Rules [1266]. Echo [2220]. Echocardiographic [831]. Ecologies [1519]. Economic [197]. Economical [1697]. ECRC [990]. Edge [1053, 168, 1115, 1667, 815, 852, 146]. Edge-Coloring [1667]. Edge-Colorings [146]. Edge-Face [168]. Edges [60, 853, 804]. Edit [402]. Editing [818, 711, 683, 2184]. Editor [2185]. Editors [1969]. Education [930]. Educational [1972]. EES [1806]. EF [210]. effect [70]. Effective [380, 391, 1224, 122, 1592, 2117, 1615, 1113]. Effects [1332, 955, 1893, 1172, 1171]. Efficiency [54, 1935, 70]. Efficient [1121, 1214, 1626, 1134, 1084, 379, 329, 435, 2046, 798, 127, 760, 2158, 1419, 2004, 1720, 156, 1164, 276, 2055, 434, 1562, 630, 1896, 1953, 1193, 39, 294, 546, 430, 131, 542, 1943, 312, 192, 1826, 2145, 446, 1922, 1192, 1227, 1661, 2204, 322, 581, 1100, 981, 421, 2282, 2268, 465, 2315, 1113, 736, 1076, 1964]. Efficiently [159, 1983, 1942]. Effort [1380]. EGI [1549]. Egomotion [709]. Elastic [703, 819]. Elasticity [797, 591]. Election [759]. ELECTRE [1649]. Electrocardiographic [667]. Electronic [903, 322, 2263]. Element [1538]. Elementary [563, 557, 2124]. Elements [356, 354]. Elimination [388, 804, 2053]. ELLA [1314]. Elliott [952]. Elliott-MacMahon [952]. Ellipses [863]. Elliptic [291, 2265]. Elliptical [650]. Elongated [841, 663]. ELS [1943]. ELS-Programs [1943]. Embedded [1987]. Embedding [732, 1305]. Embeddings [139, 423, 1729, 421]. Emedded [940]. Emerge [90]. Emergence [1346]. Emergent [96, 92, 75]. EMIR02 [1058]. Emotion [2226]. Empirical [14, 1768, 120, 1862]. Empirically [1175]. Empty [2308, 160]. Emulation [422, 1020]. Enaction [1358]. Encoding [89, 1310]. encountered [1304]. Encryption [305, 301, 1789, 1811]. End [2025, 2041, 2057]. End-to-End [2041]. ended [84]. Energy [688, 2202, 1797]. Engine [1274, 909]. engineer [1806]. Engineering [1910, 1674, 498, 1833, 783, 1918, 2121, 266, 1017, 1596, 2015, 497, 1037, 1485, 1371, 1065, 1610, 1362, 2086, 1000, 2247, 2103, 2101, 1883]. Engineers [1355]. Engines [428]. English [1844, 2074]. Enhance [1452]. Enhanced [1078, 2093]. Enhancement [857, 606, 625]. Enlargement [857]. Entailment [1169]. Entity [1992, 2115, 2096, 2093, 2106]. Entity-Relationship [1992, 2115, 2096, 2093]. Entries [378]. Entropy [1511]. Enumerating [150]. Enumeration [61, 1309]. Environment [1742, 1067, 113, 1359, 1758, 1441, 1016, 1753, 1822, 344, 1041, 346, 1024, 1015, 1596, 1825, 947, 1606, 1926, 980, 1456, 1770, 1326, 1526, 1345]. Environmental [585, 1033]. Environments [1470, 1396, 1978, 463, 1334, 1614, 781, 1014]. Epipolar [702, 1533]. Epistemic [1405]. Equal [155, 480]. Equality [1949, 2145]. Equality-Testing [2145]. Equation [505]. Equational [377, 501, 2290]. Equations [951, 361, 399, 364, 1463, 241, 353]. Equivalence [159, 252, 172, 714, 2161]. Equivalences [263, 1727]. Equivalent [513, 2083, 1117]. Erlang [1297]. erroneous [1571]. Error [2278, 226, 932, 1311]. Error-Correcting [2278]. Errors [2281]. Esablishment [2259]. Escrow [305, 303, 302, 734]. ESP [2169]. ESP-MC [2169]. Essential [1031]. Establishment [2038, 2121, 2175]. Estelle [334]. Estelle-specified [334]. Estimates [700]. Estimating [698, 674, 658, 618, 2077, 1770, 587, 2201]. Estimation [860, 810, 609, 1631, 669, 1546, 863, 2224, 607, 617, 709, 2199, 2232, 2221, 803, 1543, 635, 702, 644, 900, 1835].

ET-LOTOS [267]. ETF [1138]. Ethics [1251]. Euclidean [2309]. Eulerian [1704]. Eurocrypt [307]. Evaluate [637]. Evaluating [1055, 811, 1850, 1354]. Evaluation [337, 1278, 1240, 2305, 2318, 995, 1205, 363, 1038, 1590, 1773, 1195, 1018, 763, 342, 2023, 1943, 1255, 346, 2080, 479, 1928, 1607, 1774, 994, 1210, 1964, 682]. Evaluations [1239, 1059, 1185, 393, 999, 1382]. Evaluator [1034, 1194]. Even [2091, 1563]. Event [346, 1815, 259]. Event-oriented [346]. Events [1939, 2089]. Evolution [894, 577, 1395, 2102, 1376, 1399, 74]. Evolutionary [1519, 1587, 780, 2063, 1517, 1673, 1524, 76, 85, 83]. Evolving [1910, 1039, 1490, 84, 88]. Exact [1004, 190, 1805]. Example [1407]. Examples [227, 1429, 241, 1250, 1503, 1563, 1564]. Exceptions [1505, 1445]. Exceptions-Based [1505]. Exchange [1551, 47, 2015, 291]. Exclusion [140]. Executable [1207]. Execution [1129, 409, 1262, 98, 420, 1372, 344, 2017, 1226, 2137]. Executions [914]. Exemplification [1484]. Exercise [752]. Exhaustive [1672]. Exist [29]. Existential [24]. existing [1305]. Exits [378]. Experience [1911]. Experiences [2130, 1597, 1825, 69, 1762, 1760]. Experiment [2169, 317, 1399, 1872]. Experimental [1240, 1590, 838, 420, 683, 1945, 1662, 77]. Experimentation [189, 2066]. Experiments [716, 1812, 54, 482, 246, 448, 1860]. Expert [1056, 1082, 1054, 1927, 1412, 1422]. expertise [1841]. Experts [1455]. Explained [1240, 1241]. Explaining [2132]. Explanation [1272, 1161, 1422, 1877]. Explanation-Based [1161]. Explicit [1916, 1418, 1181, 653, 1560]. exploit [1805]. Exploitation [1584, 578]. Exploiting [1185, 428, 1452]. Exploration [1061]. Exploratory [1378]. Exploring [493, 1326, 1527]. explosion [2062]. Explosive [317]. Exponential [528, 537, 2295]. Express [351]. Expressing [206]. Expression [2225, 2305, 511, 1206, 393, 300, 2227, 877]. Expression-Or-Parallel [1206]. Expressions [1205, 257, 1070, 133]. Expressive [2300, 220, 2288]. Expressiveness [287]. ExSel [589]. Extended [671, 1435, 2281, 566, 1889, 2157, 1058, 2096, 1066, 1654]. Extending [1469, 594, 1906, 1266, 152, 912, 486]. Extensibility [1558]. Extension [1084, 487, 1052, 2090, 1161, 2119, 352]. Extensional [1444, 1480]. Extensions [1193, 2294, 762, 1463, 2321]. Exterior [2199]. External [1104, 1346, 1621]. External-Memory [1104]. Extract [901, 589]. Extracting [873, 636]. Extraction [2231, 827, 204, 1062, 1478, 662, 2190, 1065, 687]. Extrapolation [1010]. Extremal [1730]. Extremely [1735]. Eye [2171].

F [1568, 1228]. F-Horn [1568]. Face [168, 595, 631]. Facial [2227, 820]. Facilitating [1628]. Facilities [2085, 1031, 1747]. Factorization [1147]. Factors [1381, 59]. Fading [2266]. Failure [1161, 2156]. Fair [270, 91, 304]. False [804]. Falsifiability [543]. Families [851]. Fantasy [1546]. Fast [2225, 1129, 292, 889, 49, 1730, 665, 1089, 1186, 383, 1095, 821, 50, 1697, 1088, 1125, 890, 1726, 933, 288, 290, 1155, 291, 1723, 178, 934, 1289, 2178, 900, 664, 1807]. Faster [1738, 61, 761, 269]. Fastmap [908]. Fat [1101, 1087]. Fat-Tree [1087]. Fault [944, 2076, 141, 456, 996, 1997, 1840]. Fault-Free [456]. Fault-Tolerance [944]. Fault-Tolerant [141, 1997]. Faults [754, 457, 1715, 456, 1869]. Faulty [1089, 2002, 456]. Favourite [1768]. FD [983]. FEAL [1801]. Feasibility [200]. Feasible [202, 1111]. Feasibly [215]. Featherweight [441]. Feature [676, 2229, 658, 660, 2193, 598, 242, 692, 2202, 1369, 1875, 1856]. feature-based [1875]. Features [337, 2231, 1622, 802, 1260, 1475, 662, 1330, 881, 696, 1073, 817, 850, 820, 877, 1863]. Federated [2114, 1770, 2113]. Feedback [1709, 1796, 1647]. FEM [1142]. FEM-Meshes [1142]. FERSA [2248]. Few [1984, 60]. FFTs [455]. Fiberings [117]. Fibonacci [1808, 327]. Fibrational [256]. Fibre [825]. Field [1134, 1493, 2214, 585, 2176, 615]. Fields [112, 932, 844, 1717, 1796]. File [2045, 1041, 348, 350, 1774, 2316]. Files [1773]. Filling [812]. Film [711]. Filter [936, 852]. Filtered [859, 803]. Filtering [1426, 893, 1424, 2235, 625]. Filters [707, 816, 815, 937]. Final [1557, 2318]. Financial [1898, 2120]. Find [835]. Finding [1705, 2272, 1736, 697, 1896, 1706, 1733, 1920, 1983, 820, 137, 178, 1103, 1731]. Fine [1821, 340, 466]. Fine-Grain [1821]. Fine-Grained [466, 340]. Fingerprinting [323]. Finitary [251]. Finite [272, 326, 540, 289, 722, 218, 192, 932, 1643, 620, 531, 401, 223, 354, 1717, 135, 133, 1784, 2059, 1796, 1570, 228]. Finitism [200]. Finsler [1643]. FIR [936, 937]. Firing [561, 485]. First [1239, 535, 1450, 958, 1737, 221, 1174, 468, 1828, 23, 804, 1475, 2127]. First-Order [535, 958, 221, 23]. Fish [1882, 2171]. Fish-Eye [2171]. Fitness [1513]. Fits [681]. Fitting [650]. Five [162]. Fixed [1644]. Fixpoint [254, 253]. Fixpoint-Analysis [253]. Flash [1300]. Flash-Memory [1300]. Flat [655]. Flexible [1552, 432, 1847]. Floating [726]. Floodlight [30]. Floodlights [1734]. Floor [784]. Floundering [1199]. Flow [361, 669, 586, 1720, 1220, 1223, 1224, 2043, 1926, 895, 1879]. Flows [2022, 2108, 1094]. Fluid [361, 1141]. Fluidity [95]. Fluorescent [670]. fly [1293]. Focus [626]. Fold [2185]. Following [1942]. Font [872]. Force [2175]. Force-Driven [2175]. forecasting [1879]. Foreground [841]. Forests [1741, 2003]. Forgeries [2183]. Form [874, 1062, 521, 861]. Formal [487, 328, 1314, 1908, 1821, 1491, 498, 470, 1363, 1833, 271, 473, 724, 420, 495, 1373, 483, 1374, 729, 3, 1832, 731, 472, 399, 1261, 475, 450, 1931, 239, 1824, 1306, 1308]. Formalising [2064]. Formalism [1454, 1372, 1351]. Formalisms [1675]. Formalization [1560, 1823, 738]. Formalizing [2165, 1499, 2155, 1392]. Formally [1305]. Format [482, 876]. Formation [94, 2185, 93]. Formations [894]. Formats [1338]. formedness [522]. Forms [327, 648, 869]. Formula [998]. Formulae [572]. Formulas [543, 142, 545, 404, 2055, 1568]. Formulation [1134, 1156, 728]. Formulations [2153, 1675]. Fortran [338]. Forward [968, 1037]. foundation [1308]. Foundations [773]. foundedness [509]. Four [317, 838, 1724]. Four-Vertex-Connectivity [1724]. Fourier [347, 700, 1545, 933, 934]. FPGA [920, 922, 931, 927, 936, 910, 907, 935, 913, 912, 934, 906, 930]. FPGA-Based [920, 935, 912, 934]. FPGAs [938, 914, 942, 926, 924, 923, 919, 933, 937, 918]. FPLD [932]. Fractal [890, 888]. Frame [1434, 2044, 2148, 2247, 1025]. Frame-Based [1025]. Frames [1431, 702]. Framework [1203, 2284, 1336, 1205, 1363, 474, 1476, 2099, 2098, 98, 1395, 1474, 2215, 967, 1384, 1893, 472, 1543, 707, 440, 587, 1650, 807, 103, 589, 1655, 1489, 87, 1797]. Free [554, 563, 386, 758, 1919, 556, 741, 1226, 2154, 469, 456, 55, 1114, 2062, 1797]. Free-Word-Order [1919]. Frege [211]. French [489]. Frequency [1129, 654, 396, 2242]. Fresh [1383]. Friendly [1555]. Front [2203, 2057]. front-end [2057]. Frontier [527]. FTAPE [996]. Fully [257, 1095, 1097]. Fully-Dynamic [1095]. Function [715, 525, 1240, 205, 1124, 1725, 2275, 775, 1852, 70]. Functional [1203, 1217, 412, 1180, 1220, 1221, 725, 464, 1285, 1222, 1206, 2131, 2136, 1375, 1184, 1196, 440, 1266, 1286, 405, 1233, 1550, 1887, 2129, 1212, 1197, 2140, 2132, 2130, 2125, 2137, 2128, 2135, 2124]. Functionality [1261]. Functions [1505, 289, 161, 321, 1173, 159, 2098, 387, 1167, 1073, 1726, 1555, 302, 1523, 288, 733, 1894, 1931, 1787, 1788, 1570, 1791, 1800, 1810]. Fundamental [1673]. Fundamentals [1923]. Further [357]. Fusion [1928, 644, 1776]. Future [483, 1591, 1602, 1614, 2262, 1355, 1763, 1620]. Fuzzy [1469, 814, 2238, 1470, 1890, 2200, 1634, 1032, 834, 1896, 1057, 1898, 652, 1900, 1892, 772, 1886, 1888, 1889, 1895, 1924, 1927, 1891, 1887, 1885, 898, 1689, 1879, 1894]. Fuzzy-Neuro [814]. FuzzyBase [1057].

G [168]. G01 [1143]. GA [1465, 1513, 1525, 1897, 1398]. GA/TS [1398]. Gabor [859]. Gadgets [1196]. Gaining [1024]. Galaxy [897]. Game [1153, 1877, 84]. Games [219, 248, 124, 1870, 77]. Garbage [1295, 1291, 1285, 1293, 1296, 1287, 1294, 1192, 1298, 1290, 1286, 753, 1297]. Gas [1522]. GASBOR [82]. Gauss [2053, 1717, 664]. Gaussian [1887]. GC [1292]. GENA [1236]. General [2000, 316, 1156, 1442, 952, 110, 1132, 1958, 179, 589, 1042, 1881]. Generalisation [612, 1331, 1330]. Generalised [1699, 2149, 2274]. Generalization [13, 1384, 1161, 1979]. Generalized [2312, 149, 1488, 539, 2184, 888, 1997, 1677]. Generated [836, 165]. Generating [1236, 1193, 302, 1969]. Generation [1214, 1406, 987, 1217, 1002, 1453, 1608, 431, 940, 1467, 1334, 6, 1468, 1429, 2268, 9, 974, 1870]. Generational [1295, 1286]. Generator [2296, 2057]. Generators [2253, 1808, 1784, 1786]. Generic [99, 1177, 1374, 1815, 1445, 590, 1542]. Genetic [1510, 1896, 1898, 941, 81, 683, 809, 1893, 1895, 1501, 1523, 787, 1151, 1527, 1518, 1521, 1511, 1894, 878, 1526, 71, 2160, 72, 78, 82, 89, 73, 79, 80]. Genetic-Fuzzy [1898]. Genome [1668, 1695]. Geodesic [1100]. GEOFF [2185]. Geographic [1104, 1030, 2241, 1329, 1328]. Geographically [1755, 1326]. Geographically-Distributed [1755]. Geography [1321]. Geological [1055]. Geometric [1671, 1102, 54, 1702, 1099, 27, 39, 69, 106, 405, 664]. Geometrical [505, 610, 2185, 578]. Geometry [653, 1537]. Geons [840]. Gestalten [1248]. Gesture [2223]. GF [932]. Gibbs [2220, 585]. GIMS [1914]. Ginjoshu [1892]. GIS [1344, 1040, 1914, 1328]. Given [2268]. Glasgow [1234]. Glia [1402]. Global [760, 965, 1131, 954, 708, 1293, 1294, 437]. Goal [1489, 1201, 1439, 1854]. Goal-Directed [1201]. Goals [101, 1242, 1395]. Goedel [492]. Going [36, 1899]. Gollmann [299]. Good [10, 227, 1003, 2261, 104]. GoslerP [247]. Gossiping [391]. GP [70]. Gradient [853, 2157]. Grading [885]. Grain [1821]. Grained [466, 340]. Gram [1106]. Grammar [1371]. Grammars [554, 1414, 386, 1413, 1178, 1957]. Grammatical [1561]. Granularity [868]. Graph [1995, 177, 1990, 283, 798, 363, 139, 55, 1986, 1985, 1781, 1164, 126, 1136, 1739, 12, 61, 1992, 44, 2192, 547, 551, 1989, 558, 1371, 1719, 137, 562, 1713, 1579]. Graph-Theoretical [1992]. Graphical [1769, 988, 488, 2112, 806, 1196]. Graphics [873, 2170]. Graphs [10, 47, 1984, 1709, 1270, 1477, 63, 161, 2011, 127, 381, 16, 168, 1707, 1741, 2003, 180, 1732, 19, 870, 390, 2008, 2283, 1987, 384, 162, 11, 18, 17, 1098, 799, 526, 1740, 2009, 2010, 1710, 1095, 22, 190, 1114, 1982, 1132, 6, 1093, 2006, 21, 561, 2282, 24, 1997, 9, 145, 23, 20, 619, 738, 89]. GRASP [1151]. Gray [630, 878]. Gray-Scale [878]. Great [501]. Greedy [1710]. Grey [640, 854, 654, 604]. Grey-level [854]. Grey-Scale [604]. Grid [127, 1732, 634, 1520]. Grids [1141, 391, 2177, 421]. Gröbner [116, 176]. Ground [518, 1404]. Group [180, 665]. Grouping [582, 2187, 1267, 578, 800]. Groups [2037, 111, 2224, 165, 578, 1784]. Groupwork [784]. Grow [1419]. Grow-And-Learn [1419]. Growing [584, 385]. Growth [1991]. Guarantees [2043]. Guarded [971, 1553, 969]. GUI [1978]. Guidance [704]. guide [2052]. Guided [2174, 1081, 236, 1860]. Guidelines [591]. GUIs [1197]. Gupta [1484]. Gupta-Belnap [1484].

Habits [2117]. Hair [2225]. Halftoning [867]. Halldorsson [186]. Hamilton [47]. Hamiltonian [1732, 165]. Hamming [423, 148]. Hand [2223]. Handle [1504, 112]. Handling [1758, 2111, 459, 1779, 1521, 1369]. Hands [2222]. HandTalker [2223]. Handwriting [881]. Handwritten [2238, 882, 2193, 1937]. Harari [2276]. Hard [1718, 402, 2295, 132, 2003]. Harder [170]. hardness [2315]. Hardware [1910, 917, 941, 915, 927, 916, 1298, 366, 918, 1314, 1304, 1308, 1315]. Hardware-Assisted [1298]. Harmony [1450]. Hash [2275, 1814, 288, 1800, 1810, 1313]. Hashing [290]. Haskell [491, 1234]. Hasse [1011]. Haystack [1672]. HCTRs [513]. HDL [922]. HDL-Based [922]. Head [828, 675, 546, 632]. Headed [1077]. health [1841]. Heap [1722]. Hearing [1506]. Heart [624]. Hedge [546]. Hedonic [102]. Helical [2208]. Hellman [2249]. Helmholtz [364]. Help [485, 1250, 2279]. Herbrand [209]. Heterogeneous [1758, 2041, 684, 463, 444, 2048, 1955]. Heuristic [2302, 1070, 1246, 848]. Heuristics [957]. Hidden [640, 1631, 321, 891]. Hierarchical [860, 2198, 1004, 1339, 739, 508, 1637, 1015, 424, 1457, 1257, 2204, 2056, 1445, 1894, 1651, 1878]. Hierarchically [1008]. Hierarchies [203, 528, 2309, 152, 213]. Hierarchy [2311, 665, 922, 172, 449, 1348, 395, 214]. High [1708, 327, 889, 884, 1912, 830, 1890, 2200, 922, 630, 915, 2172, 2020, 849, 913, 623, 78, 1787, 1305]. High-Level [1912]. High-Order [884]. High-Performance [915, 2172]. High-Speed [913]. Higher [1916, 680, 1227, 1184, 998, 733, 34, 1212, 1798]. Higher-Order [1916, 1227, 1184, 1212]. Highly [671, 141, 2117, 1550, 1997]. Highways [1776]. Hillclimbing [1518]. Hints [499]. Histogram [44, 613]. Historical [1326]. History [1917, 1299]. History-Based [1299]. History-Oriented [1917]. Hive [1986]. HLS [820]. HMM [881]. HOL [716, 719, 720, 726, 730, 714, 1556, 735, 737]. HOL90 [725]. HOLCF [733]. Home [36]. Homogeneous [442]. Homogeneously [2011, 384, 1456]. homophonic [1807]. Honest [313]. Honeycomb [548]. Hong [152]. Hong-Kung [152]. Hopfield [795, 817, 1932]. Horn [1568]. Hot [1998, 1107]. Hot-Potato [1998, 1107]. Hough [1124, 902, 2237]. Housing [2073]. HT [671]. Hull [812]. Human [675, 1428, 1918, 1333, 1176, 2070, 1354, 2227, 624, 631, 2222]. Human-Centered [1428]. Human-Computer [675, 1918]. Human-Subjects [1354]. HYbrid [1642, 1631, 1633, 1636, 1635, 1496, 1637, 1638, 1639, 1640, 1641, 1886, 1643, 1645, 424, 2074, 1646, 1648, 1652, 822, 1653, 1654, 1655, 1656, 1398, 79, 80]. Hydra [2135]. Hydrographic [2241]. Hyper [1729]. Hyper-Rings [1729]. Hypercube [1129, 1997]. Hypercubes [1134, 754, 1729, 456]. Hypergraph [1995, 78]. Hypergraphs [2005, 331]. Hypermedia [1976, 647, 1602, 1694, 1979]. Hyperrectangles [2146]. Hypertext [1975, 474, 1047]. Hypertexts [1063]. HYTECH [1642, 2052]. HyTime [1045].

I/O [53, 1703, 1067, 54, 2019, 2050, 1556, 943]. I/O-Algorithms [53]. I/O-automata [2050]. I/O-Complexity [1703]. IDA [1076]. Ideal [116, 2127]. Ideals [374]. Ideas [735, 1665]. IDEFIX [776]. IDEFO [775]. Identical [1727]. Identifiable [320]. Identification [828, 2084, 231, 181, 593, 2276, 872]. Identifying [1726, 1583]. Identity [1480, 2264, 2113]. if [1301]. IFO2 [1065]. Ignorance [1405]. IGOR [1237]. II [597]. III [915]. Illuminance [2205]. Illuminated [865]. Illuminating [31]. Illumination [1734, 30, 713]. Illustrative [491]. Image [647, 813, 802, 894, 676, 798, 608, 884, 584, 666, 830, 2220, 942, 819, 657, 667, 826, 831, 641, 678, 2216, 643, 805, 708, 2232, 890, 2221, 857, 598, 1058, 1925, 467, 880, 849, 1039, 600, 645, 816, 888, 895, 694, 892, 1774, 901, 807, 359, 794, 1399, 2201, 856, 603, 796, 898, 2245, 625, 682, 693, 855, 1838]. Image-Intensity [666]. Image-Processing [813]. Images [640, 827, 836, 828, 609, 893, 854, 795, 1062, 834, 651, 662, 630, 2206, 585, 705, 696, 2208, 2212, 865, 824, 845, 859, 2180, 646, 655, 896, 844, 2234, 2202, 846, 793, 2210, 878, 604, 822, 825, 599]. Imaging [2217, 644, 900]. Imai [307]. Imitate [2070]. immune [1851]. Immunized [310]. Impact [1380, 1602, 823]. Impacts [1614]. Impediments [271]. Imperative [337]. Imperfectly [1240, 1241]. Implement [1497]. Implementation [1622, 1916, 16, 2045, 419, 788, 1186, 1946, 18, 387, 1182, 1773, 1603, 959, 1206, 948, 931, 821, 1023, 367, 932, 364, 69, 686, 933, 1947, 981, 447, 1930, 934, 1029, 735, 1269, 1210]. Implementations [1067, 1180, 1234]. Implemented [938]. Implementing [1087, 490, 432, 354, 2249]. Implicational [543]. Implications [1338, 2249, 135]. Importance [2214, 790]. Imprecise [2230]. Improve [357, 1852, 1853]. Improved [67, 1270, 312, 186, 982, 642, 1800, 644, 1313]. Improvement [1229, 888, 790, 1360]. Improving [875, 1342, 2067, 656, 478, 300, 1517]. Improvise [1359]. Imputation [1060]. In-Place [847, 1721, 542, 58]. inappropriate [1581]. Including [1429]. Incomplete [1423, 1849]. Inconsistency [1933]. Inconsistent [2114]. Incorporating [1032]. Increase [275]. Increasing [1431, 369]. incredible [1872]. Incremental [2069, 1986, 1380, 131, 397, 1922, 574, 2287, 1567, 1249, 1370]. Independence [128, 532, 1956]. Independent [63, 1725, 2223, 1710, 147, 541, 972]. Indeterminate [2288]. Index [324, 2066]. Indexing [2231, 886, 2198, 1186, 1696, 1871, 1880, 1839]. Indicator [832]. Indices [1247]. Indirect [1292]. Individual [2091, 1716, 1778]. Induced [1741, 2003, 1983]. Inducing [1162, 2133]. induct [2133]. Induction [1238, 1239, 1684, 254, 523, 1572]. Inductive [230, 247, 2165, 1566, 226, 245, 727, 2155, 1444, 1650, 1249, 120, 238, 2160, 1865, 1582]. Industrial [1968, 2028, 1355, 991]. Industry [483, 791, 930]. Inequalities [376]. Inequations [951]. Inescapable [1480]. Inexact [883]. Inference [230, 1189, 1433, 247, 229, 1220, 226, 730, 547, 237, 1212, 1169, 523, 1865, 2320, 1582, 1797, 1561, 1574]. Inferring [1571]. Infinite [371, 249, 133]. influence [86]. Informal [495]. Informatics [1695]. Information [1551, 1104, 1346, 1630, 1082, 1423, 1478, 1833, 2122, 1030, 234, 2114, 1905, 684, 1889, 2080, 1334, 1337, 806, 2190, 366, 1714, 1955, 1326, 1485, 1776, 1080, 774, 1690, 1000, 1422, 1328, 2318]. Informative [1690, 832]. Inherent [744, 2237]. Inheritance [1952, 1227, 485, 1951]. Inhomogeneously [865]. Initial [618]. Injection [996]. Innovation [271]. Input [426]. Input-Dependent [426]. INRECA [1865]. Insertion [704]. Insertions [2184]. Inspection [905, 902, 837, 2243]. Instance [2082, 1937, 1884]. Instance-Based [1937, 1884]. Instances [170]. Instead [2089]. Instruction [1131, 2078]. Instructional [1468]. Instrumentation [1362]. Integer [157, 960, 953, 928, 1117]. Integral [1889]. Integrated [2091, 495, 1373, 1504, 2121, 1589, 2105, 1010, 2033, 2077, 1412, 800, 1865, 82, 1857]. Integrating [1238, 1239, 1478, 2098, 1136, 1918, 1873, 1176, 1001, 1861, 2106, 1370]. Integration [1834, 1160, 1338, 1064, 1764, 2094, 4, 1948, 1978, 1024, 999, 2115, 849, 1585, 1246, 1929, 1624, 1846]. Integrity [2318, 1162, 1276]. Intelligence [1428, 1157, 2070]. Intelligent [1440, 343, 777, 1614, 1623, 1616, 1081, 87]. Intensity [666]. Intensive [1379]. Intentional [1444]. Intentions [1435, 1395]. Inter [1397, 1363]. Inter-Layer [1397]. Inter-level [1363]. Interact [1534]. Interacting [884, 1242, 75]. Interaction [1973, 1974, 675, 921, 1918, 1978, 96, 2066, 1619]. Interactions [1508, 1474, 1451]. Interactive [1203, 2039, 1359, 988, 311, 2067, 2226, 895, 2177, 105, 2060]. Interchange [482]. Interconnection [912, 451, 450]. Interface [2213, 647, 1974, 908, 1590, 1918, 488, 779, 1980, 1329, 1652, 1628, 735, 1885, 939, 1076]. Interfaces [988, 2067, 1196]. Interfacing [725]. Interleaved [1203]. Interleaving [98]. Interlocking [1416]. Intermediate [987, 109, 595, 622]. Internal [1346]. Internally [1698]. Internet [2027, 1020, 1694]. Internetwork [1771]. Interoperation [1025]. Interpolation [1342, 142, 2179]. Interpretation [1479, 1218, 2200, 834, 1442, 1372, 1641, 648, 122, 1228, 7, 645, 1680, 1211, 1235]. Interpreting [2191]. Interprocedural [1214]. Intersection [172, 160]. Interval [1996, 526, 748, 962, 1093, 1427, 1466]. Interval-Based [1427, 1466]. Intervals [371, 183]. InterViews [1970]. intractability [1805]. Intrinsic [208]. Introducing [1907, 2125, 1615]. Introduction [1618, 1904, 1762]. introspective [1871, 1853]. Invariant [639, 580, 1536, 630, 881, 579, 650, 713, 842]. Invariants [987, 1530, 2143, 843, 677, 673]. Inversion [644]. Inverting [691]. Investigating [1150]. Investigation [683, 1891, 89, 1864]. Invited [2124]. Involving [2184, 1353]. IOs [1598]. IRO [1954]. IRO-DB [1954]. Irregular [1145, 1016, 1143, 1126, 1130, 1132, 1133]. Irregularly [1129]. Isabelle [716, 728, 490, 1556]. Isabelle/HOL [1556]. ISDN [1627]. Isometries [888]. Isomorphic [549]. Isomorphism [1984, 19, 1739, 15]. Isophotes [629]. Issue [133]. Issues [360, 1938, 1635, 1468, 1585, 1047, 735, 2014, 1677, 1580]. Iterated [149, 88, 77]. Iterative [676, 1510, 1136, 2227, 349, 1006, 353].

JACK [1315]. Japanese [876]. Jitter [2034]. Job [983, 1398]. Job-Shop [983]. Jobs [32]. Join [2158, 1275, 2159]. Joint [673]. JTMS [1390]. Jump [1657, 1186]. Jumping [1506].

K-64 [309, 1799]. K-means [2211]. k-SR [2254]. kappa [155]. KBED [1053]. KBMSs [1078]. KBS [1843]. KDD [1059, 1580]. Kernel [1022, 1277]. Kernels [349]. Key [306, 1681, 310, 309, 2250, 303, 297, 302, 2259, 151, 307, 2263, 629, 291, 734, 2265, 2249]. Key-Schedule [309]. Keying [2277]. keystream [1808]. Kind [164]. Kit [491, 930]. KIV [1830]. KIV-Approach [1830]. KLIC [447]. Knowledge [1251, 747, 647, 1493, 1053, 1443, 1494, 362, 313, 1162, 1418, 1453, 1454, 2064, 22, 1163, 1892, 1455, 653, 1060, 1061, 1980, 1484, 897, 1456, 1498, 181, 1883, 1951, 898, 1956, 438, 1874, 2061, 1881, 1862]. Knowledge-Base [1956]. Knowledge-Based [1053, 1892, 1980, 438]. Knowledge-Bases [1951]. known [1794]. Kohonen [657]. Korean [1393]. Korso [1819, 1833]. Krylov [1930]. KSR [448]. KSR-1 [448]. Kung [152].

Labeled [2149]. Labeling [913, 1113]. Labelled [1436, 1437]. Laboratory [246, 790]. Lambda [1415, 1230, 2149, 1181, 1555, 1560]. lambda-Calculus [2149, 1181, 1560]. lambda-Terms [1555]. Landscapes [1510]. Language [337, 487, 1477, 338, 418, 1424, 1491, 362, 1030, 1946, 1180, 1822, 725, 464, 1206, 1150, 1176, 1377, 2112, 1467, 2083, 134, 1334, 1158, 469, 1028, 351, 1559, 1485, 1460, 1564, 1823, 1458, 224, 1316, 1756, 1297, 1391, 9, 1076, 1677, 1314, 2127, 2125, 1573]. Language-Free [469]. Languages [552, 1912, 277, 1217, 739, 1475, 1819, 33, 389, 220, 1640, 1285, 1222, 1919, 556, 1948, 1225, 1178, 2150, 2153, 1286, 181, 1233, 1932, 239, 1713, 1212, 263, 238, 235, 236, 1563]. Larchant [753]. Larchant-RDOSS [753]. Large [751, 1379, 2272, 317, 430, 545, 1840, 442, 1942, 1625, 1289, 1006, 2130]. Large-Scale [442, 1006, 1840]. Largest [2308]. Late [252]. later [1799, 1810]. Lattice [1974]. Lattice-Based [1974]. Lattices [1340]. Law [1728]. Layer [1397, 367]. Layered [1051, 1419]. Layers [2061]. Layout [2046, 874, 156, 2047, 2246]. Layouts [1120]. Lazy [1204, 1220, 1285, 1196, 1286, 1931]. Leader [759]. Leafary [131]. Leakage [314]. Leaks [1286]. Leapfrogging [2141]. Learn [1419, 1413, 1881]. Learnable [159]. Learning [1440, 232, 1565, 1505, 1569, 1471, 2069, 835, 1519, 14, 158, 1419, 1396, 1242, 228, 1441, 1871, 227, 233, 1896, 1577, 240, 1472, 1877, 1563, 1853, 1688, 1501, 242, 1568, 1570, 1447, 1525, 241, 1859, 1445, 1165, 1894, 1451, 1249, 631, 239, 1713, 1399, 120, 238, 225, 930, 236, 1937, 1503, 1843, 1469, 1578, 1634, 1579, 1562, 1573, 1567, 1564, 1842, 1884, 71, 244]. Learning-Goals [1242]. Least [681]. Leaves [2002]. Left [516, 10]. Left-Linear [516]. Legacy [1037]. Legal [1063]. Lemma [330, 1294]. Length [67, 212, 949, 2278, 1388, 1697, 353]. Length-Limited [1697]. Lenses [2171]. LESD [1381]. Lesions [829]. Less [1721]. Level [640, 1912, 2200, 654, 444, 892, 2024, 1069, 268, 854, 1363, 1305, 1193, 432]. Lexical [1163]. Libraries [939, 1870]. Library [1021, 1027, 988, 184, 805]. Life [2090, 1658]. Light [207, 2228, 1822]. Lighting [655]. Lightweight [1197]. like [302]. Limit [197, 1583]. Limitations [1243, 1673, 1875]. Limited [37, 1164, 233, 1697]. Linda [1150]. Line [1104, 2270, 194, 871, 1985, 663, 678, 881, 26, 40, 843, 2199, 195, 2193, 2192, 2183, 612, 649, 869, 2197, 42, 875]. Line-to-Line [2199]. Linear [951, 327, 2251, 1664, 1698, 321, 2253, 2272, 1721, 1736, 19, 55, 1101, 56, 952, 1050, 1781, 1330, 1739, 207, 962, 1795, 220, 1733, 455, 516, 112, 696, 279, 301, 1801, 2143, 628, 2309, 518, 864, 1645, 44, 1198, 1463, 300, 846, 953, 822, 353, 1103, 163, 1273, 1802, 1563, 1788, 1570, 1792]. Linear-Time [1103, 1721, 44]. linearity [1791]. Lines [183, 666, 899, 611, 634, 1349]. Linguistic [1478, 1354]. Linkage [1737]. Lip [2248]. Lip-Synchronous [2248]. List [425, 189]. Lists [1048, 1517, 1123, 1578]. liveness [1317]. Liver [639, 824]. LL [310, 1922]. Load [752, 2029, 28, 424, 1146, 1526, 1108]. Loading [1056]. Local [37, 1995, 68, 965, 2293, 549, 192, 954, 859, 64, 700, 1294, 1198, 576, 1517, 2202, 637, 629, 852, 248, 1392, 1859]. Locality [1198]. Localization [2198, 676, 1352]. Localized [1268]. Locally [1493, 2283, 2058]. Location [1332, 802, 1988, 182]. Locations [1106]. Lock [1078]. Lofar [899]. Log [153, 588, 586, 527, 173]. Log-Polar [588, 586]. Log-Time [153]. Logarithmic [1086, 936]. Logic [2163, 1481, 1048, 247, 287, 1173, 1415, 1681, 1049, 1202, 976, 1449, 1711, 944, 1435, 1417, 218, 1186, 1420, 2099, 500, 1405, 911, 1274, 1057, 207, 946, 528, 501, 652, 1900, 1206, 410, 344, 948, 112, 1888, 573, 134, 221, 947, 1924, 1198, 1523, 1462, 1939, 1184, 2162, 1444, 1509, 1387, 411, 982, 733, 1438, 1208, 734, 2191, 1201, 1427, 1466, 1828, 24, 1936, 23, 2160, 1369, 1566, 2054, 1567, 1581]. Logic-Based [2191]. Logic-Oriented [1939]. Logical [1421, 1405, 1500, 1900, 201, 117, 1502]. Logics [1433, 1166, 222, 1404, 1436]. Logspace [535, 1700]. LOLITA [2130]. Long [2229]. Long-Term [2229]. Look [1396, 1383]. Looking [2279]. Loop [330, 430, 432, 433, 627]. Loops [431, 1130]. Loosely [1329]. Loss [272, 2027]. Lossless [822]. LOTOS [267]. Low [1129, 808, 830, 420, 1710, 2232, 632, 444, 2258, 359]. Low-Cost [808, 359]. Low-Weight [2258]. Lower [243, 552, 25, 1702, 537, 544, 551, 186, 532, 553]. LSE [506]. LSign [1231]. LU [436]. Lucas [318]. Lucas-Based [318]. Lung [2209, 2208]. Lungs [697]. LUT [919]. LVQ [801, 646].

MAC [288]. Maceral [2224]. MacGuffin [1790]. Mach [2023]. Machine [1572, 911, 1896, 488, 779, 1688, 1187, 934, 359, 253, 1135, 1899, 1663, 1304, 1872, 1842]. machine-assisted [1304]. Machines [232, 552, 153, 1089, 1153, 546, 231, 444, 350, 401, 193, 1309, 66, 1582, 87]. Machining [39]. MacMahon [952]. MacRad [1838]. Macrodynamics [1975]. Macrostructure [1979]. MACs [289, 288]. Magic [166]. Magnetic [828, 2220, 2210, 625]. MAI02L [103]. Mail [876]. Main [2255, 1942]. Mainstream [406]. Maintain [2070]. Maintaining [383, 1766]. Maintenance [1054, 1380, 1254, 1520, 1392, 2079, 1843]. Major [1338, 2224]. Majority [551]. Make [1170]. Making [1032, 388, 1898, 1892, 2074]. malicious [1569]. Malignness [171]. Malleability [1994]. Man [779]. Managed [1615]. Management [920, 494, 360, 1743, 2022, 1611, 1594, 1260, 2111, 1605, 1279, 1280, 2114, 1753, 2102, 2016, 1589, 1604, 1300, 1744, 1597, 1748, 1754, 806, 1598, 1761, 1760, 1745, 1747, 2024, 1069, 1610, 774, 1763, 1615, 1370]. Managerial [1032]. Managers [1619]. Managing [2088, 141, 1599, 2113]. Manifesto [1260]. Manifolds [1643]. Manipulation [1703, 340, 466, 998, 1969]. Manuscripts [2191, 2245]. Many [1727, 1575]. Map [691, 2241, 1349, 1113, 72, 1016]. map-trees [72]. Mapped [1067]. Mapping [1134, 1493, 435, 592, 1139, 445, 2096, 919, 2195, 1135, 713]. Mappings [522, 1785]. Maps [1470, 722, 679]. Mark [1295, 1293, 1292]. Mark-Sweep [1295, 1293]. marker [1864]. marker-passing [1864]. Markers [2098]. Market [1430]. Markets [2120]. Marking [425]. Markings [1532]. Markov [640, 540, 585, 1157, 844, 822, 1005, 1006]. Markov/Gibbs [585]. Markup [1677]. Marrying [1650]. MARS [2046]. Masks [1897]. Massively [1012, 808, 2065, 356, 1143, 364, 358]. MASTAC [354]. Master [1092, 452]. Master-Slave [452]. Match [1077]. Matching [860, 703, 2144, 2198, 836, 1738, 1183, 798, 1102, 184, 498, 883, 819, 799, 621, 1739, 2068, 41, 1150, 557, 792, 198, 2192, 620, 581, 622, 21, 2303, 2036, 1723, 1106, 800, 619, 861]. Matching-Time [1739]. Matchings [46]. Material [1072]. Materialized [1780]. Mathema [1441]. Mathematica [493]. Mathematical [1674, 107, 469, 118, 523, 1677, 2131]. Mathematics [1559]. Matrices [27, 2261, 167, 1803]. Matrix [1128]. Matroid [120]. Matsumoto [307]. Maurer [2259]. Max [962, 1712]. MAXCLIQUE [186]. Maximal [2144, 56, 162, 147, 1733, 2294, 957, 636, 456]. Maximization [544]. Maximum [1094, 136, 1926, 1511, 178]. MBC [358]. MBC-100 [358]. MC [2169]. McCurley [2265]. McGuffin [1809]. MCM [451]. MCM-Based [451]. MCR [2240, 877]. Mcube [450]. MD4 [1804]. MDS [2255]. MDx [288]. MDx-MAC [288]. Mean [1134, 124]. Means [1270, 2211]. Measure [534, 380, 1167, 604, 1079]. Measurement [670, 657, 1020]. Measurements [2018]. Measures [1325]. Measuring [1450, 1365, 2032, 712, 996, 1325]. Meat [905]. Mechanically [1194]. Mechanics [2262]. Mechanism [1505, 2082, 420, 1267, 994]. Mechanisms [2231, 1494, 1453, 1761]. Mechanized [734]. Mechanizing [714, 475]. mEDA [352]. mEDA-2 [352]. Media [2019, 2049, 2032, 644, 2035, 1776]. Medial [1736]. Median [382, 803]. Medical [609, 1833, 662, 684, 1927]. Medicine [9]. Meet [2267]. Meeting [2021]. Meetings [1073]. Meets [1918, 1279, 2170]. Meldable [49]. Membership [181, 1894, 1564]. Memoing [1210]. Memorable [2207].

Memories [2067]. Memory

[37, 1104, 1067, 355, 540, 718, 1243, 1086, 1164, 434, 233, 2064, 1087, 758, 1300, 708, 1019, 426, 1294, 414, 2195, 2291, 1942, 349, 152, 753, 1005, 1148, 1885, 2274, 1851, 1878]. memory-based [1878]. Memory-Limited [1164]. Memory-Mapped [1067]. MEMSY [365]. Mental [101, 500]. Merging [1406, 1752, 542]. Meristematic [1991]. Mesh [2158, 1850]. Mesh-Connected [2158]. Meshes [422, 1143, 1087, 1088, 2001, 562, 1135, 2196, 1142]. Meshing [2203]. Message [289, 25, 2155, 290]. Mesure [904]. Meta [2100, 1386]. Meta-Reasoning [1386]. Metacircular [464]. Metafinite [216]. metallic [1837]. Metaphor [1512, 1593, 1497]. Metaphors [1627]. Metatheory [1554]. Meteosat [894]. Method [802, 721, 874, 584, 739, 1492, 176, 1432, 1041, 824, 365, 809, 1455, 2194, 141, 636, 655, 364, 302, 998, 1813, 1928, 2277, 354, 568, 2032, 1522, 1006, 876, 336, 2246, 163, 693, 2236, 930, 1876]. Methodolgy [926]. Methodological [1239]. Methodology [1364, 470, 1918, 1592, 694, 1367, 1934, 2051]. Methods [2225, 289, 1908, 814, 1127, 1342, 498, 1510, 1833, 271, 483, 1992, 2101, 2180, 469, 117, 1930, 475, 120, 1824, 768, 682]. Metonymy [10]. Metric [274, 1364, 1859, 1884]. Metrical [38]. Metrics [1381, 1378]. METU [1022]. Micronesia [1348]. Microprocessors [356]. Microscope [2202]. Microsystems [940]. Mid [892]. Mid-Level [892]. MIDAS [1076]. Migration [37, 1774]. MIKE [1375]. Millennium [1326]. MIMD [914, 1150, 1148]. Min [2304, 962]. Min-Cut [2304]. Mind [1170]. Minimal [1200, 2251, 2144, 161, 229, 2254, 2143, 1400, 919, 2256]. Minimization [688, 1797]. Minimizing [2305, 1930]. Minimum [757, 2272, 183, 885, 162, 62, 2309, 58, 175, 89]. Minimum-Redundancy [58].

Minimum-Weight [2272]. minor [1579]. Miranda [2126]. Mirror [674]. Miss [1299]. Missing [880]. Mixed [1083, 1958]. Mixing [367]. MNP [187]. MOA [1289]. Mobile [1121, 1160, 1002, 41, 1895, 254, 1601, 613, 2171, 750]. Modal [575, 2053, 1404, 1544, 1438, 1885, 2051]. Modal-calculus [2053]. Modality [1515]. Mode [1190]. Moded [1226]. Model [2166, 337, 1431, 1767, 1204, 1084, 328, 1965, 1109, 2313, 802, 814, 2091, 1487, 1423, 884, 1364, 1684, 1215, 100, 1262, 249, 1417, 1446, 2059, 1032, 1442, 778, 276, 1320, 434, 18, 1496, 989, 2085, 1175, 1372, 585, 1351, 1373, 216, 1530, 2102, 1504, 1500, 678, 1150, 594, 1195, 2170, 191, 2280, 2101, 278, 250, 373, 2083, 221, 2157, 2221, 1334, 574, 1400, 1058, 2205, 2093, 1623, 1456, 1762, 1927, 1815, 642, 223, 454, 1617, 152, 590, 407, 1676, 248, 1361, 1331, 596, 710, 1941, 974, 1846, 1849, 1301, 2053, 2050, 85, 2056]. Model-Based [989, 1351, 2170, 642, 1846]. model-building [85]. Model-Checking [373, 2056]. Model-Equivalent [2083]. Modelers [2117]. Modeling [776, 1344, 1976, 1630, 1950, 2108, 771, 1435, 2099, 2098, 1635, 989, 1358, 690, 772, 2097, 2083, 1019, 2227, 117, 1498, 2100, 1650, 1550, 2079, 1689, 1654, 1655, 2317]. Modelling [672, 597, 1420, 1613, 1020, 1045, 1508, 775, 1462, 334, 1001, 1485, 977, 624, 1756]. Models [640, 836, 1674, 215, 205, 257, 2122, 500, 1991, 641, 1534, 1693, 1337, 1354, 332, 1207, 2116, 1612, 2152, 1525, 1014, 1683, 1247, 1662, 589, 822, 118, 1361, 1362, 2106, 1691, 768, 1899, 1663]. Moderate [1719, 1148]. Modes [1332, 1078, 1190, 1545]. modifiable [83]. Modifying [2218]. Modular [2005, 1917, 942, 2004, 1817, 567, 2016, 278, 520]. Modules [2168, 1748]. Modulo [550, 2269, 1797]. Moffat [1096]. Molecular [198]. Moment [650, 2236]. Moment-Based [650]. Moments [639, 843, 998, 664]. MONA [2054].

Monaco [446]. Monadic [219, 2054].

Money [1897]. Monitoring [675, 463, 1616]. Monocular [635]. Monoids [379]. Monotone [1124, 1725]. Monotonic [1389, 1650, 1726]. monotonicity [1573]. Montague [1957]. Moore [1303]. Morishima [167]. Morphing [684]. Morphisms [149]. Morphological [1076]. MOSES [2120]. Motion [2225, 810, 609, 608, 2229, 588, 607, 701, 652, 891, 1543, 892, 615, 593, 34, 823, 599, 2319]. Motion-Based [593]. Motzkin [150]. Moulded [689]. Mouse [639]. Movements [1176]. Moving [361, 2173, 627]. MPEG [2026]. MPEG2 [2048]. MPI [419]. MPP [406]. MPx [359]. MR [827]. MRF [884, 891]. MRI [2211, 832]. MST [885, 333]. MTF [644]. Mu [279]. Mu-Calculus [279]. Muller [1523]. Multi [2266, 2038, 873, 99, 1470, 363, 1419, 2220, 1397, 1217, 296, 1441, 1450, 2306, 748, 2151, 138, 1605, 1193, 678, 1087, 2168, 786, 1416, 1597, 96, 1998, 1246, 1452, 1069, 1458, 1522, 862, 1776, 1885, 351]. Multi-Action [2151]. Multi-Agent [99, 1470, 1397, 1441, 1450, 786, 1416, 96, 1452, 1458]. Multi-Dimensional [2266, 748, 1087, 1998]. Multi-domain [1597]. Multi-Echo [2220]. Multi-Layered [1419]. Multi-Level [1069, 1193]. Multi-media [1776]. Multi-Modal [1885]. Multi-Model [678]. Multi-object [2168]. Multi-objective [1522]. Multi-Part [873]. Multi-Party [2038, 296]. Multi-Polygonal [862]. Multi-Scale [2306]. Multi-Service [1605]. Multi-SIMD [351]. Multi-Step [1246]. Multi-Threaded [363, 1217]. Multi-Weighted [138]. Multiattribute [350]. Multibody [2213]. Multicast [2039, 2037, 2000, 1018, 2036, 1625]. Multicomputers [457]. Multicut [1111]. Multidatabase [1025]. Multienvironment [1977]. Multigrid [1127, 364]. Multilevel [1336, 1766, 1427, 1466]. Multilinguality [1468]. Multimedia [1359, 1692, 2045, 1018, 2017, 1693, 2015, 1913, 2033, 1617, 2024, 994, 1610]. Multimodal [684]. Multiparadigm [1205]. multipermutations [1804]. Multiple [1471, 378, 550, 1974, 833, 1952, 1050, 98, 1492, 2173, 1465, 1455, 685, 1043, 1386, 1869, 1801, 1839]. Multiplierless [936]. Multipliers [729]. Multiprocessor [365, 2157, 452, 448]. Multiprocessors [434, 330, 1131, 708]. Multiresolution [1337, 2211, 796]. Multiscale [2235, 662, 817, 606]. Multisearch [1084]. Multiset [1178, 1198]. MultiSite [1755]. Multithreading [357, 416]. Multiwave [451]. Multiwindow [1014]. MUM [2279]. Muscle [825]. Museum [1074]. Music [1064, 2191]. must [1301]. Mutation [1893, 73].

n [527, 1626, 1016]. N-MAP [1016]. n050 [136]. n050/02 [136]. Nagoya [1896]. Naive [1321, 1591]. Name [1229, 2301]. Narrowing [1204, 506, 1182, 1184, 975]. NARX [1903]. National [1520]. native [1844]. Natural [1477, 1424, 786, 1467, 1334, 1158, 1460, 1211, 1458, 9, 1076, 74]. Natural-Semantics-Based [1211]. Naturals [952]. Navier [361]. Navigable [1473]. Navigation [1332, 1327, 1496, 839, 1333, 2171]. Navigational [1972]. NC [1741, 2003, 39, 1700, 224]. Near [2281, 1115, 2228]. Near-Optimal [1115]. Nearest [1245, 1855]. Nearly [229]. Need [1356, 517, 785, 1804]. Needed [1744]. Needles [1672]. Needs [784]. Negated [159]. Negation [18, 1958, 1161]. Neighbor [1245, 1855]. Neighborhood [44]. Neighbourhood [1139, 1515]. Neighbourhood-Constrained [1139]. Nested [1078, 1066, 1569]. Net [258, 389, 334, 1683]. Nets [1008, 386, 1007, 273, 260, 2092, 1654].

Network [337, 2165, 1595, 494, 16, 197, 657, 2094, 630, 1291, 1120, 445, 2280, 1920, 1888, 1604, 688, 2241, 1623, 2239, 1521, 2024, 450, 1576, 1881, 83]. Network-Based [2024]. Networked [1913]. Networking [2042]. Networks [2304, 1493, 1159, 1004, 795, 1473, 1519, 277, 423, 473, 2293, 1605, 2041, 1009, 1903, 749, 50, 2020, 954, 618, 2043, 1889, 817, 646, 1601, 1107, 2162, 444, 2234, 816, 2262, 815, 244, 1498, 978, 2210, 1625, 179, 451, 1659, 468, 548, 1522, 1901, 1887, 2040, 86, 1576]. Networks-Zero-Copy [2020]. Neural [328, 1446, 473, 2094, 630, 445, 1888, 688, 817, 2234, 816, 815, 244, 2239, 2210, 1659, 1887, 661, 86, 1576, 1881]. Neuro [814, 1890, 1634, 1897]. Neuro-Fuzzy [1890, 1634]. Neuro-Money [1897]. Neurobiology [830]. Neurocomputing [1448]. Neurons [1402]. Neuropathy [1256]. Next [1608]. NFS [317]. NG [415]. Nirmani [2077]. NLP [1480]. NN [1525]. No [1735, 1600, 2062]. Node [2156]. Nodule [2209]. Noise [893, 1504, 850, 1569, 70]. Noisy [844, 2184, 1574]. Non [715, 1389, 1408, 1995, 68, 1940, 609, 1347, 830, 1217, 833, 1475, 1548, 745, 2168, 1480, 1943, 1375, 742, 1650, 521, 1461, 160, 2196, 855, 1844, 2059, 1490, 1316]. Non-Atomic [1408]. Non-Blocking [745]. Non-Convex [2196]. Non-deterministic [2168, 1316]. Non-Diffident [1461]. Non-Duplicating [521]. Non-empty [160]. Non-extensional [1480]. non-finite [2059]. Non-first-order [1475]. Non-Functional [1375]. Non-Monotonic [1389, 1650]. non-native [1844]. Non-Oblivious [1995]. Non-Overlapping [68]. Non-Parametric [1548, 855]. Non-Polyhedral [833]. Non-Primitive [715]. Non-Rigid [609]. Non-Spatial [1347]. Non-stochastic [1631]. Non-Stratified [1943]. Non-Strict [1217]. Non-Terminating [1940]. non-Trivial [1490]. Non-Uniform [830]. Nonconvex [457]. Nondeterminism [275]. Nondeterministic [2313, 257]. Nonfrontal [2217]. Nonlinear [1257]. nonlinearity [1787]. Nonmonotonic [530, 1404, 1403]. Nonregular [552]. Nonstationary [1526]. Nonuniform [537]. Nonvisual [1332]. Normal [587, 521, 1439]. Normalization [1781, 1552, 511, 1964]. Normalizing [975]. Normative [101, 1590]. Norms [101]. Not-So-Nearly-Minimal-Size [229]. Notation [1951]. Note [2275, 1641]. Notes [920]. Nothing [2159]. Notion [534, 274, 2149, 538, 1199]. Notions [523]. Novel [2044, 21]. NP [188, 2003, 187, 219, 125, 1700, 547]. NP-Complete [125, 547]. NP-Hard [2003]. NPO [544]. NR [1782]. Nuclear [2121]. Nuclei [639]. NUMA [2157]. Number [1666, 755, 168, 936, 144, 1391]. Numbers [550, 202]. Numerical [367]. Nuts [1738].

O [1067, 54, 2019, 136, 1556, 943]. O-Algorithms [53]. O-automata [2050]. O-Complexity [1703]. OASIS [1028]. OBDD. [147]. Object [2166, 1332, 1976, 1356, 2082, 1965, 672, 836, 2088, 1547, 132, 1038, 2099, 1030, 1022, 11, 2072, 1961, 372, 1045, 2085, 1385, 1352, 1918, 1280, 2084, 1530, 1772, 2109, 763, 1365, 1036, 1948, 1015, 1959, 1963, 1816, 2083, 653, 2080, 1060, 1612, 2117, 1528, 627, 1779, 1329, 1537, 351, 1928, 764, 1929, 1942, 2100, 892, 2113, 2090, 2191, 1966, 1529, 486, 807, 1029, 1531, 862, 2118, 900, 2081, 1269, 673, 2106, 2188, 1964, 1956, 2182, 1542, 2168]. Object- [1332]. Object-Instance-Based [2082]. Object-Orientated [1329]. Object-Orientation [1332]. Object-Oriented [1976, 2082, 1965, 132, 1038, 1030, 1022, 11, 1961, 1045, 1385, 1918, 1280, 763, 1365, 1036, 1015, 1959, 1816, 2083, 2080, 1060, 351, 1928, 764, 1929, 2100, 892, 2191, 1966, 807, 1029, 1269, 2188, 1964, 1956]. Object-Relationship [2088]. Object-Z [486]. Objectbase [1769]. Objectbases [1931]. Objective [1364, 1522]. Objects [751, 1051, 835, 13, 496, 2091, 1950, 1443, 917, 597, 101, 1536, 833, 1101, 1352, 1960, 758, 2173, 1978, 1535, 891, 817, 685, 1046, 2096, 1490, 858, 1680, 2107, 2219, 1066, 1550, 1653, 1615, 1331, 596, 840, 2179, 2131]. Obligations [1363]. Oblivious [1995, 295, 296, 297]. Observable [1157]. Observation [252, 2292]. observational [1319]. Observer [627]. Observers [1507]. Observing [2297]. Obstacle [701]. Obstructions [60]. Occluded [817]. Occluding [602]. OCR [2239]. Odd [1735]. Odin [1757]. ODMG [1961]. ODMG-93 [1961]. ODP [1621]. Off [875, 416, 2183]. Off-Line [2183, 875]. Off-the-Shelf [416]. Offline [1194]. oil [177]. Old [2245, 1561]. omega [1316, 113]. omega-automata [1316]. Omniscience [1405, 201]. OMT [1612]. On-Board [929, 1840]. On-Line [2270, 871, 881, 195, 2193, 2192, 42]. On-the-fly [1293]. Onco [1082]. Onco-Pathology [1082]. One [1295, 535, 546, 1097, 1799, 1802, 1562, 1810]. one-counter [1562]. One-Way [535]. Online [41, 1118]. Only [2025, 246]. onto [445, 923, 919]. Ontology [2, 22]. OO [2098, 2121, 2119]. OODB [1946]. OODBMS [765]. OODBPL [1023]. OP [72]. Open [1291, 1596, 2015, 1914, 1452, 1610, 84]. Operands [353]. Operating [2016]. operation [2316]. Operational [768, 1319]. Operations [507, 708, 9, 1328]. Operative [2069]. Operator [577, 1836, 74]. Operators [1516, 847, 1323, 1644, 1521, 1511, 1066]. Opportunities [475, 1451]. Optical [1109, 669, 586, 872]. Optimal [422, 53, 747, 1084, 552, 1722, 1085, 158, 427, 1730, 863, 1124, 423, 2281, 1115, 1139, 1352, 330, 1120, 393, 797, 591, 708, 1714, 241, 436, 2001, 137]. Optimality [1215]. Optimisation [1523, 1522, 1868, 1873, 79, 80]. Optimization [1671, 230, 525, 187, 27, 1962, 2095, 344, 957, 1963, 426, 1070, 1151, 1521, 1117, 1461, 2175]. Optimization-Based [957]. Optimizations [448]. Optimize [1278]. Optimized [1105, 1897]. Optimizing [1220, 964, 110, 1233]. Optimum [1335, 1793]. Option [801]. Or-parallelism [411]. Oracles [539, 1574]. Order [1916, 535, 884, 218, 630, 528, 958, 1919, 143, 221, 1227, 1184, 733, 960, 1828, 1212, 23, 804, 1317, 1475, 2054, 1798]. Order-Sorted [143]. Orderable [2011, 384]. Ordered [1703, 1049, 1340, 579, 1577]. Ordering [966, 565, 2214, 1523, 756, 1341]. Orderings [509, 1221, 2294]. Orders [564, 17, 1717]. ordinate [1619]. Organisation [1601]. Organising [816, 1346]. Organization [1061]. Organizational [95]. Organizations [1451]. Organized [1326]. Orientated [1329]. Orientation [1332, 605, 39, 2199, 900]. Oriented [1976, 2082, 1965, 1917, 132, 1038, 2099, 1030, 1022, 11, 2072, 1961, 1016, 1045, 1385, 1918, 1280, 2109, 763, 1365, 1036, 1033, 1015, 1959, 1816, 2083, 2080, 1060, 910, 1939, 1779, 351, 1928, 2202, 764, 1929, 2100, 892, 2191, 1966, 807, 1029, 1269, 928, 1430, 2188, 1964, 1956, 1318, 346, 2051]. Orthogonal [1734, 2147, 127, 665, 511, 1745, 1568]. Other [1510, 925, 34]. OTI [988]. our [2262]. Outerplanar [168, 383]. Outlines [1532]. Outperforms [1510]. Output [212, 147, 1647]. Output-Size [147]. Overcome [1672]. Overhead [420]. Overlapping [68, 445, 516]. Overview [1742, 1530, 775]. Oz [1187, 407, 1676].

P [534]. P4s [1984]. PAC [1578, 244]. PAC-Learning [244]. Packages [903]. Packet [2027, 2040]. Packet-Switched [2040]. Packing [808, 1091, 136]. Packings [155]. PAG [1214]. Page [37]. Pair [55, 131]. Pairs [2304, 1096, 2007]. Pandiagonal [166]. paper [2124]. Paraconsistent [1417]. Paradigm [1145, 529, 1572, 1561]. Paradigms [1501, 2285]. Parallel [524, 326, 1708, 327, 1084, 2144, 461, 212, 1147, 338, 425, 203, 435, 1012, 808, 2065, 811, 1140, 363, 16, 1912, 1085, 2158, 415, 1089, 1156, 2004, 1124, 412, 1145, 1011, 1154, 1720, 156, 1142, 339, 386, 356, 1016, 368, 1143, 439, 445, 946, 1141, 2010, 369, 464, 1034, 157, 1206, 331, 1153, 1277, 367, 1041, 130, 346, 365, 809, 1716, 446, 433, 2307, 348, 364, 686, 1925, 1146, 437, 459, 1661, 333, 350, 2048, 1151, 443, 442, 1152, 1155, 335, 1112, 451, 465, 1662, 359, 1035, 1209, 2164, 1149, 1148, 413, 1013, 353, 1135, 1899, 1663, 710, 163]. Parallel [193, 358, 1273, 418, 80]. Parallel/Distributed [193]. Parallelising [458]. Parallelism [408, 1185, 410, 517, 2062, 467, 400, 428, 406, 351, 275, 411]. Parallelization [361, 431, 1130, 1944, 15, 354, 438]. Parallelizing [337, 1166, 1128, 438]. Parameter [863, 2221, 2071]. Parameterised [277]. Parameterized [2092]. Parameters [588, 797, 591]. Parametric [1143, 1548, 399, 589, 1524, 855]. Parametrization [665, 1533]. Parental [1514]. Parity [2258]. PARIX [421]. PARMA [350]. Parser [1426, 1393]. Parsers [2287]. Parsimonious [1425, 1464]. Parsing [554, 1919, 1922, 1178]. Part [873, 2286, 1459]. Part-of-Speech [1459]. Partial [747, 744, 564, 1194, 1195, 59, 1077, 1774, 1731, 1317]. partial-order [1317]. Partiality [2293]. Partially [1443, 1157]. Particle [614, 2177, 75]. Particles [1146]. Partition [2006]. Partitioning [1741, 2003, 1217, 1783, 433, 1135, 928, 78]. Partitions [1101, 2146]. Partly [1293].

Parts [1536]. Party [2038, 1758, 296, 319, 953]. Pass [1295]. Passing [2299, 2301, 1864]. Passive [1109]. Past [1406]. Patch [709]. Patches [1536]. Path [2302, 1394, 35, 2281, 2008, 2283, 1333, 1096, 1070, 1624, 1335, 2007]. Pathology [1082]. Paths [47, 150, 1085, 383, 1704, 145]. Pattern [868, 605, 1102, 184, 694, 2303, 2036, 237, 2246]. Pattern-Matching [2036]. Patterns [841, 835, 1343, 385, 612, 1565, 1877]. paying [10]. Payoff [124]. PB [544]. PB-complete [544]. pC [1010]. PDL [1476]. PDL-based [1476]. Pedagogical [501]. Pei [417]. People [648, 1033]. PEPP [346]. Perception [1541, 1347]. Perceptron [1419]. Perfect [316, 2257, 41]. Perform [1505]. Performance [1288, 1051, 995, 1012, 1378, 1890, 565, 1038, 2029, 1016, 1773, 1605, 950, 915, 656, 369, 1018, 1017, 342, 2172, 2020, 1600, 2043, 999, 916, 293, 2018, 1010, 1607, 1810, 1774, 462, 1934, 1035, 1353, 1209, 2164, 1362, 1013, 1000, 693]. Performance-Oriented [1016]. Periodic [2253, 2034]. Periodically [149]. Periodicity [149, 555]. Periods [1717]. Permutation [2158, 2010]. Permutation-Based [2158]. permutations [1785]. Perpetual [314]. PerPreT [1012]. PERs [1188]. Persistent [1067, 753]. person [77]. Personal [1357, 1600, 1623, 1607]. Perspective [1428, 1673, 1328, 599, 76, 1883]. Perspectives [1355]. Pertaining [665]. Petri [1008, 258, 386, 1007, 389, 273, 394, 334, 260, 1683, 2092, 1654]. PFFM [167]. PFSLib [348]. Phase [667, 955, 859, 700]. PHdPHd1dK [1009]. Phenomena [1446, 1003]. philosophers [2056]. Philosophical [1677]. Photometric [836]. PhotoPix [807]. Physical [1487, 1068, 408, 191, 2317]. Physics [1546, 1547, 1543, 1411, 1550, 52, 533]. Physics-Based [1547, 1543, 1411, 1550].

Picks [1490]. pieces [1839]. Piecewise [841, 846]. Piecewise-Linear [846]. Pipe [1521]. Pipeline [729, 731]. Pipelined [2159]. Pipelines [1783]. Pipelining [1136, 357]. Pixel [849]. PKZIP [1794]. Place [847, 1721, 542, 58]. plaintext [1794]. Planar [381, 2003, 128, 2283, 383, 56, 162, 126, 1639, 1733, 2307, 1119, 634, 2282, 43, 2179, 738]. Planarity [1098]. Planarization [1987]. Plane [54, 586, 1097, 1601]. Planes [681]. Planned [1520]. Planner [1491, 1257, 1878]. Planning [1495, 109, 1242, 1274, 98, 767, 1351, 1409, 2078, 1751, 1074, 92, 117, 837, 34, 769, 1427, 1466, 791, 774, 1848, 2319]. Plans [110]. Plant [826, 615]. Plants [1647]. Plastic [689]. Platform [366, 913]. Plausibilistic [845]. Play [90]. Playback [2029]. Player [2026]. Playout [2046]. PLD [921]. Pohnpei [1348]. Point [128, 1730, 1900, 726, 594, 2228, 637]. Pointers [1067]. Points [1116, 382, 1644, 1100, 579, 2181, 2236]. Poisson [364]. Polar [588, 586, 594, 26]. Polarized [2301]. Policies [540]. Policy [1005]. Poly [1514]. Poly-Parental [1514]. Polyadic [255]. Polygon [1698, 35, 1736, 1100, 1538, 1103]. Polygonal [2186, 2187, 862]. Polygons [382, 1124, 1699, 30, 130, 677]. Polygons-Beyond [1699]. Polyhedra [31]. Polyhedral [833, 1119]. Polymorphic [1219, 255, 1208, 1234, 2290]. Polynesia [1347]. Polynomial [212, 35, 19, 1219, 1388, 374, 444, 981, 2295, 169, 1719, 213]. Polynomials [164, 2315]. Population [489, 1893]. Portable [461, 446, 1233, 1210]. Portuguese [1457, 1459]. Pose [685, 632, 635]. Position [657]. Positions [579]. Positive [1726, 559, 238, 1563]. Possibility [761]. Possible [328]. Postirradiation [697]. Potato [1998, 1107]. Potential [1887]. Power [920, 2300, 219, 658, 220, 396, 2288, 673].

Powers [2269]. PQL [1377]. Practical [1251, 1338, 721, 54, 1711, 1944, 106]. Practice [1908, 743, 2121, 266, 1661, 2054]. Pragmatic [1551, 7, 8]. Pragmatics [1]. PRAM [426, 400]. PRAM-Programs [426]. Praxis [1688]. Pre [1626]. Pre-Design [1626]. Preattentive [707]. Precomputing [295]. Precondition [1495]. Predicate [1471, 274, 569, 24]. Predicates [717, 760, 1953, 437, 235]. Predicted [2012]. Prediction [894, 1012, 1474, 1448, 822, 1013]. Predictive [2173, 1245]. Predictors [1378]. Preemption [2150]. Preference [1333, 1404]. Preferential [1169]. Prefetching [430, 1299]. Preliminary [1479, 1433, 1638, 1472]. Preorder [565, 515]. Prepaid [2263]. PREPARE [1604]. Preprocessing [845]. Presburger [1213]. Presence [2281, 2156, 1569]. Presentations [217]. Presenting [1454, 1422]. Preservation [223]. Preserve [171]. Preserving [507, 522]. Pretty [266]. Preventive [1500]. Primal [199]. Primal-Dual [199]. Primary [2231, 1540]. Prime [2269]. Primer [1669]. Primes [317]. Primitive [715, 387]. Primitives [2233, 635, 2288, 2150, 2318]. PRINCE [1907]. Principles [306, 1619, 2125]. Print [879]. Printed [648, 877]. Prints [1027]. priori [872]. Priorities [1481]. Priority [48, 49, 1155]. prisoner [88, 77]. Privacy [1588]. Private [296]. Proactive [314]. Probabilistic [2313, 621, 233, 699, 2237, 231, 618, 398, 1573, 2162, 265, 264, 668, 1313, 2058]. Problem [524, 230, 1708, 67, 698, 1709, 1160, 1434, 63, 2011, 1423, 194, 2281, 1092, 56, 1492, 549, 1739, 1740, 138, 1872, 941, 1071, 1718, 147, 543, 191, 26, 1118, 131, 173, 190, 192, 628, 1465, 1926, 561, 1696, 953, 1515, 2007, 1149, 1391, 1113, 713, 72, 89]. Problems [1995, 188, 1664, 1482, 1802, 1516, 2003, 187, 1304, 156, 384, 2308, 1126, 27, 1141, 979, 557, 1472, 1904, 544, 108, 1988, 182, 364, 104, 551, 1695, 1119, 1517, 1246, 1518, 978, 1133, 562, 1935, 1128, 985, 956, 1111, 1868, 1873, 81, 1876, 1850, 1797, 2315]. Procedure [563, 12, 981, 268]. Procedures [719, 1483]. Process [370, 1359, 1630, 283, 811, 285, 2029, 2151, 1753, 262, 2102, 1017, 1892, 999, 1980, 1498, 1412, 1411, 997, 1360]. Processes [540, 378, 2299, 563, 565, 249, 1177, 386, 1358, 1157, 254, 265, 1727, 2059, 2058]. Processess [2103]. Processing [1332, 1104, 1477, 1771, 813, 808, 1140, 830, 942, 1145, 921, 657, 356, 1954, 1953, 1141, 369, 331, 1041, 805, 365, 843, 1596, 708, 1158, 1925, 1039, 121, 366, 935, 443, 1774, 1201, 451, 807, 359, 1458, 1538, 918, 1979, 1402, 1524, 1273]. Processor [810, 25, 908, 911, 330, 357, 749, 1072, 195, 2033, 1133, 1901]. Processor-Time [25]. Processors [416, 333]. Procol [571]. Producer [977]. Producer/Consumer [977]. Product [1056, 550, 1365, 916]. Production [770, 1886, 1832, 787, 769, 791, 1398]. Products [2120]. PROFIL [1837]. Profiles [623]. Program [2163, 204, 1681, 1821, 229, 1378, 1177, 340, 466, 246, 344, 1377, 440, 465, 1685, 2137, 1581]. Programmable [943]. Programmes [986]. Programming [929, 1682, 188, 247, 418, 1449, 739, 1435, 1417, 1420, 339, 458, 434, 989, 979, 1640, 343, 777, 1383, 1179, 108, 1225, 2083, 1501, 969, 1924, 1462, 1184, 2162, 1444, 980, 497, 411, 982, 993, 1527, 1438, 986, 1117, 953, 407, 1676, 1209, 2164, 991, 1936, 1197, 928, 1430, 2160, 1956, 72, 2130, 2131, 2125, 2137, 2136, 2128, 2129, 2124]. Programs [1203, 1129, 995, 435, 1216, 1202, 976, 1186, 339, 458, 1220, 439, 12, 1221, 946, 245, 410, 1943, 346, 1921, 1226, 110, 1501, 963, 437, 1010, 1958, 960, 1649, 1208, 1651, 1367, 214, 1964, 553, 2132, 1567, 1943, 426]. PROGRES [1371]. Progress [1333, 1298]. Progressive [2204, 953]. Project [1381, 2122, 1326]. Projection [16, 1234, 20]. Projective [677]. projects [1835]. Proliferative [829]. Prolog [1236, 1777, 1226, 448, 459, 1237]. PROMAL [1438]. Promoting [1418]. Proof [113, 1908, 717, 418, 109, 1363, 249, 1388, 1552, 12, 476, 110, 1558, 568, 2148, 1828, 211, 1168, 1304, 2139, 2051]. Proofs [204, 2165, 313, 1167, 2155, 532, 181, 523, 737, 2315]. Propagation [25, 1275, 445, 118, 1392, 710, 974]. Proper [158, 922]. Properties [63, 329, 91, 1216, 281, 273, 867, 545, 398, 114, 1174, 1792, 520, 1649, 1361, 336, 1317, 2058]. Property [197, 521]. Proposal [1896, 686, 852, 1504]. Proposed [357]. Propositional [122, 1387, 532]. Propositions [115]. ProSet [1150]. ProSet-Linda [1150]. Prospect [346]. Prospects [1284]. Protecting [2262]. PROTEUS [1756]. Protocol [720, 2286, 428, 2031, 734, 1302]. Protocols [140, 306, 718, 99, 311, 1018, 293, 1019, 297, 334, 429, 750, 1042]. Prototype [940, 1618, 936, 1234]. Prototyping [1150, 927, 1024, 947, 2105, 895, 918]. Prove [2264]. Prover [721, 1303]. Provide [170, 472, 1761]. Providing [1327, 1591]. Proving [281, 107, 1407, 1436, 106, 105]. Provision [2039]. Proximity [1323, 45]. Pruning [866]. PSCS [1622]. Pseudo [302]. Pseudo-Random [302]. Pseudorandom [289, 1784]. PTL [280]. Public [306, 315, 2103, 313, 310, 1604, 2259, 151, 307, 2263]. Public-Key [151, 2263]. Pullback [1990]. Pumping [330]. Pupil [675]. Pure [543]. Purpose [460]. Pursuing [627]. Pursuit [104]. Pushdown [398]. Putnam [981]. PVM [352]. PVS [2286]. Pyramidal [812]. Pyramids [890]. PYTHAGORAS [1977].

Q [1469]. Q-learning [1469]. QoS [2042, 2022, 1613, 2021, 2023, 2048]. QoS-Based [2042]. QoS-Control [2023].

QPN [1008]. QPN-Tool [1008].

Quadrangulation [1735].

Quadrangulations [43]. Qualifications [1219]. Qualitative [1344, 1339, 1322, 460, 1323, 1324, 1409, 789, 1174, 1410, 1411, 1247]. Quality [905, 889, 1379, 1381, 1357, 1600, 857, 1382, 1165, 1866]. Quality-Based [1165]. Quantification [1460]. Quantified [437]. Quantitative [538, 1382]. Quantization [889, 2216, 823, 2178]. Quantum [321, 1657, 1638, 297, 2262]. Quasi [1530, 167]. Quasi-Invariants [1530]. Quasi-Morishima [167]. Queries [132, 159, 158, 1725, 1953, 1057, 959, 1963, 1925, 1966, 1353, 1565, 1564]. Query [1030, 1946, 1954, 2084, 220, 725, 1962, 2112, 1963, 1766, 1947, 1201, 224, 1964, 1273]. Querying [1048, 2085, 1955]. Quest [1948]. Question [1031]. Questioning [2066]. Questions [1031]. Queue [144]. Queueing [1008, 1004, 997]. Queues [48, 49, 1009, 1155]. Quiescence [447]. Quo [1914, 1663]. Quorum [140, 28, 141]. Quorum-Based [140].

R [964]. R0ik [203]. R1040 [2318]. Race [419, 1622, 2318]. Radiation [488]. Radio [2030]. Radiography [2209]. Radiology [1838, 2244]. Radiosity [2203, 2204]. RAID [2045]. RAID-3 [2045]. Ramification [1686]. Random [526, 585, 539, 302, 844, 2183, 145, 1785, 66]. Randomized [38, 131, 397, 186]. Randomly [754]. Range [65, 2158, 845, 846, 1353]. Range-Join [2158]. Rankable [170]. Ranked [530]. Ranking [47, 1088]. Rapid [1584, 927, 2105, 1947, 918]. Rare [1447]. Rate [2021, 2017, 2044, 632, 2049]. Rate-Based [2017]. Rates [949]. Ratio [1709]. Rational [931, 308]. Rationale [1376, 1750]. Ratios [680]. Ray [2206]. rays [697]. RBF [1894]. RBF-Fuzzy [1894].

RC5 [301, 1789]. RCL [2292]. RDBMS [765]. RDOSS [753]. Re [2103, 1385, 2101]. Re-Architecturing [1385]. Re-engineering [2103, 2101]. Reachability [1378, 268]. Reacting [367]. Reaction [697]. Reactive [2166, 2284, 409, 1264, 1921, 2150, 1815, 1649, 1362]. Reading [2189]. Real [1295, 2038, 345, 2026, 2230, 942, 2214, 2013, 572, 652, 821, 2016, 112, 2023, 843, 250, 373, 742, 2015, 467, 1298, 1490, 895, 1909, 1658, 1297, 596, 943, 2182, 1562]. Real-Time [1295, 2038, 2026, 942, 2013, 821, 2016, 2023, 373, 742, 2015, 467, 1298, 895, 1909, 1297, 943, 2182, 345]. Real-Tune [2230, 2214]. Real-World [596]. Realisation [1606]. Realising [921]. Realistic [400]. Realization [1280, 2256]. Realizations [1600]. Realizing [1814]. Rearrangements [1668]. Reason [10]. Reasoner [1238, 1867]. Reasoners [1507]. Reasoning [1055, 1389, 1845, 1239, 1040, 1241, 2069, 119, 721, 1950, 1487, 739, 1243, 1421, 2073, 1252, 1476, 2072, 276, 1265, 1322, 2068, 1323, 1499, 1319, 1254, 530, 1172, 2075, 1256, 789, 1174, 1410, 645, 2071, 1171, 1246, 1686, 1650, 1341, 1165, 1386, 1452, 1249, 1887, 1486, 1951, 1690, 1403, 1259, 1439, 2285, 1865, 1846, 1835, 1847, 1871, 1872, 1873, 1853, 1879, 1841, 1857, 1859]. reasons [1845]. Receiver [294]. Receptive [1493]. Recognition [640, 703, 868, 2238, 1547, 883, 882, 1536, 1985, 679, 667, 630, 1532, 2223, 1530, 881, 1886, 824, 560, 817, 685, 595, 2193, 2192, 1528, 2190, 2184, 1529, 1531, 1897, 900, 673, 661, 2222, 2247, 2182, 1542, 2240, 877, 1937, 1875]. Recognition-By-Alignment [1531]. Recognizing [2189, 1479, 1535]. Recombination [1514]. Recombinative [1516]. Reconfigurability [949]. Reconfigurable [1147, 917, 944, 946, 950, 945, 948, 916, 947, 939]. Reconstructing [2179]. Reconstruction [315, 602, 884, 711, 678, 705, 643, 653, 2241, 600, 846].

Recording [737]. Recovering [1282]. Recovery [1548, 1960, 2155]. Rectangles [158]. Rectangular [2219]. Rectangulating [130]. Rectilinear [1730, 382, 671, 130]. Recurrence [238]. Recurrence-Term [238]. Recurrent [1888, 2269, 1659]. Recursion [1219, 2134]. Recursions [212, 1050, 1781]. Recursive [715, 387, 1553, 1679, 1555, 559, 1979, 214, 235, 236, 1273, 71]. RED [2137, 2040]. RED-VBR [2040]. REDOC [915]. Reduce [14, 1515]. Reducibilities [539]. Reducibility [518, 558]. Reducing [2298, 983, 2187]. Reduction [283, 363, 228, 511, 1232, 146]. Reductions [549, 389, 2295]. Redundancy [58, 603]. Reed [2260, 2252, 1523]. Reference [871, 1819, 1352, 1603, 8]. References [1942]. refine [1871]. Refinement [718, 274, 286, 15, 275, 1427, 1466]. Refining [471, 1354]. Reflectance [691, 2201]. Reflecting [1582]. Reform [448]. Reformulation [1503]. refurbishment [1839]. Refuses [1284]. Refutation [2161]. Refutation-completeness [2161]. refute [1581]. Regenerative [1005]. Regime [1931]. Region [584, 913, 793, 1212]. Region-Based [793]. Regions [639, 2190]. Register [1131]. registers [1796]. Registration [849]. Regular [376, 277, 257, 1126, 273, 59, 923, 1208, 1951]. Regular-Factors [59]. Regularization [587]. Regulations [1259]. Reification [1386]. Reinventing [1428]. Reiter [1434]. Reject [801]. Rejuvenation [1780]. Related [2308, 575, 1119, 21, 930]. Relating [410]. Relation [524, 1320, 1960]. Relational [1051, 995, 1917, 965, 1961, 1454, 1057, 962, 1777, 762, 2192, 2093, 1065, 2086, 619, 133, 1834]. Relations [1278, 1175, 515, 1354, 1066, 1353, 133, 1310, 71]. Relationship [1666, 2088, 745, 1992, 2115, 1598, 2096, 2093, 2081, 2106]. Relationships [1011, 1753, 2097]. Relative [2281, 511, 44]. relatively [1876]. Relativized [222]. Relaxation [621, 699, 668]. Relaxed [1019]. Reliable [617, 2180]. relocation [1841]. Remarks [318]. Remindings [1250]. Remote [1616]. Rendering [596, 2196]. Rendezvous [1263]. Reordered [1231]. Reordering [1275]. Repealed [1728]. replacement [73]. Replay [419, 420, 1257]. Replenishment [2229]. Replicated [141]. Replication [1931, 2047]. Report [1333, 2318]. Reporting [1033]. Repositories [1045, 2111]. Repository [1772, 2116]. Representation [1251, 1040, 1540, 1541, 1338, 1494, 1339, 597, 1536, 1390, 2306, 1418, 1492, 663, 1548, 792, 2097, 681, 1058, 1528, 1043, 1545, 1444, 1484, 1537, 694, 892, 901, 1531, 1550, 1843, 1856, 2319]. Representational [1340, 8]. Representations [525, 886, 427, 17, 1532, 1330, 2254, 1549, 2104, 1539, 1535, 1544, 622]. Representing [1482, 1839]. Requirements [1584, 498, 1045, 1253, 2127, 4, 766, 1375, 1000]. Requiring [2183]. Resampling [2178]. Research [1692, 986, 1763, 768]. Reservation [2013]. Reservations [2012]. Reserving [2014]. Resharing [1121]. Residential [2073]. Residual [1930]. Resolution [868, 830, 1962, 1513, 849, 2161, 623, 1933, 1199]. Resolvable [154]. Resolving [1952, 2115, 1463]. Resonance [828, 2220, 2210, 625]. Resource [1740, 2016, 2024, 1838]. Resources [2167, 2014]. Respect [565]. Response [1108]. Responsibility [2118]. Restarting [392]. Restoration [641, 844]. Restricted [760, 218, 514, 1713]. Restrictions [981, 1391]. Restructuring [2087]. Result [479]. Results [48, 357, 1472, 575, 1174, 401, 532, 468]. Rethinking [1428]. Retrieval [2065, 1974, 819, 2072, 1105, 1602, 2044, 1077, 1058, 1782, 15, 1690, 1391, 1834, 1851, 1879, 1838, 1839, 1860, 1864]. Retrieve [1248, 1882]. Retrieving [1854]. Reusability [1163]. Reusable [1826]. Reuse [1364, 1418, 1817, 2075, 1862, 1308]. Reusing [2287]. Revealing [1874]. Reverse [858, 1037, 1065, 2086, 1806]. Review [494, 1283]. Revision [1389, 1488, 2, 530, 1502]. Revisited [1949, 1481, 1818, 304, 440, 1199]. Revisiting [745]. revolutionary [1572]. Rewrite [506, 507, 245, 512, 514, 516]. Rewriting [376, 1990, 379, 2147, 1050, 246, 515, 143, 518, 1923, 520, 151, 1989, 558, 521, 522]. Rialto [2016]. Rich [873, 266]. Ridge [666]. Ridge-Valley [666]. Rights [480]. Rigid [609, 174, 597]. Rigorous [2087]. Rigorously [494]. Ring [2266, 449, 456]. Rings [1729, 759, 449, 560]. RIPE [2318]. Rippling [504]. RISC [357, 416]. Risk [1032]. RNLC [1713]. Road [2190]. Robot [1490, 1648, 2171, 704, 84]. Robotic [1494, 1413]. Robotics [839]. Robots [2174, 1492, 1895, 1513, 1490]. Robust [1546, 1547, 882, 607, 709, 705, 1645, 1158, 850, 599, 88]. Robustness [534, 306, 1891]. Role [929, 2039, 1406, 2106]. Roles [1486]. Rosser [2271, 151, 521]. Rotated [883]. Rotation [630, 702]. Rough [660, 120]. Round [1020, 1802, 1791]. Round-Trip [1020]. Rounds [755]. Route [1333, 1848]. Routing [51, 2037, 1996, 1519, 526, 1999, 748, 1465, 424, 1998, 1107, 46, 1625, 82, 83]. Routings [1997]. RPL [1936]. RQL [1936]. RSA [292, 293]. RT [1361]. RTL [1471, 919]. Ruby [924]. Rule [113, 1270, 1940, 119, 1938, 1262, 1059, 1775, 824, 1072, 683, 1523, 1267, 1268, 616, 1852]. Rule-Based [824, 1852]. Rule-Driven [1072]. Rulebase [1260]. Rules [1271, 1272, 1274, 1896, 1892, 432, 1939, 1411, 1445, 1894, 1941, 1861, 1266]. Run [1940, 1224, 1130, 465, 1128]. Run-Time [1224, 1130, 465, 1128, 1940]. Runtime [446]. rural [1841]. Russia [368].

S [1802, 1791]. S-box [1802]. S-boxes [1791]. S012 [210]. S3.mp [429]. SA [1361]. SA/RT [1361]. Safe [1195, 960]. SAFER [309, 1799, 1804]. Safety [1959, 1909, 1361, 133]. Safety-Critical [1909]. SAI [1549]. Sake [1892]. Salesman [42, 941]. Salespersons [2264]. Samplesort [2141]. SAR [893, 2221]. SAT [1712, 981]. Satan [1682]. Satellite [895, 1774, 898]. Satisfaction [957, 64, 96, 404, 956, 1482, 1876, 1858]. Satisfiability [191]. Satisfy [1274]. Scalability [1934, 413]. Scalable [2039, 810, 2022, 2025, 1660, 2048]. Scale [1338, 666, 2306, 630, 583, 442, 1625, 878, 852, 604, 1006, 2130, 1840]. Scale-Space [666]. Scales [193]. Scaling [955]. Scan [443]. Scattering [644]. Scenario [1163]. Scenarios [1373, 117]. Scene [2200, 600]. Scenes [835, 833, 655, 858, 631]. Schedule [309]. Scheduler [1069]. Schedules [330]. Scheduling [524, 2046, 2167, 194, 767, 1136, 330, 983, 979, 1131, 453, 1137, 1716, 2043, 1520, 708, 195, 2034, 787, 2033, 444, 32, 454, 452, 335, 994, 2035, 193, 1398]. Schema [1765, 2088, 1044, 1764, 2095, 2080, 2096, 1065, 199, 2087, 1846]. schema-based [1846]. Schemas [1961, 2115, 480]. Schemata [4]. Scheme [597, 2138, 344, 2250, 322, 1947, 307, 1133, 2265, 1289, 2276, 2197]. Schemes [316, 1996, 311, 2293, 1999, 748, 1553, 294, 1432, 320, 181]. SCI [1302]. Science [1657, 2139, 2125, 1883]. Scientific [1075, 1912, 1136, 1935, 1580]. SCM [1759]. Scope [1138, 108]. Script [875, 1760, 2197]. Script-Based [1760]. SDL [1606, 1609]. Seamless [415]. seamlessly [1865]. Search [1995, 2302, 48, 14, 362, 1124, 1164, 1274, 691, 1105, 549, 983, 1153, 1472, 29, 64, 1400, 1528, 1672, 300, 1517, 1151, 1525, 1673, 1246, 702, 848, 985, 956, 1873]. Searching [1793, 65, 1512, 174, 2002, 1725, 1699, 869]. Second [218, 528, 2054, 2128]. second-order [2054]. Secret [1121, 315, 316, 2270, 314, 320]. sections [1837]. Sector [1592]. Secure [292, 323, 311, 2093, 1766, 181, 734, 1069, 2318]. Security [1587, 293, 1588, 297, 2093, 299, 1586, 1080, 2249]. Seduct [1828]. Seeding [1515]. Seeing [1506]. SEER [2093]. Segment [795, 2277]. Segmentation [875, 669, 584, 818, 1124, 2224, 834, 607, 826, 585, 583, 690, 859, 1543, 846, 642, 793, 892, 2210, 649, 794, 2211, 832, 856, 796, 898, 682, 855]. Segmentation/Reconstruction [846]. Segmenting [840]. Segments [1104, 649, 650]. Seismic [365]. Selected [2233, 1831, 790]. Selecting [871]. Selection [1342, 1721, 1352, 27, 1333, 660, 692, 694, 74]. Selective [671]. Self [938, 2302, 1109, 757, 1634, 838, 758, 759, 1601, 2066, 907, 816, 1081, 633, 2296, 930, 1346, 1582, 83]. Self-Adjusting [2302]. Self-Calibration [838, 633]. self-confident [1582]. Self-Guided [1081]. Self-Learning [930, 1634]. self-modifiable [83]. Self-Organisation [1601]. Self-Organising [816, 1346]. Self-Questioning [2066]. Self-Simulation [1109]. Self-Stabilization [758]. Self-Stabilizing [757, 759]. Self-Testing [2296]. Self-Timed [907]. Self-Validating [938]. SEM [705]. Semantic [1976, 1426, 1183, 711, 566, 1962, 1195, 1963, 2152]. Semantics [1495, 487, 377, 1434, 283, 267, 258, 1202, 1262, 2099, 1318, 252, 1372, 2110, 1404, 405, 1211, 264, 1651, 1, 571, 1169, 1941, 1319]. Semi [874, 1340, 403, 1439]. Semi-Automatic [874]. Semi-Lattices [1340]. Semi-Normal [1439]. Semi-Traces [403]. Semicontinuous [866]. Semidefinite [188]. Semilattices [326]. Semistructured [1955]. Sender [294]. Sense [1901]. Sensitive [1923, 2040]. Sensitivities [147]. Sensitivity [1432]. Sensor [938, 837].

Sentence [6]. Sentences [1479]. Separating [1847, 40, 347, 181]. Separations [539]. Sequence [2253, 176, 2216, 2277, 164, 1571]. Sequences [860, 886, 609, 894, 2257, 831, 614, 198, 2193, 600, 2269]. Sequential [1216, 1177, 157, 221, 519, 576, 2291, 1112, 2252, 1311]. Sequentiality [516]. Serial [729, 2179, 937]. Series [2144, 959, 1448]. Server [292, 2045, 293, 2049, 1625]. Server-Aided [292, 293]. Servers [2023, 2048, 2047]. Serves [929]. Service [2039, 2012, 839, 1605, 1618, 1600, 1597, 1606, 5, 1619, 997]. Services [1620, 1611, 1613, 1590, 1605, 1608, 1591, 1592, 1612, 1609, 1623, 1586, 1617, 1627, 1610]. Session [1617]. Set [716, 1200, 1709, 63, 2253, 147, 134, 1644, 1557, 1927, 1066]. Set-Term [134]. Set-Valued [1644]. Sets [128, 1730, 183, 500, 2098, 1710, 541, 125, 660, 2145, 1924, 2146, 1814, 169, 1268, 579, 239, 43, 1951, 120, 492, 1689, 1369]. Seven [2117]. Several [1067]. SGML [1045, 482]. SGML/HyTime [1045]. Shading [2228, 628, 629]. Shadows [617]. Shafer [616]. Shallow [1478]. Shamir [298]. Shape [1339, 2306, 1548, 2228, 628, 1543, 590, 2201, 2320]. Shape-from-Shading [628]. Shaped [2179]. Shapes [671, 663, 688]. Shaping [2021]. Shared [355, 2167, 1086, 1087, 758, 1978, 2159, 1294, 443, 2291, 753]. Shared-Nothing [2159]. Sharing [1121, 315, 316, 2270, 314, 453, 320, 302, 1513, 520, 282]. Shelf [416]. shift [1796]. shifts [1572]. Shop [983, 784, 1398]. Shop-Floor-Area [784]. Short [894, 1552, 791]. Short-Term [894]. Shortcomings [762]. Shortest [67, 35, 1085, 383, 1096, 2007, 145]. Should [66, 1744, 1728]. Shuffle [47]. Shuffle-Exchange [47]. Side [1093]. Sides [406]. Signal [1160, 2235, 706, 935, 1538, 918, 1524]. Signals [1630, 830]. Signature [597, 871, 311, 1773, 1041, 2183]. Signatures [292, 580, 871, 319, 646, 115]. Significant [1350, 1422]. SIMD [1131, 1153, 455, 351]. similarities [1581]. Similarity [959, 1859]. SIMON [791]. Simone [263]. SIMPLE [346, 377, 1698, 418, 35, 1578, 382, 1736, 585, 397, 1922, 1726, 302, 1100, 637, 1103, 146, 2314]. Simpler [62]. Simplification [2294, 1523, 2186, 1276, 970, 2055]. Simulated [1151, 1518, 79, 80]. Simulating [1575]. Simulation [272, 770, 355, 1631, 1109, 1595, 1044, 1089, 473, 1002, 460, 357, 453, 367, 896, 1623, 1001, 442, 1655, 1850]. Simulations [1086]. Simulator [1015]. Simultaneous [557]. Single [194, 1472, 1884]. Single-Vehicle [194]. Sink [1097, 1882]. Situ [626]. Situated [92, 2066, 970]. Situation [1408, 1449]. Situations [112]. Size [860, 229, 1101, 147, 1093, 1719]. Sizing [1893]. Skeleton [1125]. Skeletons [866, 1127, 412]. Skeptical [1168]. Sketches [819, 2098]. Skew [580, 1094]. Skew-Symmetric [1094]. Skill [501]. skills [2139]. Skin [809, 623]. Skip [1123]. SKL [1431]. Skylight [2205]. Slave [452]. SLDNF [1199]. SLDNF-Resolution [1199]. Sliced [1773]. Sliding [1289]. Small [1085, 1983, 1942, 148]. Smallest [1706, 1724]. Smalltalk [1980]. SMART2 [995]. SMD [657]. Smooth [2306, 2218]. SMP [406]. SMP/MPP/ [406]. sn [2295]. sn-Reductions [2295]. Snake [656, 797]. Snakes [591, 642]. SNePS [1425, 1464]. Social [1347, 1348, 1452]. Societies [1397, 1450]. Society [1451]. Soft [2260]. Software [1742, 1755, 1767, 461, 1910, 1746, 1379, 1812, 1674, 1752, 1424, 2025, 498, 470, 1833, 1418, 1136, 1817, 1918, 1753, 1374, 1017, 478, 1357, 430, 1383, 1033, 463, 1744, 1972, 1366, 2075, 1382, 366, 469, 1813, 1977, 1760, 1830, 978, 1355, 1675, 2292, 1824, 928, 1000, 1835]. Software-Heterogeneous [463].

Software-Intensive [1379]. Solid [597]. Solomon [2260, 2252]. Solution [929, 973, 1434, 1406, 361, 1517, 353, 861]. Solutions [2203, 435, 2058]. Solve [1037]. Solver [988, 123]. Solvers [973, 349]. Solving [951, 505, 1482, 1423, 1492, 962, 1141, 979, 1472, 108, 364, 562, 1006, 985, 956, 1872, 1873, 1876, 1465]. Some [318, 920, 1702, 1510, 156, 1775, 1753, 266, 1875, 544, 210, 545, 195, 481, 1913, 2256, 401, 1930, 1935, 2274, 27, 1874, 1527]. Sometimes [1092]. Sophisticated [97]. Sort [1821]. Sorted [66, 27, 143]. Sorting [157, 1088, 876, 2196]. Sorts [405]. Source [1097]. Sources [2228]. Space [2305, 13, 1347, 14, 1102, 666, 1101, 582, 667, 1492, 527, 2216, 583, 614, 1340, 453, 759, 1697, 2226, 221, 45, 598, 2146, 581, 820, 1286, 1930, 1335, 1289, 397]. Space-Economical [1697]. Spaces [1512, 1494, 7, 1680]. Span [2253]. Spanning [2142, 1732, 885, 17, 138, 61, 62, 2309, 89]. spareC [1023]. Sparse [2304, 1147, 1741, 541, 845, 2261, 349, 1128, 1570]. Spatial [1040, 1346, 1336, 1347, 1342, 830, 1343, 1252, 829, 1327, 1322, 1323, 1175, 1351, 1348, 1334, 1354, 598, 1350, 1174, 2242, 1035, 1353, 1331, 2319]. Spatio [607, 701, 806, 1345]. Spatio-Temporal [607, 701, 806, 1345]. Spatiotemporal [608]. Special [460, 1627]. Special-Purpose [460]. Specialised [1516]. Specific [940, 1382]. Specification [2284, 1008, 1631, 494, 2091, 2167, 1621, 474, 495, 959, 2043, 1832, 1825, 472, 1261, 2107, 1778, 1823, 1367, 1263, 120, 1967, 2135]. Specification-Based [1367]. Specifications [1424, 473, 1236, 478, 2168, 484, 481, 1207, 479, 502, 1779, 493, 1929, 405, 499]. specified [334]. Specifying [1377, 1366]. Speckle [893]. Spectral [1330]. Spectrum [1820, 1818, 1820]. Speech [814, 1886, 2273, 1459]. Speed [25, 808, 565, 1388, 357, 916, 913, 78].

Speed-Ups [1388]. Speedy [1010].

Spherical [1549]. SPIN [281]. SPLASH [941]. Spline [2235, 1539, 2179]. Splines [176]. Splitting [1406]. SPN [1005]. Spots [809]. Spreading [611, 1931]. SQL [1052, 1768]. SQL-Statements [1768]. SQL-Usage [1768]. SQL0* [1782]. SQL0*/NR [1782]. Squad [561]. Square [681, 888]. Squares [166]. SR [2254]. SS [1600]. SS-No [1600]. SSADM4 [1907]. Stability [1635, 1891, 673]. Stabilization [740, 758]. Stabilizing [757, 759]. Stable [1941, 1876]. Stable-Model [1941]. Stack [1222, 144]. Staff [790]. Standard [929, 356, 579, 1677]. Star [1109]. StarT [415]. StarT-NG [415]. State [272, 1631, 1692, 1309, 249, 192, 1643, 1754, 769, 1005, 2059, 2062]. State-of-the-Art [1754]. Statecharts [1373]. Statements [1768]. States [2089]. Static [1769, 2037, 1225, 708, 336, 1284]. Stationary [1003]. Statistical [2123, 1546, 2199, 694, 587, 659]. Status [346, 1750]. Stay [1284]. Steady [1005]. Steam [2076]. Steerable [707]. Steiner [2011, 1730, 190]. Step [1681, 853, 1037, 1246]. Stepwise [1384, 1921]. Stereo [860, 836, 2173, 705, 634, 695, 861]. Stereoscopic [2206]. Still [822, 1284]. Stochastic [1631, 1510, 1007, 1413, 1518, 693]. Stochastic/Non [1631]. Stochastic/Non-stochastic [1631]. Stock [1149, 81]. Stokes [361]. Stopwatch [1649]. Storage [51, 1216, 546, 2044, 2259, 1774, 2047, 1289, 1283]. Store [1758]. Stored [1456]. Straight [194]. Strapping [592]. Strategic [2077]. Strategies [770, 1440, 90, 1641, 1944, 1146, 241, 93, 1035, 1273, 88, 1836, 1877]. Strategy [1974, 1514, 578, 84]. Stratified [1943, 960]. Stream [2258, 2031, 1794, 1795]. Streamed [414]. Streams [994]. Street [36]. Streets [1699]. Strict [200, 1217, 1233]. Strictness [2149, 1232, 1235]. String [560, 620, 151, 2184, 1106]. String-Rewriting [151]. Strings [679]. Strong [744, 552, 1552, 516, 544, 2124]. Strong-Partial [744]. strongly [1563]. Structural [2238, 2115, 2187, 898, 1328, 877, 1503]. Structure [1917, 116, 185, 2010, 1724, 26, 1701, 350, 1894, 687, 2247, 625, 599]. Structured [1129, 1950, 1216, 1372, 1046, 1412, 1854, 1856]. Structures [525, 2213, 1632, 174, 218, 17, 577, 1679, 256, 1340, 217, 530, 574, 332, 1192, 919, 5, 259, 15, 349, 912, 2134, 1788, 2128, 784]. Structuring [474, 1163, 472, 481, 244]. student [2140]. students [2139, 2133]. Studies [1831, 585, 1862]. Study [1160, 1240, 1051, 1771, 1378, 1833, 1483, 491, 1818, 1380, 1832, 700, 429, 1624, 450, 1035, 1360, 2120, 77]. Studying [1893]. Sub [841, 849, 2007]. Sub-Cubic [2007]. Sub-Patterns [841]. Sub-Pixel [849]. Subcode [2252]. Subcube [456]. Subcube-Ring [456]. Subdefinite [123, 118]. Subexpression [432]. Subexpression-level [432]. Subgraph [1708, 56, 15, 175]. Subgraphs [1705, 1733, 1983]. Subjects [1354]. Subpixel [587]. Subquadratic [1737]. Subsequence [2184]. Subset [1320, 280, 1318]. Subsets [160]. Substitution [1916]. Substitutions [1181, 2184, 1560]. Substructural [1436]. Subsumption [14, 2161]. Subsymbolic [121, 1402]. Subsystems [781]. Subtype [1219]. Subtypes [484]. Subtyping [372]. such [1868, 2070]. Suffices [172]. Sugeno [1889]. Suggestion [1517]. Suitability [935]. Summation [261]. Super [1388, 849, 1902, 446, 359]. Super-Polynomial [1388]. Supercomputing [945]. Superpolynomial [553]. Superposition [503].

Supersampling [2194]. Superscalar [1935]. Superstring [67]. Supertrees [1706]. Supervised [1894, 682]. Supervision [1412, 1836]. Supply [1522]. Support [1075, 914, 1908, 1595, 2073, 778, 2094, 1264, 1253, 1009, 1374, 2016, 1031, 1751, 1073, 1972, 1980, 2118, 1025, 2040, 1834, 1866, 1314, 1837, 1841, 1345]. Supported [1080]. Supporting [337, 1755, 1973, 1052, 1051, 2044, 1376, 1778, 1269, 2106, 1846]. Supports [2251, 275]. Surface [602, 705, 896, 2186, 623, 659, 1538, 687, 2179, 2320]. Surfaces [592, 666, 1987, 2306, 1532, 1143, 2218, 2201]. Surrogate [1766]. Survey [1261, 1927, 1290, 1283]. Sustainable [95]. Sweep [1295, 54, 1293, 1292]. switchbox [82]. Switched [2040]. Symbiosis [1480]. Symbiotic [1394]. Symbol [1160]. Symbolic [1629, 973, 886, 1493, 1144, 984, 1320, 699, 107, 1393, 1307, 531, 622, 1647, 123, 901, 1361, 1402, 1310, 1301]. Symbols [873]. Symmetric [718, 1094, 1107]. Symmetry [580, 605, 1684, 654, 637]. Synchronization [339, 403, 561, 2031, 2032]. Synchronized [994]. Synchronous [325, 744, 431, 2248, 1362, 1303]. Synchronous-Asynchronous [325]. Syntactic [1426, 1752, 1559]. Syntax [1555, 1, 571]. Synthesis [1505, 1629, 724, 2257, 946, 245, 246, 1921, 1656, 1305, 1581]. Synthesising [858]. Synthesized [717, 2222]. Synthesizing [2213, 572, 2285]. System [2166, 1056, 905, 1051, 1027, 1053, 418, 719, 1674, 494, 1082, 1547, 2025, 128, 1758, 1757, 1833, 903, 1634, 1827, 1260, 1064, 340, 466, 1453, 1059, 1253, 1508, 1280, 740, 2109, 2172, 2173, 2016, 4, 1600, 365, 446, 463, 936, 303, 2080, 255, 348, 1256, 806, 1146, 910, 1376, 1228, 1463, 2195, 2074, 1761, 2077, 1037, 816, 1760, 1258, 1945, 442, 1774, 2024, 1891, 1894, 992, 359, 1458, 211, 791, 1209, 2164, 1885, 2222, 1250, 2079, 1076, 1422, 943, 358, 872, 1398, 1834, 1865, 1866, 1851, 2316, 1838, 1857, 1881, 1842]. System-Level [2024]. Systematic [1411]. Systems [272, 1440, 1104, 1629, 1281, 2082, 1940, 2284, 1631, 1769, 2313, 1908, 1595, 506, 920, 1829, 1469, 345, 1632, 1012, 808, 1379, 507, 99, 721, 360, 2147, 1633, 409, 1397, 771, 788, 1497, 474, 249, 1038, 2073, 1011, 1420, 1054, 1636, 1030, 921, 1442, 1635, 743, 778, 1002, 471, 1139, 310, 305, 1817, 1279, 2114, 962, 511, 572, 1898, 512, 2064, 1905, 1071, 1638, 1639, 1034, 1874, 266, 1772, 1223, 284, 780, 763, 1365, 28, 281, 514, 38, 2155, 516, 1602, 1886, 143, 786, 1036, 1643, 518, 999, 250, 373, 2028, 1645, 2157, 2159, 472, 1296, 2070]. Systems [1694, 1913, 302, 1661, 1462, 756, 576, 520, 151, 897, 1766, 1779, 998, 907, 935, 282, 1927, 1815, 1081, 1646, 443, 558, 764, 2291, 1909, 1607, 1485, 452, 1650, 1747, 2113, 291, 1452, 782, 1652, 1069, 807, 769, 2292, 521, 774, 1392, 1756, 918, 413, 1013, 353, 522, 997, 269, 975, 135, 523, 1042, 1025, 1654, 1967, 1727, 163, 2182, 1655, 2285, 2060, 1317, 2318, 2061, 2051, 1880, 2317, 1840, 1883, 75]. Systolic [391, 78].

T [351]. Table [10]. Tableau [1232]. Tabled [1185]. Tables [1186, 66, 1814, 869, 1792]. Tabu [1151, 1873]. Tactical [90]. Tagging [1459]. Tailorable [1971]. Takaoka [1096]. Take [536]. Taking [406]. Talisman [1458]. Tangent [803]. Tangents [40]. Tape [1774]. Target [341]. Targeted [946]. Targeting [1609]. Task [1242, 38, 1525, 351, 454, 704, 193, 1861]. Task-Tree [193]. Tasks [2167, 98, 1716, 1413, 1846, 1867, 1874]. Tatzelwurm [1827]. Tautologies [1388]. Taxonomy [1180]. TCM [2266]. TEA [2321, 1811]. Teaching [813, 2126, 501, 2135, 497, 2130, 2137]. Teams [232, 231, 1575]. Technical [1580, 1834]. Technique [53, 2073, 2224, 921, 2176, 110, 626, 900, 2179].

Techniques [1671, 1478, 495, 417, 2215, 643, 1978, 1393, 1290, 837, 644, 1128, 438, 1858, 79, 80]. Technological [1584]. Technologies [1608]. Technology [271, 2094, 927, 1033, 1906, 1642]. Telecommunication [1620, 1590, 1591, 1614, 910]. Telecommunication-Oriented [910]. Telecommunications [1612, 1586]. Teleoperating [1613]. Telephone [2030]. Telesensation [1776]. Temperature [938]. Templates [1664, 593]. Temporal [1344, 1048, 287, 1173, 2312, 1491, 607, 276, 701, 2111, 1499, 1474, 2110, 954, 2112, 806, 1939, 849, 1276, 1427, 1466, 1345]. Temporally [1326]. Temporary [1286]. Temporized [1268]. Teramac [909]. Term [717, 506, 894, 2147, 2229, 509, 1221, 246, 143, 134, 520, 558, 791, 521, 522, 238, 1856]. Terminating [1940, 214, 975]. Termination [976, 508, 388, 1167, 1221, 512, 514, 1989, 2156, 510]. Terminological [1165, 1386]. Terms [1748, 1555, 1856]. Ternary [2097, 1523]. Terrain [1337]. Tertiary [1774]. Tesseral [1040]. Test [929, 19, 479]. Testability [192]. Testbeds [104]. Testing [270, 132, 518, 2145, 1354, 2034, 2296, 1715, 477, 1367]. Testpilot [1767]. Tests [478]. Text [640, 1473, 1478, 1105, 1046, 2279]. text-text [2279]. Texts [184, 1468, 2303, 1459]. Texture [639, 827, 592, 904, 605, 2233, 792, 859, 2234, 606]. Texture-Mapping [592]. Textured [793, 2182]. Textures [850]. Their [1892, 2110, 867, 484, 2264, 374, 1680, 670, 57, 2009, 1602, 790]. Themes [1264]. Theorem [721, 209, 109, 1050, 107, 1407, 555, 2161, 1436, 1557, 106, 668, 105, 664, 1303]. Theorems [223, 160]. Theoretic [1213]. Theoretical [1992]. Theories [505, 377, 723, 208, 115, 1509, 1502, 1902, 1439].

Theory [1666, 716, 2000, 1765, 1968, 1338, 380, 722, 206, 1449, 1417, 743, 724, 566, 216, 1668, 1679, 266, 1009, 1981, 1688, 1661, 1444, 1484, 1557, 1927, 259, 223, 1650, 275, 120, 2247, 616, 2132, 87]. Therapy [488]. Thinning [878, 2181, 710]. Third [1758]. Third-Party [1758]. Thoughts [266]. Threaded [363, 1217]. Threads [2034, 441]. Three [31, 2283, 1749, 1991, 1831, 1534, 354, 129]. Three-Dimensional [31, 1991, 1534, 354]. Threshold [243, 1725, 320, 319, 655]. tiered [1843]. Tight [425]. Tightly [1069]. tile [177]. Tiling [404]. Tillich [2275]. Time [1295, 698, 370, 212, 2038, 1698, 25, 1004, 153, 35, 2026, 194, 1736, 942, 55, 1219, 2013, 1492, 1739, 572, 959, 652, 1285, 2216, 614, 537, 453, 1733, 1224, 1137, 2110, 821, 2016, 279, 2023, 1130, 2309, 843, 250, 373, 742, 2015, 136, 1192, 467, 1545, 1229, 1298, 444, 1448, 32, 895, 1909, 1649, 2295, 169, 712, 465, 1128, 1013, 1235, 997, 1297, 1103, 1108, 943, 2182, 1940, 345, 1721, 1562, 44]. Time-bounded [1649]. Time-Series [959]. Time-Space [2216]. Time-To-Contact [712]. Time-Varying [2110]. Timed [1629, 287, 286, 573, 907, 1312]. TimeNET [1007]. Times [1195]. Timetable [1517]. Timetabling [1516, 1518]. Timing [1312, 268]. TINA [1594, 1618, 1619, 1617]. TINA-C [1594]. tiny [1811]. Tissue [885, 832]. TkWinHOL [730]. TLA [728]. TLP0S [246]. TM [908]. TMN [1595, 1621, 1587, 1596, 1588, 1585, 1624, 1599]. TNC [211]. Toeplitz [149]. Tolerance [944, 996]. Tolerant [141, 1997]. Tolerating [456]. Tongan [1347]. Tool [1759, 770, 247, 1271, 345, 995, 1012, 1272, 1068, 885, 1011, 1236, 491, 501, 1009, 1374, 1642, 730, 502, 1010, 1237, 1745, 1628, 996, 1689, 1837, 2137, 1008]. Tool-Kit [491]. Toolkit [1970]. Tools [461, 1692, 2169, 458, 1064, 679, 925, 777, 2101, 108, 1694, 895, 1269, 1844, 2057]. Top [2118, 2285]. Top-Down [2118, 2285]. Topological [232, 610, 842]. Topologies [336, 1135]. Topology [57, 1322, 1669]. TOPYDE [1068]. Tori [1998]. Tortoise [2311]. Total [161, 168, 756]. Tour [1092, 236]. Trace [379, 2168, 264, 280, 1727]. Trace-Based [264]. Traces [403]. Tracing [2206, 2177]. Tracker [2173]. Tracking [592, 2229, 675, 614, 2172, 2173, 615, 659, 862]. Tracking-Based [615]. Tractable [196, 967, 629]. Traction [1891]. Tradeoffs [25]. Trading [5]. Traffic [424, 896]. Trail [1732]. train [1840]. Training [86, 2183, 790, 1576]. Trajectory [1332]. Trajectory-Centered [1332]. Transaction [1771, 1751, 1959, 1069, 1691, 1941]. Transaction-Conscious [1941]. Transactions [1078, 1258]. Transfer [295, 271, 296, 1602, 297]. Transform [588, 902, 513, 2237, 347, 692, 1123, 612, 934]. Transformation [1829, 2147, 2289, 417, 2095, 1943, 1072, 1198, 858, 522, 796]. Transformational [1826]. Transformations [1781, 162, 171, 432, 123]. Transformers [2163, 274, 569]. Transforming [1310, 1961]. Transforms [665, 933]. Transient [1007, 1003]. transistors [2135]. Transition [955, 2051]. transitions [2061]. Transitive [1097]. Translation [1415, 630, 1947, 1229]. Translations [1397]. Transmission [2230, 749, 2277]. Transmissions [50]. Transparent [1973]. Transportation [97, 1118]. Transpositions [2184]. Transputer [16]. Traveling [42, 2264]. Travelling [1092, 941]. Traverdi [1829]. TRDL [1485]. Treat [10]. Treating [1242]. Treatment [1375]. Tree [53, 2011, 757, 156, 17, 387, 138, 658, 1087, 1153, 1920, 62, 2309, 136, 242, 692, 661, 193, 1565, 89]. Tree-Packing [136]. Tree-Wavelet [692]. Trees [422, 2302, 1048, 2305, 48, 57, 1730, 885, 2002, 1110, 1091, 61, 1818, 1340, 59, 131, 29, 45, 547, 265, 46, 179, 1731, 72, 2142, 144]. Treewidth [1996, 1085]. Trend [786]. Trends [355, 1687, 782]. Triadic [3]. Triangles [1540]. Triangular [1143]. Triangulated [1331]. Triangulation [601, 1103, 175]. Triangulations [128, 621, 2307]. Tribulation [740]. Triggering [1270]. Trip [1020]. Triple [55, 1086, 696, 1114]. Triple-free [55, 1114]. Triple-Logarithmic [1086]. Triplet [673]. Triumph [740]. Trivial [1490]. TROLL [1822]. TROM [2166]. TRSs [513]. Truly [1084, 329, 1301]. Truncated [1798]. Trust [1230]. Trusted [319]. TS [1398]. TSL [1026]. TSP [1868, 1805]. TSP2Opt [1994]. TTP [1588]. TTP-based [1588]. Tumor [824]. Tumours [827, 623]. Tune [2230, 2214]. Tuning [1288]. Turbines [2076]. Turing [552, 153, 546, 401, 1902]. Tutorial [985]. Tutoring [1440, 1442]. Tutte [1719]. Two [1433, 111, 158, 184, 30, 1991, 495, 1704, 391, 1534, 26, 2180, 2232, 1077, 333, 401, 702, 761, 1149, 599, 1843, 1908, 1244]. Two- [1991]. Two-Dimensional [184, 495, 391, 1534, 2232, 401, 1149]. Two-Floodlight [30]. Two-Headed [1077]. Two-Line [26]. two-tiered [1843]. Type [487, 1189, 212, 206, 1222, 1223, 255, 405, 213, 559]. Type-Based [1222]. Typed [1960, 1226, 1555]. Types [805, 115, 1208, 2290, 492]. Typical [545, 1171].

Ultrasonic [824, 2234]. ULYSSES [1974]. Unambiguous [1372]. Unbalanced [197, 1016]. Unbounded [1632, 294]. Uncertainty [1409, 459, 1410]. Unconstrained [2161]. Undefined [1678]. undergraduates [2125]. Underlying [1434]. Underspecification [1678]. Understanding [666, 1054, 2130, 2241, 1460, 901, 825]. Underwater [902, 844, 900]. Undirected [1709, 1724]. Uneven [655]. Uni [180]. Uni-Directional [180]. UniCalc [123]. Unicast [2000]. Unicast-Based [2000]. Unidirectional [449]. Unification [1200, 1916, 1681, 394, 1226, 134, 20]. Unification-Free [1226]. Uniform [1990, 170, 757, 830, 759, 891, 1024, 195, 1188, 1884]. Uniformity [538, 1791]. Uniformly [222]. Unifying [1336, 474, 967, 2083]. Unions [158, 1565]. Unique [254, 521]. Uniqueness [1189, 2104]. Unit [2274]. Units [1971]. Universal [751, 2019, 171]. Unknown [36, 137]. Unobservable [618]. Unplug [2030]. Unstructured [1134, 1142, 1016, 1141, 2177, 1135]. Unsupervised [584, 1894, 855]. Unsupervised/Supervised [1894]. Untangling [1743]. Update [189]. Updates [1274]. Updating [1332, 2145, 1337]. Ups [1388]. UPT [1618, 1624]. Usability [1592]. Usage [2305, 1768]. Use [2123, 875, 1910, 949, 1347, 2169, 473, 922, 1772, 2110, 1886, 1033, 653, 484, 110, 2116, 1661, 1517, 935, 816, 1619, 791, 1259, 1868]. Used [115, 2243]. Usefulness [998]. User [1052, 988, 1611, 819, 1590, 921, 1953, 1918, 488, 2016, 1555, 1980, 1196, 1628, 1328, 2052]. User-Centric [2016]. User-Defined [1953]. Users [1591, 1586, 1628, 1844]. Using [640, 639, 703, 674, 1278, 672, 1040, 914, 836, 289, 609, 893, 602, 1044, 676, 494, 795, 14, 884, 1419, 2220, 362, 1026, 597, 1497, 1124, 588, 2021, 1274, 1322, 1220, 1492, 1532, 1045, 989, 2111, 234, 1280, 1221, 585, 962, 1530, 1960, 2121, 1150, 594, 2260, 478, 2286, 1534, 1852, 1031, 131, 2228, 1888, 927, 2137, 1777, 859, 1015, 916, 936, 817, 685, 2193, 1413, 890, 636, 1501, 114, 2259, 467, 1257, 844, 858, 1609, 2184, 2234, 895, 5, 2202, 1646, 2239, 913, 2263, 820, 1881, 1208, 1624, 1248, 1232, 2036, 712, 2171, 456, 2092]. Using [1106, 822, 1897, 1263, 2089, 2178, 1887, 2081, 1756, 918, 1361, 1360, 9, 2246, 664, 438, 2236, 133, 1937, 2225, 1565, 72, 647, 1802, 78, 277, 2200, 1837, 1319, 1849, 2138, 2215, 1801, 4, 1879, 1312, 707, 1523, 1854, 121, 89, 1858, 1884]. Utilisation [1227]. UTT [1554]. UU [387]. UU-Algorithm [387].

v2 [1028]. Vadetis [1663]. Vadis [1914]. Validated [1175]. Validating [938]. Validation [420, 1662]. Valley [666]. Valleys [611]. Valuation [2073]. Value [2299, 109, 364]. Value-Passing [2299]. Valued [1644]. Values [618, 2071]. Variability [1756]. Variable [966, 949, 2278, 2098, 2044, 1523, 2049, 997]. Variable-Length [949, 2278]. Variables [951, 1632, 518, 618]. Variance [830]. Variations [1264]. Various [551]. Varying [2110]. VBR [2040]. VC [1982]. VC-Dimensions [1982]. VDM [476]. Vector [889, 853, 2176, 2178, 1935]. Vectorization [439, 347]. Vectorized [843]. Vehicle [194, 1465, 1081]. Velocity [2176]. Verb [1457]. Verifiable [1121]. Verification [1829, 1632, 2312, 2169, 871, 903, 1827, 1831, 1605, 262, 726, 729, 62, 1816, 1832, 731, 2183, 1646, 1830, 734, 1651, 261, 1685, 1317, 2057, 1318, 1307, 1302, 1313]. verification-oriented [1318]. Verified [1182]. Verifier [329, 313]. Verifiers [1700]. Verify [1361]. Verifying [277, 1315, 1194, 2286, 429, 1649, 1303]. Verilog [925]. Versatile [2085]. Version [1433, 1745, 1370, 1369]. Versioning [1749]. Versus [1275, 356, 1126, 1332, 233, 1112]. Vertex [2144, 1709, 1724, 1107, 2282]. Vertex-Connectivity [1724]. Vertex-Symmetric [1107]. Vertices [60]. Very [632, 2261, 736]. Vessels [2212]. VHDL [1318, 1319, 926, 927, 732]. VHDL-Based [927]. VHDL’87 [1320]. via [592, 580, 2098, 27, 656, 660, 64, 46, 702]. Viable [1636]. Vibration [1545]. Video [2231, 1122, 2039, 949, 2022, 2025, 2230, 818, 2021, 2029, 2170, 2044, 2028, 632, 2048, 2047, 2040]. Video-on-Demand [2039, 2044, 2047]. Vienna [338]. View [2123, 744, 597, 2094, 495, 1900, 26, 1024, 2015, 1031, 1073]. View-Synchronous [744]. Viewpoint [1355]. Viewpoints [1052, 496]. Views [1965, 2198, 1044, 838, 1377, 685, 595, 1780, 576, 1029, 2092, 1967]. Virtual [723, 1326]. Visibility [1337, 129]. Visible [1698]. Vision [2174, 905, 847, 1494, 675, 911, 1687, 839, 1150, 1176, 2170, 2173, 1539, 708, 707, 444, 1537, 359, 2273]. Vision-Based [675, 839]. Visited [305]. Visits [1074]. VISPAT [342]. Visual [689, 1271, 814, 2229, 1446, 1030, 2173, 618, 1334, 1178, 887, 980, 694, 837, 2243, 704, 2320]. Visualization [1359, 1272, 342, 2212, 2176, 502]. Visualizing [1991]. VITAL [1271]. VLSI [810]. Vocabulary [2066]. Volume [2196]. VOODOO [1745]. Voronoi [397, 1100, 34, 1103]. VPN [1603, 1598]. vs [1346, 1546, 808, 2065, 54, 313, 1405, 266]. Vulnerability [1997]. VuSystem [2018].

W [486]. Wait [758, 741]. Wait-Free [758, 741]. Waiver [1360]. Walk [547]. WARPmemory [462]. Warrant [303]. WASA [1075]. Watchman [35]. Watermeters [2243]. Watershed [686, 610, 796]. Wave [710, 2037]. Wavelet [646, 692, 1148]. Wavelets [893, 706]. Way [535, 112]. Weak [252, 528, 263, 129, 956]. Weak-Commitment [956]. Weakest [1495]. Weakly [1698, 975]. Weakness [309]. Weaving [2242]. Web [1327, 1915]. Weight [2272, 2254, 2258, 2268, 148, 175]. Weighted [1709, 1707, 138, 643, 145]. Weighting [2216, 1863]. Weights [243, 1116]. Welcoming [1902]. Weld [902]. Well [509, 1226, 522]. Well-formedness [522]. Well-foundedness [509]. Well-Moded [1226]. Well-Typed [1226]. Where [1395, 1899, 2267]. which [1698, 1510, 73]. While [1472, 431]. Wide [1327, 1915]. Widgets [1970]. Width [864]. Widths [2001]. Will [266]. Window [194, 730, 2195]. Wireless [1601]. Within [60, 2081, 1315, 1474, 44, 714]. Without [40, 1325, 951, 10, 1572, 81, 319, 2183, 2296, 1286, 73, 596]. Witness [549]. Witness-Isomorphic [549]. Word [2238, 882, 1919, 518]. Words [873, 2272, 149, 870, 555, 2268]. Work [2108, 1753, 1778, 1080]. workers [1841]. Workflow [1075, 360]. Workflow-Based [1075]. Workload [1014]. Workpiece [39]. Workstation [2029, 348]. Workstations [462]. World [1546, 1487, 408, 1534, 1658, 596, 1327, 1915]. Wormhole [468]. Worst [1722, 1088]. Worst-Case [1088]. Wrappers [1947]. Writing [499, 1844]. Written [1424]. WWW [1082, 1915].

X [2091, 697]. X-rays [697]. XC6200 [908]. XOR [289]. XTEA [2322].

year [1799, 2128, 1810].

Z [2284, 494, 474, 500, 482, 491, 478, 488, 490, 472, 481, 479, 502, 486, 480, 492]. Z-into-Haskell [491]. Zemor [2275]. Zero [313, 2020]. Zero-Knowledge [313]. ZF [716, 1557]. ZM4 [346]. ZM4/SIMPLE/PEPP [346].

References

Sowa:1995:SSP

[1] J. F. Sowa. Syntax, semantics, and pragmatics of contexts. Lecture Notes in Computer Science, 954:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foo:1995:OR

[2] N. Foo. Ontology revision. Lecture Notes in Computer Science, 954:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehmann:1995:TAF

[3] F. Lehmann and R. Wille. A triadic approach to formal concept analysis. Lecture Notes in Computer Science, 954:32–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamath:1995:AID

[4] R. Y. Kamath and W. R. Cyre. Automatic integration of digital system requirements using schemata. Lecture Notes in Computer Science, 954:44–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puder:1995:STU

[5] A. Puder, S. Markwitz, and F. Gudermann. Service trading using conceptual structures. Lecture Notes in Computer Science, 954:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicolov:1995:SGC

[6] N. Nicolov, C. Mellish, and G. Ritchie. Sentence generation from conceptual graphs. Lecture Notes in Computer Science, 954:74–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulin:1995:DSP

[7] B. Moulin. Discourse spaces: A pragmatic interpretation of contexts. Lecture Notes in Computer Science, 954:

89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulin:1995:PRA

[8] B. Moulin. A pragmatic representational approach of context and reference in discourses. Lecture Notes in Computer Science, 954:105–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1995:UCG

[9] J. C. Wagner, R. H. Baud, and J.-R. Scherrer. Using the conceptual graphs operations for natural language generation in medicine. Lecture Notes in Computer Science, 954:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amghar:1995:TLP

[10] T. Amghar, F. Gayral, and B. Levrat. Table 10 left without paying the bill! A good reason to treat metonymy with conceptual graphs. Lecture Notes in Computer Science, 954:129–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellis:1995:OOC

[11] G. Ellis. Object-oriented conceptual graphs. Lecture Notes in Computer Science, 954:144–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghosh:1995:DPP

[12] Bikash Chandra Ghosh and V. Wuwongse. A direct proof procedure for definite conceptual graph programs. Lecture Notes in Computer Science, 954:158–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bournaud:1995:CCC

[13] I. Bournaud and J.-G. Ganascia. Conceptual clustering of complex objects: A generalization space based approach. Lecture Notes in Computer Science, 954:173–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Champesme:1995:UES

[14] M. Champesme. Using empirical subsumption to reduce the search space in learning. Lecture Notes in Computer Science, 954:188–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roberts:1995:NPS

[15] J. D. Roberts. A new parallelization of subgraph isomorphism refinement for classification and retrieval of conceptual structures. Lecture Notes in Computer Science, 954:202–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1995:TNI

[16] A. Chan and P. Kocura. Transputer network implementation of a parallel projection algorithm for conceptual graphs. Lecture Notes in Computer Science, 954:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fall:1995:STR

[17] A. Fall. Spanning tree representations of graphs and orders in conceptual structures. Lecture Notes in Computer Science, 954:232–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esch:1995:IMC

[18] J. Esch and R. Levinson. An implementation model for contexts and negation in conceptual graphs. Lecture Notes in Computer Science, 954:247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cogis:1995:LDC

[19] O. Cogis and O. Guinaldo. A linear descriptor for conceptual graphs and a class for polynomial isomorphism test. Lecture Notes in Computer Science, 954:263–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willems:1995:PUC

[20] M. Willems. Projection and unification for conceptual graphs. Lecture Notes in Computer Science, 954:278–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poole:1995:NAM

[21] J. Poole and J. A. Campbell. A novel algorithm for matching conceptual and related graphs. Lecture Notes in Computer Science, 954:293–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoede:1995:OKG

[22] C. Hoede. On the ontology of knowledge graphs. Lecture Notes in Computer Science, 954:308–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wermelinger:1995:CGF

[23] M. Wermelinger. Conceptual graphs and first-order logic. Lecture Notes in Computer Science, 954:323–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandenBerg:1995:EGD

[24] H. Van den Berg. Existential graphs and dynamic predicate logic. Lecture Notes in Computer Science, 954:338– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bilardi:1995:LBP

[25] G. Bilardi and F. P. Preparata. Lower bounds to processor-time tradeoffs under bounded speed message propagation. Lecture Notes in Computer Science, 955:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaromczyk:1995:TLC

[26] J. W. Jaromczyk and M. Kowaluk. The two-line center problem from a polar view: a new algorithm and data structure. Lecture Notes in Computer Science, 955:13–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glozman:1995:SGS

[27] A. Glozman, K. Kedem, and G. Shpitalnik. On some geometric selection and optimization problems via sorted matrices. Lecture Notes in Computer Science, 955:26–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzman:1995:LBQ

[28] R. Holzman, Y. Marcus, and D. Peleg. Load balancing in quorum systems. Lecture Notes in Computer Science, 955:38–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kroell:1995:BDS

[29] B. Kroell and P. Widmayer. Balanced distributed search trees do not exist. Lecture Notes in Computer Science, 955:50–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estivill-Castro:1995:TFI

[30] V. Estivill-Castro and J. Urrutia. Twofloodlight illumination of convex polygons. Lecture Notes in Computer Science, 955:62–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1995:CAI

[31] G. Das and M. T. Goodrich. On the complexity of approximating and illuminating three-dimensional convex polyhedra. Lecture Notes in Computer Science, 955:74–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phillips:1995:SJA

[32] C. Phillips, C. Stein, and J. Wein. Scheduling jobs that arrive over time. Lecture Notes in Computer Science, 955:86–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frandsen:1995:DAD

[33] Gudmund Skovbjerg Frandsen, T. Husfeldt, P. Bro Miltersen, and T. Rauhe. Dynamic algorithms for the dyck languages. Lecture Notes in Computer Science, 955:98–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sharir:1995:AHD

[34] M. Sharir. Arrangements in higher dimensions: Voronoi diagrams, motion planning, and other applications. Lecture Notes in Computer Science, 955: 109–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1995:CSW

[35] S. Carlsson and H. Jonsson. Computing a shortest watchman path in a simple polygon in polynomial time. Lecture Notes in Computer Science, 955: 122–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez-Ortiz:1995:GHT

[36] A. Lopez-Ortiz and S. Schuierer. Going home through an unknown street. Lecture Notes in Computer Science, 955: 135–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1995:PML

[37] S. Albers and H. Koga. Page migration with limited local memory capacity. Lecture Notes in Computer Science, 955:147–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Irani:1995:RAM

[38] S. Irani and S. Seiden. Randomized algorithms for metrical task systems. Lecture Notes in Computer Science, 955:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1995:EGA

[39] P. Gupta, R. Janardan, J. Majhi, and T. Woo. Efficient geometric algorithms for workpiece orientation in 4- and 5 axis NC machining. Lecture Notes in Computer Science, 955:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirkpatrick:1995:CCT

[40] D. Kirkpatrick and J. Snoeyink. Computing common tangents without a separating line. Lecture Notes in Computer Science, 955:183–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grove:1995:OPM

[41] E. F. Grove, M.-Y. Kao, P. Krishnan, and J. Scott Vitter. Online perfect matching and mobile computing. Lecture Notes in Computer Science, 955:194–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ausiello:1995:CAL

[42] G. Ausiello, E. Feuerstein, S. Leonardi, and L. Stougie. Competitive algorithms for the on-line traveling salesman. Lecture Notes in Computer Science, 955:206–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toussaint:1995:QPS

[43] G. Toussaint. Quadrangulations of planar sets. Lecture Notes in Computer Science, 955:218–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lingas:1995:LTC

[44] A. Lingas and A. Mukhopadhyay. A linear-time construction of the relative neighborhood graph within a histogram. Lecture Notes in Computer Science, 955:228–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liotta:1995:CPD

[45] G. Liotta and G. Di Battista. Computing proximity drawings of trees in the 3-dimensional space. Lecture Notes in Computer Science, 955:239–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roberts:1995:RTM

[46] A. Roberts, A. Symvonis, and L. Zhang. Routing on trees via matchings. Lecture Notes in Computer Science, 955:251–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Annexstein:1995:RAH

[47] F. S. Annexstein and E. A. Kuchko. A ranking algorithm for Hamilton paths in shuffle-exchange graphs. Lecture Notes in Computer Science, 955:263– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyar:1995:ARC

[48] J. Boyar, R. Fagerberg, and K. S. Larsen. Amortization results for chromatic search trees, with an application to priority queues. Lecture Notes in Computer Science, 955:270–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodal:1995:FMP

[49] Gerth Stoelting Brodal. Fast meldable priority queues. Lecture Notes in Computer Science, 955:282–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kagaris:1995:CFD

[50] D. Kagaris, G. E. Pantziou, S. Tragoudas, and C. D. Zaroliagis. On the computation of fast data transmissions in networks with capacities and delays. Lecture Notes in Computer Science, 955:291–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baruah:1995:RDW

[51] S. Baruah and G. Pruesse. Routing with delays when storage is costly. Lecture Notes in Computer Science, 955: 303–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitanyi:1995:AAP

[52] P. Vitanyi and M. Li. Algorithmic arguments in physics of computation. Lecture Notes in Computer Science, 955:315–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arge:1995:BTN

[53] L. Arge. The buffer tree: A new technique for optimal I/O-algorithms. Lecture Notes in Computer Science, 955: 334–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiang:1995:EPE

[54] Y.-J. Chiang. Experiments on the practical I/O efficiency of geometric algorithms: Distribution sweep vs. plane sweep. Lecture Notes in Computer Science, 955:346–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corneil:1995:CDP

[55] D. G. Corneil, S. Olariu, and L. Stewart. Computing a dominating pair in an asteroidal triple-free graph in linear time. Lecture Notes in Computer Science, 955:358–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djidjev:1995:LAM

[56] H. N. Djidjev. A linear algorithm for the maximal planar subgraph problem. Lecture Notes in Computer Science, 955:369–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Callahan:1995:TBT

[57] P. Callahan, M. T. Goodrich, and K. Ramaiyer. Topology B-trees and their applications. Lecture Notes in Computer Science, 955:381–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moffat:1995:PCM

[58] A. Moffat and J. Katajainen. In-place calculation of minimum-redundancy codes. Lecture Notes in Computer Science, 955:393–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaller:1995:RFC

[59] D. Kaller, A. Gupta, and T. Shermer. Regular-factors in the complements of partial k-trees. Lecture Notes in Computer Science, 955:403–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattell:1995:OWF

[60] K. Cattell, M. J. Dinneen, and M. R. Fellows. Obstructions to within a few vertices or edges of acyclic. Lecture Notes in Computer Science, 955:415– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hariharan:1995:FEA

[61] R. Hariharan, S. Kapoor, and V. Kumar. Faster enumeration of all spanning trees of a directed graph. Lecture Notes in Computer Science, 955: 428–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1995:SMS

[62] V. King. A simpler minimum spanning tree verification algorithm. Lecture Notes in Computer Science, 955: 440–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1995:API

[63] P. Berman and T. Fujito. On approximation properties of the independent set problem in degree 3 graphs. Lecture Notes in Computer Science, 955: 449–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1995:ACS

[64] Hoong Chuin Lau. Approximation of constraint satisfaction via local search. Lecture Notes in Computer Science, 955:461–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersson:1995:DRS

[65] A. Andersson and K. Swanson. On the difficulty of range searching. Lecture Notes in Computer Science, 955: 473–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fich:1995:TSS

[66] F. Fich and P. Bro Miltersen. Tables should be sorted (on random access machines). Lecture Notes in Computer Science, 955:482–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armen:1995:ILB

[67] C. Armen and C. Stein. Improved length bounds for the shortest superstring problem. Lecture Notes in Computer Science, 955:494–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bafna:1995:NOL

[68] V. Bafna, B. Narayanan, and R. Ravi. Non-overlapping local alignments. Lecture Notes in Computer Science, 955: 506–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehlhorn:1995:EIG

[69] K. Mehlhorn. Experiences with the implementation of geometric algorithms. Lecture Notes in Computer Science, 955:518–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:EFN

[70] J. Y. B. Lee and P. C. Wong. The effect of function noise, on GP efficiency. Lecture Notes in Computer Science, 956:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whigham:1995:GAL

[71] P. A. Whigham and R. I. McKay. Genetic approaches to learning recursive relations. Lecture Notes in Computer Science, 956:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aytekin:1995:AGP

[72] T. Aytekin, E. E. Korkmaz, and H. A. Guevennir. An application of genetic programming to the 4-OP problem using map-trees. Lecture Notes in Computer Science, 956:28–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rowe:1995:DRG

[73] J. Rowe and I. East. Direct replacement: a genetic algorithm without mutation which avoids deception. Lecture Notes in Computer Science, 956:41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

TuanPham:1995:CEN

[74] Q. Tuan Pham. Competitive evolution: a natural approach to operator selection. Lecture Notes in Computer Science, 956:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:ECC

[75] Z. Zhang, S. Bai, and G.-J. Li. Emergent collective computational abilities in interacting particle systems. Lecture Notes in Computer Science, 956:61–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michalewicz:1995:PEC

[76] Z. Michalewicz. A perspective on evolutionary computation. Lecture Notes in Computer Science, 956:73–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1995:ESP

[77] X. Yao and P. Darwen. An experimental study of N-person iterated prisoner’s dilemma games. Lecture Notes in Computer Science, 956:90–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1995:SAH

[78] H. Chan and P. Mazumder. A systolic architecture for high speed hypergraph partitioning using genetic algorithms. Lecture Notes in Computer Science, 956:109–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1995:DHO

[79] K. P. Wong and S. Y. W. Wong. Development of hybrid optimisation techniques based on genetic algorithms and simulated annealing. Lecture Notes in Computer Science, 956:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1995:DPH

[80] K. P. Wong and S. Y. W. Wong. Development of parallel hybrid optimisation techniques based on genetic algorithms and simulated annealing. Lecture Notes in Computer Science, 956: 155–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinterding:1995:GAC

[81] R. Hinterding and L. Khan. Genetic algorithms for cutting stock problems: with and without contiguity. Lecture Notes in Computer Science, 956:166– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lienig:1995:GGA

[82] J. Lienig and K. Thulasiraman. GAS-BOR: A genetic algorithm for switchbox routing in integrated circuits. Lecture Notes in Computer Science, 956: 187–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seunarine:1995:CSM

[83] D. Seunarine and E. Eberbach. The calculus of self-modifiable algorithm based evolutionary computer network routing. Lecture Notes in Computer Science, 956:201–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sugiyama:1995:ERS

[84] T. Sugiyama, T. Kido, and M. Nakanishi. Evolving robot strategy for open ended game. Lecture Notes in Computer Science, 956:225–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pan:1995:EAA

[85] Z. Pan, L. Kang, J. He, and Y. Liu. An evolutionary approach to adaptive model-building. Lecture Notes in Computer Science, 956:236–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Machado:1995:TNN

[86] A. M. C. Machado and M. F. M. Campos. Training neural networks with influence diagrams. Lecture Notes in Computer Science, 956:245–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindley:1995:BTI

[87] C. A. Lindley. A behavioural theory of intelligent machines as a framework for the analysis of adaptation. Lecture Notes in Computer Science, 956:257– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darwen:1995:ERS

[88] P. J. Darwen and X. Yao. On evolving robust strategies for iterated prisoner’s dilemma. Lecture Notes in Computer Science, 956:276–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piggott:1995:EGG

[89] P. Piggott and F. Suraweera. Encoding graphs for genetic algorithms: an investigation using the minimum spanning tree problem. Lecture Notes in Computer Science, 956:305–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drogoul:1995:WAP

[90] A. Drogoul. When ants play chess (or can strategies emerge from tactical behaviours?). Lecture Notes in Computer Science, 957:13–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkhard:1995:HDA

[91] H.-D. Burkhard. How to define agent properties — or: What is a fair agent? Lecture Notes in Computer Science, 957:28–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagao:1995:EPC

[92] K. Nagao, K. Hasida, and T. Miyata. Emergent planning: A computational architecture for situated behaviour. Lecture Notes in Computer Science, 957:42–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shehory:1995:CFA

[93] O. Shehory and S. Kraus. Coalition formation among autonomous agents: Strategies and complexity. Lecture Notes in Computer Science, 957:56–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ketchpel:1995:CFA

[94] S. P. Ketchpel. Coalition formation among autonomous agents. Lecture Notes in Computer Science, 957:73–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glance:1995:OFS

[95] N. S. Glance and B. A. Huberman. Organizational fluidity and sustainable cooperation. Lecture Notes in Computer Science, 957:89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1995:ECS

[96] J. Liu and K. Sycara. Emergent constraint satisfaction through multiagent coordinated interaction. Lecture Notes in Computer Science, 957:107– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1995:SDT

[97] K. Fischer, N. Kuhn, H. J. Mueller, and J. P. Mueller. Sophisticated and distributed: The transportation domain. Lecture Notes in Computer Science, 957:122–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ephrati:1995:FIE

[98] E. Ephrati and J. S. Rosenschein. A framework for the interleaving of execution and planning for dynamic tasks by multiple agents. Lecture Notes in Computer Science, 957:139–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burmeister:1995:GCC

[99] B. Burmeister, A. Haddadi, and K. Sundermeyer. Generic, configurable, cooperation protocols for multi-agent systems. Lecture Notes in Computer Science, 957:157–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Correa:1995:AAA

[100] M. Correa and H. Coelho. Around the architectural agent approach to model conversations. Lecture Notes in Computer Science, 957:172–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conte:1995:NMO

[101] R. Conte and C. Castelfranchi. Norms as mental objects — from normative beliefs to normative goals. Lecture Notes in Computer Science, 957:186– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourgine:1995:HAC

[102] P. Bourgine. The hedonic agent: A constructivist approach of abductive capacities. Lecture Notes in Computer Science, 957:199–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steiner:1995:CFM

[103] D. Steiner, A. Burt, M. Kolb, and C. Leri. The conceptual framework of MAI02L. Lecture Notes in Computer Science, 957:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manela:1995:DGP

[104] M. Manela and J. A. Campbell. Designing good pursuit problems as testbeds for distributed. Lecture Notes in Computer Science, 957:231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ueberberg:1995:ITP

[105] J. Ueberberg. Interactive theorem proving and computer algebra. Lecture Notes in Computer Science, 958:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rege:1995:PAG

[106] A. Rege and J. Canny. A practical algorithm for geometric theorem proving. Lecture Notes in Computer Science, 958:10–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Homann:1995:CTP

[107] K. Homann and J. Calmet. Combining theorem proving and symbolic mathematical computing. Lecture Notes in Computer Science, 958:18–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapitonova:1995:TSP

[108] Y. V. Kapitonova, A. A. Letichevsky, V. A. Volkov, and M. S. L’vov. Tools for solving problems in the scope of algebraic programming. Lecture Notes in Computer Science, 958:30–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chippendale:1995:PPI

[109] M. Chippendale. Planning a proof of the intermediate value theorem. Lecture Notes in Computer Science, 958: 48–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madden:1995:GTA

[110] P. Madden and I. Green. A general technique for automatically optimizing programs through the use of proof plans. Lecture Notes in Computer Science, 958:64–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1995:DTG

[111] G. Butler. Datalog and two groups and C++. Lecture Notes in Computer Science, 958:80–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jumpertz:1995:LLR

[112] P. Jumpertz. Linear logic and real closed fields: a way to handle situations dynamically. Lecture Notes in Computer Science, 958:93–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1995:PEA

[113] S. Baker and A. Smaill. A proof environment for arithmetic with the Omega rule. Lecture Notes in Computer Science, 958:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Missura:1995:UCP

[114] S. A. Missura and A. Weber. Using commutativity properties for controlling coercions. Lecture Notes in Computer Science, 958:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Missura:1995:TSP

[115] S. A. Missura. Theories = signatures + propositions used as types. Lecture Notes in Computer Science, 958:144– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collart:1995:ISG

[116] S. Collart and D. Mall. The ideal structure of Gröbner base computations. Lecture Notes in Computer Science, 958:156–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfalzgraf:1995:MCA

[117] J. Pfalzgraf, K. Stokkermans, and U. C. Sigmund. Modeling cooperating agents scenarios by deductive planning methods and logical fiberings. Lecture Notes in Computer Science, 958:167– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Telerman:1995:PMC

[118] V. Telerman. Propagation of mathematical constraints in subdefinite models. Lecture Notes in Computer Science, 958:191–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buendgen:1995:CCA

[119] R. Buendgen. Combining computer algebra and rule based reasoning. Lecture Notes in Computer Science, 958: 209–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1995:ASE

[120] S. Tsumoto and H. Tanaka. Algebraic specification of empirical inductive learning methods based on rough sets and matroid theory. Lecture Notes in Computer Science, 958:224– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nettleton:1995:SPU

[121] D. J. Nettleton and R. Garigliano. Subsymbolic processing using adaptive algorithms. Lecture Notes in Computer Science, 958:244–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laita:1995:IPB

[122] L. M. Laita, L. De Ledesma, E. Roanes-Lozano, and E. Roanes-Macias. An interpretation of the propositional Boolean algebra as a k-algebra effective calculus. Lecture Notes in Computer Science, 958:255–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Semenov:1995:SCS

[123] A. Semenov, A. Babichev, and A. Leshchenko. Subdefinite computations and symbolic transformations in the UniCalc solver. Lecture Notes in Computer Science, 958:264–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zwick:1995:CMP

[124] U. Zwick and M. S. Paterson. The complexity of mean payoff games. Lecture Notes in Computer Science, 959:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwama:1995:ACS

[125] K. Iwama and S. Miyazaki. Approximation of coNP sets by NP-Complete sets. Lecture Notes in Computer Science, 959:11–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feng:1995:HDP

[126] Q.-W. Feng, R. F. Cohen, and P. Eades. How to draw a planar clustered graph. Lecture Notes in Computer Science, 959:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calamoneri:1995:EOG

[127] T. Calamoneri and R. Petreschi. An efficient orthogonal grid drawing algorithm for cubic graphs. Lecture Notes in Computer Science, 959:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1995:CIS

[128] S.-W. Cheng and Y.-F. Xu. Constrained independence system and triangulations of planar point sets. Lecture Notes in Computer Science, 959: 41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1995:TDW

[129] C. Wang and B. Zhu. Three dimensional weak visibility: Complexity and applications. Lecture Notes in Computer Science, 959:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1995:RRP

[130] Sung Kwon Kim. Rectangulating rectilinear polygons in parallel. Lecture Notes in Computer Science, 959:61–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamakoti:1995:ERI

[131] V. Kamakoti, K. Krithivasan, and C. Pandu Rangan. Efficient randomized incremental algorithm for the closest pair problem using leafary trees. Lecture Notes in Computer Science, 959:71–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1995:TCO

[132] E. P. F. Chan and R. Van der Meyden. Testing containment of objectoriented conjunctive queries is 0p2-Hard. Lecture Notes in Computer Science, 959:81–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:CIR

[133] R. Zhang. Computing infinite relations using finite expressions: a new approach to the safety issue in relational databases. Lecture Notes in Computer Science, 959:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lima:1995:STU

[134] Seung Jin Lima and Y.-K. Ng. Setterm unification in a logic database language. Lecture Notes in Computer Science, 959:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wild:1995:CFC

[135] M. Wild. Computations with finite closure systems and implications. Lecture Notes in Computer Science, 959:111– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lingas:1995:MTP

[136] A. Lingas. Maximum tree-packing in time O(n050/02). Lecture Notes in Computer Science, 959:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1995:OAF

[137] W. Shi and D. B. West. Optimal algorithms for finding connected components of an unknown graph. Lecture Notes in Computer Science, 959:131– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganley:1995:MWS

[138] J. L. Ganley, M. J. Golin, and J. S. Salowe. The multi-weighted spanning tree problem. Lecture Notes in Computer Science, 959:141–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:AGE

[139] J. Chen. Algorithmic graph embeddings. Lecture Notes in Computer Science, 959:151–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agrawal:1995:AQB

[140] D. Agrawal, O. Egecioglu, and A. El Abbadi. Analysis of quorum-based protocols for distributed (k + 1)- exclusion. Lecture Notes in Computer Science, 959:161–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1995:HFT

[141] X. Lin and M. E. Orlowska. A highly fault-tolerant quorum consensus method for managing replicated data. Lecture Notes in Computer Science, 959:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1995:CCI

[142] G. Huang. Constructing craig interpolation formulas. Lecture Notes in Computer Science, 959:181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawabe:1995:COS

[143] Y. Kawabe and N. lshii. Currying of order-sorted term rewriting systems. Lecture Notes in Computer Science, 959:191–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rengarajan:1995:SQN

[144] S. Rengarajan and C. E. Veni Madhavan. Stack and queue number of 2-Trees. Lecture Notes in Computer Science, 959:203–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walley:1995:SPR

[145] S. K. Walley and H. H. Tan. Shortest paths in random weighted graphs. Lecture Notes in Computer Science, 959: 213–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1995:SRC

[146] X. Zhou and T. Nishizeki. Simple reduction of f-colorings to edgecolorings. Lecture Notes in Computer Science, 959:223–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayase:1995:OSS

[147] K. Hayase, K. Sadakane, and S. Tani. Output-size sensitivities of OBDD. construction through maximal independent set problem. Lecture Notes in Computer Science, 959:229–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tromp:1995:SWB

[148] J. Tromp, L. Zhang, and Y. Zhao. Small weight bases for Hamming codes. Lecture Notes in Computer Science, 959:235–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cassaigne:1995:TWG

[149] J. Cassaigne and J. Karhumaeki. Toeplitz words, generalized periodicity and periodically iterated morphisms. Lecture Notes in Computer Science, 959:244–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barcucci:1995:CEC

[150] E. Barcucci, A. Del Lungo, E. Pergola, and R. Pinzani. A construction for enumerating k-coloured Motzkin paths. Lecture Notes in Computer Science, 959:254–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oleshchuk:1995:PKC

[151] V. A. Oleshchuk. On public-key cryptosystem based on Church–Rosser string-rewriting systems. Lecture Notes in Computer Science, 959:264– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Savage:1995:EHK

[152] J. E. Savage. Extending the Hong-Kung model to memory hierarchies. Lecture Notes in Computer Science, 959:270–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1995:LTA

[153] L. Cai and J. Chen. On log-time alternating Turing machines of alternating depth κ. Lecture Notes in Computer Science, 959:282–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurosawa:1995:NBA

[154] K. Kurosawa and S. Kageyama. New bound for affine resolvable designs and its application to authentication codes. Lecture Notes in Computer Science, 959:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lubachevsky:1995:DPK

[155] B. D. Lubachevsky and R. L. Graham. Dense packings of 3kappa(kappa + 1) + 1 equal disks in a circle for kappa = 1,2,3,4, alibi 5. Lecture Notes in Computer Science, 959:303–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz:1995:EPA

[156] J. Diaz, A. Gibbons, G. Pantziou, and M. Serna. Efficient parallel algorithms for some tree layout problems. Lecture Notes in Computer Science, 959:313–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1995:CAP

[157] Y. Han and X. Shen. Conservative algorithms for parallel and sequential integer sorting. Lecture Notes in Computer Science, 959:324–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:OAP

[158] Z. Chen. An optimal algorithm for proper learning of unions of two rectangles with queries. Lecture Notes in Computer Science, 959:334–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:DNC

[159] Z. Chen. Disjunctions of negated counting functions are efficiently learnable with equivalence queries. Lecture Notes in Computer Science, 959:344– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1995:NEC

[160] S. Wu. Non-empty cross-3-intersection theorems of subsets. Lecture Notes in Computer Science, 959:350–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bo:1995:CMT

[161] Y. Bo. Convexity of minimal total dominating functions in graphs. Lecture Notes in Computer Science, 959:357– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duparc:1995:TMP

[162] J. H. Duparc and P. Rolland. Transformations for maximal planar graphs with minimum degree five. Lecture Notes in Computer Science, 959:366– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

You:1995:APM

[163] Z. You and C. Wang. An asynchronous parallel method for linear systems. Lecture Notes in Computer Science, 959:372–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:KSP

[164] X. Zhang. On a kind of sequence of polynomials. Lecture Notes in Computer Science, 959:379–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meng:1995:HCG

[165] J. Meng. Hamiltonian cycles in 2generated Cayley digraphs of Abelian groups. Lecture Notes in Computer Science, 959:384–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1995:PMS

[166] C.-X. Xu and Z.-W. Lu. Pandiagonal magic squares. Lecture Notes in Computer Science, 959:388–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1995:PQM

[167] H. Wang and Z.-Y. You. PFFM and quasi-morishima matrices. Lecture Notes in Computer Science, 959:392– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1995:EFT

[168] C. F. Chang, J. X. Chang, X. C. Liu, and P. C. B. Lam. Edge-face total chromatic number of outerplanar graphs with cap delta(G) = 6. Lecture Notes in Computer Science, 959:396–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuler:1995:SCP

[169] R. Schuler and T. Yamakami. Sets computable in polynomial time on average. Lecture Notes in Computer Science, 959:400–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belanger:1995:RDD

[170] J. Belanger and J. Wang. Rankable distributions do not provide harder instances than uniform distributions. Lecture Notes in Computer Science, 959:410–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1995:TPM

[171] K. Kobayashi. Transformations that preserve malignness of universal distributions. Lecture Notes in Computer Science, 959:420–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemaspaandra:1995:ISB

[172] L. A. Hemaspaandra and J. Rothe. Intersection suffices for Boolean hierarchy equivalence. Lecture Notes in Computer Science, 959:430–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelley:1995:LCA

[173] D. Kelley, P.-J. Wan, and Q. Yang. A 3/2 log 3-Competitive algorithm for the counterfeit coin problem. Lecture Notes in Computer Science, 959:436– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1995:SRD

[174] S. Carlsson and J. Chen. Searching rigid data structures. Lecture Notes in Computer Science, 959:446–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:BSM

[175] B.-T. Yang. A better subgraph of the minimum weight triangulation. Lecture Notes in Computer Science, 959:452– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gao:1995:SDM

[176] W. Gao and B. Yin. Sequence decomposition method for computing a Gröbner basis and its application to bivariate splines. Lecture Notes in Computer Science, 959:456–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bai:1995:BAO

[177] L. Bai, P. M. Yamakawa, H. Ebara, and H. Nakano. A broadcasting algorithm oil tile arrangement graph. Lecture Notes in Computer Science, 959: 462–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shiau:1995:FMF

[178] S.-H. Shiau and C.-B. Fang. A fast maximum finding algorithm on broadcast communication. Lecture Notes in Computer Science, 959:472–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shastri:1995:BGN

[179] A. Shastri. Broadcasting in general networks I: Trees. Lecture Notes in Computer Science, 959:482–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chern:1995:UDA

[180] S.-C. Chern, T.-C. Tuan, and J.-S. Jwo. Uni-directional alternating group graphs. Lecture Notes in Computer Science, 959:490–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakurai:1995:SPK

[181] K. Sakurai. On separating proofs of knowledge from proofs of membership of languages and its application to secure identification schemes. Lecture Notes in Computer Science, 959:496– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krumke:1995:CLP

[182] S. O. Krumke, H. Noltemeier, S. S. Ravi, and M. V. Marathe. Compact location problems with budget and communication constraints. Lecture Notes in Computer Science, 959: 510–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1995:MDS

[183] S.-W. Cheng, M. Kaminski, and S. Zaks. Minimum dominating sets of intervals on lines. Lecture Notes in Computer Science, 959:520–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1995:TDP

[184] Y. Choi and T. W. Lam. Twodimensional pattern matching on a dynamic library of texts. Lecture Notes in Computer Science, 959:530–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzi:1995:SAC

[185] P. Crescenzi, V. Kann, R. Silvestri, and L. Trevisan. Structure in approximation classes. Lecture Notes in Computer Science, 959:539–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peinado:1995:ILB

[186] M. Peinado. Improved lower bounds for the randomized Boppana-Halldorsson algorithm for MAXCLIQUE. Lecture Notes in Computer Science, 959:549– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1995:MCN

[187] Q. Cheng and H. Zhu. MNP: A class of NP optimization problems. Lecture Notes in Computer Science, 959:559– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bacik:1995:SPA

[188] R. Bacik and S. Mahajan. Semidefinite programming and its applications to NP problems. Lecture Notes in Computer Science, 959:566–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hui:1995:AEL

[189] L. C.-K. Hui and K.-Y. Lam. Analysis and experimentation on list update algorithms. Lecture Notes in Computer Science, 959:576–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khoury:1995:EBB

[190] B. N. Khoury and P. M. Pardalos. An exact branch and bound algorithm for the Steiner problem in graphs. Lecture Notes in Computer Science, 959:582–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1995:PMS

[191] W. Huang, W. Li, W. Lu, and Y. Zhang. A physical model for the satisfiability problem. Lecture Notes in Computer Science, 959:591–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1995:EAL

[192] Sam Myo Kim and R. McNaughton. An efficient algorithm for local testability problem of finite state automata. Lecture Notes in Computer Science, 959: 597–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1995:STT

[193] X. Yu and M. Yung. Scheduling tasktree with additive scales on parallel/ distributed machines. Lecture Notes in Computer Science, 959:607–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1995:SVS

[194] C.-I. Chan and G. H. Young. Singlevehicle scheduling problem on a straight line with time window constraints. Lecture Notes in Computer Science, 959:617–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1995:LAS

[195] R. Li and L. Shi. An on-line algorithm for some uniform processor scheduling. Lecture Notes in Computer Science, 959:627–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeavons:1995:ACT

[196] P. Jeavons and D. Cohen. An algebraic characterization of tractable constraints. Lecture Notes in Computer Science, 959:633–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:1995:LPU

[197] J. Ding, C. Qing, and G. Song. Limit property of unbalanced development in economic network. Lecture Notes in Computer Science, 959:643–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1995:MCS

[198] T. Jiang. Matching and comparing sequences in molecular biology. Lecture Notes in Computer Science, 959:648– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vazirani:1995:PDS

[199] V. V. Vazirani. Primal-dual schema based approximation algorithms. Lecture Notes in Computer Science, 959: 650–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardone:1995:SFF

[200] F. Cardone. Strict finitism and feasibility. Lecture Notes in Computer Science, 960:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parikh:1995:LO

[201] R. Parikh. Logical omniscience. Lecture Notes in Computer Science, 960:22–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sazonov:1995:FN

[202] V. Y. Sazonov. On feasible numbers. Lecture Notes in Computer Science, 960:30–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloch:1995:PHR

[203] S. Bloch. On parallel hierarchies and R0ik. Lecture Notes in Computer Science, 960:52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1995:PEC

[204] U. Berger and H. Schwichtenberg. Program extraction from classical proofs. Lecture Notes in Computer Science, 960:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clote:1995:CMF

[205] P. Clote. Computational models and function algebras. Lecture Notes in Computer Science, 960:98–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Constable:1995:ECC

[206] R. L. Constable. Expressing computational complexity in constructive type theory. Lecture Notes in Computer Science, 960:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girard:1995:LLL

[207] J.-Y Girard. Light linear logic. Lecture Notes in Computer Science, 960: 145–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leivant:1995:ITC

[208] D. Leivant. Intrinsic theories and computational complexity. Lecture Notes in Computer Science, 960:177–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buss:1995:HT

[209] S. R. Buss. On Herbrand’s theorem. Lecture Notes in Computer Science, 960:195–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krajicek:1995:SCC

[210] J. Krajicek and P. Pudlak. Some consequences of cryptographical conjectures for S012 and EF. Lecture Notes in Computer Science, 960:210–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takeuti:1995:FPS

[211] G. Takeuti. Frege proof system and TNC. Lecture Notes in Computer Science, 960:221–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellantoni:1995:CPT

[212] S. J. Bellantoni. Characterizing parallel time by type 2 recursions with polynomial output length. Lecture Notes in Computer Science, 960:253–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seth:1995:TPH

[213] A. Seth. Type 2 polynomial hierarchies. Lecture Notes in Computer Science, 960:269–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wainer:1995:HTR

[214] S. S. Wainer. The hierarchy of terminating recursive programs over N. Lecture Notes in Computer Science, 960: 281–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cenzer:1995:FCM

[215] D. Cenzer and J. B. Remmel. Feasibly categorical models. Lecture Notes in Computer Science, 960:300–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graedel:1995:MMT

[216] E. Graedel and Y. Gurevich. Metafinite model theory. Lecture Notes in Computer Science, 960:313–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khoussainov:1995:APS

[217] B. Khoussainov and A. Nerode. Automatic presentations of structures. Lecture Notes in Computer Science, 960: 367–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawar:1995:RSO

[218] A. Dawar. A restricted second order logic for finite structures. Lecture Notes in Computer Science, 960:393– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fagin:1995:CPM

[219] R. Fagin. Comparing the power of monadic NP games. Lecture Notes in Computer Science, 960:414–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grumbach:1995:LCQ

[220] S. Grumbach, J. Su, and C. Tollu. Linear constraint query languages: Expressive power and complexity. Lecture Notes in Computer Science, 960: 426–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindell:1995:CSS

[221] S. Lindell. A constant-space sequential model of computation for firstorder logic. Lecture Notes in Computer Science, 960:447–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makowsky:1995:LCR

[222] J. A. Makowsky and Y. B. Pnueli. Logics capturing relativized complexity classes uniformly. Lecture Notes in Computer Science, 960:463–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosen:1995:PTF

[223] E. Rosen and S. Weinstein. Preservation theorems in finite model theory. Lecture Notes in Computer Science, 960:480–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suciu:1995:QLN

[224] D. Suciu and V. Breazu-Tannen. A query language for NC. Lecture Notes in Computer Science, 960:503– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiehagen:1995:LC

[225] R. Wiehagen and T. Zeugmann. Learning and consistency. Lecture Notes in Computer Science, 961:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1995:EDI

[226] R. Freivalds, E. B. Kinber, and R. Wiehagen. Error detecting in inductive inference. Lecture Notes in Computer Science, 961:25–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1995:LGE

[227] R. Freivalds, E. B. Kinber, and R. Wiehagen. Learning from good examples. Lecture Notes in Computer Science, 961:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daley:1995:TRA

[228] R. Daley and B. Kalyanasundaram. Towards reduction arguments for FINite learning. Lecture Notes in Computer Science, 961:63–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Case:1995:NMS

[229] J. Case, M. Suraj, and S. Jain. Notso-nearly-minimal-size program inference. Lecture Notes in Computer Science, 961:76–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1995:OPI

[230] A. Ambainis. Optimization problem in inductive inference. Lecture Notes in Computer Science, 961:96–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1995:ITP

[231] S. Jain and A. Sharma. On identification by teams and probabilistic machines. Lecture Notes in Computer Science, 961:108–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apsitis:1995:TCC

[232] K. Apsitis. Topological considerations in composing teams of learning machines. Lecture Notes in Computer Science, 961:146–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1995:PVD

[233] R. Freivalds, E. B. Kinber, and C. H. Smith. Probabilistic versus deterministic memory limited learning. Lecture Notes in Computer Science, 961:155– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasarch:1995:CUI

[234] W. I. Gasarch, M. G. Pleszkoch, and M. Velauthapillai. Classification using information. Lecture Notes in Computer Science, 961:162–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiehagen:1995:CRP

[235] R. Wiehagen, C. H. Smith, and T. Zeugmann. Classifying recursive predicates and languages. Lecture Notes in Computer Science, 961:174– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeugmann:1995:GTA

[236] T. Zeugmann and S. Lange. A guided tour across the boundaries of learning recursive languages. Lecture Notes in Computer Science, 961:190–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shinohara:1995:PI

[237] T. Shinohara and S. Arikawa. Pattern inference. Lecture Notes in Computer Science, 961:259–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1995:ILR

[238] P. Watson. Inductive learning of recurrence-term languages from positive data. Lecture Notes in Computer Science, 961:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takada:1995:LFL

[239] Y. Takada. Learning formal languages based on control sets. Lecture Notes in Computer Science, 961:316–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Globig:1995:LCB

[240] C. Globig and S. Wess. Learning in case-based classification algorithms. Lecture Notes in Computer Science, 961:340–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Posthoff:1995:OSL

[241] C. Posthoff and M. Schlosser. Optimal strategies — learning from examples — Boolean equations. Lecture Notes in Computer Science, 961:363–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehlsam:1995:FCD

[242] G. Mehlsam, H. Kaindl, and W. Barth. Feature construction during tree learning. Lecture Notes in Computer Science, 961:391–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albrecht:1995:LBD

[243] A. Albrecht. On lower bounds for the depth of threshold circuits with weights from 0 − 1, 0, +1. Lecture Notes in Computer Science, 961:404–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pippig:1995:SNN

[244] E. Pippig. Structuring neural networks and PAC-Learning. Lecture Notes in Computer Science, 961:417–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldammer:1995:ISR

[245] U. Goldammer. Inductive synthesis of rewrite programs. Lecture Notes in Computer Science, 961:435–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grieser:1995:TTR

[246] G. Grieser. TLP0S — a term rewriting laboratory (not only) for experiments in automatic program synthesis. Lecture Notes in Computer Science, 961:467–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beick:1995:GLP

[247] H.-R. Beick and V. Stankov. GoslerP — A logic programming tool for inductive inference. Lecture Notes in Computer Science, 961:482–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stirling:1995:LMC

[248] C. Stirling. Local model checking games. Lecture Notes in Computer Science, 962:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dam:1995:CPS

[249] M. Dam. Compositional proof systems for model checking infinite state processes. Lecture Notes in Computer Science, 962:12–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laroussinie:1995:CMC

[250] F. Laroussinie and K. G. Larsen. Compositional model checking of real time systems. Lecture Notes in Computer Science, 962:27–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montanari:1995:CBF

[251] U. Montanari and M. Pistore. Checking bisimilarity for finitary π-calculus. Lecture Notes in Computer Science, 962: 42–56, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1995:WLC

[252] G.-L. Ferrari, U. Montanari, and P. Quaglia. The weak late π-calculus semantics as observation equivalence. Lecture Notes in Computer Science, 962:57–71, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steffen:1995:FAM

[253] B. Steffen, A. Classen, M. Klein, and J. Knoop. The fixpoint-analysis machine. Lecture Notes in Computer Science, 962:72–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1995:UFI

[254] H. Lin. Unique fixpoint induction for mobile processes. Lecture Notes in Computer Science, 962:88–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1995:PTS

[255] X. Liu and D. Walker. A polymorphic type system for the polyadic πcalculus. Lecture Notes in Computer Science, 962:103–116, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermida:1995:FCS

[256] C. Hermida and J. Power. Fibrational control structures. Lecture Notes in Computer Science, 962:117–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1995:FAM

[257] F. Corradini, R. De Nicola, and A. Labella. Fully abstract models for nondeterministic regular expressions. Lecture Notes in Computer Science, 962: 130–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busi:1995:PNS

[258] N. Busi and R. Gorrieri. A Petri net semantics for π-calculus. Lecture Notes in Computer Science, 962:145–159, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rensink:1995:CTD

[259] A. Rensink. A complete theory of deterministic event structures. Lecture Notes in Computer Science, 962:160– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1995:CBC

[260] M. Nielsen, L. Priese, and V. Sassone. Characterising behavioural congruences for Petri nets. Lecture Notes in Computer Science, 962:175– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaandrager:1995:VDS

[261] F. Vaandrager. Verification of a distributed summation algorithm. Lecture Notes in Computer Science, 962:190– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groote:1995:CPV

[262] J. F. Groote and M. P. A. Sellink. Confluence for process verification. Lecture Notes in Computer Science, 962:204–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ulidowski:1995:AWE

[263] I. Ulidowski. Axiomatisations of weak equivalences for De Simone languages. Lecture Notes in Computer Science, 962:219–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Segala:1995:CTB

[264] R. Segala. A compositional tracebased semantics for probabilistic automata. Lecture Notes in Computer Science, 962:234–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nuthez:1995:ATP

[265] M. Nuthez, D. De Frutos, and L. Llana. Acceptance trees for probabilistic processes. Lecture Notes in Computer Science, 962:249–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harel:1995:WPW

[266] D. Harel. Will I be pretty, will I be rich?: Some thoughts on theory vs. practice in systems engineering. Lecture Notes in Computer Science, 962:264–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryans:1995:TDS

[267] J. Bryans, J. Davies, and S. Schneider. Towards a denotational semantics for ET-LOTOS. Lecture Notes in Computer Science, 962:269–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1995:RAP

[268] F. Wang. Reachability analysis at procedure level through timing coincidence. Lecture Notes in Computer Science, 962:284–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1995:FAS

[269] W. Vogler. Faster asynchronous systems. Lecture Notes in Computer Science, 962:299–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brinksma:1995:FT

[270] E. Brinksma, A. Rensink, and W. Vogler. Fair testing. Lecture Notes in Computer Science, 962:313– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Craigen:1995:FMT

[271] D. Craigen. Formal methods technology transfer: Impediments and innovation. Lecture Notes in Computer Science, 962:328–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdulla:1995:DSB

[272] P. A. Abdulla and M. Kindahl. Decidability of simulation and bisimulation between loss channel systems and finite state systems. Lecture Notes in Computer Science, 962:333–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jancar:1995:CRP

[273] P. Jancar and F. Moller. Checking regular properties of Petri nets. Lecture Notes in Computer Science, 962:348– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonsangue:1995:MPT

[274] M. M. Bonsangue, J. N. Kok, and E. De Vink. Metric predicate transformers: Towards a notion of refinement for concurrency. Lecture Notes in Computer Science, 962:363–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siegel:1995:RTS

[275] M. Siegel. A refinement theory that supports both “decrease of nondeterminism” and “increase of parallelism”. Lecture Notes in Computer Science, 962:378–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerson:1995:MCE

[276] E. A. Emerson. Model checking and efficient automation of temporal reasoning. Lecture Notes in Computer Science, 962:393–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1995:VPN

[277] E. M. Clarke, O. Grumberg, and S. Jha. Verifying parameterised networks using abstraction and regular languages. Lecture Notes in Computer Science, 962:395–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kupferman:1995:CBM

[278] O. Kupferman and M. Y. Vardi. On the complexity of branching modular model checking. Lecture Notes in Computer Science, 962:408–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaivola:1995:ALT

[279] R. Kaivola. Axiomatising linear time mu-calculus. Lecture Notes in Computer Science, 962:423–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thiagarajan:1995:TCS

[280] P. S. Thiagarajan. A trace consistent subset of PTL. Lecture Notes in Computer Science, 962:438–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzmann:1995:PPC

[281] G. J. Holzmann. Proving properties of concurrent systems with SPIN. Lecture Notes in Computer Science, 962: 453–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Philippou:1995:SDC

[282] A. Philippou and D. Walker. On sharing and determinacy in concurrent systems. Lecture Notes in Computer Science, 962:456–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brock:1995:PSG

[283] S. Brock and G. Ostheimer. Process semantics of graph reduction. Lecture Notes in Computer Science, 962:471– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hennessy:1995:BCB

[284] M. Hennessy and J. Rathke. Bisimulations for a calculus of broadcasting systems. Lecture Notes in Computer Science, 962:486–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DArgenio:1995:DCP

[285] P. R. D’Argenio and S. Mauw. Delayed choice for process algebra with abstraction. Lecture Notes in Computer Science, 962:501–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ceraens:1995:CCT

[286] K. Ceraens. CTR: A calculus of timed refinement. Lecture Notes in Computer Science, 962:516–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1995:TLT

[287] A. Bouajjani and Y. Lakhnech. Temporal logic + timed automata: Expressiveness and decidability. Lecture Notes in Computer Science, 962:531– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preneel:1995:MMB

[288] Bart Preneel and Paul C. van Oorschot. MDx-MAC and building fast MACs from hash functions. Lecture Notes in Computer Science, 963: 1–14, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630001.
pdf.

Bellare:1995:XMN

[289] Mihir Bellare, Roch Guérin, and Phillip Rogaway. XOR MACs: New methods for message authentication using finite pseudorandom functions. Lecture Notes in Computer Science, 963: 15–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630015.
pdf.

Rogaway:1995:BHA

[290] Phillip Rogaway. Bucket hashing and its application to fast message authentication. Lecture Notes in Computer Science, 963:29–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630029.
pdf.

Schroeppel:1995:FKE

[291] Richard Schroeppel, Hilarie Orman, Sean O’Malley, and Oliver Spatscheck. Fast key exchange with elliptic curve systems. Lecture Notes in Computer Science, 963:43– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630043.
pdf.

Beguin:1995:FSA

[292] Philippe Béguin and Jean-Jacques Quisquater. Fast server-aided RSA signatures secure against active attacks. Lecture Notes in Computer Science, 963:57–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0963/09630057.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0963/09630057.
pdf.

Lim:1995:SPS

[293] Chae Hoon Lim and Pil Joong Lee. Security and performance of server-aided RSA computation protocols. Lecture Notes in Computer Science, 963:70– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630070.
pdf.

Halevi:1995:ECS

[294] Shai Halevi. Efficient commitment schemes with bounded sender and unbounded receiver. Lecture Notes in Computer Science, 963:84– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630084.
pdf.

Beaver:1995:POT

[295] Donald Beaver. Precomputing oblivious transfer. Lecture Notes in Computer Science, 963:97–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630097.
pdf.

Crepeau:1995:COT

[296] Claude Crépeau, Jeroen van de Graaf, and Alain Tapp. Committed oblivious transfer and private multi-party computation. Lecture Notes in Computer Science, 963:110– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630110.
                                        pdf.

Mayers:1995:SQO

[297] Dominic Mayers. On the security of the quantum oblivious transfer and key distribution protocols. Lecture Notes in Computer Science, 963:124– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630124.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630124.
                                        pdf.

Theobald:1995:HBS

[298] Thorsten Theobald. How to break Shamir’s asymmetric basis. Lecture Notes in Computer Science, 963:136– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630136.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630136.
                                        pdf.

Park:1995:SGC

[299] Sang-Joon Park, Sang-Jin Lee, and Seung-Cheol Goh. On the security of the gollmann cascades. Lecture Notes in Computer Science, 963:148– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630148.
pdf.

Ohta:1995:ISA

[300] Kazuo Ohta, Shiho Moriai, and Kazumaro Aoki. Improving the search algorithm for the best linear expression. Lecture Notes in Computer Science, 963:157–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630157.
pdf.

Kaliski:1995:DLC

[301] Burton S. Kaliski Jr. and Yiqun Lisa Yin. On differential and linear cryptanalysis of the RC5 encryption algorithm. Lecture Notes in Computer Science, 963:171–183, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630171.
pdf.

Micali:1995:SMG

[302] Silvio Micali and Ray Sidney. A simple method for generating and sharing pseudo-random functions, with applications to Clipper-like key escrow systems. Lecture Notes in Computer Science, 963:185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630185.
                                        pdf.

Lenstra:1995:KES

[303] Arjen K. Lenstra, Peter Winkler, and Yacov Yacobi. A key escrow system with warrant bounds. Lecture Notes in Computer Science, 963:197– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630197.
                                        pdf.

Kilian:1995:FCR

[304] Joe Kilian and Tom Leighton. Fair cryptosystems, revisited. Lecture Notes in Computer Science, 963:208– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630208.
                                        pdf.

Frankel:1995:EES

[305] Yair Frankel and Moti Yung. Escrow encryption systems visited: Attacks, analysis and designs. Lecture Notes in Computer Science, 963:222– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630222.
pdf.

Anderson:1995:RPP

[306] Ross Anderson and Roger Needham. Robustness principles for public key protocols. Lecture Notes in Computer Science, 963:236–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630236.
pdf.

Patarin:1995:CMI

[307] Jacques Patarin. Cryptanalysis of the Matsumoto and Imai public key scheme of Eurocrypt ’88. Lecture Notes in Computer Science, 963:248– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630248.
pdf.

Klapper:1995:CBA

[308] Andrew Klapper and Mark Goresky. Cryptanalysis based on 2-Adic rational approximation. Lecture Notes in Computer Science, 963:262–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630262.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630262.
                                        pdf.

Knudsen:1995:KSW

[309] Lars R. Knudsen. A key-schedule weakness in SAFER K-64. Lecture Notes in Computer Science, 963:274– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630274.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630274.
                                        pdf.

Frankel:1995:CIL

[310] Yair Frankel and Moti Yung. Cryptanalysis of the immunized LL public key systems. Lecture Notes in Computer Science, 963:287–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630287.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630287.
                                        pdf.

Cramer:1995:SSS

[311] Ronald Cramer and Ivan Bjerre Damgård. Secure signature schemes based on interactive protocols. Lecture Notes in Computer Science, 963: 297–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/0963/09630297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630297.
pdf.

Kilian:1995:IEA

[312] Joe Kilian. Improved efficient arguments. Lecture Notes in Computer Science, 963:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0963/09630311.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0963/09630311.pdf.

Damgaard:1995:HVV

[313] Ivan Bjerre Damgård, Oded Goldreich, Tatsuaki Okamoto, and Avi Wigderson. Honest verifier vs. dishonest verifier in public coin zeroknowledge proofs. Lecture Notes in Computer Science, 963:325–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630325.
pdf.

Herzberg:1995:PSS

[314] Amir Herzberg, Stanislaw Jarecki, Hugo Krawczyk, and Moti Yung. Proactive secret sharing or: How to cope with perpetual leakage. Lecture Notes in Computer Science, 963:339– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/0963/09630339.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630339.
                                        pdf.

Beimel:1995:SSP

[315] Amos Beimel and Benny Chor. Secret sharing with public reconstruction. Lecture Notes in Computer Science, 963:353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0963/09630353.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630353.
                                        pdf.

Blakley:1995:GPS

[316] G. R. Blakley and Gregory A. Kabatianski. On general perfect secret sharing schemes. Lecture Notes in Computer Science, 963:367–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0963/09630367.
                                        pdf.

Dodson:1995:NFL

[317] Bruce Dodson and Arjen K. Lenstra. NFS with four large primes: An explosive experiment. Lecture Notes in Computer Science, 963:372–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0963/09630372.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/0963/09630372.
pdf.

Bleichenbacher:1995:SRL

[318] Daniel Bleichenbacher, Wieb Bosma, and Arjen K. Lenstra. Some remarks on Lucas-based cryptosystems. Lecture Notes in Computer Science, 963:386– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630386.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630386.
pdf.

Langford:1995:TDS

[319] Susan K. Langford. Threshold DSS signatures without a trusted party. Lecture Notes in Computer Science, 963: 397–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630397.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630397.
pdf.

Kurosawa:1995:CIT

[320] Kaoru Kurosawa, Satoshi Obana, and Wakaha Ogata. t-cheater identifiable (k, n) threshold secret sharing schemes. Lecture Notes in Computer Science, 963:410–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630410.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/0963/09630410.
pdf.

Boneh:1995:QCH

[321] Dan Boneh and Richard J. Lipton. Quantum cryptanalysis of hidden linear functions. Lecture Notes in Computer Science, 963:424–437, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630424.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630424.
pdf;   http://theory.stanford.edu/
~dabo/papers/quantum.ps.gz.

Okamoto:1995:EDE

[322] Tatsuaki Okamoto. An efficient divisible electronic cash scheme. Lecture Notes in Computer Science, 963:438– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0963/09630438.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0963/09630438.
pdf.

Boneh:1995:CSF

[323] Dan Boneh and James Shaw. Collusionsecure fingerprinting for digital data. Lecture Notes in Computer Science, 963:452–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0963/09630452.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0963/09630452.
pdf.

Anonymous:1995:AI

[324] Anonymous. Author index. Lecture Notes in Computer Science, 963:467– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/0963/0963indx.pdf.

Achasova:1995:SAC

[325] S. M. Achasova. Synchronousasynchronous cellular computations. Lecture Notes in Computer Science, 964:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agibalov:1995:PCF

[326] G. P. Agibalov. Parallel computations and finite automata on semilattices. Lecture Notes in Computer Science, 964:7–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anisimov:1995:LFF

[327] A. V. Anisimov. Linear Fibonacci forms and parallel algorithms for high dimension arithmetic. Lecture Notes in Computer Science, 964:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bandman:1995:CNC

[328] O. L. Bandman. Cellular-neural computations. formal model and possible applications. Lecture Notes in Computer Science, 964:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bianchi:1995:EVT

[329] A. Bianchi, S. Coluccini, P. Degano, and C. Priami. An efficient verifier of truly concurrent properties. Lecture Notes in Computer Science, 964:36–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasperoni:1995:OLS

[330] F. Gasperoni, U. Schwiegelshohn, and J. Turek. Optimal loop scheduling on multiprocessors: A pumping lemma for p-processor schedules. Lecture Notes in Computer Science, 964:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmann:1995:PDP

[331] P. Hartmann. Parallel and distributed processing of cellular hypergraphs. Lecture Notes in Computer Science, 964: 57–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markova:1995:CMC

[332] V. Markova and S. Piskunov. Computer models of 3D cellular structures. Lecture Notes in Computer Science, 964:70–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nepomniaschaya:1995:CTM

[333] A. S. Nepomniaschaya. Comparison of two MST algorithms for associative parallel processors. Lecture Notes in Computer Science, 964:85–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nepomniaschy:1995:PNM

[334] V. A. Nepomniaschy, G. I. Alekseev, A. V. Bystrov, and T. G. Churina. Petri net modelling of Estelle-specified communication protocols. Lecture Notes in Computer Science, 964:94–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnor:1995:DSP

[335] B. Schnor. Dynamic scheduling of parallel applications. Lecture Notes in Computer Science, 964:109–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vesselovski:1995:MAC

[336] G. Vesselovski and M. Kupriyanova. A method for analyzing combinatorial properties of static connecting topologies. Lecture Notes in Computer Science, 964:117–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adamovich:1995:CIL

[337] A. I. Adamovich. cT: An imperative language with parallelizing features supporting the computation model “autotransformation of the evaluation network”. Lecture Notes in Computer Science, 964:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:1995:VFA

[338] S. Benkner. Vienna Fortran 90 — an advanced data parallel language. Lecture Notes in Computer Science, 964: 142–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doroshenko:1995:PAS

[339] A. E. Doroshenko. Programming abstracts for synchronization and communication in parallel programs. Lecture Notes in Computer Science, 964:157– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evstigneev:1995:PMSa

[340] V. A. Evstigneev and V. N. Kasyanov. A program manipulation system for fine-grained architectures. Lecture Notes in Computer Science, 964:163– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hochberger:1995:CCD

[341] C. Hochberger, R. Hoffmann, and S. Waldschmidt. Compilation of CDL for different target architectures. Lecture Notes in Computer Science, 964: 169–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hondroudakis:1995:PEV

[342] A. Hondroudakis, R. Procter, and K. Shanmugam. Performance evaluation and visualization with VIS-PAT. Lecture Notes in Computer Science, 964:180–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huesler:1995:DPI

[343] R. Huesler, H. Vonder Muehll, A. Gunzinger, and G. Troester. Dataparallel programming with intelligent communication. Lecture Notes in Computer Science, 964:186–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurson:1995:OSE

[344] A. R. Hurson and B.-U. Jun. Optimization scheme on execution of logic program in a dataflow environment. Lecture Notes in Computer Science, 964: 204–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borshchev:1995:CTD

[345] A. V. Borshchev, Y. G. Karpov, and V. V. Roudakov. COVERS — A tool for the design of real-time concurrent systems. Lecture Notes in Computer Science, 964:219–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klar:1995:SPZ

[346] R. Klar and P. Dauphin. Status and prospect of ZM4/SIMPLE/PEPP: an event-oriented evaluation environment for parallel and distributed programs. Lecture Notes in Computer Science, 964:234–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klimova:1995:SDD

[347] O. V. Klimova. The separating decomposition of discrete Fourier transform and vectorization of its calculation. Lecture Notes in Computer Science, 964:241–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludwig:1995:PPF

[348] T. Ludwig and S. Lamberts. PFSLib — A parallel file system for workstation clusters. Lecture Notes in Computer Science, 964:246–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saad:1995:DSC

[349] Y. Saad and A. V. Malevsky. Data structures, computational, and communication kernels for distributed memory sparse iterative solvers. Lecture Notes in Computer Science, 964: 252–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oezerdim:1995:PMF

[350] A. Oezerdim, M. O. Uenalir, O. Dikenelli, and E. Ozkarahan. PARMA: A multiattribute file structure for parallel database machines. Lecture Notes in Computer Science, 964:258–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pic:1995:OOL

[351] M. M. Pic, H. Essafi, M. Viala, and L. Nicolas. T++: An objectoriented language to express task and data parallelism on Multi-SIMD computers. Lecture Notes in Computer Science, 964:273–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vlassov:1995:MEP

[352] V. Vlassov, H. Ahmed, and L.-E. Thorelli. mEDA-2: An extension of PVM. Lecture Notes in Computer Science, 964:288–293, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vazhenin:1995:PIS

[353] A. Vazhenin and V. Morozov. Parallel iterative solution of systems of linear equations with dynamically changed length of operands. Lecture Notes in Computer Science, 964:294–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Royak:1995:PCC

[354] M. Royak, E. Shurina, Y. Soloveichik, and V. Malyshkin. Parallelization of computer code MASTAC threedimensional finite elements method implementing. Lecture Notes in Computer Science, 964:304–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barriga:1995:NTS

[355] L. Barriga and R. Ayani. New trends in simulation of distributed shared memory architectures. Lecture Notes in Computer Science, 964:314–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Etiemble:1995:SMV

[356] D. Etiemble and C. Germain. Standard microprocessors versus custom processing elements for massively parallel architectures. Lecture Notes in Computer Science, 964:320–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goossens:1995:FPM

[357] B. Goossens and D. T. Vu. Further pipelining and multithreading to improve RISC processor speed. A proposed architecture and simulation results. Lecture Notes in Computer Science, 964:326–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zabrodin:1995:MPC

[358] A. V. Zabrodin, V. K. Levin, and V. V. Korneev. The massively parallel computer system MBC-100. Lecture Notes in Computer Science, 964: 341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srini:1995:DSM

[359] V. P. Srini. DFS-Super MPx: Low-cost parallel processing system for machine vision and image processing. Lecture Notes in Computer Science, 964:356– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bussler:1995:AID

[360] C. Bussler, S. Jablonski, T. Kirsche, and H. Schuster. Architectural issues of distributed workflow management systems. Lecture Notes in Computer Science, 964:370–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bessonov:1995:PSN

[361] O. Bessonov, V. Brailovskaya, V. Polezhaev, and B. Roux. Parallelization of the solution of 3D Navier–Stokes equations for fluid flow in a cavity with moving covers. Lecture Notes in Computer Science, 964:385–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1995:DSK

[362] P. Ciancarini and P. Mancini. Distributing search and knowledge using a coordination language. Lecture Notes in Computer Science, 964:400– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caudal:1995:DEM

[363] F. Caudal and B. Lecussan. Design and evaluation of a multi-threaded architecture for parallel graph reduction. Lecture Notes in Computer Science, 964:411–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mamedova:1995:IMM

[364] I. G. Mamedova. Implementation of the multigrid method for solving the boundary-value problems for the Poisson and Helmholtz equations on the massively parallel computers. Lecture Notes in Computer Science, 964:427– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kremlev:1995:PSD

[365] A. Kremlev, O. Monakhov, and T. Thiel. Parallel seismic data processing method for MEMSY multiprocessor system. Lecture Notes in Computer Science, 964:434–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ozkarahan:1995:HSP

[366] E. Ozkarahan. Hardware and software platform for information processing. Lecture Notes in Computer Science, 964:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kessy:1995:NSR

[367] E. Kessy, A. Stoukov, and D. Vandromme. Numerical simulation of reacting mixing layer with a parallel implementation. Lecture Notes in Computer Science, 964:453–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fet:1995:PCR

[368] Y. I. Fet and D. A. Pospelov. Parallel computing in russia. Lecture Notes in Computer Science, 964:464–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haendler:1995:EAP

[369] W. Haendler. Early approaches to parallel processing: Increasing performance and dependability. Lecture Notes in Computer Science, 964:477– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1995:DTP

[370] J. C. M. Baeten and J. A. Bergstra. Discrete time process algebra with abstraction. Lecture Notes in Computer Science, 965:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaochen:1995:DCI

[371] Z. Chaochen, D. V. Hung, and L. Xiaoshan. A duration calculus with infinite intervals. Lecture Notes in Computer Science, 965:16–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1995:DBO

[372] K. Fisher and J. C. Mitchell. A delegation-based object calculus with subtyping. Lecture Notes in Computer Science, 965:42–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1995:MCR

[373] K. G. Larsen, P. Pettersson, and W. Yi. Model-checking for real-time systems. Lecture Notes in Computer Science, 965:62–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1995:PIT

[374] E. W. Mayr. On polynomial ideals, their complexity, and applications. Lecture Notes in Computer Science, 965:89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amadio:1995:CCC

[375] R. M. Amadio, L. Leth, and B. Thomsen. From a concurrent λ-calculus to the π-calculus. Lecture Notes in Computer Science, 965:106–115, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antimirov:1995:RRI

[376] V. Antimirov. Rewriting regular inequalities. Lecture Notes in Computer Science, 965:116–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1995:SAS

[377] G. Barthe. A simple abstract semantics for equational theories. Lecture Notes in Computer Science, 965:126– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergstra:1995:PME

[378] J. A. Bergstra and G. Stefanescu. Processes with multiple entries and exits. Lecture Notes in Computer Science, 965:136–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertol:1995:ERC

[379] M. Bertol. Efficient rewriting in cograph trace monoids. Lecture Notes in Computer Science, 965:146–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calude:1995:ECM

[380] C. Calude and M. Zimand. Effective category and measure in abstract complexity theory. Lecture Notes in Computer Science, 965:156–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaboud:1995:APC

[381] T. Chaboud. About planar Cayley graphs. Lecture Notes in Computer Science, 965:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chepoi:1995:CMP

[382] V. D. Chepoi and F. F. Dragan. On Condorcet and median points of simple rectilinear polygons. Lecture Notes in Computer Science, 965:181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djidjev:1995:FAM

[383] H. N. Djidjev, G. E. Pantziou, and C. D. Zaroliagis. Fast algorithms for maintaining shortest paths in outerplanar and planar digraphs. Lecture Notes in Computer Science, 965:191– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragan:1995:DPH

[384] F. F. Dragan and F. Nicolai. rdomination problems on homogeneously orderable graphs. Lecture Notes in Computer Science, 965:201– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand:1995:GPC

[385] B. Durand and J. Mazoyer. Growing patterns in 1D cellular automata. Lecture Notes in Computer Science, 965: 211–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esparza:1995:PNC

[386] J. Esparza. Petri nets, commutative context-free grammars, and basic parallel processes. Lecture Notes in Computer Science, 965:221–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fassbender:1995:IUA

[387] H. Fassbender. Implementation of a UU-Algorithm for primitive recursive tree functions. Lecture Notes in Computer Science, 965:233–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1995:DEM

[388] M. C. F. Ferreira and H. Zantema. Dummy elimination: Making termination easier. Lecture Notes in Computer Science, 965:243–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gronewold:1995:CPN

[389] A. Gronewold and H. Fleischhack. Computing Petri net languages by reductions. Lecture Notes in Computer Science, 965:253–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeHaas:1995:CG

[390] E. De Haas. Categorial graphs. Lecture Notes in Computer Science, 965: 263–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1995:ESA

[391] J. Hromkovic, R. Klasing, D. Pardubska, and W. Unger. Effective systolic algorithms for gossiping in cycles and two-dimensional grids. Lecture Notes in Computer Science, 965:273–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jancar:1995:RA

[392] P. Jancar, F. Mraz, M. Platek, and J. Vogel. Restarting automata. Lecture Notes in Computer Science, 965: 283–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kessler:1995:OCE

[393] C. W. Kessler and T. Rauber. Optimal contiguous expression DAG evaluations. Lecture Notes in Computer Science, 965:293–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klaudel:1995:CUP

[394] H. Klaudel and E. Pelz. Communication as unification in the Petri box calculus. Lecture Notes in Computer Science, 965:303–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kudlek:1995:DCC

[395] M. Kudlek and A. Mateescu. Distributed catenation and Chomsky hierarchy. Lecture Notes in Computer Science, 965:313–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kummer:1995:PFC

[396] M. Kummer and F. Stephan. The power of frequency computation. Lecture Notes in Computer Science, 965: 323–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Le:1995:RIC

[397] N.-M. Le. Randomized incremental construction of simple abstract Voronoi diagrams in 3-Space. Lecture Notes in Computer Science, 965:333–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macarie:1995:PPP

[398] I. I. Macarie and M. Ogihara. Properties of probabilistic pushdown automata. Lecture Notes in Computer Science, 965:343–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makanin:1995:FPE

[399] G. S. Makanin, H. Abdulrab, and M. N. Maksimenko. Formal parametric equations. Lecture Notes in Computer Science, 965:353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niedermeier:1995:PTR

[400] R. Niedermeier and P. Rossmanith. PRAM’s towards realistic parallelism: BRAM’s. Lecture Notes in Computer Science, 965:363–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersen:1995:SRC

[401] H. Petersen. Some results concerning two-dimensional Turing machines and finite automata. Lecture Notes in Computer Science, 965:374–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pighizzini:1995:HHC

[402] G. Pighizzini. How hard is to compute the edit distance. Lecture Notes in Computer Science, 965:383–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reinhardt:1995:SST

[403] K. Reinhardt. On the synchronization of semi-traces. Lecture Notes in Computer Science, 965:393–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Remila:1995:TBS

[404] E. Remila. Tiling with bars and satisfaction of Boolean formulas. Lecture Notes in Computer Science, 965:404– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schewe:1995:FSD

[405] K.-D. Schewe. Functional sorts in data type specifications; A geometric approach to semantics. Lecture Notes in Computer Science, 965:424–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulos:1995:MPT

[406] G. Papadopoulos. Mainstream parallelism: Taking sides on the SMP/ MPP/ cluster debate. Lecture Notes in Computer Science, 966:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smolka:1995:OPMa

[407] G. Smolka. The Oz programming model. Lecture Notes in Computer Science, 966:5–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engquist:1995:PCA

[408] B. Engquist. Parallelism in computational algorithms and the physical world. Lecture Notes in Computer Science, 966:9–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caspi:1995:EDR

[409] P. Caspi and A. Girault. Execution of distributed reactive systems. Lecture Notes in Computer Science, 966:15–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermenegildo:1995:RDP

[410] M. V. Hermenegildo and M. Carro. Relating data-parallelism and (and-) parallelism in logic programs. Lecture Notes in Computer Science, 966:27–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pontelli:1995:DBP

[411] E. Pontelli and G. Gupta. On the duality between or-parallelism and andparallelism in logic programming. Lecture Notes in Computer Science, 966: 43–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darlington:1995:FSP

[412] J. Darlington, Y.-K. Guo, Hing Wing To, and J. Yang. Functional skeletons for parallel coordination. Lecture Notes in Computer Science, 966:55–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Unrau:1995:SDD

[413] R. C. Unrau, M. Stumm, and O. Krieger. On the scalability of demand-driven parallel systems. Lecture Notes in Computer Science, 966: 69–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKee:1995:BMB

[414] S. A. McKee, W. A. Wulf, and T. C. Landon. Bounds on memory bandwidth in streamed computations. Lecture Notes in Computer Science, 966: 83–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiou:1995:SND

[415] D. Chiou, B. S. Ang, R. Greiner, and Arvind. StarT-NG: Delivering seamless parallel computing. Lecture Notes in Computer Science, 966:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maquelin:1995:CBM

[416] O. C. Maquelin, H. H. J. Hum, and G. R. Gao. Costs and benefits of multithreading with off-the-shelf RISC processors. Lecture Notes in Computer Science, 966:117–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Genaud:1995:TTP

[417] S. Genaud, E. Violard, and G.-R. Perrin. Transformation techniques in pei. Lecture Notes in Computer Science, 966:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouge:1995:CPS

[418] L. Bouge and D. Cachera. On the completeness of a proof system for a simple data-parallel programming language. Lecture Notes in Computer Science, 966:143–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clemencon:1995:IRD

[419] C. Clemencon, J. Fritscher, M. J. Meehan, and R. Ruehl. An implementation of race detection and deterministic replay with MPI. Lecture Notes in Computer Science, 966:155–166, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fagot:1995:FEV

[420] A. Fagot and J. Chassin de Kergommeaux. Formal and experimental validation of a low overhead execution replay mechanism. Lecture Notes in Computer Science, 966:167–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roemke:1995:EEG

[421] T. Roemke, M. Roettger, U.-P. Schroeder, and J. Simon. On efficient embeddings of grids into grids in PARIX. Lecture Notes in Computer Science, 966:181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Achilles:1995:OEM

[422] A.-C. Achilles. Optimal emulation of meshes on meshes of trees. Lecture Notes in Computer Science, 966:193– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1995:OEH

[423] S. K. Das and A. Mao. Optimal embeddings in the Hamming cube networks. Lecture Notes in Computer Science, 966:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1995:HAR

[424] Z. Liu. Hierarchical adaptive routing under hybrid traffic load. Lecture Notes in Computer Science, 966:217– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhatt:1995:TBP

[425] S. N. Bhatt, G. Bilardi, K. T. Herley, and G. Pucci. Tight bounds on parallel list marking. Lecture Notes in Computer Science, 966:231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loewe:1995:OPP

[426] W. Loewe. Optimization of PRAM-Programs with input-dependent memory access. Lecture Notes in Computer Science, 966:243–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:OCA

[427] L. Chen. Optimal circular arc representations. Lecture Notes in Computer Science, 966:255–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nowatzyk:1995:EPC

[428] A. Nowatzyk, G. Aybay, M. Browne, and E. Kelly. Exploiting parallelism in cache coherency protocol engines. Lecture Notes in Computer Science, 966: 269–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pong:1995:VDD

[429] F. Pong, A. Nowatzyk, G. Aybay, and M. Dubois. Verifying distributed directory-based cache coherence protocols: S3.mp, a case study. Lecture Notes in Computer Science, 966:287– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1995:ESD

[430] S.-J. Hwang and M.-S. Park. Efficient software data prefetching for a loop with large arrays. Lecture Notes in Computer Science, 966:301–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griebl:1995:GSC

[431] M. Griebl and J.-F. Collard. Generation of synchronous code for automatic parallelization of while loops. Lecture Notes in Computer Science, 966:315– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kulkarni:1995:IFC

[432] D. Kulkarni, M. Stumm, and R. C. Unrau. Implementing flexible computation rules with subexpression-level loop transformations. Lecture Notes in Computer Science, 966:327–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeFur:1995:APA

[433] M. Le Fur, J.-L. Pazat, and F. Andre. An array partitioning analysis for parallel loop distribution. Lecture Notes in Computer Science, 966:351–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erzmann:1995:MEP

[434] A. Erzmann, M. Hadeler, and C. Mueller-Schloer. A model for efficient programming of dynamic applications on distributed memory multiprocessors. Lecture Notes in Computer Science, 966:367–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouvry:1995:ESM

[435] P. Bouvry, J. Chassin de Kergommeaux, and D. Trystram. Efficient solutions for mapping parallel programs. Lecture Notes in Computer Science, 966:379–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauber:1995:ODD

[436] T. Rauber and G. Ruenger. Optimal data distributions for LU decomposition. Lecture Notes in Computer Science, 966:391–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minas:1995:DQG

[437] M. Minas. Detecting quantified global predicates in parallel programs. Lecture Notes in Computer Science, 966: 403–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:UKB

[438] C.-T. Yang, S.-S. Tseng, C.-D. Chuang, and W.-C. Shih. Using knowledgebased techniques for parallelization on parallelizing compilers. Lecture Notes in Computer Science, 966:417– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Germain:1995:AVC

[439] C. Germain and F. Delaplace. Automatic vectorization of communications for data-parallel programs. Lecture Notes in Computer Science, 966: 429–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pouzet:1995:PCR

[440] M. Pouzet. The program compaction revisited: the functional framework. Lecture Notes in Computer Science, 966:441–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sloman:1995:FTA

[441] B. Sloman and T. Lake. Featherweight threads and ANDF compilation of concurrency. Lecture Notes in Computer Science, 966:457–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Romein:1995:PBS

[442] J. W. Romein and H. E. Bal. Parallel N-body simulation on a large-scale homogeneous distributed system. Lecture Notes in Computer Science, 966: 473–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahm:1995:APS

[443] E. Rahm and T. Stoehr. Analysis of parallel scan processing in shared disk database systems. Lecture Notes in Computer Science, 966:485–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolan:1995:PTS

[444] A. R. Nolan and B. Everding. Polynomial time scheduling of low level computer vision algorithms on networks of heterogeneous machines. Lecture Notes in Computer Science, 966:501– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girau:1995:MNN

[445] B. Girau. Mapping neural network back-propagation onto parallel computers with computation/ communication overlapping. Lecture Notes in Computer Science, 966:513– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larson:1995:SMP

[446] J. S. Larson, B. C. Massey, and E. Tick. Super Monaco: Its portable and efficient parallel runtime system. Lecture Notes in Computer Science, 966:527– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rokusawa:1995:QDD

[447] K. Rokusawa, A. Nakase, and T. Chikayama. Quiescence detection in a distributed KLIC implementation. Lecture Notes in Computer Science, 966:539–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindgren:1995:COR

[448] T. Lindgren, J. Bevemyr, and H. Millroth. Compiler optimizations in reform Prolog: Experiments on the KSR-1 multiprocessor. Lecture Notes in Computer Science, 966:553–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaseemuddin:1995:BRA

[449] M. Jaseemuddin and Z. G. Vranesic. Bidirectional ring: An alternative to the hierarchy of unidirectional rings. Lecture Notes in Computer Science, 966:567–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singhvi:1995:FSM

[450] N. K. Singhvi and K. Ghose. A formal study of the mcube interconnection network. Lecture Notes in Computer Science, 966:579–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shionoya:1995:MIN

[451] S. Shionoya, T. Aoki, and T. Higuchi. Multiwave interconnection networks for MCM-Based parallel processing. Lecture Notes in Computer Science, 966:593–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sahni:1995:SMS

[452] S. Sahni. Scheduling master-slave multiprocessor systems. Lecture Notes in Computer Science, 966:611–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hori:1995:TSS

[453] A. Hori, Y. Ishikawa, J. Nolte, and H. Konaka. Time space sharing scheduling: A simulation analysis. Lecture Notes in Computer Science, 966: 623–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rost:1995:ASM

[454] J. Rost, F.-J. Markus, and L. Yan-Hua. “agency scheduling” — A model for dynamic task scheduling. Lecture Notes in Computer Science, 966:635–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johannesson:1995:FLS

[455] M. Johannesson. FFTs on a linear SIMD array. Lecture Notes in Computer Science, 966:649–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sheu:1995:TFF

[456] J.-P. Sheu and Y.-S. Chen. Tolerating faults in faulty hypercubes using maximal fault-free subcube-ring. Lecture Notes in Computer Science, 966: 661–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chalasani:1995:CMN

[457] S. Chalasani and R. V. Boppana. Communication in multicomputers with nonconvex faults. Lecture Notes in Computer Science, 966:673–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doroshenko:1995:PPA

[458] A. E. Doroshenko and A. B. Godlevsky. Parallelising programs with algebraic programming tools. Lecture Notes in Computer Science, 966:687–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molnar:1995:PPU

[459] K. Molnar. Parallel Prolog with uncertainty handling. Lecture Notes in Computer Science, 966:691–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Friedl:1995:SPC

[460] G. Friedl, M. Platzner, and B. Rinner. A special-purpose coprocessor for qualitative simulation. Lecture Notes in Computer Science, 966:695–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnes:1995:PST

[461] C. Barnes and C. Wadsworth. Portable software tools for parallel architectures. Lecture Notes in Computer Science, 966:699–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siegelin:1995:BPW

[462] C. Siegelin, U. Finger, and C. O’Donnell. Boosting the performance of workstations through WARPmemory. Lecture Notes in Computer Science, 966:703– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laurentowski:1995:MSS

[463] A. Laurentowski, J. Szymaszek, A. Uszok, and K. Zielinski. A monitoring system for software-heterogeneous distributed environments. Lecture Notes in Computer Science, 966:707– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hains:1995:MDP

[464] G. Hains and J. Mullins. A metacircular data-parallel functional language. Lecture Notes in Computer Science, 966:711–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silc:1995:ERT

[465] J. Silc and B. Robic. Efficient run-time program allocation on a parallel coprocessor. Lecture Notes in Computer Science, 966:715–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evstigneev:1995:PMSb

[466] V. A. Evstigneev and V. N. Kasyanov. A program manipulation system for fine-grained architectures. Lecture Notes in Computer Science, 966:719– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moravie:1995:RTI

[467] P. Moravie, H. Essafi, C. Lambert-Nebout, and J.-L. Basille. Realtime image compression using dataparallelism. Lecture Notes in Computer Science, 966:723–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smai:1995:CCW

[468] A.-H. Smai. Congestion control in wormhole networks: First results. Lecture Notes in Computer Science, 966: 727–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parnas:1995:LFM

[469] D. L. Parnas. Language-free mathematical methods for software design. Lecture Notes in Computer Science, 967:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciaccia:1995:FAS

[470] P. Ciaccia, P. Ciancarini, and W. Penzo. A formal approach to software design: The clepsydra methodology. Lecture Notes in Computer Science, 967:5–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edmond:1995:RDS

[471] D. Edmond. Refining database systems. Lecture Notes in Computer Science, 967:25–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luck:1995:SZS

[472] M. Luck and M. D’Inverno. Structuring a Z specification to provide a formal framework for autonomous agent systems. Lecture Notes in Computer Science, 967:47–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLimaMachado:1995:UFS

[473] P. D. De Lima Machado and S. L. Meira. On the use of formal specifications in the design and simulation of artificial neural networks. Lecture Notes in Computer Science, 967:63–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DInverno:1995:SSZ

[474] M. D’Inverno and M. Priestley. Structuring specification in Z to build a unifying framework for hypertext systems. Lecture Notes in Computer Science, 967:83–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rushby:1995:MFM

[475] J. Rushby. Mechanizing formal methods: Opportunities and challenges. Lecture Notes in Computer Science, 967:105–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hughes:1995:APV

[476] A. P. Hughes and A. A. Donnelly. An algebraic proof in VDM. Lecture Notes in Computer Science, 967:114– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stepney:1995:TA

[477] S. Stepney. Testing as abstraction. Lecture Notes in Computer Science, 967: 137–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoercher:1995:IST

[478] H.-M. Hoercher. Improving software tests using Z specifications. Lecture Notes in Computer Science, 967:152– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikk:1995:CZS

[479] E. Mikk. Compilation of Z specifications into C for automatic test result evaluation. Lecture Notes in Computer Science, 967:167–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valentine:1995:ERS

[480] S. H. Valentine. Equal rights for schemas in Z. Lecture Notes in Computer Science, 967:183–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacDonald:1995:SZS

[481] A. MacDonald and D. Carrington. Structuring Z specifications: Some choices. Lecture Notes in Computer Science, 967:203–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

German:1995:EZI

[482] D. M. German and D. D. Cowan. Experiments with the Z interchange format and SGML. Lecture Notes in Computer Science, 967:224–236 (or 224– 233??), 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1995:FFM

[483] J. A. Hall, D. L. Parnas, N. Plat, and J. Rushby. The future of formal methods in industry. Lecture Notes in Computer Science, 967:237–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liskov:1995:STU

[484] B. Liskov and J. M. Wing. Specifications and their use in defining subtypes. Lecture Notes in Computer Science, 967:245–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strulo:1995:HFC

[485] B. Strulo. How firing conditions help inheritance. Lecture Notes in Computer Science, 967:264–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1995:EWO

[486] G. Smith. Extending W for object-Z. Lecture Notes in Computer Science, 967:276–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bancroft:1995:FSL

[487] P. Bancroft and I. Hayes. A formal semantics for a language with type extension. Lecture Notes in Computer Science, 967:299–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacky:1995:ZCG

[488] J. Jacky and J. Unger. From Z to code: A graphical user interface for a radiation therapy machine. Lecture Notes in Computer Science, 967:315–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernard:1995:FPC

[489] P. Bernard and G. Laffitte. The French population census for 1990. Lecture Notes in Computer Science, 967:334– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kraan:1995:IZI

[490] I. Kraan and P. Baumann. Implementing Z in isabelle. Lecture Notes in Computer Science, 967:355–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goodman:1995:ZHT

[491] H. S. Goodman. The Z-into-Haskell tool-kit: An illustrative case study. Lecture Notes in Computer Science, 967:374–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

West:1995:TSG

[492] M. M. West. Types and sets in goedel and Z. Lecture Notes in Computer Science, 967:389–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reilly:1995:ESM

[493] C. Reilly. Exploring specifications with Mathematica. Lecture Notes in Computer Science, 967:408–422, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryant:1995:UZR

[494] T. Bryant, A. Evans, L. Semmens, and R. Milovanovic. Using Z to rigorously review a specification of a network management system. Lecture Notes in Computer Science, 967:423–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

France:1995:TDV

[495] R. B. France and M. M. Larrondo-Petrie. A two-dimensional view of integrated formal and informal specification techniques. Lecture Notes in Computer Science, 967:434–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowman:1995:VO

[496] H. Bowman, J. Derrick, and M. Steen. Viewpoints and objects. Lecture Notes in Computer Science, 967:449– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parnas:1995:TPE

[497] D. L. Parnas. Teaching programming as engineering. Lecture Notes in Computer Science, 967:471–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciaccia:1995:CFM

[498] P. Ciaccia and P. Ciancarini. A course on formal methods in software engineering: Matching requirements with design. Lecture Notes in Computer Science, 967:482–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wing:1995:HWS

[499] J. M. Wing. Hints for writing specifications. Lecture Notes in Computer Science, 967:497–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dean:1995:MMZ

[500] N. Dean. Mental models of Z: I — sets and logic. Lecture Notes in Computer Science, 967:498–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gries:1995:ELG

[501] D. Gries. Equational logic: A great pedagogical tool for teaching a skill in logic. Lecture Notes in Computer Science, 967:508–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikusiak:1995:ZBT

[502] L. Mikusiak, V. Vojtek, J. Hasaralejko, and J. Hanzelova. Z browser — A tool for visualization of Z specifications. Lecture Notes in Computer Science, 967:510–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1995:ACS

[503] L. Bachmair and H. Ganzinger. Associative-commutative superposition. Lecture Notes in Computer Science, 968:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1995:CR

[504] D. A. Basin and T. Walsh. A calculus for rippling. Lecture Notes in Computer Science, 968:15–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balbiani:1995:ESG

[505] P. Balbiani. Equation solving in geometrical theories. Lecture Notes in Computer Science, 968:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bockmayr:1995:LND

[506] A. Bockmayr and A. Werner. LSE narrowing for decreasing conditional term rewrite systems. Lecture Notes in Computer Science, 968:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buendgen:1995:PCR

[507] R. Buendgen. Preserving confluence for rewrite systems with built-in operations. Lecture Notes in Computer Science, 968:71–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1995:HT

[508] N. Dershowitz. Hierarchical termination. Lecture Notes in Computer Science, 968:89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1995:WFT

[509] M. C. F. Ferreira and H. Zantema. Well-foundedness of term orderings. Lecture Notes in Computer Science, 968:106–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gelis:1995:NCA

[510] J.-M. Gelis. A new characterisation of AC-Termination and application. Lecture Notes in Computer Science, 968:124–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glauert:1995:RNO

[511] J. Glauert and Z. Khasidashvili. Relative normalization in orthogonal expression reduction systems. Lecture Notes in Computer Science, 968:144– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gramlich:1995:TCC

[512] B. Gramlich. On termination and confluence of conditional rewrite systems. Lecture Notes in Computer Science, 968:166–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hintermeier:1995:HTC

[513] C. Hintermeier. How to transform canonical decreasing HCTRs into equivalent canonical TRSs. Lecture Notes in Computer Science, 968:186– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoot:1995:TRD

[514] C. Hoot. Termination for restricted derivations and conditional rewrite systems. Lecture Notes in Computer Science, 968:206–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inverardi:1995:RPR

[515] P. Inverardi. Rewriting for preorder relations. Lecture Notes in Computer Science, 968:223–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jouannaud:1995:SSL

[516] J.-P. Jouannaud and W. Sadfi. Strong sequentiality of left-linear overlapping rewrite systems. Lecture Notes in Computer Science, 968:235–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennaway:1995:CBC

[517] R. Kennaway. A conflict between call-by-need computation and parallelism. Lecture Notes in Computer Science, 968:247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kucherov:1995:CTG

[518] G. Kucherov and M. Rusinowitch. The complexity of testing ground reducibility for linear word rewriting systems with variables. Lecture Notes in Computer Science, 968:262–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mori:1995:CCC

[519] A. Mori and Y. Matsumoto. Coherence for Cartesian closed categories: A sequential approach. Lecture Notes in Computer Science, 968:276–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohlebusch:1995:MPC

[520] E. Ohlebusch. Modular properties of constructor-sharing conditional term rewriting systems. Lecture Notes in Computer Science, 968:296–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toyama:1995:CRP

[521] Y. Toyama and M. Oyamaguchi. Church–Rosser property and unique normal form property of non-duplicating term rewriting systems. Lecture Notes in Computer Science, 968:316–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verheul:1995:TTR

[522] J. C. Verheul and P. G. Kluit. The transformation of term rewriting systems based on well-formedness preserving mappings. Lecture Notes in Computer Science, 968:332–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wirth:1995:ANI

[523] C.-P. Wirth and K. Becker. Abstract notions and inference systems for proofs by mathematical induction. Lecture Notes in Computer Science, 968: 353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adler:1995:SPC

[524] M. Adler, J. W. Byers, and R. M. Karp. Scheduling parallel communication: The h-relation problem. Lecture Notes in Computer Science, 969:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnborg:1995:DSB

[525] S. Arnborg. Decomposable structures, Boolean function representations, and optimization. Lecture Notes in Computer Science, 969:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flammini:1995:CIR

[526] M. Flammini, J. Van Leeuwen, and A. Marchetti-Spaccamela. The complexity of interval routing on random graphs. Lecture Notes in Computer Science, 969:37–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geffert:1995:BAL

[527] V. Geffert. Bridging across the log(n) space frontier. Lecture Notes in Computer Science, 969:50–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gottlob:1995:SOL

[528] G. Gottlob, N. Leone, and H. Veith. Second order logic and the weak exponential hierarchies. Lecture Notes in Computer Science, 969:66–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmanis:1995:CPC

[529] J. Hartmanis. On the computing paradigm and computational complexity. Lecture Notes in Computer Science, 969:82–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehmann:1995:RSN

[530] D. Lehmann. Ranked structures in nonmonotonic reasoning and belief revision. Lecture Notes in Computer Science, 969:93–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perrin:1995:SDF

[531] D. Perrin. Symbolic dynamics and finite automata. Lecture Notes in Computer Science, 969:94–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Razborov:1995:LBP

[532] A. A. Razborov. Lower bounds for propositional proofs and independence results in bounded arithmetic. Lecture Notes in Computer Science, 969:105– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitanyi:1995:PNC

[533] P. Vitanyi. Physics and the new computation. Lecture Notes in Computer Science, 969:106–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allender:1995:MPR

[534] E. Allender and M. Strauss. Measure on P: Robustness of the notion. Lecture Notes in Computer Science, 969:129–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burtschick:1995:CCC

[535] H.-J. Burtschick. Comparing counting classes for logspace, one-way logspace, and first-order. Lecture Notes in Computer Science, 969:139–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damm:1995:ATA

[536] C. Damm and M. Holzer. Automata that take advice. Lecture Notes in Computer Science, 969:149–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Homer:1995:NLB

[537] S. Homer and S. Mocas. Nonuniform lower bounds for exponential time classes. Lecture Notes in Computer Science, 969:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufmann:1995:QNU

[538] S. Kaufmann and M. Kummer. On a quantitative notion of uniformity. Lecture Notes in Computer Science, 969: 169–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merkle:1995:SRO

[539] W. Merkle and Y. Wang. Separations by random oracles and “almost” classes for generalized reducibilities. Lecture Notes in Computer Science, 969:179–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beauquier:1995:CFM

[540] D. Beauquier, D. Burago, and A. Slissenko. On the complexity of finite memory policies for Markov decision processes. Lecture Notes in Computer Science, 969:191–200, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmeister:1995:DSA

[541] T. Hofmeister and H. Lefmann. Derandomization for sparse approximations and independent sets. Lecture Notes in Computer Science, 969:201–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katajainen:1995:AEP

[542] J. Katajainen, T. Pasanen, and G. Titan. Asymptotically efficient in-place merging. Lecture Notes in Computer Science, 969:211–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heusch:1995:CFP

[543] P. Heusch. The complexity of the falsifiability problem for pure implicational formulas. Lecture Notes in Computer Science, 969:221–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kann:1995:SLB

[544] V. Kann. Strong lower bounds on the approximability of some NPO PBcomplete maximization problems. Lecture Notes in Computer Science, 969: 227–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lefmann:1995:STP

[545] H. Lefmann and P. Savicky. Some typical properties of large AND/OR Boolean formulas. Lecture Notes in Computer Science, 969:237–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huehne:1995:HES

[546] M. Huehne. The hedge: An efficient storage device for Turing machines with one head. Lecture Notes in Computer Science, 969:247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maruyama:1995:GIW

[547] O. Maruyama and S. Miyano. Graph inference from a walk for trees of bounded degree 3 is NP-complete. Lecture Notes in Computer Science, 969: 257–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stojmenovic:1995:HN

[548] I. Stojmenovic. Honeycomb networks. Lecture Notes in Computer Science, 969:267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1995:WIR

[549] S. Fischer, L. Hemaspaandra, and L. Torenvliet. Witness-isomorphic reductions and the local search problem. Lecture Notes in Computer Science, 969:277–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertram-Kretzberg:1995:MPM

[550] C. Bertram-Kretzberg and T. Hofmeister. Multiple product modulo arbitrary numbers. Lecture Notes in Computer Science, 969:288–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinel:1995:LBM

[551] C. Meinel and S. Waack. Lower bounds for the majority communication complexity of various graph accessibility problems. Lecture Notes in Computer Science, 969:299–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertoni:1995:SOL

[552] A. Bertoni, C. Mereghetti, and G. Pighizzini. Strong optimal lower bounds for Turing machines that accept nonregular languages. Lecture Notes in Computer Science, 969:309– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zak:1995:SLB

[553] S. Zak. A superpolynomial lower bound for (1, +k(n))-branching programs. Lecture Notes in Computer Science, 969:319–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breveglieri:1995:DPA

[554] L. Breveglieri, A. Cherubini, and S. Crespi-Reghizzi. Deterministic parsing for augmented context-free grammars. Lecture Notes in Computer Science, 969:326–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mignosi:1995:PTW

[555] F. Mignosi, A. Restivo, and S. Salemi. A periodicity theorem on words and applications. Lecture Notes in Computer Science, 969:337–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hotz:1995:NAA

[556] G. Hotz and G. Pitsch. A new approach to analyse coupled-context-free languages. Lecture Notes in Computer Science, 969:349–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermann:1995:CCS

[557] M. Hermann and P. G. Kolaitis. Computational complexity of simultaneous elementary matching problems. Lecture Notes in Computer Science, 969: 359–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rao:1995:GRT

[558] M. R. K. Krishna Rao. Graph reducibility of term rewriting systems. Lecture Notes in Computer Science, 969:371–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Urzyczyn:1995:PRT

[559] P. Urzyczyn. Positive recursive type assignment. Lecture Notes in Computer Science, 969:382–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kranakis:1995:SRA

[560] E. Kranakis, D. Krizanc, and F. L. Luccio. String recognition on anonymous rings. Lecture Notes in Computer Science, 969:392–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roka:1995:FSS

[561] Z. Roka. The firing squad synchronization problem on Cayley graphs. Lecture Notes in Computer Science, 969: 402–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1995:SCG

[562] J. F. Sibeyn and M. Kaufmann. Solving cheap graph problems on meshes. Lecture Notes in Computer Science, 969: 412–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkart:1995:EBD

[563] O. Burkart, D. Caucal, and B. Steffen. An elementary bisimulation decision procedure for arbitrary context-free processes. Lecture Notes in Computer Science, 969:423–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauget:1995:CPO

[564] S. Bauget and P. Gastin. On congruences and partial orders. Lecture Notes in Computer Science, 969:434– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1995:PPO

[565] F. Corradini, R. Gorrieri, and M. Roccetti. Performance preorder: Ordering processes with respect to speed. Lecture Notes in Computer Science, 969: 444–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1995:TST

[566] W. Ferreira and M. Hennessy. Towards a semantic theory of CML (extended abstract). Lecture Notes in Computer Science, 969:454–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huguet:1995:MCD

[567] S. Huguet and A. Petit. Modular constructions of distributing automata. Lecture Notes in Computer Science, 969:467–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sangiorgi:1995:PMB

[568] D. Sangiorgi. On the proof method for bisimulation. Lecture Notes in Computer Science, 969:479–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1995:TCP

[569] C. Martin. Towards a calculus of predicate transformers. Lecture Notes in Computer Science, 969:489–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1995:AAC

[570] M. Abadi and S. Merz. An abstract account of composition. Lecture Notes in Computer Science, 969:499–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderGoot:1995:SSP

[571] R. Van der Goot and A. De Bruin. Syntax and semantics of Procol. Lecture Notes in Computer Science, 969:509– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godskesen:1995:SDF

[572] J. C. Godskesen and K. G. Larsen. Synthesizing distinguishing formulae for real time systems. Lecture Notes in Computer Science, 969:519–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laroussinie:1995:TAL

[573] F. Laroussinie, K. G. Larsen, and C. Weise. From timed automata to logic — and back. Lecture Notes in Computer Science, 969:529–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makowsky:1995:IMC

[574] J. A. Makowsky and E. V. Ravve. Incremental model checking for decomposable structures. Lecture Notes in Computer Science, 969:540–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janin:1995:AMC

[575] D. Janin and I. Walukiewicz. Automata for the modal μ-calculus and related results. Lecture Notes in Computer Science, 969:552–562, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niebert:1995:CLV

[576] P. Niebert. A v-calculus with local views for systems of sequential agents. Lecture Notes in Computer Science, 969:563–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feinsilver:1995:OCA

[577] P. Feinsilver and R. Schott. An operator calculus approach to the evolution of dynamic data structures. Lecture Notes in Computer Science, 969: 574–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGool:1995:GGS

[578] L. Van Gool, T. Moons, and M. Proesmans. Groups for grouping: A strategy for the exploitation of geometrical constraints. Lecture Notes in Computer Science, 970:1–8, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suesse:1995:ISP

[579] H. Suesse, K. Voss, and I. Rothe. Invariant standard positions of ordered sets of points. Lecture Notes in Computer Science, 970:9–16, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruckstein:1995:SSD

[580] A. M. Bruckstein and D. Shaked. Skew symmetry detection via invariant signatures. Lecture Notes in Computer Science, 970:17–24, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pajdla:1995:EMS

[581] T. Pajdla and L. Van Gool. Efficient matching of space curves. Lecture Notes in Computer Science, 970:25–32, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubuc:1995:CSC

[582] B. Dubuc and S. W. Zucker. A complexity space for curve grouping. Lecture Notes in Computer Science, 970: 33–40, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henkel:1995:SSS

[583] R. D. Henkel. Segmentation in scale space. Lecture Notes in Computer Science, 970:41–48, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiavetta:1995:URG

[584] F. Chiavetta and V. Di Gesu. An unsupervised region growing method for 3-D image segmentation. Lecture Notes in Computer Science, 970:49–56, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimelfarb:1995:SIE

[585] G. Gimel’farb and N. Kovalevskaya. Segmentation of images for environmental studies using a simple Markov/ Gibbs random field model. Lecture Notes in Computer Science, 970:57–64, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daniilidis:1995:OFC

[586] K. Daniilidis and V. Krueger. Optical flow computation in the log-polar plane. Lecture Notes in Computer Science, 970:65–72, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ricquebourg:1995:SRF

[587] Y. Ricquebourg and P. Bouthemy. A statistical regularization framework for estimating normal displacements along contours with subpixel accuracy. Lecture Notes in Computer Science, 970: 73–81, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daniilidis:1995:CDM

[588] K. Daniilidis. Computation of 3-D-motion parameters using the log-polar transform. Lecture Notes in Computer Science, 970:82–89, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stricker:1995:EGF

[589] M. Stricker and A. Leonardis. ExSel++: A general framework to extract parametric models. Lecture Notes in Computer Science, 970:90–97, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shen:1995:GDS

[590] X. Shen and D. Hogg. Generic 3-D shape model: Acquisitions and applications. Lecture Notes in Computer Science, 970:98–105, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1995:GCO

[591] O. V. Larsen, P. Radeva, and E. Marti. Guidelines for choosing optimal parameters of elasticity for snakes. Lecture Notes in Computer Science, 970:106– 113, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bray:1995:TST

[592] A. Bray. Tracking surfaces via texturemapping: A boot-strapping approach. Lecture Notes in Computer Science, 970:114–121, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnoerr:1995:MBI

[593] C. Schnoerr and W. Peckar. Motionbased identification of deformable templates. Lecture Notes in Computer Science, 970:122–129, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heap:1995:EPD

[594] T. Heap and D. Hogg. Extending the point distribution model using polar coordinates. Lecture Notes in Computer Science, 970:130–137, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lew:1995:IVF

[595] M. S. Lew, A. C. She, and T. S. Huang. Intermediate views for face recognition. Lecture Notes in Computer Science, 970:138–145, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werner:1995:RRW

[596] T. Werner, R. D. Hersch, and V. Hlavac. Rendering real-world objects without 3-D model. Lecture Notes in Computer Science, 970:146–153, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cole:1995:MDR

[597] P. A. R. Cole and M. S. Khan. Modelling 3-D rigid solid objects using the view signature II representation scheme. Lecture Notes in Computer Science, 970:154–161, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matas:1995:SFS

[598] J. Matas and J. Kittler. Spatial and feature space clustering: Applications in image analysis. Lecture Notes in Computer Science, 970:162–173, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:ARA

[599] Z. Zhang. An automatic and robust algorithm for determining motion and structure from two perspective images. Lecture Notes in Computer Science, 970:174–181, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1995:DSR

[600] H. V. Nguyen and M. Hanajik. 3-D scene reconstruction from image sequences. Lecture Notes in Computer Science, 970:182–189, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartley:1995:T

[601] R. I. Hartley and P. Sturm. Triangulation. Lecture Notes in Computer Science, 970:190–197, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyer:1995:DSR

[602] E. Boyer and M. O. Berger. 3-D surface reconstruction using occluding contours. Lecture Notes in Computer Science, 970:198–205, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volden:1995:IRC

[603] E. Volden, G. Giraudon, and M. Berthod. Image redundancy and classification. Lecture Notes in Computer Science, 970:206–213, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Starovoitov:1995:TMD

[604] V. V. Starovoitov. Towards a measure of diversity between grey-scale images. Lecture Notes in Computer Science, 970:214–221, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chetverikov:1995:POT

[605] D. Chetverikov. Pattern orientation and texture symmetry. Lecture Notes in Computer Science, 970:222–229, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weickert:1995:MTE

[606] J. Weickert. Multiscale texture enhancement. Lecture Notes in Computer Science, 970:230–237, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duc:1995:STR

[607] B. Duc, P. Schroeter, and J. Biguen. Spatio-temporal robust motion estimation and segmentation. Lecture Notes in Computer Science, 970:238–245, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caplier:1995:NSA

[608] A. Caplier and F. Luthon. A new spatiotemporal approach for image analysis application to motion detection. Lecture Notes in Computer Science, 970:246–253, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benayoun:1995:DNR

[609] S. Benayoun and N. Ayache. Dense non-rigid motion estimation in sequences of medical images using differential constraints. Lecture Notes in Computer Science, 970:254–261, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Najman:1995:TGC

[610] L. Najman and R. Vaillant. Topological and geometrical corners by watershed. Lecture Notes in Computer Science, 970:262–269, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piegay:1995:CLD

[611] E. Piegay, N. Selmaoui, and C. Leschi. Crest lines detection by valleys spreading. Lecture Notes in Computer Science, 970:270–277, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pridmore:1995:DTL

[612] T. P. Pridmore and S. V. Ablameyko. The distance transform for line patterns: Generalisation and development. Lecture Notes in Computer Science, 970:278–285, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maravall:1995:MDB

[613] D. Maravall, J. A. Sanandres, and L. Baumela. Mobile detection based on histogram difference. Lecture Notes in Computer Science, 970:286–293, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hering:1995:PTS

[614] F. Hering, D. Wierzimok, and B. Jaehne. Particle tracking in space time sequences. Lecture Notes in Computer Science, 970:294–301, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchiz:1995:PTB

[615] J. M. Sanchiz, F. Pla, J. A. Marchant, and R. Brivot. Plant tracking-based motion analysis in a crop field. Lecture Notes in Computer Science, 970:302– 309, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zouhal:1995:ARB

[616] L. M. Zouhal and T. Denoeux. An adaptive k-NN rule based on Dempster-Shafer theory. Lecture Notes in Computer Science, 970:310–317, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elder:1995:SDR

[617] J. Elder and S. W. Zucker. Shadows, defocus and reliable estimation. Lecture Notes in Computer Science, 970:318– 325, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwoh:1995:EIV

[618] C.-K. Kwoh and D. F. Gillies. Estimating the initial values of unobservable variables in visual probabilistic networks. Lecture Notes in Computer Science, 970:326–333, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1995:RMA

[619] R. C. Wilson and E. R. Hancock. Relational matching with active graphs. Lecture Notes in Computer Science, 970:334–341, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melichar:1995:ASM

[620] B. Melichar. Approximate string matching by finite automata. Lecture Notes in Computer Science, 970:342– 349, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finch:1995:MDT

[621] A. M. Finch, R. C. Wilson, and E. R. Hancock. Matching Delaunay triangulations by probabilistic relaxation. Lecture Notes in Computer Science, 970:350–358, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinz:1995:AMI

[622] A. Pinz, M. Prantl, and H. Ganster. Affine matching of intermediate symbolic representations. Lecture Notes in Computer Science, 970:359–367, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ross:1995:ACS

[623] T. Ross, H. Handels, J. Kreusch, and H. Busche. Automatic classification of skin tumours with high resolution surface profiles. Lecture Notes in Computer Science, 970:368–375, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strauss:1995:DMH

[624] E. Strauss and P. Burger. 4 dimensional modelling of the human heart. Lecture Notes in Computer Science, 970:376–383, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:SAA

[625] G. Z. Yang, P. Burger, D. N. Firmin, and S. R. Underwood. Structure adaptive anisotropic filtering for magnetic resonance image enhancement. Lecture Notes in Computer Science, 970:384– 391, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scholz:1995:SDC

[626] T. Scholz, B. Jaehne, H. Suhr, and G. Wehnert. In situ determination of cell concentration in bioreactors with a new depth from focus technique. Lecture Notes in Computer Science, 970:392–399, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nordlund:1995:CLP

[627] P. Nordlund and T. Uhlin. Closing the loop: Pursuing a moving object by a moving observer. Lecture Notes in Computer Science, 970:400–407, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kozera:1995:ALS

[628] R. Kozera. An algorithm for a linear shape-from-shading problem. Lecture Notes in Computer Science, 970:408– 415, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sara:1995:IKT

[629] R. Sara. Isophotes: the key to tractable local shading analysis. Lecture Notes in Computer Science, 970:416–423, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foltyniewicz:1995:EHO

[630] R. Foltyniewicz. Efficient high order neural network for rotation, translation and distance invariant recognition of gray scale images. Lecture Notes in Computer Science, 970:424–431, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sung:1995:LHF

[631] K.-K. Sung and T. Poggio. Learning human face detection in cluttered scenes. Lecture Notes in Computer Science, 970:432–439, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez:1995:HPC

[632] R. Lopez and T. S. Huang. Head pose computation for very low bitrate video coding. Lecture Notes in Computer Science, 970:440–447, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quan:1995:SCA

[633] L. Quan and R. Mohr. Self-calibration of an affine camera. Lecture Notes in Computer Science, 970:448–455, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Polanski:1995:SCP

[634] A. Polanski, K. Wojciechowski, and A. Borek. Stereo calibration by planar grid lines. Lecture Notes in Computer Science, 970:456–463, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neubauer:1995:MPE

[635] R. Neubauer and K. Voss. Monocular pose estimation of circular primitives. Lecture Notes in Computer Science, 970:464–471, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lohmann:1995:NME

[636] G. Lohmann. A new method of extracting closed contours using maximal discs. Lecture Notes in Computer Science, 970:472–479, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quereda:1995:SAE

[637] J. M. I. Quereda and M. B. Gomez. A simple algorithm to evaluate the local symmetry at each point of a closed contour. Lecture Notes in Computer Science, 970:480–487, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sloboda:1995:BA

[638] F. Sloboda and B. Zat’ko. On boundary approximation. Lecture Notes in Computer Science, 970:488–495, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albregtsen:1995:TCM

[639] F. Albregtsen, H. Schulerud, and L. Yang. Texture classification of mouse liver cell nuclei using invariant moments of consistent regions. Lecture Notes in Computer Science, 970:496– 502, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aas:1995:TRG

[640] K. Aas, L. Eikvil, and T. Andersen. Text recognition from grey level images using hidden Markov models. Lecture Notes in Computer Science, 970:503– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grigorian:1995:NMI

[641] A. Grigorian. New models of image restoration. Lecture Notes in Computer Science, 970:509–514, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radeva:1995:IMS

[642] P. Radeva and E. Marti. An improved model of snakes for model-based segmentation. Lecture Notes in Computer Science, 970:515–520, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kazantsev:1995:WBT

[643] I. Kazantsev. The weighted backprojection techniques of image reconstruction. Lecture Notes in Computer Science, 970:521–525, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shkvarko:1995:FBE

[644] Y. V. Shkvarko and A. S. Netjukhailo. Fusion of Bayesian estimation and MTF inversion techniques for improved array imaging in scattering media. Lecture Notes in Computer Science, 970: 526–531, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perner:1995:CBR

[645] P. Perner. Case based reasoning for image interpretation. Lecture Notes in Computer Science, 970:532–537, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Livens:1995:CCI

[646] S. Livens, P. Scheunders, G. Van de Wouwer, and D. Van Dyck. Classification of corrosion images by wavelet signatures and LVQ networks. Lecture Notes in Computer Science, 970:538– 543, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balev:1995:KAI

[647] B. Balev, A. Bloemer, and O. Dehning. Knowledge acquisition for image analysis using hypermedia interface. Lecture Notes in Computer Science, 970:544–549, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ihle:1995:IPF

[648] T. Ihle, H. Schirmer, and S. Fuchs. Interpretation of printed forms for blind people. Lecture Notes in Computer Science, 970:550–555, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmid:1995:ASB

[649] G. Schmid, L. Altamirano Robles, and W. Eckstein. Automatic segmentation of boundaries in line segments and circular arcs. Lecture Notes in Computer Science, 970:556–561, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voss:1995:MBI

[650] K. Voss, H. Suesse, and R. Neubauer. Moment-based invariant fitting of elliptical segments. Lecture Notes in Computer Science, 970:562–567, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiGesu:1995:CBI

[651] V. Di Gesu, S. Mantaci, and G. Tortorici. Compression of binary images based on covering. Lecture Notes in Computer Science, 970:568–573, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustin:1995:MDF

[652] V. Gustin and M. Cufer. Motion detection with fuzzy logic in real time. Lecture Notes in Computer Science, 970:574–579, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liedtke:1995:UEK

[653] C.-E. Liedtke, O. Grau, and S. Growe. Use of explicit knowledge for the reconstruction of 3-D object geometry. Lecture Notes in Computer Science, 970:580–587, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gofman:1995:DGL

[654] Y. Gofman and N. Kiryati. Detecting grey level symmetry: The frequency domain approach. Lecture Notes in Computer Science, 970:588–593, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luchowski:1995:NMT

[655] L. Luchowski. A new method to threshold images of flat binary scenes under uneven lighting. Lecture Notes in Computer Science, 970:594–599, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunn:1995:ISP

[656] S. R. Gunn and M. S. Nixon. Improving snake performance via a dual active contour. Lecture Notes in Computer Science, 970:600–605, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duin:1995:SPM

[657] R. P. W. Duin and E. T. G. Hoek. SMD position measurement by a Kohonen network compared with image processing. Lecture Notes in Computer Science, 970:606–611, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gracia:1995:EFD

[658] I. Gracia, F. Pla, F. J. Ferri, and P. Garcia. Estimating feature discriminant power in decision tree classifiers. Lecture Notes in Computer Science, 970:612–617, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sharp:1995:SST

[659] N. G. Sharp and E. R. Hancock. Statistical surface tracking. Lecture Notes in Computer Science, 970:618–624, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuncheva:1995:FSR

[660] L. I. Kuncheva and R. K. Kounchev. On feature selection via rough sets. Lecture Notes in Computer Science, 970:625–630, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voracek:1995:TNC

[661] J. Voracek. Tree neural classifier for character recognition. Lecture Notes in Computer Science, 970:631–636, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fidrich:1995:MEF

[662] M. Fidrich and J.-P. Thirion. Multiscale extraction of features from medical images. Lecture Notes in Computer Science, 970:637–642, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frucci:1995:LRE

[663] M. Frucci and A. Marcelli. Line representation of elongated shapes. Lecture Notes in Computer Science, 970:643– 648, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:FCD

[664] L. Yang, F. Albregtsen, and T. Taxt. Fast computation of 3-D geometric moments using a discrete Gauss’ theorem. Lecture Notes in Computer Science, 970:649–654, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chernov:1995:GAH

[665] V. M. Chernov. On the group algebras’ hierarchy pertaining to the parametrization of fast algorithms of discrete orthogonal transforms. Lecture Notes in Computer Science, 970:655– 660, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chinveeraphan:1995:URV

[666] S. Chinveeraphan, R. Takamatsu, and M. Sato. Understanding of ridge-valley lines on image-intensity surfaces in scale-space. Lecture Notes in Computer Science, 970:661–667, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fainzilberg:1995:CAR

[667] L. Fainzilberg and T. Potapova. Computer analysis and recognition of cognitive phase space electrocardiographic image. Lecture Notes in Computer Science, 970:668–673, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoddart:1995:NAP

[668] A. J. Stoddart, M. Petrou, and J. Kittler. A new algorithm for probabilistic relaxation based on the Baum Eagon theorem. Lecture Notes in Computer Science, 970:674–679, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bors:1995:SEO

[669] A. G. Bors and I. Pitas. Segmentation and estimation of the optical flow. Lecture Notes in Computer Science, 970:680–685, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boecker:1995:ADF

[670] W. Boecker, H.-W. Gantenberg, W.-U. Mueller, and C. Streffer. Automated detection of fluorescent cells and measurement of their DNA-Content. Lecture Notes in Computer Science, 970: 686–691, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cucchiara:1995:HSH

[671] R. Cucchiara and F. Filicori. A highly selective HT based algorithm for detecting extended, almost rectilinear shapes. Lecture Notes in Computer Science, 970:692–698, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bayro-Corrochano:1995:OMC

[672] E. Bayro-Corrochano and G. Sommer. Object modelling and collision avoidance using Clifford algebra. Lecture Notes in Computer Science, 970:699– 704, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veillon:1995:JIT

[673] F. Veillon, L. Quan, and P. Sturm. Joint invariants of a triplet of coplanar conics: Stability and discriminating power for object recognition. Lecture Notes in Computer Science, 970:705– 710, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnspang:1995:UMC

[674] J. Arnspang, H. Nielsen, M. Christensen, and K. Henriksen. Using mirror cameras for estimating depth. Lecture Notes in Computer Science, 970:711– 716, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colombo:1995:CHT

[675] C. Colombo, A. Del Bimbo, and S. De Magistris. Combining head tracking and pupil monitoring in vision-based human-computer interaction. Lecture Notes in Computer Science, 970:717– 722, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BrentSeales:1995:ACC

[676] W. Brent Seales and D. Eggert. Activecamera calibration using iterative image feature localization. Lecture Notes in Computer Science, 970:723–728, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suk:1995:PIP

[677] T. Suk and J. Flusser. The projective invariants for polygons. Lecture Notes in Computer Science, 970:729– 734, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haindl:1995:MMI

[678] M. Haindl and S. Simberova. A multimodel image line reconstruction. Lecture Notes in Computer Science, 970: 735–740, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eikvil:1995:TAR

[679] L. Eikvil, K. Aas, and M. Holden. Tools for automatic recognition of character strings in maps. Lecture Notes in Computer Science, 970:741–746, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kropatsch:1995:THD

[680] W. G. Kropatsch. Towards higher decimation ratios. Lecture Notes in Computer Science, 970:747–752, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klette:1995:RDP

[681] R. Klette, I. Stojmenovic, and J. Zunic. A representation of digital planes by least square fits. Lecture Notes in Computer Science, 970:753–758, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:SAE

[682] L. Yang, F. Albregtsen, T. Loennestad, and P. Groettum. A supervised approach to the evaluation of image segmentation methods. Lecture Notes in Computer Science, 970:759–765, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuncheva:1995:EIE

[683] L. I. Kuncheva and Y. K. Yotzov. Experimental investigation on editing for the k-NN rule through a genetic algorithm. Lecture Notes in Computer Science, 970:766–771, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jankovic:1995:HMM

[684] V. Jankovic, E. Ruzicky, and E. Groeller. Heterogeneous morphing of multimodal medical information. Lecture Notes in Computer Science, 970:772– 777, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leonardis:1995:RPD

[685] A. Leonardis, S. Kovacic, and F. Pernus. Recognition and pose determination of 3-D objects using multiple views. Lecture Notes in Computer Science, 970:778–783, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meijster:1995:PIP

[686] A. Meijster and J. B. T. M. Roerdink. A proposal for the implementation of a parallel watershed algorithm. Lecture Notes in Computer Science, 970:790– 795, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turner:1995:BED

[687] M. Turner and E. R. Hancock. Bayesian extraction of differential surface structure. Lecture Notes in Computer Science, 970:790–795, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Law:1995:NNE

[688] T. Law, H. Itoh, and H. Seki. A neural network energy minimization approach to approximation of 2dimensional shapes. Lecture Notes in Computer Science, 970:796–801, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anzalone:1995:VDD

[689] A. Anzalone and A. Machi. Visual detection of defects in moulded plastic drippers. Lecture Notes in Computer Science, 970:802–807, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogs:1995:SM

[690] A. Hoogs and R. Bajcsy. Segmentation modeling. Lecture Notes in Computer Science, 970:808–813, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faure:1995:IRM

[691] F. Faure. Inverting the reflectance map with binary search. Lecture Notes in Computer Science, 970:814–819, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palmer:1995:FST

[692] P. L. Palmer, N. Fatemi-Ghomi, and M. Petrou. Feature selection for the tree-wavelet transform. Lecture Notes in Computer Science, 970:820–825, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1995:PCD

[693] S. Yu and K. Weigl. Performance comparison of a deterministic and a stochastic method for image classification. Lecture Notes in Computer Science, 970:826–831, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pudil:1995:ASM

[694] P. Pudil, J. Novovicova, F. Ferri, and J. Kittler. Advances in the statistical methodology for the selection of image descriptors for visual pattern representation and classification. Lecture Notes in Computer Science, 970:832– 837, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sturm:1995:ASC

[695] P. Sturm and L. Quan. Affine stereo calibration. Lecture Notes in Computer Science, 970:838–843, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kadyrov:1995:TFL

[696] A. Kadyrov. Triple features for linear distorted images. Lecture Notes in Computer Science, 970:844–849, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chmielewski:1995:FPR

[697] L. Chmielewski, E. Chmielewska, M. Sklodowski, and W. Cudny. Finding postirradiation reaction in lungs from digitized X-rays. Lecture Notes in Computer Science, 970:850–855, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnspang:1995:ETC

[698] J. Arnspang, K. Henriksen, and R. Stahr. Estimating time to contact with curves, avoiding calibration and aperture problem. Lecture Notes in Computer Science, 970:856–861, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hatef:1995:CPR

[699] M. Hatef and J. Kittler. Constraining probabilistic relaxation with symbolic attributes. Lecture Notes in Computer Science, 970:862–867, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maki:1995:LFP

[700] A. Maki, L. Bretzner, and J.-O. Eklundh. Local Fourier phase and disparity estimates: An analytical study. Lecture Notes in Computer Science, 970:868–873, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fornland:1995:DOD

[701] P. Fornland. Direct obstacle detection and motion from spatio-temporal derivatives. Lecture Notes in Computer Science, 970:874–879, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roy:1995:DER

[702] S. Roy and I. J. Cox. Direct estimation of rotation from two frames via epipolar search. Lecture Notes in Computer Science, 970:880–887, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrianasy:1995:DCR

[703] F. Andrianasy and M. Milgram. Dynamic character recognition using an elastic matching. Lecture Notes in Computer Science, 970:888–893, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ungureanu:1995:VRG

[704] D. Ungureanu, L. Van Gool, and T. Moons. Visual robot guidance for an insertion task. Lecture Notes in Computer Science, 970:894–899, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janova:1995:RSR

[705] D. Janova and J. Jan. Robust surface reconstruction from stereo SEM images. Lecture Notes in Computer Science, 970:900–905, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kautsky:1995:AWS

[706] J. Kautsky and R. Turcajova. Adaptive wavelets for signal analysis. Lecture Notes in Computer Science, 970:906– 911, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michaelis:1995:CFP

[707] M. Michaelis, R. Herpers, and G. Sommer. A common framework for preattentive and attentive vision using steerable filters. Lecture Notes in Computer Science, 970:912–919, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:SGS

[708] C. Lee, Y.-F. Wang, and T. Yang. Static global scheduling for optimal computer vision and image processing operations on distributed-memory multiprocessors. Lecture Notes in Computer Science, 970:920–925, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herwig:1995:RPC

[709] C. Herwig and H.-O. Carmesin. Robust patch concept for egomotion estimation. Lecture Notes in Computer Science, 970:926–931, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xia:1995:PTA

[710] F. Xia. Parallel thinning algorithm based on the wave propagation model. Lecture Notes in Computer Science, 970:932–937, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corridoni:1995:FER

[711] J. M. Corridoni and A. Del Bimbo. Film editing reconstruction and semantic analysis. Lecture Notes in Computer Science, 970:938–943, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seales:1995:MTC

[712] W. Brent Seales. Measuring timeto-contact using active camera control. Lecture Notes in Computer Science, 970:944–949, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiemker:1995:CCP

[713] R. Wiemker. The color constancy problem: An illumination invariant mapping approach. Lecture Notes in Computer Science, 970:950–955, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohamed:1995:MCE

[714] Otmane Ait Mohamed. Mechanizing acalculus equivalence within HOL. Lecture Notes in Computer Science, 971: 1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agerholm:1995:NPR

[715] S. Agerholm. Non-primitive recursive function definitions. Lecture Notes in Computer Science, 971:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agerholm:1995:EZS

[716] S. Agerholm and M. Gordon. Experiments with ZF set theory in HOL and isabelle. Lecture Notes in Computer Science, 971:32–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Black:1995:AST

[717] P. E. Black and P. J. Windley. Automatically synthesized term denotation predicates: A proof aid. Lecture Notes in Computer Science, 971:46–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodeveiz:1995:RSM

[718] J.-P. Bodeveiz and M. Filali. On the refinement of symmetric memory protocols. Lecture Notes in Computer Science, 971:58–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulton:1995:CDP

[719] R. J. Boulton. Combining decision procedures in the HOL system. Lecture Notes in Computer Science, 971:75–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brackin:1995:DCP

[720] S. H. Brackin. Deciding cryptographic protocol adequacy with HOL. Lecture Notes in Computer Science, 971:90–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busch:1995:PMR

[721] H. Busch. A practical method for reasoning about distributed systems in a theorem prover. Lecture Notes in Computer Science, 971:106–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collins:1995:TFM

[722] G. Collins and D. Syme. A theory of finite maps. Lecture Notes in Computer Science, 971:122–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curzon:1995:VT

[723] P. Curzon. Virtual theories. Lecture Notes in Computer Science, 971:138– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eisenbiegler:1995:ATD

[724] D. Eisenbiegler and R. Kumar. An automata theory dedicated towards formal circuit synthesis. Lecture Notes in Computer Science, 971:154–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunter:1995:IHF

[725] E. L. Gunter and L. Libkin. Interfacing HOL90 with a functional database query language. Lecture Notes in Computer Science, 971:170–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1995:FPV

[726] J. Harrison. Floating point verification in HOL. Lecture Notes in Computer Science, 971:186–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1995:IDA

[727] J. Harrison. Inductive definitions: Automation and application. Lecture Notes in Computer Science, 971:200– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalvala:1995:FTI

[728] S. Kalvala. A formulation of TLA in isabelle. Lecture Notes in Computer Science, 971:214–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1995:FVS

[729] Jang Dae Kim and S.-K. Chin. Formal verification of serial pipeline multipliers. Lecture Notes in Computer Science, 971:229–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laangbacka:1995:TTW

[730] T. Laangbacka, R. Ruksenas, and J. Von Wright. TkWinHOL: A tool for window inference in HOL. Lecture Notes in Computer Science, 971:245– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loewenstein:1995:FVC

[731] P. N. Loewenstein. Formal verification of counterflow pipeline architecture. Lecture Notes in Computer Science, 971:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reetz:1995:DEV

[732] R. Reetz. Deep embedding VHDL. Lecture Notes in Computer Science, 971: 277–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Regensburger:1995:HHO

[733] F. Regensburger. HOLCF: Higher order logic of computable functions. Lecture Notes in Computer Science, 971: 293–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schubert:1995:MLS

[734] T. Schubert and S. Mocas. A mechanized logic for secure key escrow protocol verification. Lecture Notes in Computer Science, 971:308–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Syme:1995:NIH

[735] D. Syme. A new interface for HOL — ideas, issues and implementation. Lecture Notes in Computer Science, 971: 324–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welinder:1995:VEC

[736] M. Welinder. Very efficient conversions. Lecture Notes in Computer Science, 971:340–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1995:RCH

[737] W. Wong. Recording and checking HOL proofs. Lecture Notes in Computer Science, 971:353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamamoto:1995:FPG

[738] M. Yamamoto, S.-Y. Nishizaki, M. Hagiya, and Y. Toda. Formalization of planar graphs. Lecture Notes in Computer Science, 971:369–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cui:1995:HMR

[739] Z. Cui, B. R. Becker, M. R. Heckman, and K. Levitt. A hierarchical method for reasoning about distributed programming languages. Lecture Notes in Computer Science, 971:385–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gouda:1995:TTS

[740] M. G. Gouda. The triumph and tribulation of system stabilization. Lecture Notes in Computer Science, 972:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jayanti:1995:WFC

[741] P. Jayanti. Wait-free computing. Lecture Notes in Computer Science, 972: 19–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeLann:1995:RTN

[742] G. Le Lann. On real-time and non realtime distributed computing. Lecture Notes in Computer Science, 972:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dolev:1995:TPD

[743] D. Dolev. Theory and practice in distributed systems. Lecture Notes in Computer Science, 972:71–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babaoglu:1995:ICS

[744] O. Babaoglu, R. Davoli, L. Giachini, and P. Sabattini. The inherent cost of strong-partial view-synchronous communication. Lecture Notes in Computer Science, 972:72–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerraoui:1995:RRB

[745] R. Guerraoui. Revisiting the relationship between non-blocking atomic commitment and consensus. Lecture Notes in Computer Science, 972:87–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ricciardi:1995:DDC

[746] A. Ricciardi. Dissecting distributed coordination. Lecture Notes in Computer Science, 972:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Awerbuch:1995:OBP

[747] B. Awerbuch, S. Kutten, Y. Mansour, and D. Peleg. Optimal broadcast with partial knowledge. Lecture Notes in Computer Science, 972:116–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flammini:1995:MDI

[748] M. Flammini, G. Gambosi, U. Nanni, and R. B. Tan. Multi-dimensional interval routing schemes. Lecture Notes in Computer Science, 972:131– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakoby:1995:DTP

[749] A. Jakoby and R. Reischuk. Data transmission in processor networks. Lecture Notes in Computer Science, 972:145–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spirakis:1995:DPA

[750] P. Spirakis, B. Tampakas, and H. Antonopoulou. Distributed protocols against mobile eavesdroppers. Lecture Notes in Computer Science, 972: 160–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1995:UCL

[751] J. H. Anderson and M. Moir. Universal constructions for large objects. Lecture Notes in Computer Science, 972: 168–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arora:1995:LBE

[752] A. Arora and M. G. Gouda. Load balancing: an exercise in constrained convergence. Lecture Notes in Computer Science, 972:183–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shapiro:1995:LRD

[753] M. Shapiro and P. Ferreira. Larchant-RDOSS: a distributed shared persistent memory and its garbage collector. Lecture Notes in Computer Science, 972:198–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bao:1995:BHR

[754] F. Bao, Y. Igarashi, and K. Katano. Broadcasting in hypercubes with randomly distributed Byzantine faults. Lecture Notes in Computer Science, 972:215–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borcherding:1995:NAR

[755] M. Borcherding. On the number of authenticated rounds in Byzantine agreement. Lecture Notes in Computer Science, 972:230–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moser:1995:TOA

[756] L. E. Moser and P. M. Melliar-Smith. Total ordering algorithms for asynchronous Byzantine systems. Lecture Notes in Computer Science, 972:242– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butelle:1995:USS

[757] F. Butelle, C. Lavault, and M. Bui. A uniform self-stabilizing minimum diameter tree algorithm. Lecture Notes in Computer Science, 972:257–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoepman:1995:SSW

[758] J.-H. Hoepman, M. Papatriantafilou, and P. Tsigas. On self-stabilization in wait-free shared memory objects. Lecture Notes in Computer Science, 972: 273–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Itkis:1995:DCS

[759] G. Itkis, C. Lin, and J. Simon. Deterministic, constant space, selfstabilizing leader election on uniform rings. Lecture Notes in Computer Science, 972:288–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chase:1995:EDR

[760] C. M. Chase and V. K. Garg. Efficient detection of restricted classes of global predicates. Lecture Notes in Computer Science, 972:303–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoller:1995:FPD

[761] S. D. Stoller and F. B. Schneider. Faster possibility detection by combining two approaches. Lecture Notes in Computer Science, 972:318–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kroha:1995:SER

[762] P. Kroha. Shortcomings and extensions of relational DBMS. Lecture Notes in Computer Science, 973:50–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heumer:1995:EOO

[763] C. Heumer, G. Kappel, S. Rausch-Schott, and W. Retschitzegger. Evaluation of object-oriented database systems. Lecture Notes in Computer Science, 973:70–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rausch-Schott:1995:AOO

[764] K. G. Rausch-Schott, W. Retschitzegger, A. M. Tjoa, and S. Vieweg. Active object-oriented database systems for CIM applications. Lecture Notes in Computer Science, 973:96–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kroha:1995:DBR

[765] P. Kroha. Databdrige between RDBMS and OODBMS. Lecture Notes in Computer Science, 973:132–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappel:1995:DRC

[766] G. Kappel and S. Vieweg. Database requirements of CIM applications. Lecture Notes in Computer Science, 973: 136–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finke:1995:PS

[767] G. Finke. Planning and scheduling. Lecture Notes in Computer Science, 973:165–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Widmer:1995:ORM

[768] M. Widmer. Operational research models and methods in CIM. Lecture Notes in Computer Science, 973:179– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sriharan:1995:PPC

[769] V. Sriharan and J. J. Kanet. Production planning and control systems — state of the art and new directions. Lecture Notes in Computer Science, 973:195–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adelsberger:1995:STD

[770] H. H. Adelsberger and C. Hohendorf. Simulation — A tool for developing advanced production strategies. Lecture Notes in Computer Science, 973:217– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conen:1995:DMC

[771] W. Conen. Dynamic modeling of CIM systems. Lecture Notes in Computer Science, 973:237–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horacek:1995:FMC

[772] P. Horacek. Fuzzy modeling and control. Lecture Notes in Computer Science, 973:257–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matousek:1995:FCC

[773] V. Matousek. Foundations of computer communications. Lecture Notes in Computer Science, 973:289–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Traunmueller:1995:IMI

[774] R. Traunmueller. Information management and information systems planning. Lecture Notes in Computer Science, 973:325–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joergensen:1995:OFM

[775] F. Joergensen. Overview of function modelling — IDEFO. Lecture Notes in Computer Science, 973:340–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adelsberger:1995:DMI

[776] H. H. Adelsberger and F. Koerner. Data modeling with IDEFIX. Lecture Notes in Computer Science, 973:355– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacak:1995:APC

[777] W. Jacak and B. Kapsammer. Automatic programming in CIM based on intelligent tools. Lecture Notes in Computer Science, 973:392–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drazan:1995:MBD

[778] P. J. Drazan. Model based decision support systems. Lecture Notes in Computer Science, 973:418–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kouba:1995:MMI

[779] Z. Kouba and T. Vlcek. Man — machine interface for CIM. Lecture Notes in Computer Science, 973:427– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hensgens:1995:DES

[780] J. Hensgens. Designing for evolutionary systems. Lecture Notes in Computer Science, 973:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preucil:1995:ASC

[781] I. Preucil. Active subsystems for CIM environments. Lecture Notes in Computer Science, 973:454–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slavik:1995:CST

[782] P. Slavik. CAD systems: Trends and developments. Lecture Notes in Computer Science, 973:478–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fohn:1995:CE

[783] S. M. Fohn, A. Greef, R. E. Young, and P. O’Grady. Concurrent engineering. Lecture Notes in Computer Science, 973:493–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Specht:1995:GSF

[784] D. Specht and F. Fehler. Groupwork in the shop-floor-area needs decentral CIM-Structures and components. Lecture Notes in Computer Science, 973:506–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marik:1995:DCN

[785] V. Marik, J. Lazansky, and O. Stepankova. Does CIM need AI? Lecture Notes in Computer Science, 973:517– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keilmann:1995:MAS

[786] K.-P. Keilmann. Multi-agent systems: A natural trend in CIM. Lecture Notes in Computer Science, 973:548– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neubauer:1995:PSG

[787] M. Neubauer. Production scheduling and genetic algorithms. Lecture Notes in Computer Science, 973:563– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Csonto:1995:ISD

[788] J. Csonto. Implementation of systems with declarative constraints. Lecture Notes in Computer Science, 973:583– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marvan:1995:QRC

[789] I. Marvan and O. Stepankova. Qualitative reasoning and CIM. Lecture Notes in Computer Science, 973:594– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Starzacher:1995:CAS

[790] P. Starzacher, G. Quirchmayr, and R. Wagner. Comparison and analysis of selected CIM-Laboratory concepts and their importance for the improvement of staff training. Lecture Notes in Computer Science, 973:613–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tjoa:1995:UAD

[791] A. M. Tjoa, R. R. Wagner, C. Gierlinger, and C. Kaiser. The use of the active database SIMON for a short term production planning system in the automotive industry. Lecture Notes in Computer Science, 973:638–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1995:TAR

[792] A. K. Jain and K. Karu. Texture analysis: Representation and matching. Lecture Notes in Computer Science, 974:

3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rouquet:1995:RBS

[793] C. Rouquet and P. Bonton. Regionbased segmentation of textured images. Lecture Notes in Computer Science, 974:11–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Succi:1995:NAI

[794] A. Succi and V. Torre. A new approach to image segmentation. Lecture Notes in Computer Science, 974:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campadelli:1995:UHN

[795] P. Campadelli, D. Medici, and R. Schettini. Using Hopfield networks to segment color images. Lecture Notes in Computer Science, 974:25–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wegner:1995:WTM

[796] S. Wegner, T. Harms, J. H. Builtjes, and H. Oswald. The watershed transformation for multiresolution image segmentation. Lecture Notes in Computer Science, 974:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1995:BOE

[797] O. V. Larsen, P. Radeva, and E. Marti. Bounds on the optimal elasticity parameters for a snake. Lecture Notes in Computer Science, 974:37–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bunke:1995:EAG

[798] H. Bunke and B. T. Messmer. Efficient attributed graph matching and its application to image analysis. Lecture Notes in Computer Science, 974:45–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finch:1995:MDG

[799] A. M. Finch and E. R. Hancock. Matching Delaunay graphs. Lecture Notes in Computer Science, 974:56–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1995:IAG

[800] R. C. Wilson and E. R. Hancock. An integrated approach to grouping and matching. Lecture Notes in Computer Science, 974:62–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordella:1995:ARO

[801] L. P. Cordella, C. De Stefano, C. Sansone, and M. Vento. An adaptive reject option for LVQ classifiers. Lecture Notes in Computer Science, 974:68–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaszka:1995:MBM

[802] T. Blaszka and R. Deriche. A model based method for characterization and location of curved image features. Lecture Notes in Computer Science, 974: 77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matas:1995:ECT

[803] J. Matas, Z. Shao, and J. Kittler. Estimation of curvature and tangent direction by median filtered differencing. Lecture Notes in Computer Science, 974:83–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziou:1995:AEF

[804] D. Ziou and S. Tabbone. Adaptive elimination of false edges for first order detectors. Lecture Notes in Computer Science, 974:89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koelma:1995:IPL

[805] D. Koelma and A. Smeulders. An image processing library based on abstract image data-types in C++. Lecture Notes in Computer Science, 974: 97–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maruyama:1995:STD

[806] M. Maruyama, T. Teraoka, T. Tamada, and S. Nishida. Spatio-temporal data management for a 3D graphical information system. Lecture Notes in Computer Science, 974:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SouzaSol:1995:POO

[807] A. A. Souza Sol and A. De Albuquerque Araujo. PhotoPix: an objectoriented framework for digital image processing systems. Lecture Notes in Computer Science, 974:109–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broggi:1995:DPV

[808] A. Broggi. Data packing vs. processing speed on low-cost massively parallel systems. Lecture Notes in Computer Science, 974:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kutics:1995:NPM

[809] A. Kutics and M. Date. A new parallel method based on a genetic approach for determination and classification of skin spots. Lecture Notes in Computer Science, 974:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baglietto:1995:VSP

[810] P. Baglietto, M. Maresca, A. Migliaro, and M. Migliardi. A VLSI scalable processor array for motion estimation. Lecture Notes in Computer Science, 974:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cantoni:1995:EDA

[811] V. Cantoni, L. Cinque, S. Levialdi, and L. Lombardi. Evaluating digital angles by a parallel diffusion process. Lecture Notes in Computer Science, 974:133–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albanesi:1995:PAC

[812] M. G. Albanesi, M. Ferretti, and L. Zangrandi. A pyramidal approach to convex hull and filling algorithms. Lecture Notes in Computer Science, 974: 139–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biancardi:1995:CBA

[813] A. Biancardi and M. Pini. A code based approach to image-processing teaching. Lecture Notes in Computer Science, 974:145–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bothe:1995:VSM

[814] H. H. Bothe. A visual speech model based on fuzzy-neuro methods. Lecture Notes in Computer Science, 974: 153–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinho:1995:EDF

[815] A. J. Pinho and L. B. Almeida. Edge detection filters based on artificial neural networks. Lecture Notes in Computer Science, 974:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petkov:1995:UCF

[816] N. Petkov. Use of cortical filters and neural networks in a self-organising image classification system. Lecture Notes in Computer Science, 974:165– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:OOR

[817] J.-S. Lee, C.-H. Chen, Y.-N. Sun, and G.-S. Tseng. Occluded objects recognition using multiscale features and Hopfield neural networks. Lecture Notes in Computer Science, 974:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corridoni:1995:AVS

[818] J. M. Corridoni and A. Del Bimbo. Automatic video segmentation through editing analysis. Lecture Notes in Computer Science, 974:179–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelBimbo:1995:IRE

[819] A. Del Bimbo and P. Pala. Image retrieval by elastic matching of user sketches. Lecture Notes in Computer Science, 974:185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranefall:1995:FFF

[820] P. Ranefall, B. Nordin, and E. Bengtsson. Finding facial features using an HLS colour space. Lecture Notes in Computer Science, 974:191–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1995:NFR

[821] J. Jiang. A new and fast real-time implementation of 2-D DCT. Lecture Notes in Computer Science, 974:197– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takamura:1995:HLC

[822] S. Takamura and M. Takagi. A hybrid lossless compression of still images using Markov models and linear prediction. Lecture Notes in Computer Science, 974:203–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stiller:1995:IAQ

[823] C. Stiller. The impact of aliasing and quantization on motion compensation. Lecture Notes in Computer Science, 974:209–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kovalev:1995:RBM

[824] V. A. Kovalev. Rule-based method for tumor recognition in liver ultrasonic images. Lecture Notes in Computer Science, 974:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taylor:1995:UMF

[825] C. C. Taylor, M. R. Faghihi, and I. L. Dryden. An understanding of muscle fibre images. Lecture Notes in Computer Science, 974:223–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1995:NPC

[826] G. Fernandez, M. Kunt, and J.-P. Zryd. A new plant cell image segmentation algorithm. Lecture Notes in Computer Science, 974:229–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Batista:1995:ETM

[827] J. Batista and R. Kitney. Extraction of tumours from MR images of the brain by texture and clustering. Lecture Notes in Computer Science, 974: 235–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bello:1995:AIB

[828] F. Bello and R. Kitney. Automatic identification of brain contours in magnetic resonance images of the head. Lecture Notes in Computer Science, 974:241–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DellaMea:1995:ASA

[829] V. Della Mea and C. A. Beltrami. Analysis of the spatial arrangement of cells in the proliferative breast lesions. Lecture Notes in Computer Science, 974:247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chitti:1995:AIP

[830] Y. Chitti. Application of image processing in neurobiology: Detection of low signals with high spatial resolution and a non-uniform variance. Lecture Notes in Computer Science, 974:253– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giachetti:1995:CAA

[831] A. Giachetti, M. Cappello, G. Gigli, and V. Torre. Computer assisted analysis of echocardiographic image sequences. Lecture Notes in Computer Science, 974:258–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vinitski:1995:TSM

[832] S. Vinitski, C. Gonzalez, C. Burnett, and F. Mohamed. Tissue segmentation in MRI as an informative indicator of disease activity in the brain. Lecture Notes in Computer Science, 974:265–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costa:1995:ASC

[833] M. S. Costa and L. G. Shapiro. Analysis of scenes containing multiple nonpolyhedral 3D objects. Lecture Notes in Computer Science, 974:273–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dherete:1995:FSA

[834] P. Dherete, L. Wendling, and J. Desachy. Fuzzy segmentation and astronomical images interpretation. Lecture Notes in Computer Science, 974:281– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bischof:1995:LHF

[835] W. F. Bischof and T. Caelli. Learning how to find patterns or objects in complex scenes. Lecture Notes in Computer Science, 974:287–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellaire:1995:AMU

[836] G. Bellaire, K. Schluens, A. Mitritz, and K. Gwinner. Adaptive matching using object models generated from photometric stereo images. Lecture Notes in Computer Science, 974:293– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roberto:1995:SPT

[837] V. Roberto and E. Trucco. Sensor planning techniques and active visual inspection. Lecture Notes in Computer Science, 974:301–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Enciso:1995:ESC

[838] R. Enciso and T. Vieville. Experimental self-calibration from four views. Lecture Notes in Computer Science, 974:307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garibotto:1995:VBN

[839] G. Garibotto, M. Ilic, and S. Masciangelo. Vision-based navigation in service robotics. Lecture Notes in Computer Science, 974:313–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1995:SOG

[840] K. Wu and M. D. Levine. Segmenting 3D objects into geons. Lecture Notes in Computer Science, 974:321–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arcelli:1995:DES

[841] C. Arcelli and G. Ramella. Delineation of elongated sub-patterns in a piecewise constant foreground. Lecture Notes in Computer Science, 974:335–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xia:1995:NBC

[842] F. Xia. On a new basic concept and topological invariant. Lecture Notes in Computer Science, 974:341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lambert:1995:LMI

[843] G. Lambert and H. Gao. Line moments and invariants for real time processing of vectorized contour data. Lecture Notes in Computer Science, 974: 347–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murino:1995:RNU

[844] V. Murino, E. Frumento, and F. Gabino. Restoration of noisy underwater acoustic images using Markov random fields. Lecture Notes in Computer Science, 974:355–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krebs:1995:PPS

[845] B. Krebs, B. Korn, and F. M. Wahl. Plausibilistic preprocessing of sparse range images. Lecture Notes in Computer Science, 974:361–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puppo:1995:SRR

[846] E. Puppo. Segmentation/reconstruction of range images based on piecewiselinear approximation. Lecture Notes in Computer Science, 974:367–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caelli:1995:PCO

[847] T. Caelli and M. Ollila. In-place covariance operators for computer vision. Lecture Notes in Computer Science, 974:375–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salotti:1995:NHS

[848] M. Salotti and M. Hatimi. A new heuristic search for boundary detection. Lecture Notes in Computer Science, 974:381–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakazawa:1995:SPR

[849] Y. Nakazawa, T. Komatsu, and T. Saito. Sub-pixel registration for super high resolution image acquisition based on temporal integration. Lecture Notes in Computer Science, 974: 387–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ottonello:1995:RFT

[850] C. Ottonello, S. Pagnan, and V. Murino. Robust features for textures in additive noise. Lecture Notes in Computer Science, 974:393–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caglioti:1995:DCC

[851] V. Caglioti. Decomposing contours into curves of different families. Lecture Notes in Computer Science, 974:399– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sorrenti:1995:PLA

[852] D. G. Sorrenti. A proposal on local and adaptive determination of filter scale for edge detection. Lecture Notes in Computer Science, 974:405–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larre:1995:DSE

[853] A. Larre and E. Montseny. On the detection of step edges in algorithms based on gradient vector analysis. Lecture Notes in Computer Science, 974: 411–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bres:1995:DAG

[854] S. Bres and H. Emptoz. Directions analysis in grey-level images. Lecture Notes in Computer Science, 974:417– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zribi:1995:UNP

[855] M. Zribi and F. Ghorbel. An unsupervised and non-parametric Bayesian image segmentation. Lecture Notes in Computer Science, 974:423–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitulano:1995:BIS

[856] S. Vitulano, C. Di Ruberto, and M. Nappi. A.I. based image segmentation. Lecture Notes in Computer Science, 974:429–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marino:1995:QEI

[857] F. Marino and G. Mastronardi. Quality enhancement in image enlargement. Lecture Notes in Computer Science, 974:435–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nystroem:1995:SOS

[858] I. Nystroem and G. Borgefors. Synthesising objects and scenes using the reverse distance transformation in 2D and 3D. Lecture Notes in Computer Science, 974:441–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Landraud:1995:TSU

[859] A. M. Landraud and S. Oh Yum. Texture segmentation using local phase differences in Gabor filtered images. Lecture Notes in Computer Science, 974: 447–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Accame:1995:DES

[860] M. Accame and F. De Natale. Disparity estimation for stereo sequences based on adaptive size hierarchical block matching. Lecture Notes in Computer Science, 974:453–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:MSC

[861] Y. Zhang and J. J. Gerbrands. Matching of stereo curves — A closed form solution. Lecture Notes in Computer Science, 974:459–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tascini:1995:MPO

[862] G. Tascini, P. Puliti, and P. Zingaretti. Multi-polygonal object tracking. Lecture Notes in Computer Science, 974:465–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cui:1995:OPE

[863] Y. Cui, J. Weng, and H. Reynolds. Optimal parameter estimation of ellipses. Lecture Notes in Computer Science, 974:471–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leite:1995:LDW

[864] J. A. F. Leite and E. R. Hancock. A linear discriminator of width. Lecture Notes in Computer Science, 974:477– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kindratenko:1995:BII

[865] V. V. Kindratenko, B. A. Treiger, and P. J. M. Van Espen. Binarization of inhomogeneously illuminated images. Lecture Notes in Computer Science, 974:483–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attali:1995:PDS

[866] D. Attali, G. Sanniti di Baja, and E. Thiel. Pruning discrete and semicontinuous skeletons. Lecture Notes in Computer Science, 974:488–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karasik:1995:CPH

[867] Y. B. Karasik. On commutative properties of halftoning and their applications. Lecture Notes in Computer Science, 974:494–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bottoni:1995:CRG

[868] P. Bottoni, L. Cinque, S. Levialdi, and L. Lombardi. Combining resolution and granularity for pattern recognition. Lecture Notes in Computer Science, 974:503–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turolla:1995:LCS

[869] E. Turolla, Y. Belaid, and A. Belaid. Line and cell searching in tables or forms. Lecture Notes in Computer Science, 974:509–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Congedo:1995:DWG

[870] G. Congedo, G. Dimauro, S. Impedovo, and G. Pirlo. Describing words by graphs. Lecture Notes in Computer Science, 974:515–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Congedo:1995:SRS

[871] G. Congedo, G. Dimauro, A. M. Forte, and S. Impedovo. Selecting reference signatures for on-line signature verification. Lecture Notes in Computer Science, 974:521–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zramdini:1995:APO

[872] A. Zramdini and R. Ingold. ApOFIS: an A priori optical font identification system. Lecture Notes in Computer Science, 974:527–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burge:1995:EWM

[873] M. Burge and G. Monagan. Extracting words and multi-part symbols in graphics rich documents. Lecture Notes in Computer Science, 974:533–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bussi:1995:SAM

[874] S. Bussi and F. Mangili. A semiautomatic method for form layout description. Lecture Notes in Computer Science, 974:539–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boccignone:1995:IUC

[875] G. Boccignone, M. De Santo, A. Chianese, and A. Picariello. Improving the use of contours for off-line cursive script segmentation. Lecture Notes in Computer Science, 974:545–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsuchiya:1995:MDA

[876] T. Tsuchiya, N. Nakajima, and T. Kamimura. A method for determining address format in the automated sorting of Japanese mail. Lecture Notes in Computer Science, 974: 551–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zidouri:1995:SFM

[877] A. B. C. Zidouri, S. Chinveeraphan, and M. Sato. Structural features by MCR expression for printed Arabic character recognition. Lecture Notes in Computer Science, 974:557–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shioyama:1995:GAT

[878] T. Shioyama, A. Okumura, and Y. Aoki. Genetic algorithm for thinning gray-scale images of characters. Lecture Notes in Computer Science, 974: 563–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcu:1995:BPD

[879] G. Marcu and S. Abe. Blue-print document analysis for color classification. Lecture Notes in Computer Science, 974:569–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagy:1995:DIA

[880] G. Nagy. Document image analysis: What is missing? Lecture Notes in Computer Science, 974:577–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1995:IFH

[881] J. Hu, M. K. Brown, and W. Turin. Invariant features for HMM based online handwriting recognition. Lecture Notes in Computer Science, 974:588– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Congedo:1995:RAA

[882] G. Congedo, G. Dimauro, S. Impedovo, and G. Pirlo. A robust analytical approach for handwritten word recognition. Lecture Notes in Computer Science, 974:594–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cinque:1995:RRC

[883] L. Cinque, S. Di Zenzo, and S. Levialdi. Recognition of rotated characters by inexact matching. Lecture Notes in Computer Science, 974:600–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1995:BIR

[884] M. Chan, G. T. Herman, and E. Levitan. Bayesian image reconstruction using a high-order interacting MRF model. Lecture Notes in Computer Science, 974:609–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1995:MST

[885] H.-K. Choi, E. Bengtsson, T. Jarkrans, and J. Vasko. Minimum spanning trees (MST) as a tool for describing tissue architecture when grading bladder carcinoma. Lecture Notes in Computer Science, 974:615–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baroni:1995:SIC

[886] M. Baroni, G. Congiu, A. Del Bimbo, and A. Evangelisti. Symbolic indexing of cardiological sequences through dynamic curve representations. Lecture Notes in Computer Science, 974:621– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Netravali:1995:VC

[887] A. N. Netravali. Visual communication. Lecture Notes in Computer Science, 974:629–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Popescu:1995:GSI

[888] D. C. Popescu, A. Dimca, and H. Yan. Generalized square isometries — an improvement for fractal image coding. Lecture Notes in Computer Science, 974:637–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braccini:1995:FAH

[889] C. Braccini, F. Cocurullo, and F. Lavagetto. A fast algorithm for high quality vector quantization codebook design. Lecture Notes in Computer Science, 974:643–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1995:FFI

[890] H. Lin and A. N. Venetsanopoulos. Fast fractal image coding using pyramids. Lecture Notes in Computer Science, 974:649–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurianski:1995:HMD

[891] A. Kurianski and M. Nieniewski. Hidden MRF detection of motion of objects with uniform brightness. Lecture Notes in Computer Science, 974:657– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saito:1995:MAS

[892] T. Saito and T. Komatsu. Motion analysis and segmentation for objectoriented mid-level image representation. Lecture Notes in Computer Science, 974:663–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boeroeczky:1995:SNF

[893] L. Boeroeczky, R. Fioravanti, S. Fioravanti, and D. D. Giusto. Speckle noise filtering in SAR images using wavelets. Lecture Notes in Computer Science, 974:671–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolla:1995:PST

[894] R. Bolla, M. Marchese, C. Nobile, and S. Zappatore. Prediction of short-term evolution of cloud formations based on meteosat image sequences. Lecture Notes in Computer Science, 974:677– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Praud:1995:PIS

[895] S. Praud, P. Germain, and J. Plantier. Prototyping of interactive satellite image analysis tools using a real-time data-flow computer. Lecture Notes in Computer Science, 974:683–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marti:1995:SAI

[896] F. Marti and M. Naldi. Simulation of aircraft images for airport surface traffic control. Lecture Notes in Computer Science, 974:689–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ossola:1995:CKB

[897] J.-C. Ossola and M. Thonnat. Cooperation of knowledge based systems for galaxy classification. Lecture Notes in Computer Science, 974:697–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wendling:1995:FSS

[898] L. Wendling, M. Zehana, and J. Desachy. Fuzzy segmentation and structural knowledge for satellite image analysis. Lecture Notes in Computer Science, 974:703–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1995:DLL

[899] J.-C. Di Martino and S. Tabbone. Detection of lofar lines. Lecture Notes in Computer Science, 974:709–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tesei:1995:FOO

[900] A. Tesei, A. Trucco, and D. Zambonini. A fast object orientation estimation and recognition technique for underwater acoustic imaging. Lecture Notes in Computer Science, 974:715–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Semmar:1995:ACC

[901] N. Semmar. Applying contextual constraints to extract symbolic representation for image understanding. Lecture Notes in Computer Science, 974: 721–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greig:1995:AHT

[902] A. R. Greig. Application of the Hough transform for weld inspection underwater. Lecture Notes in Computer Science, 974:731–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darboux:1995:DAB

[903] M. Darboux and J.-M. Dinten. Development of an automated bond verification system for advanced electronic packages. Lecture Notes in Computer Science, 974:737–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruzzone:1995:CDM

[904] L. Bruzzone, F. Roli, and S. B. Serpico. Crack detection by a mesure of texture anisotropy. Lecture Notes in Computer Science, 974:743–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barni:1995:VSA

[905] M. Barni, V. Cappellini, and A. Mecocci. A vision system for automatic inspection of meat quality. Lecture Notes in Computer Science, 974:748– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stansfield:1995:DNF

[906] A. Stansfield and I. Page. The design of a new FPGA architecture. Lecture Notes in Computer Science, 975:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Payne:1995:STF

[907] R. Payne. Self-timed FPGA systems. Lecture Notes in Computer Science, 975:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Churcher:1995:XFP

[908] S. Churcher, T. Kean, and B. Wilkie. XC6200 Fastmap[TM] processor interface. Lecture Notes in Computer Science, 975:36–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snider:1995:TCC

[909] G. Snider, P. Kuekes, W. B. Culbertson, and R. J. Carter. The Teramac configurable computer engine. Lecture Notes in Computer Science, 975:44–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyazaki:1995:TOF

[910] T. Miyazaki, K. Yamada, A. Tsutsui, and H. Nakada. Telecommunicationoriented FPGA and dedicated CAD system. Lecture Notes in Computer Science, 975:54–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dunn:1995:CLP

[911] P. Dunn. A configurable logic processor for machine vision. Lecture Notes in Computer Science, 975:68–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1995:EDB

[912] P. Schulz. Extending DSP-Boards with FPGA-Based structures of interconnection. Lecture Notes in Computer Science, 975:78–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rachakonda:1995:HSR

[913] R. V. Rachakonda, P. M. Athanas, and A. L. Abbott. High-speed region detection and labeling using an FPGA based custom computing platform. Lecture Notes in Computer Science, 975:86–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaumont:1995:UFC

[914] C. Beaumont. Using FPGAs as control support in MIMD executions. Lecture Notes in Computer Science, 975:94–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerrero:1995:CHB

[915] F. Guerrero and J. M. Noras. Customised hardware based on the REDOC III algorithm for highperformance date ciphering. Lecture Notes in Computer Science, 975:104– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lawrence:1995:URH

[916] A. Lawrence, A. Kay, W. Luk, and T. Nomura. Using reconfigurable hardware to speed up product development and performance. Lecture Notes in Computer Science, 975:111–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casselman:1995:CHO

[917] S. Casselman, M. Thornburg, and J. Schewel. Creation of hardware objects in a reconfigurable computer. Lecture Notes in Computer Science, 975: 119–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turner:1995:RHP

[918] L. E. Turner and P. J. W. Graumann. Rapid hardware prototyping of digital signal processing systems using FPGAs. Lecture Notes in Computer Science, 975:129–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naseer:1995:DMM

[919] A. R. Naseer, M. Balakrishnan, and A. Kumar. Delay minimal mapping of RTL structures onto LUT based FPGAs. Lecture Notes in Computer Science, 975:139–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boemo:1995:SNP

[920] E. I. Boemo, G. Gonzalez de Rivera, S. Lopez-Buedo, and J. M. Meneses. Some notes on power management on FPGA-Based systems. Lecture Notes in Computer Science, 975:149–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimond:1995:ATR

[921] K. Dimond. An automatic technique for realising user interaction processing in PLD based systems. Lecture Notes in Computer Science, 975:158–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fields:1995:PUH

[922] C. A. Fields. The proper use of hierarchy in HDL-Based high density FPGA design. Lecture Notes in Computer Science, 975:168–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marnane:1995:CRA

[923] W. P. Marnane, C. N. Jordan, and F. J. O’Reilly. Compiling regular arrays onto FPGAs. Lecture Notes in Computer Science, 975:178–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guo:1995:CRF

[924] S. Guo and W. Luk. Compiling Ruby into FPGAs. Lecture Notes in Computer Science, 975:188–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greaves:1995:CVC

[925] D. Greaves. The CSYN verilog compiler and other tools. Lecture Notes in Computer Science, 975:198–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gschwind:1995:VDM

[926] M. Gschwind and V. Salapura. A VHDL design methodolgy for FP-GAs. Lecture Notes in Computer Science, 975:208–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khosravipour:1995:VBR

[927] M. Khosravipour and H. Gruenbacker. VHDL-Based rapid hardware prototyping using FPGA technology. Lecture Notes in Computer Science, 975:218– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weinhardt:1995:IPP

[928] M. Weinhardt. Integer programming for partitioning in software oriented codesign. Lecture Notes in Computer Science, 975:227–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahrens:1995:TSS

[929] K. Ahrens. Test standard serves dual role as on-board programming solution. Lecture Notes in Computer Science, 975:235–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zahm:1995:AMI

[930] U. Zahm, T. Hollstein, H.-J. Herpel, and N. Wehn. Advanced method for industry related education with an FPGA design self-learning kit. Lecture Notes in Computer Science, 975:241– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jebelean:1995:FIR

[931] T. Jebelean. FPGA implementation of a rational adder. Lecture Notes in Computer Science, 975:251–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klindworth:1995:FIC

[932] A. Klindworth. FPLD implementation of computation over finite fields GF(20m) with applications to error control coding. Lecture Notes in Computer Science, 975:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panneerselvam:1995:IFF

[933] G. Panneerselvam, P. J. W. Graumann, and L. E. Turner. Implementation of fast Fourier transforms and discrete cosine transforms in FP-GAs. Lecture Notes in Computer Science, 975:272–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shirazi:1995:IDF

[934] N. Shirazi, P. M. Athanas, and A. L. Abbott. Implementation of a 2-D fast Fourier transform on an FPGA-Based custom computing machine. Lecture Notes in Computer Science, 975:282– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersen:1995:ASF

[935] R. J. Petersen and B. L. Hutchings. An assessment of the suitability of FPGA-Based systems for use in digital signal processing. Lecture Notes in Computer Science, 975:293–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:FPM

[936] P. Lee. An FPGA prototype for a multiplierless FIR filter built using the logarithmic number system. Lecture Notes in Computer Science, 975:303– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turner:1995:BSF

[937] L. E. Turner, P. J. W. Graumann, and S. G. Gibb. BIT-Serial FIR filters with CSD coefficients for FP-GAs. Lecture Notes in Computer Science, 975:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atia:1995:SVT

[938] M. Atia, J. Bowles, D. W. Clarke, and M. P. Henry. A self-validating temperature sensor implemented in FP-GAs. Lecture Notes in Computer Science, 975:321–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wenban:1995:DIL

[939] A. Wenban, G. Brown, and J. O’Leary. Developing interface libraries for reconfigurable data acquisition boards. Lecture Notes in Computer Science, 975: 331–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herpel:1995:PGA

[940] H.-J. Herpel, U. Ober, and M. Glesner. Prototype generation of applicationspecific emedded controllers for microsystems. Lecture Notes in Computer Science, 975:341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1995:HGA

[941] P. Graham and B. Nelson. A hardware genetic algorithm for the travelling salesman problem on SPLASH 2. Lecture Notes in Computer Science, 975:352–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corba:1995:MAR

[942] M. Corba and Z. Ninkov. Modular architecture for real-time astronomical image processing with FP-GAs. Lecture Notes in Computer Science, 975:362–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woodward:1995:PSR

[943] D. R. Woodward, I. Page, D. C. Levy, and R. G. Harley. A programmable I/O system for real-time AC drive control applications. Lecture Notes in Computer Science, 975:370–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuddapah:1995:RLF

[944] R. Cuddapah and M. Corba. Reconfigurable logic for fault-tolerance. Lecture Notes in Computer Science, 975: 380–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guccione:1995:SRA

[945] S. A. Guccione and M. J. Gonzalez. Supercomputing with reconfigurable architectures. Lecture Notes in Computer Science, 975:389–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gokhale:1995:ASP

[946] M. Gokhale and A. Marks. Automatic synthesis of parallel programs targeted to dynamically reconfigurable logic arrays. Lecture Notes in Computer Science, 975:399–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lysaght:1995:PED

[947] P. Lysaght, H. Dick, G. McGregor, and D. McConnell. Prototyping environment for dynamically reconfigurable logic. Lecture Notes in Computer Science, 975:409–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutchings:1995:IAR

[948] B. L. Hutchings and M. J. Wirthlin. Implementation approaches for reconfigurable logic applications. Lecture Notes in Computer Science, 975:419– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brebner:1995:URV

[949] G. Brebner and J. Gray. Use of reconfigurability in variable-length code detection at video rates. Lecture Notes in Computer Science, 975:429–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guccione:1995:CPR

[950] S. A. Guccione and M. Gonzalez. Classification and performance of reconfigurable architectures. Lecture Notes in Computer Science, 975:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ajili:1995:CSL

[951] F. Ajili and E. Contejean. Complete solving of linear Diophantine equations and inequations without adding variables. Lecture Notes in Computer Science, 976:1–17, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domenjoud:1995:EMA

[952] E. Domenjoud and A. P. Tomas. From Elliott-MacMahon to an algorithm for general linear constraints on naturals. Lecture Notes in Computer Science, 976:18–35, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1995:PPP

[953] B. M. Smith, S. C. Brailsford, P. M. Hubbard, and H. P. Williams. The progressive party problem: Integer linear programming and constraint programming compared. Lecture Notes in Computer Science, 976:36–52, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koubarakis:1995:LGC

[954] M. Koubarakis. From local to global consistency in temporal constraint networks. Lecture Notes in Computer Science, 976:53–69, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gent:1995:SEC

[955] I. P. Gent, E. MacIntyre, P. Prosser, and T. Walsh. Scaling effects in the CSP phase transition. Lecture Notes in Computer Science, 976:70–87, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokoo:1995:AWC

[956] M. Yokoo. Asynchronous weakcommitment search for solving distributed constraint satisfaction problems. Lecture Notes in Computer Science, 976:88–102, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larrosa:1995:OBH

[957] J. Larrosa and P. Meseguer. Optimizationbased heuristics for maximal constraint satisfaction. Lecture Notes in Computer Science, 976:103–120, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grumbach:1995:FOD

[958] S. Grumbach and J. Su. First-order definability over constraint databases. Lecture Notes in Computer Science, 976:121–136, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldin:1995:SQT

[959] D. Q. Goldin and P. C. Kanellakis. On similarity queries for time-series data: Constraint specification and implementation. Lecture Notes in Computer Science, 976:137–153, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Revesz:1995:SSD

[960] P. Z. Revesz. Safe stratified datalog with integer order programs. Lecture Notes in Computer Science, 976:154– 169, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maher:1995:CD

[961] M. J. Maher. Constrained dependencies. Lecture Notes in Computer Science, 976:170–185, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girodias:1995:SLM

[962] P. Girodias, E. Cerny, and W. J. Older. Solving linear, min and max constraint systems using CLP based on relational interval arithmetic. Lecture Notes in Computer Science, 976:186–203, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meier:1995:DCP

[963] M. Meier. Debugging constraint programs. Lecture Notes in Computer Science, 976:204–221, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelly:1995:OCC

[964] A. D. Kelly, A. Macdonald, K. Marriott, and H. Soendergaard. An optimizing compiler for CLP(R). Lecture Notes in Computer Science, 976:222– 239, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dechter:1995:LGR

[965] R. Dechter and P. Van Beek. Local and global relational consistency. Lecture Notes in Computer Science, 976:240– 257, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bacchus:1995:DVO

[966] F. Bacchus and P. Van Run. Dynamic variable ordering in CSPs. Lecture Notes in Computer Science, 976:258– 275, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeavons:1995:UFT

[967] P. Jeavons, D. Cohen, and M. Gyssens. A unifying framework for tractable constraints. Lecture Notes in Computer Science, 976:276–291, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bacchus:1995:FCA

[968] F. Bacchus and A. Grove. On the forward checking algorithm. Lecture Notes in Computer Science, 976:292–309, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marriott:1995:CCC

[969] K. Marriott and M. Odersky. A confluent calculus for concurrent constraint programming with guarded choice. Lecture Notes in Computer Science, 976: 310–327, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Podelski:1995:SS

[970] A. Podelski and G. Smolka. Situated simplification. Lecture Notes in Computer Science, 976:328–344, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codognet:1995:GCD

[971] C. Codognet and P. Codognet. Guarded constructive disjunction: Angel or demon? Lecture Notes in Computer Science, 976:345–361, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaffanella:1995:DIA

[972] E. Zaffanella. Domain independent ask approximation in CCP. Lecture Notes in Computer Science, 976:362–379, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1995:CSC

[973] F. Baader and K. U. Schulz. On the combination of symbolic constraints, solution domains, and constraint solvers. Lecture Notes in Computer Science, 976:380–397, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:CPM

[974] J. Zhang and H. Zhang. Constraint propagation in model generation. Lecture Notes in Computer Science, 976:398–414, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werner:1995:NNW

[975] A. Werner. Normalizing narrowing for weakly terminating and confluent systems. Lecture Notes in Computer Science, 976:415–430, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colussi:1995:TCL

[976] L. Colussi, E. Marchiori, and M. Marchiori. On termination of constraint logic programs. Lecture Notes in Computer Science, 976:431–448, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simonis:1995:MPC

[977] H. Simonis and T. Cornelissens. Modelling producer/consumer constraints. Lecture Notes in Computer Science, 976:449–462, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sabin:1995:CBA

[978] D. Sabin, M. C. Sabin, R. D. Russell, and E. C. Freuder. A constraint-based approach to diagnosing software problems in computer networks. Lecture Notes in Computer Science, 976:463– 480, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerinik:1995:SCS

[979] N. Guerinik and M. Van Caneghem. Solving crew scheduling problems by constraint programming. Lecture Notes in Computer Science, 976:481–498, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paltrinieri:1995:VCP

[980] M. Paltrinieri. A visual constraintprogramming environment. Lecture Notes in Computer Science, 976:499– 514, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauzy:1995:PRS

[981] A. Rauzy. Polynomial restrictions of SAT: What can be done with an efficient implementation of the Davis and Putnam’s procedure? Lecture Notes in Computer Science, 976:515–532, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prestwich:1995:IBB

[982] S. Prestwich and S. Mudambi. Improved branch and bound in constraint logic programming. Lecture Notes in Computer Science, 976:533–548, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goltz:1995:RDS

[983] H.-J. Goltz. Reducing domains for search in CLP(FD) and its application to job-shop scheduling. Lecture Notes in Computer Science, 976:549–562, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchberger:1995:WSC

[984] B. Buchberger. What is symbolic computation? Lecture Notes in Computer Science, 976:563–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanHentenryck:1995:CSC

[985] P. Van Hentenryck. Constraint solving for combinatorial search problems: A tutorial. Lecture Notes in Computer Science, 976:564–587, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saraswat:1995:CCP

[986] V. Saraswat. The concurrent constraint programming research programmes. Lecture Notes in Computer Science, 976:588–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjoerner:1995:AGI

[987] N. Bjoerner, A. Browne, and Z. Manna. Automatic generation of invariants and intermediate assertions. Lecture Notes in Computer Science, 976:589–623, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borning:1995:OCS

[988] A. Borning and B. N. Freeman-Benson. The OTI constraint solver: A constraint library for constructing interactive graphical user interfaces. Lecture Notes in Computer Science, 976:624– 628, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fromherz:1995:MBC

[989] M. P. J. Fromherz and V. A. Saraswat. Model-based computing: Using concurrent constraint programming for modeling and model compilation. Lecture Notes in Computer Science, 976:629– 635, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meier:1995:CE

[990] M. Meier and A. Herold. CLP in ECRC. Lecture Notes in Computer Science, 976:636–639, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Touraivane:1995:CPI

[991] Touraïvane. Constraint programming and industrial applications. Lecture Notes in Computer Science, 976:640– 642, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simonis:1995:CSA

[992] H. Simonis. The CHIP system and its applications. Lecture Notes in Computer Science, 976:643–646, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puget:1995:ACP

[993] J.-F. Puget. Applications of constraint programming. Lecture Notes in Computer Science, 976:647–650, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinmetz:1995:ECS

[994] R. Steinmetz and L. C. Wolf. Evaluation of a CPU scheduling mechanism for synchronized multimedia streams. Lecture Notes in Computer Science, 977:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulos:1995:AST

[995] J. Boulos and D. Boudigue. An application of SMART2: A tool for performance evaluation of relational DataBase programs. Lecture Notes in Computer Science, 977:11–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsai:1995:MFT

[996] T. K. Tsai and R. K. Iyer. Measuring fault tolerance with the FTAPE fault injection tool. Lecture Notes in Computer Science, 977:26–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vinck:1995:QAD

[997] B. Vinck and H. Bruneel. Queueing analysis of discrete-time buffer systems with compound arrival process and variable service capacity. Lecture Notes in Computer Science, 977:41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paterok:1995:MMH

[998] M. Paterok, P. Dauphin, and U. Herzog. The method of moments for higher moments and the usefulness of formula manipulation systems. Lecture Notes in Computer Science, 977:56–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langer:1995:IPE

[999] U. Langer. Integration of performance evaluations in the design process of CPUs and computer systems. Lecture Notes in Computer Science, 977:71–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1995:IRS

[1000] L. G. Williams and C. U. Smith. Information requirements for software performance engineering. Lecture Notes in Computer Science, 977:86–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pooley:1995:IBS

[1001] R. Pooley. Integrating behavioural and simulation modelling. Lecture Notes in Computer Science, 977:102–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dunlop:1995:AGM

[1002] J. Dunlop. Assessment of 3rd generation mobile systems by simulation. Lecture Notes in Computer Science, 977: 117–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrmann:1995:HGS

[1003] C. Herrmann. How good is stationary analysis for the transient phenomena of connection admission in ATM? Lecture Notes in Computer Science, 977:135–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchholz:1995:EAA

[1004] P. Buchholz. On the exact and approximate analysis of hierarchical discrete time queueing networks. Lecture Notes in Computer Science, 977:150– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Telek:1995:SSA

[1005] M. Telek, A. Bobbio, L. Jereb, and A. Puliafito. Steady state analysis of Markov regenerative SPN with age memory policy. Lecture Notes in Computer Science, 977:165–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Touzene:1995:NIM

[1006] A. Touzene. A new iterative method for solving large-scale Markov chains. Lecture Notes in Computer Science, 977: 180–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

German:1995:TAD

[1007] R. German and J. Mitzlaff. Transient analysis of deterministic and stochastic Petri nets with TimeNET. Lecture Notes in Computer Science, 977:209– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bause:1995:QTS

[1008] F. Bause, P. Buchholz, and P. Kemper. QPN-Tool for the specification and analysis of hierarchically combined queueing Petri nets. Lecture Notes in Computer Science, 977:224–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haverkort:1995:AAN

[1009] B. R. Haverkort. Approximate analysis of networks of PHdPHd1dK queues: Theory and tool support. Lecture Notes in Computer Science, 977:239– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohr:1995:SIP

[1010] B. W. Mohr, A. D. Malony, and K. Shanmugam. Speedy: An integrated performance extrapolation tool for pC++ programs. Lecture Notes in Computer Science, 977:254–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dauphin:1995:HTA

[1011] P. Dauphin and R. Hofmann. Hasse: a tool for analyzing causal relationships in parallel and distributed systems. Lecture Notes in Computer Science, 977:269–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brehm:1995:PPP

[1012] J. Brehm, M. Madhukar, E. Smirni, and L. Dowdy. PerPreT — A performance prediction tool for massively parallel systems. Lecture Notes in Computer Science, 977:284–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGemund:1995:CTP

[1013] A. J. C. Van Gemund. Compiletime performance prediction of parallel systems. Lecture Notes in Computer Science, 977:299–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raghavan:1995:WMM

[1014] S. V. Raghavan, P. J. Joseph, and G. Haring. Workload models for multiwindow distributed environments. Lecture Notes in Computer Science, 977: 314–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1995:BHC

[1015] M. Lang, M. Stuempfle, and H. Kocher.

Building a hierarchical CAN-Simulator using an object-oriented environment. Lecture Notes in Computer Science, 977:327–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferscha:1995:POD

[1016] A. Ferscha and A. D. Malony.

Performance-oriented development of irregular, unstructured and unbalanced parallel applications in the N-MAP environment. Lecture Notes in Computer Science, 977:340–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hills:1995:PED

[1017] G. Hills, J. Rolia, and G. Serazzi.

Performance engineering of distributed software process architectures. Lecture Notes in Computer Science, 977:357– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermanns:1995:PEC

[1018] O. Hermanns. Performance evaluation of connectionless multicast protocols for cooperative multimedia applications. Lecture Notes in Computer Science, 977:372–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindemann:1995:MRM

[1019] C. Lindemann and F. Schoen. Modeling relaxed memory consistency protocols. Lecture Notes in Computer Science, 977:385–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fasbender:1995:MME

[1020] A. Fasbender and P. Davids. Measurement, modelling and emulation of Internet round-trip delays. Lecture Notes in Computer Science, 977:401– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhargava:1995:CCD

[1021] B. Bhargava and M. Annamalai. Communication costs in digital library databases. Lecture Notes in Computer Science, 978:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dogac:1995:MOO

[1022] A. Dogac, M. Altinel, C. Ozkan, and B. Arpinar. METU object-oriented DBMS kernel. Lecture Notes in Computer Science, 978:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jun:1995:DIS

[1023] Y. S. Jun and S. I. Yoo. Design and implementation of the spareC++ OODBPL. Lecture Notes in Computer Science, 978:28–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kusch:1995:GUV

[1024] J. Kusch, P. Hartel, T. Hartmann, and G. Saake. Gaining a uniform view of different integration aspects in a prototyping environment. Lecture Notes in Computer Science, 978:38–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:FBA

[1025] J. Yang and M. Papazoglou. A framebased approach for interoperation support in multidatabase systems. Lecture Notes in Computer Science, 978:48–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coenen:1995:DDD

[1026] F. P. Coenen, I. Finch, M. J. R. Shave, and T. J. M. Bench-Capon. Developing distributed database applications using TSL. Lecture Notes in Computer Science, 978:58–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boehm:1995:PCD

[1027] C. Boehm, A. Oppitz, P. Vogel, and S. Wiesener. Prints of the 17th century in a distributed digital library system. Lecture Notes in Computer Science, 978:68–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

PastorLopez:1995:OVC

[1028] O. Pastor Lopez, I. Ramos Salavert, and J. H. Canos Cerda. OASIS v2: A class definition language. Lecture Notes in Computer Science, 978:79–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SouzadosSantos:1995:DIO

[1029] C. Souza dos Santos. Design and implementation of object-oriented views. Lecture Notes in Computer Science, 978:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiLoreto:1995:VOO

[1030] F. Di Loreto, F. Ferri, F. Massari, and M. Rafanelli. A visual objectoriented query language for geographic information systems. Lecture Notes in Computer Science, 978:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kagawa:1995:CEQ

[1031] O. Kagawa, K. Katayama, S. Konomi, and Y. Kambayashi. Capturing essential questions using question support facilities in the VIEW classroom. Lecture Notes in Computer Science, 978:114–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeKorvin:1995:FMM

[1032] A. De Korvin, G. Quirchmayr, and S. Hashemi. A fuzzy model of managerial decision making incorporating risk and ambiguity aversion. Lecture Notes in Computer Science, 978:124– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krueger:1995:POS

[1033] T. Krueger, G. Kurian, A. Nair, and G. Neumann. People oriented software technology, and its use in environmental reporting. Lecture Notes in Computer Science, 978:136–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hameurlain:1995:CEP

[1034] A. Hameurlain and F. Morvan. A cost evaluator for parallel database systems. Lecture Notes in Computer Science, 978:146–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1995:PSD

[1035] K.-L. Tan and J. X. Yu. A performance study of declustering strategies for parallel spatial databases. Lecture Notes in Computer Science, 978:157–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kempe:1995:BOO

[1036] J. Kempe, W. Kowarschick, W. Kiessling, and R. Hitzelberger. Benchmarking object-oriented database systems for CAD. Lecture Notes in Computer Science, 978:167–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pernul:1995:CRF

[1037] G. Pernul and H. Hasenauer. Combining reverse with forward database engineering — a step forward to solve the legacy system dilemma. Lecture Notes in Computer Science, 978:177– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darmont:1995:PEC

[1038] J. Darmont, A. Attoui, and M. Gourgand. Performance evaluation for clustering algorithms in object-oriented database systems. Lecture Notes in Computer Science, 978:187–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naqvi:1995:AAD

[1039] W. Naqvi and S. Panyiotou. Applied active databases for evolving image processing algorithms. Lecture Notes in Computer Science, 978:197–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beattie:1995:SRG

[1040] B. Beattie, F. Coenen, T. J. M. Bench-Capon, and B. M. Diaz. Spatial reasoning for GIS using a tesseral data representation. Lecture Notes in Computer Science, 978:207–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1995:NDS

[1041] J.-K. Kim and J.-W. Chang. A new dynamic signature file method in parallel processing environment. Lecture Notes in Computer Science, 978:217– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yahata:1995:GPC

[1042] C. Yahata and M. Takizawa. General protocols for consensus in distributed systems. Lecture Notes in Computer Science, 978:227–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naja:1995:MRA

[1043] H. Naja and N. Mouaddib. The multiple representation in an architectural application. Lecture Notes in Computer Science, 978:237–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breche:1995:SSC

[1044] P. Breche, F. Ferrandina, and M. Kuklok. Simulation of schema change using views. Lecture Notes in Computer Science, 978:247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Francois:1995:SHR

[1045] P. François and P. Bazex. SGML/ HyTime repositories: Requirements and data modelling using objectoriented database concepts. Lecture Notes in Computer Science, 978:259– 271, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macleod:1995:DST

[1046] I. A. Macleod and D. Narine. A depository for structured text objects. Lecture Notes in Computer Science, 978: 272–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Signore:1995:IHD

[1047] O. Signore. Issues on hypertext design. Lecture Notes in Computer Science, 978:283–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1995:TLB

[1048] P. Becker. A temporal logic based approach for querying lists, trees, and DAG’s in databases. Lecture Notes in Computer Science, 978:293–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buccafurri:1995:DOL

[1049] F. Buccafurri, N. Leone, L. Palopoli, and P. Rullo. Disjunctive ordered logic. Lecture Notes in Computer Science, 978:303–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Du:1995:ART

[1050] X. Du and N. Ishii. An algebraic rewriting theorem of multiple linear recursions and its applications. Lecture Notes in Computer Science, 978:313– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhargava:1995:LAS

[1051] B. Bhargava and S. Goel. A layered architecture for supporting objects in a relational system: A performance study. Lecture Notes in Computer Science, 978:323–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellavia:1995:SES

[1052] G. Bellavia, D. Maio, and S. Rizzi. An SQL extension supporting user viewpoints. Lecture Notes in Computer Science, 978:334–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boucher:1995:KKB

[1053] C. Boucher, C. Daul, P. Graebling, and E. Hirsch. KBED: a knowledge-based edge detection system. Lecture Notes in Computer Science, 978:344–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debenham:1995:UES

[1054] J. Debenham. Understanding expert systems maintenance. Lecture Notes in Computer Science, 978:354–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abel:1995:ECB

[1055] M. Abel, E. B. Reategui, J. M. V. De Castilho, and J. Campbell. Evaluating case-based reasoning in a geological domain. Lecture Notes in Computer Science, 978:364–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bardis:1995:ELS

[1056] L. Bardis, G. Grigoropoulos, S. Kokkotos, and T. Loukakis. An expert loading system for chemical and product carriers. Lecture Notes in Computer Science, 978:374–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gazzotti:1995:FFL

[1057] D. Gazzotti, L. Piancastelli, C. Sartori, and D. Beneventano. FuzzyBase: A fuzzy logic aid for relational database queries. Lecture Notes in Computer Science, 978:385–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mechkour:1995:EEM

[1058] M. Mechkour. EMIR02: An extended model for image representation and retrieval. Lecture Notes in Computer Science, 978:395–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleury:1995:REK

[1059] L. Fleury, C. Djeraba, H. Briand, and J. Philippe. Rule evaluations in a KDD system. Lecture Notes in Computer Science, 978:405–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1995:OOK

[1060] A. M. J. Miller, W. A. Gray, and N. J. Fiddian. Object-oriented knowledge based imputation. Lecture Notes in Computer Science, 978:415–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Motro:1995:KOE

[1061] A. Motro and S. Goullioud. Knowledge organization for exploration. Lecture Notes in Computer Science, 978: 426–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cesarini:1995:DEF

[1062] F. Cesarini, M. Gori, S. Marinai, and G. Soda. Data extraction from form images. Lecture Notes in Computer Science, 978:438–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choquette:1995:CLH

[1063] M. Choquette, D. Poulin, and P. Bratley. Compiling legal hypertexts. Lecture Notes in Computer Science, 978: 449–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eaglestone:1995:CTI

[1064] B. Eaglestone. Composition tools integration with a music database system. Lecture Notes in Computer Science, 978:459–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soutou:1995:RDR

[1065] C. Soutou. Relational database reverse engineering: Extraction of an IFO2 schema. Lecture Notes in Computer Science, 978:469–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suekan:1995:ESO

[1066] E. Suekan and Z. M. Oezsoyoglu. Extended set operators for nested relations and complex objects. Lecture Notes in Computer Science, 978:479– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aritsugi:1995:SIP

[1067] M. Aritsugi, K. Teramoto, G. Bai, and A. Makinouchi. Several implementations of persistent pointers in a memory-mapped I/O environment. Lecture Notes in Computer Science, 978:490–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choenni:1995:TTP

[1068] S. Choenni, H. Wagterveld, H. M. Blanken, and T. Chang. TOPYDE: A tool for physical database design. Lecture Notes in Computer Science, 978: 502–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sohn:1995:TST

[1069] Y. Sohn and S. Moon. Tightly secure transaction scheduler in multilevel secure database management systems. Lecture Notes in Computer Science, 978:512–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ozkan:1995:HAO

[1070] C. Ozkan, A. Dogac, and C. Evrendilek. A heuristic approach for optimization of path expressions. Lecture Notes in Computer Science, 978:522–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gross:1995:PDD

[1071] T. Gross and R. Traunmueller. Problem dimensions in design of CSCW systems. Lecture Notes in Computer Science, 978:535–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kueng:1995:RDT

[1072] J. Kueng, R. Wagner, and W. Woess. A rule-driven transformation processor for bill of material data. Lecture Notes in Computer Science, 978:545– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:FVC

[1073] S.-H. Lee and Y. Kambayashi. Functions of VIEW conf to support dynamic features of meetings. Lecture Notes in Computer Science, 978:554–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maio:1995:CAP

[1074] D. Maio and S. Rizzi. CICERO: An assistant for planning visits to a museum. Lecture Notes in Computer Science, 978:564–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauzer-Medeiros:1995:WWB

[1075] C. Bauzer-Medeiros, G. Vossen, and M. Weske. WASA: A workflowbased architecture to support scientific database applications. Lecture Notes in Computer Science, 978:574–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winiwarter:1995:MMC

[1076] W. Winiwarter. MIDAS — the morphological component of the IDA system for efficient natural language interface design. Lecture Notes in Computer Science, 978:584–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manolopoulos:1995:PMR

[1077] Y. Manolopoulos and A. Vakali. Partial match retrieval in two-headed disks. Lecture Notes in Computer Science, 978:594–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFerreiraRezende:1995:CCN

[1078] F. De Ferreira Rezende and T. Haerder. Concurrency control in nested transactions with enhanced lock modes for KBMSs. Lecture Notes in Computer Science, 978:604–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Youssef:1995:MM

[1079] M. W. Youssef and N. Revell. A measure for measure. Lecture Notes in Computer Science, 978:614–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teufel:1995:ISC

[1080] S. Teufel, J. H. P. Eloff, K. Bauknecht, and D. Karagiannis. Information security concepts in computer supported cooperative work. Lecture Notes in Computer Science, 978:621–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preucil:1995:ISG

[1081] L. Preucil and P. Stepan. An intelligent self-guided vehicle for CIM systems. Lecture Notes in Computer Science, 978:632–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campr:1995:ISO

[1082] V. Campr, Z. Kouba, and T. Vlcek. Information system for oncopathology: From an expert system to WWW. Lecture Notes in Computer Science, 978:642–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Even:1995:MCC

[1083] S. Even, G. Itkis, and S. Rajsbaum. On mixed connectivity certificates. Lecture Notes in Computer Science, 979:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeumker:1995:TEP

[1084] A. Baeumker, W. Dittrich, and F. Meyer auf der Heide. Truly efficient parallel algorithms: c-optimal multisearch for an extension of the BSP model. Lecture Notes in Computer Science, 979:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhuri:1995:OPS

[1085] S. Chaudhuri and C. D. Zaroliagis. Optimal parallel shortest paths in small treewidth digraphs. Lecture Notes in Computer Science, 979:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czumaj:1995:SMS

[1086] A. Czumaj, F. Meyer auf der Heide, and V. Stemann. Shared memory simulations with triple-logarithmic delay. Lecture Notes in Computer Science, 979:46–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herley:1995:ISM

[1087] K. T. Herley, A. Pietracaprina, and G. Pucci. Implementing shared memory on multi-dimensional meshes and on the fat-tree. Lecture Notes in Computer Science, 979:60–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufmann:1995:BWC

[1088] M. Kaufmann, J. F. Sibeyn, and T. Suel. Beyond the worst-case bisection bound: Fast sorting and ranking on meshes. Lecture Notes in Computer Science, 979:75–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chlebus:1995:FDS

[1089] B. S. Chlebus, L. Gasieniec, and A. Pelc. Fast deterministic simulation of computations on faulty parallel machines. Lecture Notes in Computer Science, 979:89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codenotti:1995:ACD

[1090] B. Codenotti, P. Gemmell, and J. Simon. Average circuit depth and average communication complexity. Lecture Notes in Computer Science, 979: 102–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gil:1995:PT

[1091] J. Gil and A. Itai. Packing trees. Lecture Notes in Computer Science, 979: 113–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deineko:1995:STE

[1092] V. G. Deineko, R. Rudolf, and G. J. Woeginger. Sometimes travelling is easy: The master tour problem. Lecture Notes in Computer Science, 979: 128–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peer:1995:IGS

[1093] I. Pe’er and R. Shamir. Interval graphs with side (and size) constraints. Lecture Notes in Computer Science, 979: 142–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldberg:1995:MSS

[1094] A. V. Goldberg and A. V. Karzanov.

Maximum skew-symmetric flows. Lecture Notes in Computer Science, 979: 155–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1995:CFA

[1095] M. R. Henzinger and H. La Poutre.

Certificates and fast algorithms for biconnectivity in fully-dynamic graphs. Lecture Notes in Computer Science, 979:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehlhorn:1995:APS

[1096] K. Mehlhorn and V. Priebe. On the allpairs shortest path algorithm of moffat and takaoka. Lecture Notes in Computer Science, 979:185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Husfeldt:1995:FDT

[1097] T. Husfeldt. Fully dynamic transitive closure in plane dags with one source and one sink. Lecture Notes in Computer Science, 979:199–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feng:1995:PCG

[1098] Q.-W. Feng, R. F. Cohen, and P. Eades. Planarity for clustered graphs. Lecture Notes in Computer Science, 979:213–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chor:1995:GAB

[1099] B. Chor and M. Sudan. A geometric approach to betweenness. Lecture Notes in Computer Science, 979:227– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadopoulou:1995:ECG

[1100] E. Papadopoulou and D. T. Lee. Efficient computation of the geodesic Voronoi diagram of points in a simple polygon. Lecture Notes in Computer Science, 979:238–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBerg:1995:LSB

[1101] M. De Berg. Linear size binary space partitions for fat objects. Lecture Notes in Computer Science, 979:252– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chew:1995:GPM

[1102] L. P. Chew, D. Dor, A. Efrat, and K. Kedem. Geometric pattern matching in d-dimensional space. Lecture Notes in Computer Science, 979:264– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1995:FCD

[1103] Cao An Wang and F. Chin. Finding the constrained Delaunay triangulation and constrained Voronoi diagram of a simple polygon in lineartime. Lecture Notes in Computer Science, 979:280–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arge:1995:EMA

[1104] L. Arge, D. E. Vengroff, and J. S. Vitter. External-memory algorithms for processing line segments in geographic information systems. Lecture Notes in Computer Science, 979:295–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FernandesBarbosa:1995:OBS

[1105] E. Fernandes Barbosa, G. Navarro, R. Baeza-Yates, and C. Perleberg. Optimized binary search and text retrieval. Lecture Notes in Computer Science, 979:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutinen:1995:UGL

[1106] E. Sutinen and J. Tarhio. On using q-gram locations in approximate string matching. Lecture Notes in Computer Science, 979:327–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MeyeraufderHeide:1995:RBB

[1107] F. Meyer auf der Heide and C. Scheideler. Routing with bounded buffers and hot-potato routing in vertexsymmetric networks. Lecture Notes in Computer Science, 979:341–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Westbrook:1995:LBR

[1108] J. Westbrook. Load balancing for response time. Lecture Notes in Computer Science, 979:355–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthome:1995:SSP

[1109] P. Berthome, T. Duboux, T. Hagerup, and I. Newman. Self-simulation for the passive optical star model. Lecture Notes in Computer Science, 979: 369–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farach:1995:CAT

[1110] M. Farach, T. M. Przytycka, and M. Thorup. Computing the agreement of trees with bounded degrees. Lecture Notes in Computer Science, 979: 381–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1995:AAF

[1111] B. Yu and J. Cheriyan. Approximation algorithms for feasible cut and multicut problems. Lecture Notes in Computer Science, 979:394–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serna:1995:PVS

[1112] M. Serna and F. Xhafa. On parallel versus sequential approximation. Lecture Notes in Computer Science, 979: 409–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1995:EEA

[1113] F. Wagner and A. Wolff. An efficient and effective approximation algorithm for the map labeling problem. Lecture Notes in Computer Science, 979:420–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1995:ABA

[1114] T. Kloks, D. Kratsch, and H. Mueller.

Approximating the bandwidth for asteroidal triple-free graphs. Lecture Notes in Computer Science, 979:434– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubhashi:1995:NOD

[1115] D. Dubhashi and A. Panconesi. Nearoptimal distributed edge coloring. Lecture Notes in Computer Science, 979: 448–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bern:1995:CPA

[1116] M. Bern, D. Eppstein, L. Guibas, and J. Hershberger. The centroid of points with approximate weights. Lecture Notes in Computer Science, 979:460– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1995:IPO

[1117] A. S. Schulz, R. Weismantel, and G. M.

Ziegler. 0/1-Integer programming: Optimization and augmentation are equivalent. Lecture Notes in Computer Science, 979:473–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalyanasundaram:1995:OTP

[1118] B. Kalyanasundaram and K. R. Pruhs. The online transportation problem. Lecture Notes in Computer Science, 979:484–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mutzel:1995:PAP

[1119] P. Mutzel. A polyhedral approach to planar augmentation and related problems. Lecture Notes in Computer Science, 979:494–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerstel:1995:OLC

[1120] O. Gerstel, A. Wool, and S. Zaks. Optimal layouts on a chain ATM network. Lecture Notes in Computer Science, 979:508–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alou:1995:EDR

[1121] N. Alou, Z. Galil, and M. Yung. Efficient dynamic-resharing “verifiable secret sharing” against mobile adversary. Lecture Notes in Computer Science, 979:523–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1995:AVD

[1122] S. Aggarwal, J. A. Garay, and A. Herzberg. Adaptive video on demand. Lecture Notes in Computer Science, 979:538–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poblete:1995:BTA

[1123] P. V. Poblete, J. I. Munro, and T. Papadakis. The binomial transform and its application to the analysis of skip lists. Lecture Notes in Computer Science, 979:554–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damaschke:1995:OPA

[1124] P. Damaschke. An optimal parallel algorithm for digital curve segmentation using Hough polygons and monotone function search. Lecture Notes in Computer Science, 979:570–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:1995:FSC

[1125] R. Klein and A. Lingas. Fast skeleton construction. Lecture Notes in Computer Science, 979:582–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gautier:1995:RVI

[1126] T. Gautier, J. L. Roch, and G. Villard. Regular versus irregular problems and algorithms. Lecture Notes in Computer Science, 980:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botorog:1995:ASA

[1127] G. H. Botorog and H. Kuchen. Algorithmic skeletons for adaptive multigrid methods. Lecture Notes in Computer Science, 980:27–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ujaldon:1995:RTT

[1128] M. Ujaldon, S. D. Sharma, J. Saltz, and E. L. Zapata. Run-time techniques for parallelizing sparse matrix problems. Lecture Notes in Computer Science, 980:43–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auletta:1995:FEI

[1129] V. Auletta, A. Negro, and V. Scarano. Fast execution of irregularly structured programs with low communication frequency on the hypercube. Lecture Notes in Computer Science, 980:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krothapalli:1995:RTP

[1130] V. P. Krothapalli, T. Jeyaraman, and M. Giesbrecht. Run-time parallelization of irregular DOACROSS loops. Lecture Notes in Computer Science, 980:75–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hao:1995:ISG

[1131] B. Hao and D. Pearson. Instruction scheduling and global register allocation for SIMD multiprocessors. Lecture Notes in Computer Science, 980:81–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manoharan:1995:GBA

[1132] S. Manoharan. General bounds for the assignment of irregular dependency graphs. Lecture Notes in Computer Science, 980:87–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sen:1995:NSD

[1133] R. K. Sen. A new scheme for dynamic processor assignment for irregular problems. Lecture Notes in Computer Science, 980:99–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aykanat:1995:EMF

[1134] C. Aykanat and I. Haritaoglu. An efficient mean field annealing formulation for mapping unstructured domains to hypercubes. Lecture Notes in Computer Science, 980:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walshaw:1995:PMU

[1135] C. Walshaw, M. Cross, M. G. Everett, and S. Johnson. Partitioning and mapping of unstructured meshes to parallel machine topologies. Lecture Notes in Computer Science, 980:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fu:1995:ISP

[1136] C. Fu, T. Yang, and A. Gerasoulis. Integrating software pipelining and graph scheduling for iterative scientific computations. Lecture Notes in Computer Science, 980:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1995:AAT

[1137] K. Jansen and S. Oehring. Approximation algorithms for time constrained scheduling. Lecture Notes in Computer Science, 980:143–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boeres:1995:SAE

[1138] C. Boeres, G. Chochia, and P. Thanisch. On the scope of applicability of the ETF algorithm. Lecture Notes in Computer Science, 980:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franca:1995:OMN

[1139] F. M. G. Franca and L. Faria. Optimal mapping of neighbourhood-constrained systems. Lecture Notes in Computer Science, 980:165–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cantor:1995:PPD

[1140] C. R. Cantor, T. Sano, N. E. Broude, and C. L. Smith. Parallel processing in DNA analysis. Lecture Notes in Computer Science, 980:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grant:1995:SCF

[1141] P. W. Grant, M. F. Webster, and X. Zhang. Solving computational fluid dynamics problems on unstructured grids with distributed parallel processing. Lecture Notes in Computer Science, 980:187–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekmann:1995:PDU

[1142] R. Diekmann, D. Meyer, and B. Monien. Parallel decomposition of unstructured FEM-Meshes. Lecture Notes in Computer Science, 980:199–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1995:MPA

[1143] M. A. Garcia. Massively parallel approximation of irregular triangular meshes with G01 parametric surfaces. Lecture Notes in Computer Science, 980:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubeck:1995:DSC

[1144] T. Bubeck, M. Hiller, W. Kuechlin, and W. Rosenstiel. Distributed symbolic computations with DTS. Lecture Notes in Computer Science, 980:231– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1995:PPP

[1145] A. Das, L. E. Moser, and P. M.

Melliar-Smith. A parallel processing paradigm for irregular applications. Lecture Notes in Computer Science, 980:249–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miguet:1995:LBS

[1146] S. Miguet and J.-M. Pierson. Load balancing strategies for a parallel system of particles. Lecture Notes in Computer Science, 980:255–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benaini:1995:RPA

[1147] A. Benaini, D. Laiymani, and G. R.

Perrin. A reconfigurable parallel algorithm for sparse Cholesky factorization. Lecture Notes in Computer Science, 980:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uhl:1995:AWA

[1148] A. Uhl. Adapted wavelet analysis on moderate parallel distributed memory MIMD architectures. Lecture Notes in Computer Science, 980:275–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tschoeke:1995:NPA

[1149] S. Tschoeke and N. Holthoefer. A new parallel approach to the constrained two-dimensional cutting stock problem. Lecture Notes in Computer Science, 980:285–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasselbring:1995:UPL

[1150] W. Hasselbring and R. B. Fisher. Using the ProSet-Linda prototyping language for investigating MIMD algorithms for model matching in 3-D computer vision. Lecture Notes in Computer Science, 980:301–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pardalos:1995:PSC

[1151] P. M. Pardalos, L. Pitsoulis, T. Mavridou, and M. G. C. Resende. Parallel search for combinatorial optimization: Genetic algorithms, simulated annealing, tabu search and GRASP. Lecture Notes in Computer Science, 980:317– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanders:1995:BAP

[1152] P. Sanders. Better algorithms for parallel backtracking. Lecture Notes in Computer Science, 980:333–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hopp:1995:PGT

[1153] H. Hopp and P. Sanders. Parallel game tree search on SIMD machines. Lecture Notes in Computer Science, 980:349–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBruin:1995:APB

[1154] A. De Bruin, G. A. P. Kindervater, and H. W. J. M. Trienekens. Asynchronous parallel branch and bound and anomalies. Lecture Notes in Computer Science, 980:363–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanders:1995:FPQ

[1155] P. Sanders. Fast priority queues for parallel branch-and-bound. Lecture Notes in Computer Science, 980:379– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Correa:1995:PFG

[1156] R. Correa. A parallel formulation for general branch-and-bound algorithms. Lecture Notes in Computer Science, 980:395–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaelbling:1995:POM

[1157] L. P. Kaelbling, M. L. Littman, and A. R. Cassandra. Partially observable Markov decision processes for artificial intelligence. Lecture Notes in Computer Science, 981:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menzel:1995:RPN

[1158] W. Menzel. Robust processing of natural language. Lecture Notes in Computer Science, 981:19–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bricken:1995:DN

[1159] W. Bricken. Distinction networks. Lecture Notes in Computer Science, 981: 35–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bajcsy:1995:PSS

[1160] R. Bajcsy and J. Kosecka. The problem of signal and symbol integration: A study of cooperative mobile autonomous agent behaviors. Lecture Notes in Computer Science, 981:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroedl:1995:EEB

[1161] S. Schroedl. An extension of explanation-based generalization to negation as failure. Lecture Notes in Computer Science, 981:65–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Englert:1995:IIC

[1162] R. Englert. Inducing integrity constraints from knowledge bases. Lecture Notes in Computer Science, 981:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoetker:1995:DSL

[1163] W. Hoetker. Dynamic structuring of lexical knowledge in a reusability scenario. Lecture Notes in Computer Science, 981:89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eckerle:1995:EML

[1164] J. Eckerle and S. Schuierer. Efficient memory-limited graph search. Lecture Notes in Computer Science, 981:101– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnattinger:1995:QBT

[1165] K. Schnattinger, U. Hahn, and M. Klenner. Quality-based terminological reasoning for concept learning. Lecture Notes in Computer Science, 981: 113–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergmann:1995:PDL

[1166] F. W. Bergmann and J. J. Quantz.

Parallelizing description logics. Lecture Notes in Computer Science, 981:137– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giesl:1995:ATP

[1167] J. Giesl. Automated termination proofs with measure functions. Lecture Notes in Computer Science, 981: 149–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thielscher:1995:WSP

[1168] M. Thielscher. What is A skeptical proof? Lecture Notes in Computer Science, 981:161–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weydert:1995:DEP

[1169] E. Weydert. Default entailment — A preferential construction semantics for defeasible inference. Lecture Notes in Computer Science, 981:173–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanLinder:1995:AMY

[1170] B. Van Linder, W. Van der Hoek, and J.-J. C. Meyer. Actions that make you change your mind. Lecture Notes in Computer Science, 981:185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radzikowska:1995:RAA

[1171] A. Radzikowska. Reasoning about action with typical and atypical effects. Lecture Notes in Computer Science, 981:197–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lukaszewicz:1995:RAA

[1172] W. Lukaszewicz and E. Madalinska-Bugaj. Reasoning about action and change: Actions with abnormal effects. Lecture Notes in Computer Science, 981:209–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouzid:1995:TLB

[1173] M. Bouzid and A. Ligeza. Temporal logic based on characteristic functions. Lecture Notes in Computer Science, 981:221–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nebel:1995:CPQ

[1174] B. Nebel. Computational properties of qualitative spatial reasoning: First results. Lecture Notes in Computer Science, 981:233–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gapp:1995:EVM

[1175] K.-P. Gapp. An empirically validated model for computing spatial relations. Lecture Notes in Computer Science, 981:245–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herzog:1995:IVL

[1176] G. Herzog. Integrating vision and language: Towards automatic description of human movements. Lecture Notes in Computer Science, 981:257–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMoor:1995:GPS

[1177] O. De Moor. A generic program for sequential decision processes. Lecture Notes in Computer Science, 982:1–23, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marriott:1995:PVL

[1178] K. Marriott. Parsing visual languages with constraint multiset grammars. Lecture Notes in Computer Science, 982:24–25, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1995:PCC

[1179] M. P. Jones. Programming with constructor classes. Lecture Notes in Computer Science, 982:26–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Douence:1995:TTF

[1180] R. Douence and P. Fradet. Towards a taxonomy of functional language implementations. Lecture Notes in Computer Science, 982:27–44, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamareddine:1995:LCB

[1181] F. Kamareddine and A. Rios. A lambda-calculus à la De Bruijn with explicit substitutions. Lecture Notes in Computer Science, 982:45–62, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fassbender:1995:VIN

[1182] H. Fassbender. A verified implementation of narrowing. Lecture Notes in Computer Science, 982:63–80, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buetow:1995:NCS

[1183] B. Buetow, R. Giegerich, E. Ohlebusch, and S. Thesing. A new calculus for semantic matching. Lecture Notes in Computer Science, 982:81–96, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakahara:1995:CNC

[1184] K. Nakahara, A. Middeldorp, and T. Ida. A complete narrowing calculus for higher-order functional logic programming. Lecture Notes in Computer Science, 982:97–114, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freire:1995:EPT

[1185] J. Freire, R. Hu, T. Swift, and D. S.

Warren. Exploiting parallelism in tabled evaluations. Lecture Notes in Computer Science, 982:115–132, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawson:1995:DIJ

[1186] S. Dawson, C. R. Ramakrishnan, and I. V. Ramakrishnan. Design and implementation of jump tables for fast indexing of logic programs. Lecture Notes in Computer Science, 982:133–150, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehl:1995:AMO

[1187] M. Mehl, R. Scheidhauer, and C. Schulte. An abstract machine for oz. Lecture Notes in Computer Science, 982:151–168, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mycroft:1995:UPC

[1188] A. Mycroft and K. L. Solberg. Uniform PERs and comportment analysis. Lecture Notes in Computer Science, 982:169–188, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barendsen:1995:UTI

[1189] E. Barendsen and S. Smetsers. Uniqueness type inference. Lecture Notes in Computer Science, 982:189–206, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massey:1995:MCM

[1190] B. C. Massey and E. Tick. Modes of comprehension: Mode analysis of arrays and array comprehensions. Lecture Notes in Computer Science, 982:207– 222, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chin:1995:BCD

[1191] W.-N. Chin and S.-C. Khoo. Better consumers for deforestation. Lecture Notes in Computer Science, 982:223– 240, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohnen:1995:ECT

[1192] M. Mohnen. Efficient compile: Time garbage collection for arbitrary data structures. Lecture Notes in Computer Science, 982:241–258, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glueck:1995:EML

[1193] R. Glueck and J. Joergensen. Efficient multi-level generating extensions. Lecture Notes in Computer Science, 982:259–278, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hatcliff:1995:MVC

[1194] J. Hatcliff. Mechanically verifying the correctness of an offline partial evaluator. Lecture Notes in Computer Science, 982:279–298, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henglein:1995:SMB

[1195] F. Henglein and D. Sands. A semantic model of binding times for safe partial evaluation. Lecture Notes in Computer Science, 982:299–320, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noble:1995:GLF

[1196] R. Noble and C. Runciman. Gadgets:

Lazy functional components for graphical user interfaces. Lecture Notes in Computer Science, 982:321–340, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vullinghs:1995:LGF

[1197] T. Vullinghs, D. Tuijnman, and W. Schulte. Lightweight GUIs for functional programming. Lecture Notes in Computer Science, 982:341–356, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McEvoy:1995:LLL

[1198] H. McEvoy and P. H. Hartel. Local linear logic for locality consciousness in multiset transformation. Lecture Notes in Computer Science, 982:357–379, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rety:1995:NFS

[1199] J.-H. Rety. The notion of floundering for SLDNF-Resolution revisited. Lecture Notes in Computer Science, 982:380–396, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arenas-Sanchez:1995:MSU

[1200] P. Arenas-Sanchez and A. Dovier. Minimal set unification. Lecture Notes in Computer Science, 982:397–414, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimajiri:1995:GDQ

[1201] Y. Shimajiri, H. Seki, and H. Itoh. Goal-directed query processing in disjunctive logic databases. Lecture Notes in Computer Science, 982:415–430, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colussi:1995:DSC

[1202] L. Colussi, E. Marchiori, and M. Marchiori. A dataflow semantics for constraint logic programs. Lecture Notes in Computer Science, 982:431–450, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Achten:1995:FFD

[1203] P. Achten. A functional framework for deterministically interleaved interactive programs. Lecture Notes in Computer Science, 982:451–452, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arenas-Sanchez:1995:DML

[1204] P. Arenas-Sanchez and A. Gil-Luezas. A debugging model for lazy narrowing. Lecture Notes in Computer Science, 982:453–454, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burgues:1995:EEM

[1205] X. Burgues and X. Franch. Evaluation of expressions in a multiparadigm framework. Lecture Notes in Computer Science, 982:455–456, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hans:1995:EPI

[1206] W. Hans, F. Saenz, and S. Winkler. An expression-or-parallel implementation for a functional logic language. Lecture Notes in Computer Science, 982:457–458, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meriste:1995:AME

[1207] M. Meriste and J. Penjam. Attributed models of executable specifications. Lecture Notes in Computer Science, 982:459–460, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saglam:1995:ACL

[1208] H. Saglam and J. P. Gallagher. Approximating constraint logic programs using polymorphic types and regular descriptions. Lecture Notes in Computer Science, 982:461–462, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tong:1995:PDPa

[1209] B.-M. Tong and H.-F. Leung. Performance of a data-parallel concurrent constraint programming system. Lecture Notes in Computer Science, 982: 463–464, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wunderwald:1995:PIM

[1210] J. E. Wunderwald. A portable implementation of memoing evaluation. Lecture Notes in Computer Science, 982:465–466, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1995:NSB

[1211] D. A. Schmidt. Natural-semantics-based abstract interpretation. Lecture Notes in Computer Science, 983:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tofte:1995:RIH

[1212] M. Tofte. Region inference for higherorder functional languages. Lecture Notes in Computer Science, 983:19–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolper:1995:ATA

[1213] P. Wolper and B. Boigelot. An automata-theoretic approach to Presburger arithmetic constraints. Lecture Notes in Computer Science, 983:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:1995:GEI

[1214] M. Alt and F. Martin. Generation of efficient interprocedural analyzers with PAG. Lecture Notes in Computer Science, 983:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1995:OAM

[1215] R. Cleaveland, P. Iyer, and D. Yankelevich. Optimality in abstractions of model checking. Lecture Notes in Computer Science, 983:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colby:1995:DSP

[1216] C. Colby. Determining storage properties of sequential and concurrent programs with assignment and structured data. Lecture Notes in Computer Science, 983:64–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coorg:1995:PNS

[1217] S. R. Coorg. Partitioning non-strict functional languages for multi-threaded code generation. Lecture Notes in Computer Science, 983:82–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cortesi:1995:CAI

[1218] A. Cortesi, G. File, R. Giacobazzi, and C. Palamidessi. Complementation in abstract interpretation. Lecture Notes in Computer Science, 983: 100–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dussart:1995:PRS

[1219] D. Dussart, F. Henglein, and C. Mossin.

Polymorphic recursion and subtype qualifications: Polymorphic bindingtime analysis in polynomial time. Lecture Notes in Computer Science, 983: 118–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faxen:1995:OLF

[1220] K.-F. Faxen. Optimizing lazy functional programs using flow inference. Lecture Notes in Computer Science, 983:136–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giesl:1995:TAF

[1221] J. Giesl. Termination analysis for functional programs using term orderings. Lecture Notes in Computer Science, 983:154–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hannan:1995:TBA

[1222] J. Hannan. A type-based analysis for stack allocation in functional languages. Lecture Notes in Computer Science, 983:172–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heintze:1995:CFA

[1223] N. Heintze. Control-flow analysis and type systems. Lecture Notes in Computer Science, 983:189–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jagannathan:1995:EFA

[1224] S. Jagannathan and A. Wright. Effective flow analysis for avoiding run-time checks. Lecture Notes in Computer Science, 983:207–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1995:SAC

[1225] N. Kobayashi, M. Nakade, and A. Yonezawa. Static analysis of communication for asynchronous concurrent programming languages. Lecture Notes in Computer Science, 983:225– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

KrishnaRao:1995:UFE

[1226] M. R. K. Krishna Rao and R. K. Shyamasundar. Unification-free execution of well-moded and well-typed Prolog programs. Lecture Notes in Computer Science, 983:243–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohnen:1995:ECU

[1227] M. Mohnen. Efficient closure utilisation by higher-order inheritance analysis. Lecture Notes in Computer Science, 983:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monsuez:1995:SFA

[1228] B. Monsuez. System F and abstract interpretation. Lecture Notes in Computer Science, 983:279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1995:CNC

[1229] K. Nielsen and M. H. Soerensen. Callby-name CPS-Translation as a bindingtime improvement. Lecture Notes in Computer Science, 983:296–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palsberg:1995:TLC

[1230] Jens Palsberg and Peter Ørbæk. Trust in the lambda calculus. Lecture Notes in Computer Science, 983:314–330, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramachandran:1995:LR

[1231] V. Ramachandran and P. Van Hentenryck. LSign reordered. Lecture Notes in Computer Science, 983:330–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt-Schauss:1995:SAA

[1232] M. Schmidt-Schauss, S. E. Panitz, and M. Schuetz. Strictness analysis by abstract reduction using a tableau calculus. Lecture Notes in Computer Science, 983:348–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serrano:1995:BPO

[1233] M. Serrano and P. Weis. Bigloo: a portable and optimizing compiler for strict functional languages. Lecture Notes in Computer Science, 983:366– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seward:1995:BPI

[1234] J. Seward. Beyond prototype implementations: Polymorphic projection analysis for Glasgow Haskell. Lecture Notes in Computer Science, 983:382– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vedrine:1995:BTA

[1235] F. Vedrine. Binding-time analysis and strictness analysis by abstract interpretation. Lecture Notes in Computer Science, 983:400–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fecht:1995:GTG

[1236] C. Fecht. GENA — a tool for generating Prolog analyzers from specifications. Lecture Notes in Computer Science, 983:418–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nordin:1995:ITD

[1237] M. Nordin, T. Lindgren, and H. Millroth. IGOR: A tool for developing Prolog dataflow analyzers. Lecture Notes in Computer Science, 983:420– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armengol:1995:IIC

[1238] E. Armengol and E. Plaza. Integrating induction in a case-based reasoner. Lecture Notes in Computer Science, 984:3–17, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auriol:1995:IIC

[1239] E. Auriol, M. Manago, K.-D. Althoff, and S. Wess. Integrating induction and case-based reasoning: Methodological approach and first evaluations. Lecture Notes in Computer Science, 984:18–32, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bento:1995:ESE

[1240] C. Bento, J. Exposto, V. Francisco, and E. Costa. Experimental study of an evaluation function for cases imperfectly explained. Lecture Notes in Computer Science, 984:33–44, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bento:1995:RCI

[1241] C. Bento, L. Macedo, and E. Costa. Reasoning with cases imperfectly described and explained. Lecture Notes in Computer Science, 984:45–59, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cox:1995:ILG

[1242] M. T. Cox and A. Ram. Interacting learning-goals: Treating learning as a planning task. Lecture Notes in Computer Science, 984:60–74, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cunningham:1995:LMB

[1243] P. Cunningham, B. Smyth, and T. Veale. On the limitations of memory based reasoning. Lecture Notes in Computer Science, 984:75–86, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lieber:1995:CCB

[1244] J. Lieber. A criterion of comparison between two case bases. Lecture Notes in Computer Science, 984:87–100, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okamoto:1995:APA

[1245] S. Okamoto and K. Satoh. An average predictive accuracy of the nearest neighbor classifier. Lecture Notes in Computer Science, 984:101–112, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reiser:1995:CBR

[1246] C. Reiser and H. Kaindl. Casebased reasoning for multi-step problems and its integration with heuristic search. Lecture Notes in Computer Science, 984:113–125, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richards:1995:QMB

[1247] B. L. Richards. Qualitative models as a basis for case indices. Lecture Notes in Computer Science, 984:126–135, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaaf:1995:UGR

[1248] J. W. Schaaf, M. Nowak, and A. Voss.

Using gestalten to retrieve cases. Lecture Notes in Computer Science, 984: 136–150, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smyth:1995:CIC

[1249] B. Smyth and P. Cunningham. A comparison of incremental case-based reasoning and inductive learning. Lecture Notes in Computer Science, 984:151– 164, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1995:ERC

[1250] G. Weber. Examples and remindings in a case-based help system. Lecture Notes in Computer Science, 984:165– 180, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ashley:1995:CKR

[1251] K. D. Ashley and B. M. McLaren. A CBR knowledge representation for practical ethics. Lecture Notes in Computer Science, 984:181–197, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dave:1995:CBS

[1252] B. Dave, G. Schmitt, S.-G. Shih, and L. Bendel. Case-based spatial design reasoning. Lecture Notes in Computer Science, 984:198–210, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Funk:1995:CBS

[1253] P. J. Funk and D. Robertson. Casebased support for the design of dynamic system requirements. Lecture Notes in Computer Science, 984:211– 222, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kang:1995:MAC

[1254] Byeong Ho Kang and P. Compton. A maintenance approach to case-based reasoning. Lecture Notes in Computer Science, 984:223–239, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerner:1995:CBE

[1255] Y. Kerner. Case-based evaluation in computer chess. Lecture Notes in Computer Science, 984:240–254, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malek:1995:DCB

[1256] M. Malek and V. Rialle. Design of a case-based reasoning system applied to neuropathy diagnosis. Lecture Notes in Computer Science, 984:255–265, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munoz:1995:CNH

[1257] H. Muñoz, J. Paulokat, and S. Wess.

Controlling a nonlinear hierarchical planner using case replay. Lecture Notes in Computer Science, 984:266– 279, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reategui:1995:CSC

[1258] E. B. Reategui and J. A. Campbell. A classification system for credit card transactions. Lecture Notes in Computer Science, 984:280–291, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:UCB

[1259] S.-A. Yang, D. Robertson, and J. Lee.

Use of case-based reasoning in the domain of building regulations. Lecture Notes in Computer Science, 984:292– 306, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dittrich:1995:ADM

[1260] K. R. Dittrich, S. Gatziu, and A. Geppert. The active database management system manifesto: A rulebase of ADBMS features. Lecture Notes in Computer Science, 985:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paton:1995:FSA

[1261] N. W. Paton, J. Campin, A. A. A. Fernandes, and M. H. Williams. Formal specification of active database functionality: A survey. Lecture Notes in Computer Science, 985:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coupaye:1995:DSA

[1262] T. Coupaye and C. Collet. Denotational semantics for an active rule execution model. Lecture Notes in Computer Science, 985:36–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tawbi:1995:ASU

[1263] C. Tawbi, G. Jaber, and M. Dalmau. Activity specification using rendezvous. Lecture Notes in Computer Science, 985:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1995:RBS

[1264] R. Fernandez and O. Diaz. Reactive behaviour support: Themes and variations. Lecture Notes in Computer Science, 985:69–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Etzion:1995:RAB

[1265] O. Etzion. Reasoning about the behavior of active databases applications. Lecture Notes in Computer Science, 985:86–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reddi:1995:EFD

[1266] S. Reddi, A. Poulovassilis, and C. Small. Extending a functional DBPL with ECA-Rules. Lecture Notes in Computer Science, 985:101– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skoeld:1995:RCA

[1267] M. Skoeld, E. Falkenroth, and T. Risch.

Rule contexts in active databases — A mechanism for dynamic rule grouping. Lecture Notes in Computer Science, 985:119–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sturm:1995:TLR

[1268] R. Sturm, J. A. Muelle, and P. C.

Lockemann. Temporized and localized rule sets. Lecture Notes in Computer Science, 985:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1995:DIA

[1269] I. S. Thomas and A. C. Jones. Design and implementation of an active object-oriented database supporting construction of database tools. Lecture Notes in Computer Science, 985: 147–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baralis:1995:IRA

[1270] E. Baralis, S. Ceri, and S. Paraboschi.

Improved rule analysis by means of triggering and activation graphs. Lecture Notes in Computer Science, 985: 165–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benazet:1995:VVT

[1271] E. Benazet, H. Guehl, and M. Bouzeghoub. VITAL: A visual tool for analysis of rules behaviour in active databases. Lecture Notes in Computer Science, 985:182–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chakravarthy:1995:VET

[1272] S. Chakravarthy, Z. Tamizuddin, and J. Zhou. A visualization and explanation tool for debugging ECA rules in active databases. Lecture Notes in Computer Science, 985:197–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zurek:1995:SPL

[1273] T. Zurek and P. Thanisch. Strategies for parallel linear recursive query processing. Lecture Notes in Computer Science, 985:213–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Embury:1995:PCU

[1274] S. M. Embury and P. M. D. Gray. Planning complex updates to satisfy constraint rules using a constraint logic search engine. Lecture Notes in Computer Science, 985:230–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Consens:1995:CPV

[1275] M. P. Consens, A. O. Mendelzon, D. Vista, and P. T. Wood. Constant propagation versus join reordering in datalog. Lecture Notes in Computer Science, 985:245–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plexousakis:1995:CST

[1276] D. Plexousakis. Compilation and simplification of temporal integrity constraints. Lecture Notes in Computer Science, 985:260–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kersten:1995:ACP

[1277] M. L. Kersten. An active component for a parallel database kernel. Lecture Notes in Computer Science, 985:277– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baralis:1995:UDR

[1278] E. Baralis and J. Widom. Using delta relations to optimize condition evaluation in active databases. Lecture Notes in Computer Science, 985:292– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geppert:1995:DBA

[1279] A. Geppert, S. Gatziu, and K. R.

Dittrich. A designer’s benchmark for active database management systems: 007 meets the BEAST. Lecture Notes in Computer Science, 985:309– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geppert:1995:RCA

[1280] A. Geppert, M. Kradolfer, and D. Tombros. Realization of cooperative agents using an active object-oriented database management system. Lecture Notes in Computer Science, 985:327– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bailey:1995:ADA

[1281] J. Bailey, M. Georgeff, D. B. Kemp, and D. Kinny. Active databases and agent systems — A comparison. Lecture Notes in Computer Science, 985: 342–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zukunft:1995:RAD

[1282] O. Zukunft. Recovering active databases. Lecture Notes in Computer Science, 985:357–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1995:DSA

[1283] P. R. Wilson, M. S. Johnstone, M. Neely, and D. Boles. Dynamic storage allocation: A survey and critical review. Lecture Notes in Computer Science, 986:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wadler:1995:SAR

[1284] P. Wadler. Static analysis refuses to stay still: Prospects of static analysis for dynamic allocation. Lecture Notes in Computer Science, 986:117– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamilton:1995:CTG

[1285] G. W. Hamilton. Compile-time garbage collection for lazy functional languages.

Lecture Notes in Computer Science, 986:119–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roejemo:1995:GGC

[1286] N. Roejemo. Generational garbage collection without temporary space leaks for lazy functional languages. Lecture Notes in Computer Science, 986:145– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsui:1995:CGC

[1287] S. Matsui, Y. Tanaka, A. Maeda, and M. Nakanishi. Complementary garbage collector. Lecture Notes in Computer Science, 986:163–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attardi:1995:PTC

[1288] G. Attardi, T. Flagella, and P. Iglio.

Performance tuning in a customizable collector. Lecture Notes in Computer Science, 986:179–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1995:MFS

[1289] M. Suzuki, H. Koide, and M. Terashima. MOA — A fast sliding compaction scheme for a large storage space. Lecture Notes in Computer Science, 986:197–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plainfosse:1995:SDG

[1290] D. Plainfosse and M. Shapiro. A survey of distributed garbage collection techniques. Lecture Notes in Computer Science, 986:211–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1995:GCO

[1291] M. Fuchs. Garbage collection on an open network. Lecture Notes in Computer Science, 986:251–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piquer:1995:IMS

[1292] J. M. Piquer. Indirect mark and sweep: A distributed GC. Lecture Notes in Computer Science, 986:267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maeda:1995:FGG

[1293] M. Maeda, H. Konaka, Y. Ishikawa, and T. Tomokiyo. On-the-fly global garbage collection based on partly mark-sweep. Lecture Notes in Computer Science, 986:283–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matthews:1995:LDS

[1294] D. C. J. Matthews and T. Le Sergent. LEMMA: A distributed shared memory with global and local garbage collection. Lecture Notes in Computer Science, 986:297–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armstrong:1995:OPR

[1295] J. Armstrong and R. Virding. One pass real-time generational mark-sweep garbage collection. Lecture Notes in Computer Science, 986:313–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magnusson:1995:GCC

[1296] B. Magnusson and R. Henriksson.

Garbage collection for control systems. Lecture Notes in Computer Science, 986:323–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Virding:1995:GCC

[1297] R. Virding. A garbage collector for the concurrent real-time language erlang. Lecture Notes in Computer Science, 986:343–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsen:1995:PHA

[1298] K. Nilsen. Progress in hardwareassisted real-time garbage collection. Lecture Notes in Computer Science, 986:355–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phalke:1995:MHB

[1299] V. Phalke and B. Gopinath. A miss history-based architecture for cache prefetching. Lecture Notes in Computer Science, 986:381–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kjelsoe:1995:MMF

[1300] M. Kjelsoe and S. Jones. Memory management in flash-memory disks with data compression. Lecture Notes in Computer Science, 986:399–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hungar:1995:WIM

[1301] H. Hungar, O. Grumberg, and W. Damm. What if model checking must be truly symbolic. Lecture Notes in Computer Science, 987:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stern:1995:AVS

[1302] U. Stern and D. L. Dill. Automatic verification of the SCI cache coherence protocol. Lecture Notes in Computer Science, 987:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierre:1995:DVS

[1303] L. Pierre. Describing and verifying synchronous circuits with the Boyer– Moore theorem prover. Lecture Notes in Computer Science, 987:35–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curzon:1995:PEM

[1304] P. Curzon. Problems encountered in the machine-assisted proof of hardware. Lecture Notes in Computer Science, 987:56–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eisenbiegler:1995:FEE

[1305] D. Eisenbiegler and R. Kumar. Formally embedding existing high level synthesis algorithms. Lecture Notes in Computer Science, 987:71–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1995:FDC

[1306] L. G. Wang and M. Mendler. Formal design of a class of computers. Lecture Notes in Computer Science, 987:84–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McFarland:1995:SAV

[1307] M. C. McFarland and T. J. Kowalski. Symbolic analysis and verification of CPA descriptions. Lecture Notes in Computer Science, 987:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMelo:1995:FFR

[1308] A. C. V. De Melo and H. Barringer.

A foundation for formal reuse of hardware. Lecture Notes in Computer Science, 987:124–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corella:1995:SEA

[1309] F. Corella, M. Langevin, E. Cerny, and Z. Zhou. State enumeration with abstract descriptions of state machines. Lecture Notes in Computer Science, 987:146–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cabodi:1995:TBR

[1310] G. Cabodi, S. Quer, and P. Camurati. Transforming Boolean relations by symbolic encoding. Lecture Notes in Computer Science, 987:161–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wahba:1995:DED

[1311] A. Wahba and D. Borrione. Design error diagnosis in sequential circuits. Lecture Notes in Computer Science, 987:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maler:1995:TAA

[1312] O. Maler and A. Pnueli. Timing analysis of asynchronous circuits using timed automata. Lecture Notes in Computer Science, 987:189–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stern:1995:IPV

[1313] U. Stern and D. L. Dill. Improved probabilistic verification by hash compaction. Lecture Notes in Computer Science, 987:206–224, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barringer:1995:FSE

[1314] H. Barringer, G. Gough, B. Monahan, and A. Williams. Formal support for the ELLA hardware description language. Lecture Notes in Computer Science, 987:225–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNicola:1995:VHC

[1315] R. De Nicola, A. Fantechi, S. Gnesi, and S. Larosa. Verifying hardware components within JACK. Lecture Notes in Computer Science, 987:246–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tasiran:1995:LCN

[1316] S. Tasiran, R. Hojati, and R. K. Brayton. Language containment of nondeterministic omega-automata. Lecture Notes in Computer Science, 987:261– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolignano:1995:POA

[1317] D. Bolignano. A partial-order approach to the verification of concurrent systems: checking liveness properties. Lecture Notes in Computer Science, 987:278–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deharbe:1995:SVO

[1318] D. Deharbe and D. Borrione. Semantics of a verification-oriented subset of VHDL. Lecture Notes in Computer Science, 987:293–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goossens:1995:RAV

[1319] K. G. W. Goossens. Reasoning about VHDL using operational and observational semantics. Lecture Notes in Computer Science, 987:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Encrenaz:1995:SRS

[1320] E. Encrenaz. A symbolic relation for a subset of VHDL’87 descriptions and its application to symbolic model checking. Lecture Notes in Computer Science, 987:328–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egenhofer:1995:NG

[1321] M. J. Egenhofer and D. M. Mark. Naive geography. Lecture Notes in Computer Science, 988:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faltings:1995:QSR

[1322] B. Faltings. Qualitative spatial reasoning using algebraic topology. Lecture Notes in Computer Science, 988:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gahegan:1995:POQ

[1323] M. Gahegan. Proximity operators for qualitative spatial reasoning. Lecture Notes in Computer Science, 988:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hernandez:1995:QD

[1324] D. Hernandez, E. Clementini, and P. Di Felice. Qualitative distances. Lecture Notes in Computer Science, 988:45–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zimmermann:1995:MMD

[1325] K. Zimmermann. Measuring without measures: The delta-calculus. Lecture Notes in Computer Science, 988:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rennison:1995:MPC

[1326] E. Rennison and L. Strausfeld. The Millennium Project: Constructing a dynamic 3+D virtual environment for exploring geographically, temporally, and categorically organized historical information. Lecture Notes in Computer Science, 988:69–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dieberger:1995:PSN

[1327] A. Dieberger. Providing spatial navigation for the World Wide Web. Lecture Notes in Computer Science, 988:93–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuan:1995:SAG

[1328] M. Yuan and J. Albrecht. Structural analysis of geographic information and GIS operations: from a user’s perspective. Lecture Notes in Computer Science, 988:107–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peloux:1995:LCI

[1329] J.-P. Peloux and P. Rigaux. A loosely coupled interface to an objectorientated geographic database. Lecture Notes in Computer Science, 988: 123–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fritsch:1995:SRL

[1330] E. Fritsch and J. P. Lagrange. Spectral representations of linear features for generalisation. Lecture Notes in Computer Science, 988:157–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ware:1995:TSM

[1331] J. M. Ware, C. B. Jones, and G. L.

Bundy. A triangulated spatial model for cartographic generalisation of areal objects. Lecture Notes in Computer Science, 988:173–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amorim:1995:OOL

[1332] M.-A. Amorim, S. Glasauer, K. Corpinot, and A. Berthoz. Objectorientation and location updating during nonvisual navigation: The characteristics and effects of object- versus trajectory-centered processing modes. Lecture Notes in Computer Science, 988:193–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golledge:1995:PSR

[1333] R. G. Golledge. Path selection and route preference in human navigation: A progress report. Lecture Notes in Computer Science, 988:207–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maass:1995:HSI

[1334] W. Maass. How spatial information connects visual information and natural language generation in dynamic environments: Towards a computational model. Lecture Notes in Computer Science, 988:223–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stefanakis:1995:DOP

[1335] E. Stefanakis and M. Kavouras. On the determination of the optimum path in space. Lecture Notes in Computer Science, 988:241–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertolotto:1995:UFM

[1336] M. Bertolotto, L. De Floriani, and P. Marzano. A unifying framework for multilevel description of spatial data. Lecture Notes in Computer Science, 988:259–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magillo:1995:UVI

[1337] P. Magillo, L. De Floriani, and E. Bruzzone. Updating visibility information on multiresolution terrain models. Lecture Notes in Computer Science, 988:279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruegger:1995:TIS

[1338] B. P. Bruegger. Theory for the integration of scale and representation formats: Major concepts and practical implications. Lecture Notes in Computer Science, 988:297–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohn:1995:HRQ

[1339] A. G. Cohn. A hierarchical representation of qualitative shape based on connection and convexity. Lecture Notes in Computer Science, 988:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirtle:1995:RSC

[1340] S. C. Hirtle. Representational structures for cognitive space: Trees, ordered trees and semi-lattices. Lecture Notes in Computer Science, 988:327– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlieder:1995:RAO

[1341] C. Schlieder. Reasoning about ordering. Lecture Notes in Computer Science, 988:341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bucher:1995:ISA

[1342] F. Bucher and A. Vckovski. Improving the selection of appropriate spatial interpolation methods. Lecture Notes in Computer Science, 988:351–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chou:1995:SPS

[1343] Y.-H. Chou. Spatial patterns and spatial autocorrelation. Lecture Notes in Computer Science, 988:365–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1995:QCM

[1344] E. Allen, G. Edwards, and Y. Bedard. Qualitative causal modeling in temporal GIS. Lecture Notes in Computer Science, 988:397–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Story:1995:DSE

[1345] P. A. Story and M. F. Worboys. A design support environment for spatiotemporal database applications. Lecture Notes in Computer Science, 988: 413–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benenson:1995:IVE

[1346] I. Benenson and J. Portugali. Internal vs. external spatial information and cultural emergence in a self-organising city. Lecture Notes in Computer Science, 988:431–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broschart:1995:SPS

[1347] J. Broschart. The social perception of space: Non-spatial determinants of the use of directionals in tongan (polynesia). Lecture Notes in Computer Science, 988:443–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keating:1995:SCS

[1348] E. Keating. Spatial conceptualizations of social hierarchy in pohnpei, micronesia. Lecture Notes in Computer Science, 988:463–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1995:DLM

[1349] B. Smith. On drawing lines on a map. Lecture Notes in Computer Science, 988:475–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montello:1995:HSC

[1350] D. R. Montello. How significant are cultural differences in spatial cognition? Lecture Notes in Computer Science, 988:485–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glasgow:1995:FMB

[1351] J. Glasgow. A formalism for modelbased spatial planning. Lecture Notes in Computer Science, 988:501–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gapp:1995:OLS

[1352] K.-P. Gapp. Object localization: Selection of optimal reference objects. Lecture Notes in Computer Science, 988: 519–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Theodoridis:1995:RQI

[1353] Y. Theodoridis and D. Papadias. Range queries involving spatial relations: A performance analysis. Lecture Notes in Computer Science, 988: 537–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mark:1995:ERC

[1354] D. M. Mark, D. Comas, M. J. Egenhofer, and S. M. Freundschuh. Evaluating and refining computational models of spatial relations through crosslinguistic human-subjects testing. Lecture Notes in Computer Science, 988: 553–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwaertzel:1995:DPF

[1355] H. Schwaertzel. Demands and perspectives for future software engineers: An industrial viewpoint. Lecture Notes in Computer Science, 989:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bancilhon:1995:WWN

[1356] F. Bancilhon. Why we need object databases. Lecture Notes in Computer Science, 989:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Humphrey:1995:PCS

[1357] W. S. Humphrey. A personal commitment to software quality. Lecture Notes in Computer Science, 989:5–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graw:1995:DMD

[1358] G. Graw and V. Gruhn. Distributed modeling and distributed enaction of business processes. Lecture Notes in Computer Science, 989:8–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barghouti:1995:IIM

[1359] N. S. Barghouti, E. Koutsofios, and E. Cohen. Improvise: Interactive multimedia process visualization environment. Lecture Notes in Computer Science, 989:28–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Votta:1995:DPI

[1360] L. G. Votta and M. L. Zajac. Design process improvement case study using process waiver data. Lecture Notes in Computer Science, 989:44–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuya:1995:USM

[1361] J. Tuya, L. Sanchez, and J. A. Corrales. Using a symbolic model checker for verify safety properties in SA/RT models. Lecture Notes in Computer Science, 989:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ValderrutenVidal:1995:ISR

[1362] A. Valderruten Vidal, M. Vilares Ferro, and J. Gratha Gil. Instrumentation of synchronous reactive models for performance engineering. Lecture Notes in Computer Science, 989:76–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coen-Porisini:1995:FFA

[1363] A. Coen-Porisini, R. A. Kemmerer, and D. Mandrioli. A formal framework for ASTRAL inter-level proof obligations. Lecture Notes in Computer Science, 989:90–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:ORM

[1364] Y.-F. Chen, B. Krishnamurthy, and K.-P. Vo. An objective reuse metric: Model and methodology. Lecture Notes in Computer Science, 989:109– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hitz:1995:MPA

[1365] M. Hitz and B. Montazeri. Measuring product attributes of object-oriented systems. Lecture Notes in Computer Science, 989:124–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magee:1995:SDS

[1366] J. Magee, N. Dulay, S. Eisenbach, and J. Kramer. Specifying distributed software architectures. Lecture Notes in Computer Science, 989:137–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tai:1995:SBM

[1367] K.-C. Tai and R. H. Carver. A specification-based methodology for testing concurrent programs. Lecture Notes in Computer Science, 989:154– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Warren:1995:DCA

[1368] I. Warren and I. Sommerville. Dynamic configuration abstraction. Lecture Notes in Computer Science, 989: 173–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeller:1995:HVS

[1369] A. Zeller and G. Snelting. Handling version sets through feature logic. Lecture Notes in Computer Science, 989: 191–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1995:IIA

[1370] T. A. Wagner and S. L. Graham. Integrating incremental analysis with version management. Lecture Notes in Computer Science, 989:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuerr:1995:GGE

[1371] A. Schuerr, A. J. Winter, and A. Zuendorf. Graph grammar engineering with PROGRES. Lecture Notes in Computer Science, 989:219–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaskell:1995:SAF

[1372] C. Gaskell and R. Phillips. A structured analysis formalism with execution semantics to allow unambiguous model interpretation. Lecture Notes in Computer Science, 989:235–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glinz:1995:IFM

[1373] M. Glinz. An integrated formal model of scenarios based on statecharts. Lecture Notes in Computer Science, 989: 254–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heisel:1995:TSF

[1374] M. Heisel, T. Santen, and D. Zimmermann. Tool support for formal software development: A generic architecture. Lecture Notes in Computer Science, 989:272–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Landes:1995:TNF

[1375] D. Landes and R. Studer. The treatment of non-functional requirements in MIKE. Lecture Notes in Computer Science, 989:294–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monk:1995:SDR

[1376] S. Monk, I. Sommerville, J. M. Pendaries, and B. Durin. Supporting design rationale for system evolution. Lecture Notes in Computer Science, 989:307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarzabek:1995:PLS

[1377] S. Jarzabek. PQL: A language for specifying abstract program views. Lecture Notes in Computer Science, 989:324– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chamillard:1995:ESP

[1378] A. T. Chamillard. An exploratory study of program metrics as predictors of reachability analysis performance. Lecture Notes in Computer Science, 989:343–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1995:AQL

[1379] A. W. Brown, D. J. Carney, P. C.

Clements, and B. C. Meyers. Assessing the quality of large, softwareintensive systems. Lecture Notes in Computer Science, 989:384–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsia:1995:IID

[1380] P. Hsia, C.-T. Hsu, D. C. Kung, and A. T. Yaung. The impact of incremental delivery on maintenance effort: An analytical study. Lecture Notes in Computer Science, 989:405–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castell:1995:MQF

[1381] N. Castell and O. Slavkova. Metrics for quality factors in the LESD project. Lecture Notes in Computer Science, 989:423–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsumoto:1995:QES

[1382] M. J. Matsumoto. Quantitative evaluations of software quality built in with domain-specific disciplines. Lecture Notes in Computer Science, 989: 438–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jazayeri:1995:CPF

[1383] M. Jazayeri. Component programming — a fresh look at software components. Lecture Notes in Computer Science, 989:457–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koskimies:1995:DFS

[1384] K. Koskimies and H. Moessenboeck. Designing a framework by stepwise generalization. Lecture Notes in Computer Science, 989:479–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gall:1995:OOR

[1385] H. Gall, R. Kloesch, and R. Mittermeir. Object-oriented re-architecturing. Lecture Notes in Computer Science, 989: 498–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnattinger:1995:TMR

[1386] K. Schnattinger, U. Hahn, and M. Klenner. Terminological metareasoning by reification and multiple contexts. Lecture Notes in Computer Science, 990:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poli:1995:NCP

[1387] R. Poli, M. Ryan, and A. Sloman.

A new continuous propositional logic. Lecture Notes in Computer Science, 990:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1995:SPS

[1388] U. Egly. Super-polynomial speedups in proof length by new tautologies. Lecture Notes in Computer Science, 990:29–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferes:1995:BRN

[1389] J. Alferes, L. M. Pereira, T. Przymusinski, and U. Evora. Belief revision in non-monotonic reasoning. Lecture Notes in Computer Science, 990:41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dung:1995:NRJ

[1390] T. Q. Dung. A new representation of JTMS. Lecture Notes in Computer Science, 990:57–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitoria:1995:RPC

[1391] A. Vitoria, M. Mamede, and L. Monteiro. The retrieval problem in a concept language with number restrictions. Lecture Notes in Computer Science, 990:69–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trombettoni:1995:FLP

[1392] G. Trombettoni. Formalizing local propagation in constraint maintenance systems. Lecture Notes in Computer Science, 990:83–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:DPK

[1393] J.-H. Lee and G. Lee. A dependency parser of Korean based on connectionist/symbolic techniques. Lecture Notes in Computer Science, 990:95–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berlandier:1995:SAA

[1394] P. Berlandier. A symbiotic approach to arc and path consistency checking. Lecture Notes in Computer Science, 990:107–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaspar:1995:WDI

[1395] G. Gaspar and H. Coelho. Where do intentions come from?: A framework for goals and intentions adoption, derivation and evolution. Lecture Notes in Computer Science, 990:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coelho:1995:CLA

[1396] H. Coelho, A. Eusebio, and E. Costa. A closer look to artificial learning environments. Lecture Notes in Computer Science, 990:129–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coelho:1995:BMA

[1397] H. Coelho, L. Antunes, and L. Moniz.

Building multi-agent societies from descriptions to systems: Inter-layer translations. Lecture Notes in Computer Science, 990:143–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ZubizarretaAizpuru:1995:GTH

[1398] J. R. Zubizarreta Aizpuru and J. A.

Usunariz. GA/TS: A hybrid approach for job shop scheduling in a production system. Lecture Notes in Computer Science, 990:153–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teller:1995:CEE

[1399] A. Teller and M. Veloso. A controlled experiment: Evolution for learning difficult image classification. Lecture Notes in Computer Science, 990:165– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McConnell:1995:MMC

[1400] C. McConnell. Minimal model complexity search. Lecture Notes in Computer Science, 990:177–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gama:1995:CCA

[1401] J. Gama and P. Brazdil. Characterization of classification algorithms. Lecture Notes in Computer Science, 990: 189–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallace:1995:NGB

[1402] J. G. Wallace and K. Bluff. Neurons, glia and the borderline between subsymbolic and symbolic processing. Lecture Notes in Computer Science, 990: 201–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verheij:1995:ADA

[1403] B. Verheij. Arguments and defeat in argument-based nonmonotonic reasoning. Lecture Notes in Computer Science, 990:213–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nardi:1995:PSG

[1404] D. Nardi and R. Rosati. A preference semantics for ground nonmonotonic modal logics. Lecture Notes in Computer Science, 990:225–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duc:1995:LOV

[1405] Ho Ngoc Duc. Logical omniscience vs. logical ignorance on a dilemma of epistemic logic. Lecture Notes in Computer Science, 990:237–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bento:1995:RSM

[1406] C. Bento, P. Machado, and E. Costa. On the role of splitting and merging past cases for generation of a new solution. Lecture Notes in Computer Science, 990:249–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melis:1995:TPA

[1407] E. Melis. Theorem proving by analogy — A compelling example. Lecture Notes in Computer Science, 990:261– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferes:1995:NAA

[1408] J. J. Alferes, R. Li, and L. M. Pereira.

Non-atomic actions in the situation calculus. Lecture Notes in Computer Science, 990:273–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karacapilidis:1995:PUU

[1409] N. I. Karacapilidis. Planning under uncertainty: A qualitative approach. Lecture Notes in Computer Science, 990:285–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacholczyk:1995:QRU

[1410] D. Pacholczyk. Qualitative reasoning under uncertainty. Lecture Notes in Computer Science, 990:297–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reifman:1995:SCQ

[1411] J. Reifman and T. Y. C. Wei. Systematic construction of qualitative physicsbased rules for process diagnostics. Lecture Notes in Computer Science, 990: 311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quek:1995:IPS

[1412] C. Quek, P. W. Ng, and M. Pasquier. Integrated process supervision: A structured approach to expert control. Lecture Notes in Computer Science, 990:323–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lima:1995:USG

[1413] P. U. Lima and G. N. Saridis. Using stochastic grammars to learn robotic tasks. Lecture Notes in Computer Science, 990:335–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damas:1995:CCG

[1414] L. Damas and N. Moreira. Constraint categorial grammars. Lecture Notes in Computer Science, 990:347–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broda:1995:NTA

[1415] S. Broda and L. Damas. A new translation algorithm from lambda calculus into combinatory logic. Lecture Notes in Computer Science, 990:359– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kipper:1995:IMA

[1416] B. Kipper. Interlocking multi-agent and blackboard architectures. Lecture Notes in Computer Science, 990 (or 991??):371–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damasio:1995:MTP

[1417] C. V. Damasio and L. M. Pereira. A model theory for paraconsistent logic programming. Lecture Notes in Computer Science, 990 (or 991??):377–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Chamizo:1995:PSR

[1418] C. Fernandez-Chamizo, P. A. Gonzalez-Calero, and M. Gomez-Albarran. Promoting software reuse through explicit knowledge representation. Lecture Notes in Computer Science, 990 (or 991??):387–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherruel:1995:ELM

[1419] G. Cherruel, B. Solaiman, and Y. Autret. Efficient learning in multilayered perceptron using the grow-and-learn algorithm. Lecture Notes in Computer Science, 990 (or 991??):397–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAlmeidaMora:1995:MDS

[1420] I. De Almeida Mora and J. J. Alferes.

Modelling diagnosis systems with logic programming. Lecture Notes in Computer Science, 990:409–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Custodio:1995:ALA

[1421] L. M. M. Custodio and C. A. Pinto-Ferreira. Agreement: A logical approach to approximate reasoning. Lecture Notes in Computer Science, 990 (or 991??):419–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolverton:1995:PSI

[1422] M. Wolverton. Presenting significant information in expert system explanation. Lecture Notes in Computer Science, 990 (or 991??):435–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaignaud:1995:CMP

[1423] N. Chaignaud. A cognitive model of problem solving with incomplete information. Lecture Notes in Computer Science, 990 (or 991??):441–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castell:1995:FSS

[1424] N. Castell and A. Hernandez. Filtering software specifications written in natural language. Lecture Notes in Computer Science, 990 (or 991??):447–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMatos:1995:PDS

[1425] P. A. De Matos and J. P. Martins. Parsimonious diagnosis in SNePS. Lecture Notes in Computer Science, 990:457– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhattacharyya:1995:SSF

[1426] S. Bhattacharyya and S. L. Lytinen. Syntactic and semantic filtering in a chart parser. Lecture Notes in Computer Science, 990 (or 991??):465–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stephan:1995:MRPa

[1427] W. Stephan and S. Biundo. Multilevel refinement planning in an intervalbased temporal logic. Lecture Notes in Computer Science, 990:482–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1995:RRA

[1428] G. Fischer. Rethinking and reinventing artificial intelligence from the perspective of human-centered computational artifacts. Lecture Notes in Computer Science, 991:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paris:1995:ADG

[1429] C. L. Paris. Automatic documentation generation: Including examples. Lecture Notes in Computer Science, 991:12–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wellman:1995:MOP

[1430] M. P. Wellman. Market oriented programming. Lecture Notes in Computer Science, 991:26–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alvarado:1995:BIS

[1431] M. Alvarado and G. Nuthez. Belief increasing in SKL model frames. Lecture Notes in Computer Science, 991:28–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joshi:1995:SCS

[1432] A. V. Joshi, S. C. Sahasrabudhe, and K. Shankar. Sensitivity of combination schemes under conflicting conditions and a new method. Lecture Notes in Computer Science, 991:39–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becher:1995:TCL

[1433] V. Becher. Two conditional logics for defeasible inference: A comparison preliminary version. Lecture Notes in Computer Science, 991:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bedrax-Weiss:1995:USA

[1434] T. Bedrax-Weiss and L. E. Bertossi. Underlying semantics for the assessment of Reiter’s solution to the frame problem. Lecture Notes in Computer Science, 991:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DaCostaMora:1995:MIE

[1435] M. Da Costa Mora, J. G. Lopes, and H. Coelho. Modeling intentions with extended logic programming. Lecture Notes in Computer Science, 991:69–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliveira:1995:LTP

[1436] C. M. G. M. Oliveira. Labelled theorem proving for substructural logics. Lecture Notes in Computer Science, 991: 79–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RochaRiccio:1995:LA

[1437] N. C. Rocha Riccio and R. J. G. B. De Queiroz. Labelled abduction. Lecture Notes in Computer Science, 991:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodrigues:1995:PPM

[1438] O. Rodrigues and M. Benevides. PRO-MAL: Programming in modal action logic. Lecture Notes in Computer Science, 991:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaverucha:1995:GDR

[1439] G. Zaverucha and S. R. M. Veloso. A goal directed reasoning for semi-normal default theories. Lecture Notes in Computer Science, 991:112–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aimeur:1995:TNL

[1440] E. Aimeur, C. Frasson, and C. Alexe.

Towards new learning strategies in intelligent tutoring systems. Lecture Notes in Computer Science, 991:121– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBarrosCosta:1995:MLE

[1441] E. De Barros Costa, M. A. Lopes, and E. Ferneda. Mathema: A learning environment based on a multi-agent architecture. Lecture Notes in Computer Science, 991:141–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Direne:1995:GMD

[1442] A. I. Direne. A general model of dialogue interpretation for concept tutoring systems. Lecture Notes in Computer Science, 991:151–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campagnolo:1995:KBC

[1443] M. L. Campagnolo, H. Coelho, and J. H. Capelo. Knowledge based clustering of partially characterized objects. Lecture Notes in Computer Science, 991:161–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicoletti:1995:CER

[1444] M. C. Nicoletti and M. C. Monard. Constructing the extensional representation of an intentional domain theory in inductive logic programming. Lecture Notes in Computer Science, 991: 171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheffer:1995:GAL

[1445] T. Scheffer. A generic algorithm for learning rules with hierarchical exceptions. Lecture Notes in Computer Science, 991:181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCarvalho:1995:NMV

[1446] L. A. V. De Carvalho and V. L. Roitman. A neural model for the visual attention phenomena. Lecture Notes in Computer Science, 991:191–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohno-Machado:1995:LRC

[1447] L. Ohno-Machado and M. A. Musen. Learning rare categories in backpropagation. Lecture Notes in Computer Science, 991:201–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Passos:1995:AAN

[1448] E. Passos and R. Valente. An automatic adaptative neurocomputing algorithm for time series prediction. Lecture Notes in Computer Science, 991: 210–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Correa:1995:CAS

[1449] M. Correa and S. Mendes. A computational approach to situation theory based on logic programming to design cognitive agents. Lecture Notes in Computer Science, 991:219–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeOliveira:1995:MAH

[1450] F. M. De Oliveira. Measuring agreement and harmony in multi-agent societies: A first approach. Lecture Notes in Computer Science, 991:232– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shmeil:1995:DOL

[1451] M. A. H. Shmeil and E. Oliveira.

Detecting the opportunities of learning from the interactions in a society of organizations. Lecture Notes in Computer Science, 991:242–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sichman:1995:ESR

[1452] J. S. Sichman and Y. Demazeau. Exploiting social reasoning to enhance adaptation in open multi-agent systems. Lecture Notes in Computer Science, 991:253–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferneda:1995:SAD

[1453] E. Ferneda, M. E. De Souza e. Silva, and H. De Menezes Silva. A system for aiding discovery: Mechanisms for knowledge generation. Lecture Notes in Computer Science, 991:262–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fonkam:1995:CFA

[1454] M. M. Fonkam. On a composite formalism and approach to presenting the knowledge content of a relational database. Lecture Notes in Computer Science, 991:272–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Labidi:1995:MKA

[1455] S. Labidi and M. M. Gammoudi. Method for knowledge acquisition from multiple experts. Lecture Notes in Computer Science, 991:283–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Passos:1995:CMK

[1456] E. Passos, A. Sade, C. Garcez, and A. Tanaka. A conceptual model for a knowledge base homogeneously stored in a database environment. Lecture Notes in Computer Science, 991:293– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McFetridge:1995:HDP

[1457] P. McFetridge and A. Villavicencio.

A hierarchical description of the Portuguese verb. Lecture Notes in Computer Science, 991:300–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stefanini:1995:TMA

[1458] M.-H. Stefanini and Y. Demazeau.

Talisman: A multi-agent system for natural language processing. Lecture Notes in Computer Science, 991:310– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Villavicencio:1995:PST

[1459] A. Villavicencio, N. M. C. Marques, J. G. P. Lopes, and F. Villavicencio. Part-of-speech tagging for Portuguese texts. Lecture Notes in Computer Science, 991:321–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saba:1995:QCC

[1460] W. S. Saba and J.-P. Corriveau. Quantification and cognitive constraints in natural language understanding. Lecture Notes in Computer Science, 991: 331–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trombettoni:1995:NDC

[1461] G. Trombettoni, B. Neveu, P. Berlandier, and M.-C. Riff. An nondiffident combinatorial optimization algorithm. Lecture Notes in Computer Science, 991:403–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mora:1995:MDS

[1462] I. Mora and J. Alferes. Modelling diagnosis systems with logic programming. Lecture Notes in Computer Science, 991:409–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nadia:1995:CER

[1463] M. Nadia. Constructing extensions by resolving a system of linear equations. Lecture Notes in Computer Science, 991:429–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matos:1995:PDS

[1464] P. A. Matos and J. P. Martins. Parsimonious diagnosis in SNePS. Lecture Notes in Computer Science, 991:457– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krajcar:1995:GAS

[1465] S. Krajcar, D. Skrlec, B. Pribicevic, and S. Blagajac. GA approach to solving multiple vehicle routing problem. Lecture Notes in Computer Science, 991:473–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stephan:1995:MRPb

[1466] W. Stephan and S. Biundo. Multilevel refinement planning in an intervalbased temporal logic. Lecture Notes in Computer Science, 991:483–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenke:1995:NLG

[1467] N. Lenke. Natural language generation as constraint-based configuration. Lecture Notes in Computer Science, 992: 1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Not:1995:IMA

[1468] E. Not and E. Pianta. Issues of multilinguality in the automatic generation of administrative instructional texts. Lecture Notes in Computer Science, 992:13–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonarini:1995:EQL

[1469] A. Bonarini. Extending Q-learning to fuzzy classifier systems. Lecture Notes in Computer Science, 992:25–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CamargoSilva:1995:FCM

[1470] P. Camargo Silva. Fuzzy cognitive maps in multi-agent environments. Lecture Notes in Computer Science, 992: 37–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baroglio:1995:MPL

[1471] C. Baroglio and M. Botta. Multiple predicate learning with RTL. Lecture Notes in Computer Science, 992:44–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Humphrey:1995:LWS

[1472] T. Humphrey, A. Bramanti-Gregor, and H. W. Davis. Learning while solving problems in single agent search: Preliminary results. Lecture Notes in Computer Science, 992:56–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpineto:1995:ACN

[1473] C. Carpineto and G. Romano. Automatic construction of navigable concept networks characterizing text databases. Lecture Notes in Computer Science, 992:67–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gazza:1995:TPD

[1474] M. Gazza and P. Torasso. Temporal prediction: Dealing with change and interactions within a causal framework. Lecture Notes in Computer Science, 992:79–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donini:1995:NFO

[1475] F. M. Donini, D. Nardi, and R. Rosati. Non-first-order features in concept languages. Lecture Notes in Computer Science, 992:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGiacomo:1995:PBF

[1476] G. De Giacomo and M. Lenzerini. PDL-based framework for reasoning about actions. Lecture Notes in Computer Science, 992:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbero:1995:DGN

[1477] C. Barbero and V. Lombardo. Dependency graphs in natural language processing. Lecture Notes in Computer Science, 992:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciravegna:1995:ISL

[1478] F. Ciravegna and N. Cancedda. Integrating shallow and linguistic techniques for information extraction from text. Lecture Notes in Computer Science, 992:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ardissono:1995:RPS

[1479] L. Ardissono, G. Boella, and D. Sestero. Recognizing preliminary sentences in dialogue interpretation. Lecture Notes in Computer Science, 992: 139–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanellos:1995:CNE

[1480] I. Kanellos and V. H. Zaldivar-Carrillo.

Contextuality and non-extensional identity: the inescapable symbiosis in NLP. Lecture Notes in Computer Science, 992:145–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1995:PDL

[1481] G. Antoniou. Priorities in default logic revisited. Lecture Notes in Computer Science, 992:152–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennaceur:1995:BAR

[1482] H. Bennaceur. Boolean approach for representing and solving constraintsatisfaction problems. Lecture Notes in Computer Science, 992:163–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiManzo:1995:CDP

[1483] M. Di Manzo and P. Pecchiari. Composing decision procedures: The approach and a case study. Lecture Notes in Computer Science, 992:175– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orilia:1995:KRE

[1484] F. Orilia. Knowledge representation, exemplification, and the Gupta-Belnap theory of circular definitions. Lecture Notes in Computer Science, 992:187– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rumolo:1995:TLC

[1485] G. Rumolo and E. Compatangelo. TRDL: A language for conceptual modelling in information systems engineering. Lecture Notes in Computer Science, 992:199–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ThesiederDupre:1995:DRA

[1486] D. Thesieder Dupre and M. Rossotto. The different roles of abstraction in abductive reasoning. Lecture Notes in Computer Science, 992:211–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carassa:1995:CMC

[1487] A. Carassa, A. Valpiani, G. Geminiani, and S. Bandini. A cognitive model of causal reasoning about the physical world. Lecture Notes in Computer Science, 992:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragoni:1995:GAC

[1488] A. F. Dragoni, F. Mascaretti, and P. Puliti. A generalized approach to consistency based belief revision. Lecture Notes in Computer Science, 992: 231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelfranchi:1995:FDB

[1489] C. Castelfranchi, D. D’Aloisi, and F. Giacomelli. A framework for dealing with belief-goal dynamics. Lecture Notes in Computer Science, 992:237– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolfi:1995:ENT

[1490] S. Nolfi and D. Parisi. Evolving nontrivial behaviors on real robots: An autonomous robot that picks up objects. Lecture Notes in Computer Science, 992:243–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cesta:1995:FDD

[1491] A. Cesta and A. Oddi. A formal domain description language for a temporal planner. Lecture Notes in Computer Science, 992:255–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1995:MSM

[1492] C. Ferrari, E. Pagello, J. Ota, and T. Arai. A method for solving multiple autonomous robots collisions problem using space and time representation. Lecture Notes in Computer Science, 992:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blanzieri:1995:MSK

[1493] E. Blanzieri and A. Giordana. Mapping symbolic knowledge into locally receptive field networks. Lecture Notes in Computer Science, 992:267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chella:1995:KRR

[1494] A. Chella, M. Frixione, and S. Gaglio. Knowledge representation for robotic vision based on conceptual spaces and attentive mechanisms. Lecture Notes in Computer Science, 992:279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baioletti:1995:WPS

[1495] M. Baioletti, S. Marcugini, and A. Milani. A weakest precondition semantics for conditional planning. Lecture Notes in Computer Science, 992:291– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frixione:1995:CHM

[1496] M. Frixione, M. Piaggio, G. Vercelli, and R. Zaccaria. A cognitive hybrid model for autonomous navigation. Lecture Notes in Computer Science, 992:303–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAngelo:1995:UCM

[1497] A. D’Angelo. Using a chemical metaphor to implement autonomous systems. Lecture Notes in Computer Science, 992:315–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pogliano:1995:MPD

[1498] P. Pogliano and L. Riccardi. Modeling process diagnostic knowledge through causal networks. Lecture Notes in Computer Science, 992:323–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gamper:1995:FRA

[1499] J. Gamper and W. Nejdl. Formalizing reasoning about change: A temporal diagnosis approach. Lecture Notes in Computer Science, 992:335–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guida:1995:PDD

[1500] G. Guida and M. Zanella. Preventive diagnosis: Definition and logical model. Lecture Notes in Computer Science, 992:347–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Man:1995:LPD

[1501] L. W. Man and S. L. Kwong. Learning programs in different paradigms using genetic programming. Lecture Notes in Computer Science, 992:353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Semeraro:1995:RLT

[1502] G. Semeraro, F. Esposito, N. Fanizzi, and D. Malerba. Revision of logical theories. Lecture Notes in Computer Science, 992:365–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zucker:1995:REC

[1503] J.-D. Zucker. Reformulation of examples in concept learning of structural descriptions. Lecture Notes in Computer Science, 992:377–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GuerreiroNeto:1995:IMP

[1504] J. P. Guerreiro Neto and F. J. G. Moura Pires. Integrated model — a proposal to handle noise. Lecture Notes in Computer Science, 992:389– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armano:1995:EBS

[1505] G. Armano. Exceptions-based synthesis of Boolean functions as a core mechanism to perform concept learning. Lecture Notes in Computer Science, 992:395–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanLinder:1995:SBH

[1506] B. Van Linder, W. Van der Hoek, and J.-J. C. Meyer. Seeing is believing — and so are hearing and jumping. Lecture Notes in Computer Science, 992:402–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giunchiglia:1995:ARO

[1507] E. Giunchiglia, F. Giunchiglia, and L. Serafini. Agents as reasoners, observers or believers. Lecture Notes in Computer Science, 992:414–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaspari:1995:MIA

[1508] M. Gaspari. Modelling interactions in agent system. Lecture Notes in Computer Science, 992:426–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Omicini:1995:ACC

[1509] A. Omicini, E. Denti, and A. Natali. Agent coordination and control through logic theories. Lecture Notes in Computer Science, 992:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corne:1995:SCL

[1510] D. Corne and P. Ross. Some combinatorial landscapes on which a genetic algorithm outperforms other stochastic iterative methods. Lecture Notes in Computer Science, 993:1–13, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shapiro:1995:MEA

[1511] J. L. Shapiro and A. Pruegel-Bennett.

Maximum entropy analysis of genetic algorithm operators. Lecture Notes in Computer Science, 993:14–24, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bilchev:1995:ACM

[1512] G. Bilchev and I. C. Parmee. The ant colony metaphor for searching continuous design spaces. Lecture Notes in Computer Science, 993:25–39, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikami:1995:BBF

[1513] S. Mikami, Y. Kakazu, and T. C. Fogarty. Broadcast based fitness sharing GA for conflict resolution among autonomous robots. Lecture Notes in Computer Science, 993:40–47, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1995:APP

[1514] J. Smith and T. C. Fogarty. An adaptive poly-parental recombination strategy. Lecture Notes in Computer Science, 993:48–61, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swann:1995:NSR

[1515] A. J. Swann. Neighbourhood seeding to reduce problem modality. Lecture Notes in Computer Science, 993:62–74, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burke:1995:SRO

[1516] E. Burke, D. Elliman, and R. Weare. Specialised recombinative operators for timetabling problems. Lecture Notes in Computer Science, 993:75–85, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paechter:1995:ULS

[1517] B. Paechter, A. Cumming, and H. Luchian. The use of local search suggestion lists for improving the solution of timetable problems with evolutionary algorithms. Lecture Notes in Computer Science, 993:86–93, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ross:1995:CGA

[1518] P. Ross and D. Corne. Comparing genetic algorithms, simulated annealing, and stochastic hillclimbing on timetabling problems. Lecture Notes in Computer Science, 993:94–102, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carse:1995:ELC

[1519] B. Carse, T. C. Fogarty, and A. Munro.

Evolutionary learning in computational ecologies: an application to adaptive, distributed routing in communication networks. Lecture Notes in Computer Science, 993:103–116, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langdon:1995:SPM

[1520] W. B. Langdon. Scheduling planned maintenance of the National Grid. Lecture Notes in Computer Science, 993: 132–153, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Savic:1995:GOC

[1521] D. A. Savic and G. A. Walters. Genetic operators and constraint handling for pipe network optimization. Lecture Notes in Computer Science, 993:154– 165, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Surry:1995:MOA

[1522] P. D. Surry, N. J. Radcliffe, and I. D. Boyd. A multi-objective approach to constrained optimisation of gas supply networks: the COMOGA method. Lecture Notes in Computer Science, 993:166–180, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1995:TDD

[1523] J. F. Miller, P. Thomson, and P. V. G. Bradbeer. Ternary decision diagram optimisation of Reed–Muller logic functions using a genetic algorithm for variable and simplification rule ordering. Lecture Notes in Computer Science, 993:181–190, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:EAP

[1524] D. Yang and S. Flockton. An evolutionary algorithm for parametric array signal processing. Lecture Notes in Computer Science, 993:191–199, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patel:1995:CTS

[1525] M. J. Patel. Constraints on task and search complexity in GA + NN models of learning and adaptive behaviour. Lecture Notes in Computer Science, 993:200–223, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vavak:1995:LBA

[1526] F. Vavak, T. C. Fogarty, and P. Cheng.

Load balancing application of the genetic algorithm in a nonstationary environment. Lecture Notes in Computer Science, 993:224–233, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:1995:ESC

[1527] G. Robinson and P. McIlroy. Exploring some commercial applications of genetic programming. Lecture Notes in Computer Science, 993:234–262, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundy:1995:ORS

[1528] J. L. Mundy. Object recognition:

The search for representation. Lecture Notes in Computer Science, 994:19–50, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shapiro:1995:ABO

[1529] L. G. Shapiro and M. S. Costa.

Appearance-based 3D object recognition. Lecture Notes in Computer Science, 994:51–64, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gros:1995:UQI

[1530] P. Gros. Using quasi-invariants for automatic model building and object recognition: an overview. Lecture Notes in Computer Science, 994:65–76, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stockman:1995:ORR

[1531] G. Stockman. Object representation for recognition-by-alignment. Lecture Notes in Computer Science, 994:77–88, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forsyth:1995:DRR

[1532] D. Forsyth, A. Zisserman, and J. Malik. Distinctive representations for the recognition of curved surfaces using outlines and markings. Lecture Notes in Computer Science, 994:89– 100, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weiss:1995:EP

[1533] R. S. Weiss. The epipolar parametrization. Lecture Notes in Computer Science, 994:101–108, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huttenlocher:1995:UTD

[1534] D. Huttenlocher. Using two-dimensional models to interact with the threedimensional world. Lecture Notes in Computer Science, 994:109–124, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriegman:1995:RRC

[1535] D. Kriegman and J. Ponce. Representations for recognizing complex curved 3D objects. Lecture Notes in Computer Science, 994:125–138, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:1995:RIR

[1536] D. B. Cooper and Z. Lei. On representation and invariant recognition of complex objects based on patches and parts. Lecture Notes in Computer Science, 994:139–154, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petitjean:1995:AGO

[1537] S. Petitjean. Algebraic geometry and object representation in computer vision. Lecture Notes in Computer Science, 994:155–166, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taubin:1995:DSS

[1538] G. Taubin. Discrete surface signal processing: The polygon as the surface element. Lecture Notes in Computer Science, 994:167–176, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koivunen:1995:SRD

[1539] V. Koivunen and R. Bajcsy. Spline representations in 3-D vision. Lecture Notes in Computer Science, 994:177– 190, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besl:1995:TPR

[1540] P. J. Besl. Triangles as a primary representation. Lecture Notes in Computer Science, 994:191–206, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Binford:1995:BCR

[1541] T. O. Binford. Body-centered representation and perception. Lecture Notes in Computer Science, 994:207–216, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zerroug:1995:CGO

[1542] M. Zerroug and G. Medioni. The challenge of generic object recognition. Lecture Notes in Computer Science, 994:217–232, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Metaxas:1995:PBF

[1543] D. Metaxas. A physics-based framework for segmentation, shape and motion estimation. Lecture Notes in Computer Science, 994:233–248, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pentland:1995:MR

[1544] A. Pentland and S. Sclaroff. Modal representations. Lecture Notes in Computer Science, 994:249–262, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nastar:1995:TRD

[1545] C. Nastar and N. Ayache. Time representation of deformations: Combining vibration modes and Fourier analysis. Lecture Notes in Computer Science, 994:263–276, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boult:1995:PFW

[1546] T. Boult, S. D. Fenster, and T. O’Donnell. Physics in a fantasy world vs. robust statistical estimation. Lecture Notes in Computer Science, 994:277–296, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camps:1995:TRP

[1547] O. I. Camps. Towards a robust physics-based object recognition system. Lecture Notes in Computer Science, 994:297–312, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gross:1995:TNP

[1548] A. Gross and L. Latecki. Toward non-parametric digital shape representation and recovery. Lecture Notes in Computer Science, 994:313–326, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ikeuchi:1995:SRE

[1549] K. Ikeuchi and M. Hebert. Spherical representations: from EGI to SAI. Lecture Notes in Computer Science, 994:327–346, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Terzopoulos:1995:PBR

[1550] D. Terzopoulos. From physics-based representation to functional modeling of highly complex objects. Lecture Notes in Computer Science, 994:347– 359, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahn:1995:CCP

[1551] R. Ahn. Communicating contexts: a pragmatic approach to information exchange. Lecture Notes in Computer Science, 996:1–13, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geuvers:1995:SFP

[1552] H. Geuvers. A short and flexible proof of strong normalization for the calculus of constructions. Lecture Notes in Computer Science, 996:14–38, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimenez:1995:CGD

[1553] E. Gimenez. Codifying guarded definitions with recursive schemes. Lecture Notes in Computer Science, 996:39–59, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goguen:1995:MU

[1554] H. Goguen. The metatheory of UTT. Lecture Notes in Computer Science, 996:60–82, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manoury:1995:UFS

[1555] P. Manoury. A user’s friendly syntax to define recursive functions as typed lambda-terms. Lecture Notes in Computer Science, 996:83–100, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:1995:AIH

[1556] T. Nipkow and K. Slind. I/O automata in Isabelle/HOL. Lecture Notes in Computer Science, 996:101–119, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulson:1995:CFC

[1557] L. C. Paulson. A concrete final coalgebra theorem for ZF set theory. Lecture Notes in Computer Science, 996:120– 139, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pollack:1995:EPC

[1558] R. Pollack. On extensibility of proof checkers. Lecture Notes in Computer Science, 996:140–161, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranta:1995:SCL

[1559] A. Ranta. Syntactic categories in the language of mathematics. Lecture Notes in Computer Science, 996:162– 182, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saibi:1995:FLC

[1560] A. Saibi. Formalization of a lambdacalculus with explicit substitutions in Coq. Lecture Notes in Computer Science, 996:183–202, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakakibara:1995:GIO

[1561] Y. Sakakibara. Grammatical inference:

An old and new paradigm. Lecture Notes in Computer Science, 997:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fahmy:1995:ELR

[1562] A. F. Fahmy and R. Roos. Efficient learning of real time one-counter automata. Lecture Notes in Computer Science, 997:25–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koshiba:1995:LSD

[1563] T. Koshiba, E. Maekinen, and Y. Takada. Learning strongly deterministic even linear languages from positive examples. Lecture Notes in Computer Science, 997:41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakamoto:1995:LLM

[1564] H. Sakamoto. Language learning from membership queries and characteristic examples. Lecture Notes in Computer Science, 997:55–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arimura:1995:LUT

[1565] H. Arimura, H. Ishizaka, and T. Shinohara. Learning unions of tree patterns using queries. Lecture Notes in Computer Science, 997:66–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRaedt:1995:ICL

[1566] L. De Raedt and W. Van Laer. Inductive constraint logic. Lecture Notes in Computer Science, 997:80–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rao:1995:ILL

[1567] M. R. K. K. Rao. Incremental learning of logic programs. Lecture Notes in Computer Science, 997:95–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyashiro:1995:LOF

[1568] A. Miyashiro, E. Takimoto, Y. Sakai, and A. Maruoka. Learning orthogonal F-horn formulas. Lecture Notes in Computer Science, 997:110–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auer:1995:LND

[1569] P. Auer. Learning nested differences in the presence of malicious noise. Lecture Notes in Computer Science, 997: 123–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1995:LSL

[1570] A. Nakamura and S. Miura. Learning sparse linear combinations of basis functions over a finite domain. Lecture Notes in Computer Science, 997: 138–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kececioglu:1995:IDS

[1571] J. Kececioglu, M. Li, and J. Tromp. Inferring a DNA sequence from erroneous copies. Lecture Notes in Computer Science, 997:151–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Case:1995:MIR

[1572] J. Case, S. Jain, and A. Sharma. Machine induction without revolutionary paradigm shifts. Lecture Notes in Computer Science, 997:153–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1995:PLL

[1573] L. Meyer. Probabilistic language learning under monotonicity constraints. Lecture Notes in Computer Science, 997:169–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stephan:1995:NIO

[1574] F. Stephan. Noisy inference and oracles. Lecture Notes in Computer Science, 997:185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalyanasundaram:1995:STM

[1575] B. Kalyanasundaram and M. Velauthapillai. Simulating teams with many conjectures. Lecture Notes in Computer Science, 997:201–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinter:1995:CNT

[1576] C. C. Pinter. Complexity of network training for classes of neural networks. Lecture Notes in Computer Science, 997:215–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gavalda:1995:LOB

[1577] R. Gavalda and D. Guijarro. Learning ordered binary decision diagrams. Lecture Notes in Computer Science, 997: 228–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castro:1995:SPL

[1578] J. Castro and J. L. Balcazar. Simple PAC learning of simple decision lists. Lecture Notes in Computer Science, 997:239–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domingo:1995:CLM

[1579] C. Domingo and J. Shawe-Taylor. The complexity of learning minor closed graph classes. Lecture Notes in Computer Science, 997:249–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kodratoff:1995:TSI

[1580] Y. Kodratoff. Technical and scientific issues of KDD. Lecture Notes in Computer Science, 997:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadohara:1995:ALP

[1581] K. Sadohara and M. Haraguchi. Analogical logic program synthesis algorithm that can refute inappropriate similarities. Lecture Notes in Computer Science, 997:266–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jantke:1995:RSC

[1582] K. P. Jantke. Reflecting and selfconfident inductive inference machines. Lecture Notes in Computer Science, 997:282–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1995:AIC

[1583] S. Kobayashi and T. Yokomori. On approximately identifying concept classes in the limit. Lecture Notes in Computer Science, 997:298–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chester:1995:RRT

[1584] J. P. Chester and B. V. Dentskevich. Requirements for rapid technological deployment and exploitation. Lecture Notes in Computer Science, 998:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pavlou:1995:IIT

[1585] G. Pavlou and D. Griffin. Issues in the integration of IN and TMN. Lecture Notes in Computer Science, 998:12–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Popp:1995:SST

[1586] R. Popp, M. Froehlich, and N. Jefferies. Security services for telecommunications users. Lecture Notes in Computer Science, 998:28–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Endersz:1995:TSE

[1587] G. Endersz, T. Gabrielsson, F. Morast, and A. Bertsch. TMN security: An evolutionary approach. Lecture Notes in Computer Science, 998:40–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maillot:1995:TBA

[1588] D. Maillot, J. Olnes, and P. Spilling. A TTP-based architecture for TMN security and privacy. Lecture Notes in Computer Science, 998:52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karjoth:1995:IAC

[1589] G. Karjoth. Integrated access control management. Lecture Notes in Computer Science, 998:64–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denley:1995:EEN

[1590] I. Denley, B. Hill, and A. Whitefield. An experimental evaluation of a normative user interface design for the configuration of telecommunication services. Lecture Notes in Computer Science, 998:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hewson:1995:PFT

[1591] T. J. Hewson, G. Allison, and A. M.

Clarke. Providing future telecommunication services to naive users. Lecture Notes in Computer Science, 998:92–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marom:1995:UEM

[1592] R. Marom, M. Bolzoni, M. Goldberg, and D. Rotondi. Usability, an effective methodology for designing services in the agricultural sector. Lecture Notes in Computer Science, 998: 104–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Condon:1995:CCB

[1593] C. Condon and S. Keuneke. Counting the costs and benefits of metaphor. Lecture Notes in Computer Science, 998: 119–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelaFuente:1995:ATC

[1594] L. A. De la Fuente. ASTERIX: The TINA-C architecture applied to ATM connection management. Lecture Notes in Computer Science, 998:134– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bocci:1995:ANS

[1595] M. Bocci, E. Scharf, P. Georgatsos, and M. Hansen. ATM network simulation support for TMN systems. Lecture Notes in Computer Science, 998: 146–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsson:1995:ETO

[1596] R. Larsson and D. Cochrane. Engineering a TMN in an open distributed processing environment. Lecture Notes in Computer Science, 998:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewis:1995:EMD

[1597] D. Lewis, T. Tiropanis, L. H. Bjerring, and L. M. Ericsson. Experiences in multi-domain management service development. Lecture Notes in Computer Science, 998:174–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

May:1995:RBI

[1598] J. May and A. Maia. The relationship between IOs and COs in VPN charging management. Lecture Notes in Computer Science, 998:185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sartzetakis:1995:MT

[1599] S. Sartzetakis, C. Stathopoulos, V. Kalogeraki, and D. Griffin. Managing the TMN. Lecture Notes in Computer Science, 998:200–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kleier:1995:PCS

[1600] S. Kleier and C. Goerg. Personal communication system realizations: Performance and quality of service aspects on SS-no. 7. Lecture Notes in Computer Science, 998:213–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mahajan:1995:SOP

[1601] A. O. Mahajan and K. V. Lever. A selforganisation plane for distributed mobile wireless networks. Lecture Notes in Computer Science, 998:225–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karabek:1995:FHR

[1602] R. Karabek and W. Reinhardt. Future hypermedia retrieval systems and their impact on transfer systems. Lecture Notes in Computer Science, 998: 237–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaspoz:1995:VDR

[1603] J.-P. Gaspoz, C. Gbaguidi, and J. Meinkoehn. VPN on DCE: From reference configuration to implementation. Lecture Notes in Computer Science, 998:249–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kisker:1995:APN

[1604] W. Kisker and J. M. Schneider. ATM public network management in PRE-PARE. Lecture Notes in Computer Science, 998:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgatsos:1995:MSP

[1605] P. Georgatsos and D. Griffin. Management services for performance verification in broadband multi-service networks. Lecture Notes in Computer Science, 998:275–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morris:1995:SBR

[1606] C. Morris and J. Nelson. An SDL based realisation of an IN service development environment. Lecture Notes in Computer Science, 998:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritterbecks:1995:PED

[1607] A. Ritterbecks, D. Lawniczak, and N. Niebert. Performance evaluation of database concepts for personal communication systems. Lecture Notes in Computer Science, 998:309–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goebel:1995:NGD

[1608] V. Goebel, B. H. Johansen, H. C. Lochsen, and T. Plagemann. Next generation database technologies for advanced communication services. Lecture Notes in Computer Science, 998: 320–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olsen:1995:UST

[1609] A. Olsen and B. B. Noerbaek. Using SDL for targeting services to CORBA. Lecture Notes in Computer Science, 998:334–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tag:1995:EAO

[1610] M. Tag, A. Lesch, P. Fisher, and A. Gallis. An engineering approach for open multimedia services management. Lecture Notes in Computer Science, 998:347–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coates:1995:AUM

[1611] P. Coates and J. Ellman. An approach to user management of broadband services. Lecture Notes in Computer Science, 998:360–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milsted:1995:OOM

[1612] K. Milsted. OMT object models of telecommunications services. Lecture Notes in Computer Science, 998:369– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMeer:1995:QMD

[1613] J. De Meer and A. Hafid. QoS modelling of distributed teleoperating services. Lecture Notes in Computer Science, 998:380–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magedanz:1995:IIA

[1614] T. Magedanz. On the impacts of intelligent agent concepts on future telecommunication environments. Lecture Notes in Computer Science, 998: 396–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vassila:1995:IAM

[1615] A. Vassila and G. J. Knight. Introducing active managed objects for effective and autonomous distributed management. Lecture Notes in Computer Science, 998:415–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pavlou:1995:IRM

[1616] G. Pavlou, K. McCarthy, G. Mykoniatis, and J. Sanchez. Intelligent remote monitoring. Lecture Notes in Computer Science, 998:430–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruano:1995:SCM

[1617] F. Ruano and C. Aurrecoechea. Session control model for TINA multimedia services. Lecture Notes in Computer Science, 998:447–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huelamo:1995:TBA

[1618] J. Huelamo, E. Carrera, and H. Zuidweg. TINA based advanced UPT service prototype: Early introduction of TINA through the IN domain. Lecture Notes in Computer Science, 998: 458–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reyniers:1995:UAA

[1619] G. Reyniers and P. Hellemans. Use of atomic action principles to co-ordinate the interaction between TINA service managers. Lecture Notes in Computer Science, 998:468–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berndt:1995:CAF

[1620] H. Berndt, M. Chapman, P. Schoo, and I. Toennby. A comparison of architectures for future telecommunication services. Lecture Notes in Computer Science, 998:476–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAurelio:1995:EAT

[1621] M. D’Aurelio and M. Peirgigli. External access to TMN: an ODP specification. Lecture Notes in Computer Science, 998:490–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bakker:1995:IPF

[1622] H. Bakker, J. H. Heilesen, G. Schapeler, and M. Anagnostou. The implementation of PSCS features in a RACE ATM demonstrator. Lecture Notes in Computer Science, 998:492–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ostrowski:1995:SMI

[1623] M. Ostrowski and R. Coutts. Simulation model for intelligent network based personal communication services. Lecture Notes in Computer Science, 998: 495–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SantAna:1995:IPT

[1624] P. SantAna, O. Bravo, and A. Martin. Integration path of IN and TMN architectures — using UPT as a case study. Lecture Notes in Computer Science, 998:497–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sharobim:1995:DSM

[1625] H. R. Sharobim. Dedicated server multicast routing for large scale ATM networks. Lecture Notes in Computer Science, 998:500–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnold:1995:AEA

[1626] F. Arnold. Application of IS and N for efficient aircraft pre-design. Lecture Notes in Computer Science, 998: 501–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stephanidis:1995:DAM

[1627] C. Stephanidis, C. Karagiannidis, and A. Koumpis. Developing alternative metaphors for special B-ISDN services. Lecture Notes in Computer Science, 998:504–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stephanidis:1995:TBU

[1628] C. Stephanidis, Y. Mitsopoulos, A. Stergiou, and A. Koumpis. Tool based user interface construction facilitating access to users with disabilities. Lecture Notes in Computer Science, 998:506–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asarin:1995:SCS

[1629] E. Asarin, O. Maler, and A. Pnueli.

Symbolic controller synthesis for discrete and timed systems. Lecture Notes in Computer Science, 999:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birdwell:1995:CIS

[1630] J. D. Birdwell and B. C. Moore. Condensation of information from signals for process modeling and control. Lecture Notes in Computer Science, 999: 15–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benvenisbe:1995:CSS

[1631] A. Benvenisbe, B. C. Levy, E. Fabre, and P. Le Guernie. A calculus of stochastic systems for the specification simulation and hidden state estimation of hybrid stochastic/nonstochastic systems. Lecture Notes in Computer Science, 999:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1995:AVS

[1632] A. Bouajjani, R. Echahed, and R. Robbana. On the automatic verification of systems with continuous variables and unbounded discrete data structures. Lecture Notes in Computer Science, 999:61–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caines:1995:DCH

[1633] P. E. Caines and Y.-J. Wei. On dynamically consistent hybrid systems. Lecture Notes in Computer Science, 999:

86–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeClaris:1995:SLN

[1634] N. De Claris and M.-C. Su. A selflearning neuro-fuzzy system. Lecture Notes in Computer Science, 999:106– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dogruel:1995:MSI

[1635] M. Dogruel and U. Oezguener. Modeling and stability issues in hybrid systems. Lecture Notes in Computer Science, 999:118–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deshpande:1995:VCH

[1636] A. Deshpande and P. Varaiya. Viable control of hybrid systems. Lecture Notes in Computer Science, 999: 128–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godbole:1995:HHC

[1637] D. N. Godbole, J. Lygeros, and S. Sastry. Hierarchical hybrid control. Lecture Notes in Computer Science, 999: 166–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grossman:1995:HSQ

[1638] R. L. Grossman and M. Sweedler. Hybrid systems and quantum automata: Preliminary announcement. Lecture Notes in Computer Science, 999:191– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guekenheimer:1995:PHS

[1639] J. Guekenheimer and S. Johnson. Planar hybrid systems. Lecture Notes in Computer Science, 999:202–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1995:PHC

[1640] V. Gupta, R. Jagadeesan, V. Saraswat, and D. G. Bobrow. Programming in hybrid constraint languages. Lecture Notes in Computer Science, 999:226– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1995:NAI

[1641] T. A. Henzinger and P.-H. Ho. A note on abstract interpretation strategies for hybrid automata. Lecture Notes in Computer Science, 999:252–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1995:HCH

[1642] T. A. Henzinger and P.-H. Ho.

HYTECH: The Cornell HYbrid TECHnology tool. Lecture Notes in Computer Science, 999:265–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohn:1995:HSF

[1643] W. Kohn, A. Nerode, and J. B. Remmel. Hybrid systems as Finsler manifolds: Finite state control as approximation to connections. Lecture Notes in Computer Science, 999:291– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nerode:1995:CFP

[1644] A. Nerode, J. B. Remmel, and A. Yakhnis. Controllers as fixed points of setvalued operators. Lecture Notes in Computer Science, 999:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lemmon:1995:CHC

[1645] M. Lemmon, C. Bett, P. Szymanski, and P. Antsaklis. Constructing hybrid control systems from robust linear control agents. Lecture Notes in Computer Science, 999:322–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puri:1995:VHS

[1646] A. Puri and P. Varaiya. Verification of hybrid systems using abstractions. Lecture Notes in Computer Science, 999:359–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruisch:1995:CCP

[1647] J. Ruisch. Control of continuous plants by symbolic output feedback. Lecture Notes in Computer Science, 999:370– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raon:1995:HCR

[1648] A. P. Raon, H. Rischel, M. Holdgaard, and T. J. Eriksen. Hybrid control of a robot. Lecture Notes in Computer Science, 999:391–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roux:1995:VTB

[1649] O. Roux and V. Rusu. Verifying timebounded properties for ELECTRE reactive programs with stopwatch automata. Lecture Notes in Computer Science, 999:405–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarjoughian:1995:IMF

[1650] H. S. Sarjoughian and B. Zeigler. Inductive modeling: A framework marrying systems theory and non-monotonic reasoning. Lecture Notes in Computer Science, 999:417–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shyamasundar:1995:SVH

[1651] R. K. Shyamasundar and S. Ramesh.

Semantics and verification of hierarchical CRP programs. Lecture Notes in Computer Science, 999:436–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sliver:1995:ICD

[1652] J. A. Sliver, P. J. Antsaklis, and M. D. Lemmon. Interface and controller design for hybrid control systems. Lecture Notes in Computer Science, 999:462–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiltus:1995:HO

[1653] M. Tiltus and B. Egardt. Hybrid objects. Lecture Notes in Computer Science, 999:493–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:MHS

[1654] Y. Y. Yang, D. A. Linkens, and S. P. Banks. Modeling of hybrid systems based on extended coloured Petri nets. Lecture Notes in Computer Science, 999:509–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeigler:1995:DFM

[1655] B. P. Zeigler, H. S. Song, T. G. Kim, and H. Praehofer. DEVS framework for modeling, simulation, analysis, and design of hybrid systems. Lecture Notes in Computer Science, 999:529–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:SHC

[1656] Y. Zhang and A. K. Mackworth. Synthesis of hybrid constraint-based controllers. Lecture Notes in Computer Science, 999:552–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brassard:1995:QJC

[1657] G. Brassard. A quantum jump in computer science. Lecture Notes in Computer Science, 1000:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steels:1995:ALR

[1658] L. Steels and M. Tokoro. Artificial life and real world computing. Lecture Notes in Computer Science, 1000:15– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siegelmann:1995:RNN

[1659] H. T. Siegelmann. Recurrent neural networks. Lecture Notes in Computer Science, 1000:29–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McColl:1995:SC

[1660] W. F. McColl. Scalable computing. Lecture Notes in Computer Science, 1000:46–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monien:1995:EUP

[1661] B. Monien, R. Diekmann, R. Feldmann, and R. Klasing. Efficient use of parallel and distributed systems: From theory to practice. Lecture Notes in Computer Science, 1000:62–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snyder:1995:EVM

[1662] L. Snyder. Experimental validation of models of parallel computation. Lecture Notes in Computer Science, 1000: 78–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiedermann:1995:QVP

[1663] J. Wiedermann. Quo vadetis, parallel machine models? Lecture Notes in Computer Science, 1000:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bai:1995:TLA

[1664] Z. Bai, D. Day, J. Demmel, and J. Dongarra. Templates for linear algebra problems. Lecture Notes in Computer Science, 1000:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beth:1995:ABI

[1665] T. Beth, A. Klappenecker, T. Minkwitz, and A. Nueckel. The ART behind IDEAS. Lecture Notes in Computer Science, 1000:141–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adleman:1995:ANT

[1666] L. M. Adleman. Algorithmic number theory and its relationship to computational complexity. Lecture Notes in Computer Science, 1000:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakano:1995:ECA

[1667] S. Nakano, X. Zhou, and T. Nishizeki. Edge-coloring algorithms. Lecture Notes in Computer Science, 1000:172– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hannenhalli:1995:TCT

[1668] S. Hannenhalli and P. A. Pevzner. Towards a computational theory of genome rearrangements. Lecture Notes in Computer Science, 1000:184–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herlihy:1995:ATD

[1669] M. Herlihy and S. Rajsbaum. Algebraic topology and distributed computing — A primer. Lecture Notes in Computer Science, 1000:203–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anuchitanukul:1995:DB

[1670] A. Anuchitanukul, Z. Manna, and T. E.

Uribe. Differential BDDs. Lecture Notes in Computer Science, 1000:218– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1995:ATG

[1671] P. K. Agarwal and M. Sharir. Algorithmic techniques for geometric optimization. Lecture Notes in Computer Science, 1000:234–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nievergelt:1995:ANH

[1672] J. Nievergelt, R. Gasser, F. Maeser, and C. Wirth. All the needles in a haystack: Can exhaustive search overcome combinatorial chaos? Lecture Notes in Computer Science, 1000:254– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radcliffe:1995:FLS

[1673] N. J. Radcliffe and P. D. Surry. Fundamental limitations on search algorithms: Evolutionary computing in perspective. Lecture Notes in Computer Science, 1000:275–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1995:MSM

[1674] M. Broy. Mathematical system models as a basis of software engineering. Lecture Notes in Computer Science, 1000:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shaw:1995:FFS

[1675] M. Shaw and D. Garlan. Formulations and formalisms in software architecture. Lecture Notes in Computer Science, 1000:307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smolka:1995:OPMb

[1676] G. Smolka. The Oz programming model. Lecture Notes in Computer Science, 1000:324–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wood:1995:SGM

[1677] D. Wood. Standard generalized markup language: Mathematical and philosophical issues. Lecture Notes in Computer Science, 1000:344–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gries:1995:AUU

[1678] D. Gries and F. B. Schneider. Avoiding the undefined by underspecification. Lecture Notes in Computer Science, 1000:366–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harel:1995:TTR

[1679] D. Harel. Towards a theory of recursive structures. Lecture Notes in Computer Science, 1000:374–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pratt:1995:CST

[1680] V. Pratt. Chu spaces and their interpretation as concurrent objects. Lecture Notes in Computer Science, 1000:

392–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruynooghe:1995:AUK

[1681] M. Bruynooghe, M. Codish, and A. Mulkers. Abstracting unification: A key step in the design of logic program analyses. Lecture Notes in Computer Science, 1000:406–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1995:PSC

[1682] R. Anderson and R. Needham. Programming Satan’s computer. Lecture Notes in Computer Science, 1000:426– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reisig:1995:PNM

[1683] W. Reisig. Petri net models of distributed algorithms. Lecture Notes in Computer Science, 1000:441–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarke:1995:SIM

[1684] E. M. Clarke and S. Jha. Symmetry and induction in model checking. Lecture Notes in Computer Science, 1000:455–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vardi:1995:AAP

[1685] M. Y. Vardi. Alternating automata and program verification. Lecture Notes in Computer Science, 1000:471–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandewall:1995:RAA

[1686] E. Sandewall. Reasoning about actions and change with ramification. Lecture Notes in Computer Science, 1000:486– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eklundh:1995:TAV

[1687] J.-O Eklundh. Trends in active vision. Lecture Notes in Computer Science, 1000:505–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1995:CML

[1688] M. Li and P. Vitanyi. Computational machine learning in theory and praxis. Lecture Notes in Computer Science, 1000:518–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yager:1995:FST

[1689] R. R. Yager. Fuzzy sets as a tool for modeling. Lecture Notes in Computer Science, 1000:536–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanRijsbergen:1995:IRI

[1690] C. J. Van Rijsbergen. Information retrieval and informative reasoning. Lecture Notes in Computer Science, 1000: 549–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vossen:1995:DTM

[1691] G. Vossen. Database transaction models. Lecture Notes in Computer Science, 1000:560–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bulterman:1995:MAT

[1692] D. C. A. Bulterman and L. Hardman. Multimedia authoring tools: State of the art and research challenges. Lecture Notes in Computer Science, 1000:575–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaeppner:1995:CMD

[1693] T. Kaeppner and R. Steinmetz. Computational models for distributed multimedia applications. Lecture Notes in Computer Science, 1000:592–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1995:HSI

[1694] H. Maurer. Hypermedia systems as Internet tools. Lecture Notes in Computer Science, 1000:608–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyano:1995:APA

[1695] S. Miyano. Algorithmic problems arising from genome informatics. Lecture Notes in Computer Science, 1004:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimozono:1995:AAA

[1696] S. Shimozono. An approximation algorithm for alphabet indexing problem. Lecture Notes in Computer Science, 1004:2–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katajainen:1995:FSE

[1697] J. Katajainen, A. Moffat, and A. Turpin. A fast and space-economical algorithm for length-limited coding. Lecture Notes in Computer Science, 1004:12–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhattacharya:1995:CLT

[1698] B. K. Bhattacharya and A. Mukhopadhyay. Computing in linear time a chord from which a simple polygon is weakly internally visible. Lecture Notes in Computer Science, 1004:22–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Datta:1995:CSP

[1699] A. Datta, C. A. Hipke, and S. Schuierer. Competitive searching in polygonsbeyond generalised streets. Lecture Notes in Computer Science, 1004:32– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lokam:1995:LVN

[1700] S. V. Lokam, M. Mahajan, and V. Vinay. Logspace verifiers, NC, and NP. Lecture Notes in Computer Science, 1004:52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karg:1995:SAC

[1701] C. Karg and R. Schuler. Structure in average case complexity. Lecture Notes in Computer Science, 1004:62– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chien:1995:SGL

[1702] H. Chien and W. Steiger. Some geometric lower bounds. Lecture Notes in Computer Science, 1004:72–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arge:1995:COB

[1703] L. Arge. The I/O-complexity of ordered binary-decision diagram manipulation. Lecture Notes in Computer Science, 1004:82–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frank:1995:TAD

[1704] A. Frank, T. Ibaraki, and H. Nagamochi. Two arc disjoint paths in Eulerian digraphs. Lecture Notes in Computer Science, 1004:92–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asahiro:1995:FDS

[1705] Y. Asahiro and K. Iwama. Finding dense subgraphs. Lecture Notes in Computer Science, 1004:102–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1995:FSS

[1706] A. Gupta and N. Nishimura. Finding smallest supertrees. Lecture Notes in Computer Science, 1004:112–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1995:WDC

[1707] M.-S. Chang. Weighted domination of cocomparability graphs. Lecture Notes in Computer Science, 1004:122– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreev:1995:PCA

[1708] A. E. Andreev, A. Clementi, P. Crescenzi, and E. Dahlhaus. The parallel complexity of approximating the high degree subgraph problem. Lecture Notes in Computer Science, 1004: 132–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bafna:1995:CRA

[1709] V. Bafna, P. Berman, and T. Fujito. Constant ratio approximations of the weighted feedback vertex set problem for undirected graphs. Lecture Notes in Computer Science, 1004:142–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halldorsson:1995:GAI

[1710] M. M. Halldorsson and K. Yoshihara. Greedy approximations of independent sets in low degree graphs. Lecture Notes in Computer Science, 1004:152– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crossley:1995:PL

[1711] J. N. Crossley. Practical logic. Lecture Notes in Computer Science, 1004: 162–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ono:1995:AAM

[1712] T. Ono, T. Hirata, and T. Asano. An approximation algorithm for MAX 3-SAT. Lecture Notes in Computer Science, 1004:163–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tani:1995:LRR

[1713] S. Tani and K. Yamazaki. Learning of restricted RNLC graph languages. Lecture Notes in Computer Science, 1004:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadimitriou:1995:OID

[1714] C. H. Papadimitriou, S. Ramanathan, and P. V. Rangan. Optimal information delivery. Lecture Notes in Computer Science, 1004:181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santoro:1995:CTC

[1715] N. Santoro, J. Ren, and A. Nayak.

On the complexity of testing for catastrophic faults. Lecture Notes in Computer Science, 1004:188–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwon:1995:SPT

[1716] O.-H. Kwon and K.-Y. Chwa. Scheduling parallel tasks with individual deadlines. Lecture Notes in Computer Science, 1004:198–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonzurGathen:1995:OGP

[1717] J. Von zur Gathen and I. Shparlinski. Orders of Gauss periods in finite fields. Lecture Notes in Computer Science, 1004:208–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havas:1995:HPA

[1718] G. Havas and B. S. Majewski. A hard problem that is almost always easy. Lecture Notes in Computer Science, 1004:216–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sekine:1995:CTP

[1719] K. Sekine, H. Imai, and S. Tani. Computing the Tutte polynomial of a graph of moderate size. Lecture Notes in Computer Science, 1004:224–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dedorath:1995:MEP

[1720] J. Dedorath, J. Gergov, and T. Hagerup. More efficient parallel flow algorithms. Lecture Notes in Computer Science, 1004:234–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1995:LTP

[1721] S. Carlsson and M. Sundstroem. Linear-time in-place selection in less than 3n comparisons. Lecture Notes in Computer Science, 1004:244–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1995:HCO

[1722] S. Carlsson and J. Chen. Heap construction: Optimal in both worst and average cases? Lecture Notes in Computer Science, 1004:254–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1995:FAD

[1723] F. Shi. Fast approximate dictionary matching. Lecture Notes in Computer Science, 1004:264–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1995:UVC

[1724] T.-S. Hsu. Undirected vertexconnectivity structure and smallest four-vertex-connectivity augmentation. Lecture Notes in Computer Science, 1004:274–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damaschke:1995:SMF

[1725] P. Damaschke. Searching for a monotone function by independent threshold queries. Lecture Notes in Computer Science, 1004:284–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makino:1995:FSA

[1726] K. Makino and T. Ibaraki. A fast and simple algorithm for identifying 2-Monotonic positive Boolean functions. Lecture Notes in Computer Science, 1004:291–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yen:1995:DBT

[1727] H.-C. Yen, S.-T. Jian, and T.-P. Lao.

Deciding bisimulation and trace equivalences for systems with many identical processes. Lecture Notes in Computer Science, 1004:301–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preparata:1995:SAL

[1728] F. P. Preparata. Should Amdahl’s law be repealed? Lecture Notes in Computer Science, 1004:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamada:1995:EHR

[1729] Y. Hamada, A. Mei, Y. Nishitani, and Y. Igarashi. Embeddings of hyperrings in hypercubes. Lecture Notes in Computer Science, 1004:312–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1995:FAC

[1730] S.-W. Cheng and C.-K. Tang. A fast algorithm for computing optimal rectilinear Steiner trees for extremal point sets. Lecture Notes in Computer Science, 1004:322–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1995:AFC

[1731] X. Zhou and T. Nishizeki. Algorithms for finding f-colorings of partial k-trees. Lecture Notes in Computer Science, 1004:332–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1995:SCT

[1732] H.-G. Cho and A. Zelikovsky. Spanning closed trail and Hamiltonian cycle in grid graphs. Lecture Notes in Computer Science, 1004:342–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1995:LTA

[1733] W.-L. Hsu. A linear time algorithm for finding maximal planar subgraphs. Lecture Notes in Computer Science, 1004:352–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abello:1995:IOF

[1734] J. Abello, V. Estivill-Castro, T. Shermer, and J. Urrutia. Illumination with orthogonal floodlights. Lecture Notes in Computer Science, 1004:362– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1995:NQE

[1735] P. Bose and G. Toussaint. No quadrangulation is extremely odd. Lecture Notes in Computer Science, 1004:372– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chin:1995:FMA

[1736] F. Chin, J. Snoeyink, and Cao An Wang. Finding the medial axis of a simple polygon in linear time. Lecture Notes in Computer Science, 1004:382– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krznaric:1995:FSA

[1737] D. Krznaric and C. Levcopoulos. The first subquadratic algorithm for complete linkage clustering. Lecture Notes in Computer Science, 1004:392–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradford:1995:MNB

[1738] P. G. Bradford and R. Fleischer. Matching nuts and bolts faster. Lecture Notes in Computer Science, 1004: 402–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fu:1995:LMT

[1739] J. J. Fu. Linear matching-time algorithm for the directed graph isomorphism problem. Lecture Notes in Computer Science, 1004:409–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1995:RAP

[1740] S. Fujita, T. Kameda, and M. Yamashita. A resource assignment problem on graphs. Lecture Notes in Computer Science, 1004:418–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:NAPa

[1741] Z.-Z. Chen. NC algorithms for partitioning sparse graphs into induced forests with an application. Lecture Notes in Computer Science, 1004:428– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adams:1995:OCS

[1742] P. Adams and M. Solomon. An overview of the CAPITL software development environment. Lecture Notes in Computer Science, 1005:1–34, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cagan:1995:UCM

[1743] M. Cagan. Untangling configuration management. Lecture Notes in Computer Science, 1005:35–52, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leblang:1995:SCM

[1744] D. B. Leblang and P. H. Levine.

Software configuration management: Why is it needed and what should it do? Lecture Notes in Computer Science, 1005:53–60, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reichenberger:1995:VTO

[1745] C. Reichenberger. VOODOO — A tool for orthogonal version management. Lecture Notes in Computer Science, 1005:61–79, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brereton:1995:DSB

[1746] P. Brereton and P. Singleton. Deductive software building. Lecture Notes in Computer Science, 1005:81–87, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmerl:1995:DCM

[1747] B. R. Schmerl and C. D. Marlin.

Designing configuration management facilities for dynamically bound systems. Lecture Notes in Computer Science, 1005:88–100, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1995:CMT

[1748] Y.-J. Lin and S. P. Reiss. Configuration management in terms of modules. Lecture Notes in Computer Science, 1005:101–117, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estublier:1995:TDV

[1749] J. Estublier and R. Casallas. Three dimensional versioning. Lecture Notes in Computer Science, 1005:118–135, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viskari:1995:RAC

[1750] J. Viskari. A rationale for automated configuration status accounting. Lecture Notes in Computer Science, 1005:138–144, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lago:1995:TPS

[1751] P. Lago and R. Conradi. Transaction planning to support coordination. Lecture Notes in Computer Science, 1005:145–151, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buffenbarger:1995:SSM

[1752] J. Buffenbarger. Syntactic software merging. Lecture Notes in Computer Science, 1005:153–172, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godart:1995:ASR

[1753] C. Godart, G. Canals, F. Charoy, and P. Molli. About some relationships between configuration management, software process and cooperative work: The COO environment. Lecture Notes in Computer Science, 1005:173–178, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacKay:1995:SAC

[1754] S. A. MacKay. The state-of-the-art in concurrent, distributed configuration management. Lecture Notes in Computer Science, 1005:180–193, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1995:CMS

[1755] L. Allen, G. Fernandez, K. Kane, and D. Leblang. ClearCase MultiSite: Supporting geographically-distributed software development. Lecture Notes in Computer Science, 1005:194–214, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tryggeseth:1995:MSV

[1756] E. Tryggeseth, B. Gulla, and R. Conradi. Modelling systems with variability using the PROTEUS configuration language. Lecture Notes in Computer Science, 1005:216–240, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clemm:1995:OS

[1757] G. M. Clemm. The Odin system. Lecture Notes in Computer Science, 1005: 241–262, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1995:SSH

[1758] A. Christensen and T. Egge. Store — A system for handling third-party applications in a heterogeneous computer environment. Lecture Notes in Computer Science, 1005:263–276, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adams:1995:WCB

[1759] C. Adams. Why can’t I buy an SCM tool? Lecture Notes in Computer Science, 1005:278–287, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ray:1995:ESB

[1760] R. J. Ray. Experiences with a scriptbased software configuration management system. Lecture Notes in Computer Science, 1005:282–287, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parker:1995:CCC

[1761] K. Parker. Customization of a commercial CM system to provide better management mechanisms. Lecture Notes in Computer Science, 1005:289–292, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Persson:1995:ECI

[1762] A. Persson. Experiences of customization and introduction of a CM model. Lecture Notes in Computer Science, 1005:293–303, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderHoek:1995:DCM

[1763] A. Van der Hoek, D. Heimbigner, and A. L. Wolf. Does configuration management research have a future? Lecture Notes in Computer Science, 1005:305–309, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ekenberg:1995:CBS

[1764] L. Ekenberg and P. Johannesson. Conflictfreeness as a basis for schema integration. Lecture Notes in Computer Science, 1006:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biskup:1995:DSD

[1765] J. Biskup. Database schema design theory: Achievements and challenges. Lecture Notes in Computer Science, 1006:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panda:1995:MSD

[1766] B. Panda and W. Perrizo. Maintaining surrogate data for query acceleration in multilevel secure database systems. Lecture Notes in Computer Science, 1006:45–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andres:1995:CDC

[1767] F. Andres, F. Kwakkel, and M. L.

Kersten. Calibration of a DBMS cost model with the software testpilot. Lecture Notes in Computer Science, 1006:58–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poenighaus:1995:FSS

[1768] R. Poenighaus. “favourite” SQL-Statements — an empirical analysis of SQL-Usage in commercial applications. Lecture Notes in Computer Science, 1006:75–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhar:1995:SAD

[1769] S. Bhar and K. Barker. Static allocation in distributed objectbase systems: A graphical approach. Lecture Notes in Computer Science, 1006:92–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reddy:1995:EDA

[1770] M. P. Reddy and R. Y. Wang. Estimating data accuracy in a federated database environment. Lecture Notes in Computer Science, 1006:115– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhargava:1995:SDT

[1771] B. Bhargava, Y. Zhang, and S. Goel. A study of distributed transaction processing in an Internetwork. Lecture Notes in Computer Science, 1006:135– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harris:1995:UOR

[1772] W. Harris, R. McClatchey, and N. Baker. The use of an object repository in the configuration of control systems at CERN. Lecture Notes in Computer Science, 1006:153–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furuse:1995:IPE

[1773] K. Furuse, K. Asada, and A. Iizawa. Implementation and performance evaluation of compressed bit-sliced signature files. Lecture Notes in Computer Science, 1006:164–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sako:1995:PMT

[1774] K. Sako, T. Nemoto, M. Kitsuregawa, and M. Takagi. Partial migration in an 8mm tape based tertiary storage file system and its performance evaluation through satellite image processing applications. Lecture Notes in Computer Science, 1006:178–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleury:1995:SAR

[1775] L. Fleury, C. Djeraba, H. Briand, and J. Philippe. Some aspect of rule discovery in data bases. Lecture Notes in Computer Science, 1006:192–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Terashima:1995:TFM

[1776] N. Terashima. Telesensation — fusion of multi-media information and information highways. Lecture Notes in Computer Science, 1006:206–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolp:1995:RDD

[1777] M. Kolp and E. Zimanyi. Relational database design using an ER approach and Prolog. Lecture Notes in Computer Science, 1006:214–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saeki:1995:SDI

[1778] M. Saeki, S. Sureerat, and K. Yoshida.

Supporting distributed individual work in cooperative specification development. Lecture Notes in Computer Science, 1006:232–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parimala:1995:HCD

[1779] N. Parimala. Handling changes in dynamic specifications in object oriented systems. Lecture Notes in Computer Science, 1006:248–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mumick:1995:RMV

[1780] I. S. Mumick. The rejuvenation of materialized views. Lecture Notes in Computer Science, 1006:258–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Du:1995:NLR

[1781] X. Du and N. Ishii. Normalization of linear recursions based on graph transformations. Lecture Notes in Computer Science, 1006:265–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ng:1995:DRA

[1782] Y.-K. Ng and N. Qaraeen. Data retrieval and aggregates in SQL0*/ NR. Lecture Notes in Computer Science, 1006:283–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganguly:1995:PPC

[1783] S. Ganguly, A. Gerasoulis, and W. Wang. Partitioning pipelines with communication costs. Lecture Notes in Computer Science, 1006:302–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baum:1995:CCP

[1784] U. Baum and S. Blackburn. Clockcontrolled pseudorandom generators on finite groups. Lecture Notes in Computer Science, 1008:6–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chambers:1995:RMR

[1785] W. G. Chambers. On random mappings and random permutations. Lecture Notes in Computer Science, 1008: 22–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:1995:BCG

[1786] C. Ding. Binary cyclotomic generators. Lecture Notes in Computer Science, 1008:29–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobbertin:1995:CBF

[1787] H. Dobbertin. Construction of bent functions and balanced Boolean functions with high nonlinearity. Lecture Notes in Computer Science, 1008:61– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1995:ALS

[1788] X. Lai. Additive and linear structures of cryptographic functions. Lecture Notes in Computer Science, 1008: 75–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rivest:1995:REA

[1789] R. L. Rivest. The RC5 encryption algorithm. Lecture Notes in Computer Science, 1008:86–96, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaze:1995:MBC

[1790] M. Blaze and B. Schneier. The MacGuffin block cipher algorithm. Lecture Notes in Computer Science, 1008: 97–110, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

                                        counterpane.com/macguffin.html.

Nyberg:1995:BRF

[1791] K. Nyberg. S-boxes and round functions with controllable linearity and differential uniformity. Lecture Notes in Computer Science, 1008:111–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OConnor:1995:PLA

[1792] L. O’Connor. Properties of linear approximation tables. Lecture Notes in Computer Science, 1008:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1995:SOC

[1793] R. Anderson. Searching for the optimum correlation attack. Lecture Notes in Computer Science, 1008:137–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biham:1995:KPA

[1794] E. Biham and P. C. Kocher. A known plaintext attack on the PKZIP stream cipher. Lecture Notes in Computer Science, 1008:144–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golic:1995:LCS

[1795] J. D. Golic. Linear cryptanalysis of stream ciphers. Lecture Notes in Computer Science, 1008:154–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klapper:1995:FCS

[1796] A. Klapper. Feedback with carry shift registers over finite fields. Lecture Notes in Computer Science, 1008:170– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacKay:1995:FEM

[1797] D. J. C. MacKay. A free energy minimization framework for inference problems in modulo 2 arithmetic. Lecture Notes in Computer Science, 1008:179– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knudsen:1995:THO

[1798] L. R. Knudsen. Truncated and higher order differentials. Lecture Notes in Computer Science, 1008:196–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massey:1995:SKO

[1799] J. L. Massey. SAFER K-64: One year later. Lecture Notes in Computer Science, 1008:212–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rijmen:1995:ICD

[1800] V. Rijmen and B. Preneel. Improved characteristics for differential cryptanalysis of hash functions based on block ciphers. Lecture Notes in Computer Science, 1008:242–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaliski:1995:LCU

[1801] B. S. Kaliski and M. J. B. Robshaw. Linear cryptanalysis using multiple approximations and FEAL. Lecture Notes in Computer Science, 1008:249–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloecher:1995:PLC

[1802] U. Bloecher and M. Dichtl. Problems with the linear cryptanalysis of DES using more than one active S-box per round. Lecture Notes in Computer Science, 1008:265–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daemen:1995:CM

[1803] J. Daemen, R. Govaerts, and J. Vandewalle. Correlation matrices. Lecture Notes in Computer Science, 1008:275– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaudenay:1995:NMC

[1804] S. Vaudenay. On the need for multipermutations: Cryptanalysis of MD4 and SAFER. Lecture Notes in Computer Science, 1008:286–297, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucks:1995:HEI

[1805] S. Lucks. How to exploit the intractability of exact TSP for cryptography. Lecture Notes in Computer Science, 1008:298–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roe:1995:HRE

[1806] M. Roe. How to reverse engineer an EES device. Lecture Notes in Computer Science, 1008:305–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Penzhorn:1995:FHC

[1807] W. T. Penzhorn. A fast homophonic coding algorithm based on arithmetic coding. Lecture Notes in Computer Science, 1008:329–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1995:FKG

[1808] R. Anderson. On Fibonacci keystream generators. Lecture Notes in Computer Science, 1008:346–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rijmen:1995:CM

[1809] V. Rijmen and B. Preneel. Cryptanalysis of McGuffin. Lecture Notes in Computer Science, 1008:353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roe:1995:PBC

[1810] M. Roe. Performance of block ciphers and hash functions — one year later. Lecture Notes in Computer Science, 1008:359–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wheeler:1995:TTE

[1811] David J. Wheeler and Roger M. Needham. TEA, a tiny encryption algorithm. Lecture Notes in Computer Science, 1008:363–366, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.com/content/
p16916lx735m2562/. See also exten-

sions XTEA [2321] and XXTEA [2322].

Broy:1995:CSE

[1812] M. Broy and M. Wirsing. Correct software: From experiments to applications. Lecture Notes in Computer Science, 1009:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pepper:1995:MDC

[1813] P. Pepper and M. Wirsing. A method for the development of correct software. Lecture Notes in Computer Science, 1009:27–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pepper:1995:RSH

[1814] P. Pepper, R. Betschko, S. Dick, and K. Didrich. Realizing sets by hash tables. Lecture Notes in Computer Science, 1009:58–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinna:1995:EAG

[1815] M. Pinna and A. Poigne. Event automata as a generic model of reactive systems. Lecture Notes in Computer Science, 1009:74–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewerentz:1995:OOD

[1816] C. Lewerentz, T. Lindner, A. Rueping, and E. Sekerinski. On object-oriented design and verification. Lecture Notes in Computer Science, 1009:92– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gastinger:1995:DMS

[1817] S. Gastinger, R. Hennicker, and R. Stabl. Design of modular software systems with reuse. Lecture Notes in Computer Science, 1009:112–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hettler:1995:ATR

[1818] R. Hettler, D. Nazareth, F. Regensburger, and O. Slotosch. AVL trees revisited: A case study in SPEC-TRUM. Lecture Notes in Computer Science, 1009:128–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrich:1995:KRL

[1819] H.-D. Ehrich. Korso reference languages — concepts and application domains. Lecture Notes in Computer Science, 1009:150–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolter:1995:HCS

[1820] U. Wolter, K. Didrich, F. Cornelius, and M. Klar. How to cope with the spectrum of SPECTRUM. Lecture Notes in Computer Science, 1009:173– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burghardt:1995:FGS

[1821] J. Burghardt. A fine-grain sort discipline and its application to formal program construction. Lecture Notes in Computer Science, 1009:190–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1995:TLL

[1822] M. Gogolla, S. Conrad, G. Denker, and R. Herzig. TROLL light — the language and its development environment. Lecture Notes in Computer Science, 1009:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santen:1995:FAS

[1823] T. Santen, F. Kammueller, S. Jaehnichen, and M. Beyer. Formalization of algebraic specification in the development language DEVA. Lecture Notes in Computer Science, 1009:223–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonHenke:1995:CDM

[1824] F. Von Henke, A. Dold, H. Ruess, and D. Schwier. Construction and deduction methods for the formal development of software. Lecture Notes in Computer Science, 1009:239–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loeckx:1995:ESE

[1825] J. Loeckx and J. Zeyer. Experiences with a specification environment. Lecture Notes in Computer Science, 1009: 255–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krieg-Brueckner:1995:TCE

[1826] B. Krieg-Brueckner, J. Liu, H. Shi, and B. Wolff. Towards correct, efficient and reusable transformational developments. Lecture Notes in Computer Science, 1009:270–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deussen:1995:VST

[1827] P. Deussen, A. Hansmann, T. Kaeufl, and S. Klingenbeck. The verification system tatzelwurm. Lecture Notes in Computer Science, 1009:285–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stroetmann:1995:SPC

[1828] K. Stroetmann. Seduct — A proof compiler for first order logic. Lecture Notes in Computer Science, 1009:299– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bohn:1995:TTV

[1829] J. Bohn and H. Hungar. Traverdi — transformation and verification of distributed systems. Lecture Notes in Computer Science, 1009:317–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reif:1995:KAS

[1830] W. Reif. The KIV-Approach to software verification. Lecture Notes in Computer Science, 1009:339–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchss:1995:TSC

[1831] T. Fuchss, W. Reif, G. Schellhorn, and K. Stenzel. Three selected case studies in verification. Lecture Notes in Computer Science, 1009:371–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewerentz:1995:PCC

[1832] C. Lewerentz and T. Lindner. “production cell”: A comparative study in formal specification and verification. Lecture Notes in Computer Science, 1009:388–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cornelius:1995:KCS

[1833] F. Cornelius, H. Hussmann, and M. Loewe. The KORSO case study for software engineering with formal methods: A medical information system. Lecture Notes in Computer Science, 1009:417–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1995:ICB

[1834] J. R. C. Allen, D. W. R. Patterson, M. D. Mulvenna, and J. G. Hughes. Integration of case based retrieval with a relational database system in aircraft technical support. Lecture Notes in Computer Science, 1010:1–10, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bisio:1995:CES

[1835] R. Bisio and F. Malabocchia. Cost estimation of software projects through case base reasoning. Lecture Notes in Computer Science, 1010:11–22, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1995:ODA

[1836] B. Fuchs, A. Mille, and B. Chiron. Operator decision aiding by adaptation of supervision strategies. Lecture Notes in Computer Science, 1010:23–32, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geffraye:1995:PDS

[1837] F. Geffraye, J. L. Wybo, and A. Russeil. PROFIL: A decision support tool for metallic sections design using a CBR approach. Lecture Notes in Computer Science, 1010:33–42, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macura:1995:MRI

[1838] R. T. Macura and K. J. Macura.

MacRad: Radiology image resource with a case-based retrieval system. Lecture Notes in Computer Science, 1010: 43–54, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marir:1995:RIB

[1839] F. Marir and I. Watson. Representing and indexing building refurbishment cases for multiple retrieval of adaptable pieces of cases. Lecture Notes in Computer Science, 1010:55–66, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Netten:1995:LSF

[1840] B. D. Netten and R. A. Vingerhoeds.

Large-scale fault diagnosis for onboard train systems. Lecture Notes in Computer Science, 1010:67–76, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Opiyo:1995:CBR

[1841] E. T. O. Opiyo. Case-based reasoning for expertise relocation in support of rural health workers in developing countries. Lecture Notes in Computer Science, 1010:77–87, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stehr:1995:CML

[1842] J. Stehr. CBR and machine learning for combustion system design. Lecture Notes in Computer Science, 1010:98– 108, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agre:1995:KML

[1843] G. Agre. KBS maintenance as learning two-tiered domain representation. Lecture Notes in Computer Science, 1010: 109–120, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aluisio:1995:CBA

[1844] S. M. Aluisio and O. N. Oliveira. A case-based approach for developing writing tools aimed at non-native English users. Lecture Notes in Computer Science, 1010:121–132, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ashley:1995:RRC

[1845] K. D. Ashley and B. M. McLaren. Reasoning with reasons in case-based comparisons. Lecture Notes in Computer Science, 1010:133–144, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bartsch-Spoerl:1995:TIC

[1846] B. Bartsch-Spoerl. Towards the integration of case-based, schema-based and model-based reasoning for supporting complex design tasks. Lecture Notes in Computer Science, 1010:145–156, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1995:SCD

[1847] M. Brown, I. Watson, and N. Filer.

Separating the cases from the data: Towards more flexible case-based reasoning. Lecture Notes in Computer Science, 1010:157–168, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haigh:1995:RPA

[1848] K. Z. Haigh and M. Veloso. Route planning by analogy. Lecture Notes in Computer Science, 1010:169–180, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hastings:1995:CAU

[1849] J. D. Hastings, L. K. Branting, and J. A. Lockwood. Case adaptation using an incomplete causal model. Lecture Notes in Computer Science, 1010:181– 192, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurley:1995:EAC

[1850] N. Hurley. Evaluating the application of CBR in mesh design for simulation problems. Lecture Notes in Computer Science, 1010:193–204, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunt:1995:CMR

[1851] J. E. Hunt, D. E. Cooke, and H. Holstein. Case memory and retrieval based on the immune system. Lecture Notes in Computer Science, 1010:205–216, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Indurkhya:1995:UCD

[1852] N. Indurkhya and S. M. Weiss. Using case data to improve on rule-based function approximation. Lecture Notes in Computer Science, 1010:217–228, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leake:1995:LIC

[1853] D. B. Leake, A. Kinley, and D. Wilson. Learning to improve case adaptation by introspective reasoning and CBR. Lecture Notes in Computer Science, 1010:229–240, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muthoz-Avila:1995:RCS

[1854] H. Muthoz-Avila and J. Huellen. Retrieving cases in structured domains by using goal dependencies. Lecture Notes in Computer Science, 1010:241–252, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okamoto:1995:ACA

[1855] S. Okamoto and K. Satoh. An averagecase analysis of k-nearest neighbor classifier. Lecture Notes in Computer Science, 1010:253–264, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaza:1995:CTF

[1856] E. Plaza. Cases as terms: A feature term approach to the structured representation of cases. Lecture Notes in Computer Science, 1010:265–276, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Portinale:1995:AID

[1857] L. Portinale and P. Torasso. ADAPtER:

An integrated diagnostic system combining case-based and abductive reasoning. Lecture Notes in Computer Science, 1010:277–288, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Purvis:1995:AUC

[1858] L. Purvis and P. Pu. Adaptation using constraint satisfaction techniques. Lecture Notes in Computer Science, 1010:289–300, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ricci:1995:LLS

[1859] F. Ricci and P. Avesani. Learning a local similarity metric for case-based reasoning. Lecture Notes in Computer Science, 1010:301–312, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smyth:1995:EAG

[1860] B. Smyth and M. T. Keane. Experiments on adaptation-guided retrieval in case-based design. Lecture Notes in Computer Science, 1010:313–324, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Surma:1995:IRC

[1861] J. Surma and K. Vanhoof. Integrating rules and cases for the classification task. Lecture Notes in Computer Science, 1010:325–334, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Visser:1995:RKE

[1862] W. Visser. Reuse of knowledge: Empirical studies. Lecture Notes in Computer Science, 1010:335–346, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wettschereck:1995:WF

[1863] D. Wettschereck and D. W. Aha. Weighting features. Lecture Notes in Computer Science, 1010:347–358, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolverton:1995:IMP

[1864] M. Wolverton. An investigation of marker-passing algorithms for analogue retrieval. Lecture Notes in Computer Science, 1010:359–370, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auriol:1995:ISI

[1865] E. Auriol, S. Wess, M. Manago, and K.-D. Althoff. INRECA: A seamlessly integrated system based on inductive inference and case-based reasoning. Lecture Notes in Computer Science, 1010:371–380, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bakhtari:1995:DAA

[1866] S. Bakhtari and W. Oertel. DOM-ArC:

An active decision support system for quality assessment of cases. Lecture Notes in Computer Science, 1010:381– 390, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bichindaritz:1995:CBR

[1867] I. Bichindaritz. A case-based reasoner adaptive to different cognitive tasks. Lecture Notes in Computer Science, 1010:391–400, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cunningham:1995:UCO

[1868] P. Cunningham, B. Smyth, and N. Hurley. On the use of CBR in optimisation problems such as the TSP. Lecture Notes in Computer Science, 1010:401– 410, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deters:1995:CBD

[1869] R. Deters. Case-based diagnosis of multiple faults. Lecture Notes in Computer Science, 1010:411–420, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flinter:1995:AGC

[1870] S. Flinter and M. T. Keane. On the automatic generation of case libraries by chunking chess games. Lecture Notes in Computer Science, 1010:421–430, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1995:LRI

[1871] S. Fox and D. B. Leake. Learning to refine indexing by introspective reasoning. Lecture Notes in Computer Science, 1010:431–440, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goeker:1995:PSI

[1872] M. H. Goeker and H. Birkhofer. Problem solving with “the incredible machine”: An experiment in case-based reasoning. Lecture Notes in Computer Science, 1010:441–450, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grolimund:1995:ICB

[1873] S. Grolimund and J.-G. Ganascia. Integrating case-based reasoning and tabu search for solving optimisation problems. Lecture Notes in Computer Science, 1010:451–460, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanney:1995:STA

[1874] K. Hanney, M. T. Keane, B. Smyth, and P. Cunningham. Systems, tasks and adaptation knowledge: Revealing some revealing dependencies. Lecture Notes in Computer Science, 1010:461– 470, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinrichs:1995:SLF

[1875] T. R. Hinrichs. Some limitations of feature-based recognition in case-based design. Lecture Notes in Computer Science, 1010:471–480, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1995:CBM

[1876] Y. Huang and R. Miles. A case based method for solving relatively stable dynamic constraint satisfaction problems. Lecture Notes in Computer Science, 1010:481–490, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerner:1995:LSE

[1877] Y. Kerner. Learning strategies for explanation patterns: Basic game patterns with applications to chess. Lecture Notes in Computer Science, 1010: 491–500, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khemani:1995:MBH

[1878] D. Khemani and P. V. S. R. Bhanu Prasad. A memory-based hierarchical planner. Lecture Notes in Computer Science, 1010:501–509, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:CBR

[1879] R. W. Lee, R. M. Barcia, and S. K.

Khator. Case-based reasoning for cash flow forecasting using fuzzy retrieval. Lecture Notes in Computer Science, 1010:510–519, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malek:1995:CIA

[1880] M. Malek. A connectionist indexing approach for CBR systems. Lecture Notes in Computer Science, 1010:520–527, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reategui:1995:UNN

[1881] E. Reategui, J. A. Campbell, and S. Borghetti. Using a neural network to learn general knowledge in a case-based system. Lecture Notes in Computer Science, 1010:528–537, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaaf:1995:FSA

[1882] J. W. Schaaf. “fish and sink”: An anytime-algorithm to retrieve adequate cases. Lecture Notes in Computer Science, 1010:538–547, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strube:1995:KEC

[1883] G. Strube, A. Enzinger, D. Janetzko, and M. Knauff. Knowledge engineering for CBR systems from a cognitive science perspective. Lecture Notes in Computer Science, 1010:548–558, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ting:1995:TUS

[1884] Kai Ming Ting. Towards using a single uniform metric in instance-based learning. Lecture Notes in Computer Science, 1010:559–568, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ushida:1995:FAM

[1885] H. Ushida, T. Sato, T. Yamaguchi, and T. Takagi. Fuzzy associative memory system and its application to multi-modal interface. Lecture Notes in Computer Science, 1011:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasabov:1995:HCF

[1886] N. K. Kasabov. Hybrid connectionist fuzzy systems for speech recognition and the use of connectionist production systems. Lecture Notes in Computer Science, 1011:19–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teshnehlab:1995:FGP

[1887] M. Teshnehlab and K. Watanabe.

Fuzzy Gaussian potential neural networks using a functional reasoning. Lecture Notes in Computer Science, 1011:34–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khan:1995:RFL

[1888] E. Khan and F. Unal. Recurrent fuzzy logic using neural network. Lecture Notes in Computer Science, 1011:48– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:IAN

[1889] K.-M. Lee and H. Lee-Kwang. Information aggregating networks based on extended Sugeno’s fuzzy integral. Lecture Notes in Computer Science, 1011: 56–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1995:NFA

[1890] S.-B. Cho. A neuro-fuzzy architecture for high performance classification. Lecture Notes in Computer Science, 1011:67–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuster:1995:ISR

[1891] C. Schuster, T. Schmitz, and M. Hiller. Investigation of stability and robustness of a fuzzy traction control system. Lecture Notes in Computer Science, 1011:85–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honda:1995:KBR

[1892] H. Honda, T. Hanai, Y. Nishida, and I. Fukaya. Knowledge-based rules for control of the sake (ginjoshu) making process and their application in fuzzy control. Lecture Notes in Computer Science, 1011:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:FSE

[1893] M. A. Lee and H. Takagi. A framework for studying the effects of dynamic crossover, mutation. and population sizing in genetic algorithms. Lecture Notes in Computer Science, 1011: 111–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimojima:1995:USL

[1894] K. Shimojima, Y. Hasegawa, and T. Fukuda. Unsupervised/supervised learning for RBF-Fuzzy system — adaptive rules, membership functions and hierarchical structure by genetic algorithm. Lecture Notes in Computer Science, 1011:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leitch:1995:GAD

[1895] D. Leitch and P. Probert. Genetic algorithms for the development of fuzzy controllers for mobile robots. Lecture Notes in Computer Science, 1011:148– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furuhashi:1995:NAG

[1896] T. Furuhashi, Y. Miyata, K. Nakaoka, and Y. Uchikawa. A new approach to genetic based machine learning and an efficient finding of fuzzy rules — proposal of nagoya approach. Lecture Notes in Computer Science, 1011:173– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takeda:1995:NMR

[1897] F. Takeda and S. Omatu. A neuromoney recognition using optimized masks by GA. Lecture Notes in Computer Science, 1011:190–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goonatilake:1995:GFS

[1898] S. Goonatilake, J. A. Campbell, and N. Ahmad. Genetic-fuzzy systems for financial decision making. Lecture Notes in Computer Science, 1011:202– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiedermann:1995:PMM

[1899] J. Wiedermann. Parallel machine models: How they are and where are they going. Lecture Notes in Computer Science, 1012:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hajek:1995:FLL

[1900] P. Hajek. Fuzzy logic from the logical point of view. Lecture Notes in Computer Science, 1012:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tel:1995:SDP

[1901] G. Tel. Sense of direction in processor networks. Lecture Notes in Computer Science, 1012:50–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siegelmann:1995:WST

[1902] H. T. Siegelmann. Welcoming the super Turing theories. Lecture Notes in Computer Science, 1012:83–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horne:1995:WNN

[1903] B. G. Horne, H. T. Siegelmann, and C. L. Giles. What NARX networks can compute. Lecture Notes in Computer Science, 1012:95–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffery:1995:DIP

[1904] K. G. Jeffery. Database: Introduction to problems. Lecture Notes in Computer Science, 1012:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grimson:1995:DIS

[1905] J. Grimson. Distributed information systems. Lecture Notes in Computer Science, 1012:120–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paton:1995:EDT

[1906] N. W. Paton. Extending database technology. Lecture Notes in Computer Science, 1012:146–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Betts:1995:ISP

[1907] A. J. G. Betts. Introducing SSADM4+ and PRINCE. Lecture Notes in Computer Science, 1012:166–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bicarregui:1995:FMP

[1908] J. C. Bicarregui and B. M. Matthews.

Formal methods in practice: A comparison of two support systems for proof. Lecture Notes in Computer Science, 1012:184–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rischel:1995:DSC

[1909] H. Rischel, J. Cuellar, S. Moerk, and A. P. Ravn. Development of safetycritical real-time systems. Lecture Notes in Computer Science, 1012:206– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boerger:1995:WUE

[1910] E. Boerger. Why use evolving algebras for hardware and software engineering? Lecture Notes in Computer Science, 1012:236–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bac:1995:EC

[1911] C. Bac, G. Bernard, D. Conan, and Q. H. Nguyen. Experience with chorus. Lecture Notes in Computer Science, 1012:272–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:1995:HLL

[1912] B. Chapman, P. Mehrotra, and H. P. Zima. High-level languages for parallel scientific computing. Lecture Notes in Computer Science, 1012:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1995:SNA

[1913] H. Maurer. On some new aspects of networked multimedia systems. Lecture Notes in Computer Science, 1012: 315–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Limpouch:1995:QVG

[1914] A. Limpouch and K. Charvat. Quo vadis GIS: From GIS to GIMS and open GIS. Lecture Notes in Computer Science, 1012:334–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marshall:1995:WWW

[1915] V. A. Marshall. WWW — the World Wide Web. Lecture Notes in Computer Science, 1012:350–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borovansky:1995:IHO

[1916] P. Borovansky. Implementation of higher-order unification based on calculus of explicit substitution. Lecture Notes in Computer Science, 1012:363– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cappelli:1995:MHO

[1917] A. Cappelli, C. De Castro, and M. R.

Scalas. A modular history-oriented access structure for bitemporal relational databases. Lecture Notes in Computer Science, 1012:369–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gellersen:1995:SEM

[1918] H.-W. Gellersen. Software engineering meets human-computer interaction: Integrating user interface design in an object-oriented methodology. Lecture Notes in Computer Science, 1012:375–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holan:1995:PFW

[1919] T. Holan, V. Kubon, and M. Platek.

Parsing of free-word-order languages. Lecture Notes in Computer Science, 1012:379–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jennings:1995:DAF

[1920] E. Jennings. Distributed algorithm for finding a core of a tree network. Lecture Notes in Computer Science, 1012:385–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kozak:1995:SSR

[1921] P. Kozak. Stepwise synthesis of reactive programs. Lecture Notes in Computer Science, 1012:391–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1995:SEI

[1922] W. X. Li. A simple and efficient incremental LL(1) parsing. Lecture Notes in Computer Science, 1012:399–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucas:1995:FCS

[1923] S. Lucas. Fundamentals of contexsensitive rewriting. Lecture Notes in Computer Science, 1012:405–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matyska:1995:CLP

[1924] L. Matyska and H. Bures. Constraint logic programming with fuzzy sets. Lecture Notes in Computer Science, 1012:413–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meunier:1995:PPI

[1925] F. Meunier and P. Zemanek. Parallel processing of image database queries. Lecture Notes in Computer Science, 1012:419–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Motyckova:1995:MFP

[1926] L. Motyckova. Maximum flow problem in distributed environment. Lecture Notes in Computer Science, 1012: 425–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phuong:1995:FST

[1927] N. H. Phuong. Fuzzy set theory and medical expert systems: Survey and model. Lecture Notes in Computer Science, 1012:431–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pirotte:1995:FOO

[1928] A. Pirotte, T. Van den Berghe, and E. Zimanyi. The fusion object-oriented method: an evaluation. Lecture Notes in Computer Science, 1012:437–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Repaska:1995:IOO

[1929] Z. Repaska. Integration of objectoriented analysis and algebraic specifications. Lecture Notes in Computer Science, 1012:443–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rozloznik:1995:ISR

[1930] M. Rozloznik and Z. Strakos. On the implementation of some residual minimizing Krylov space methods. Lecture Notes in Computer Science, 1012:449– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1995:FLR

[1931] C. Smith and C. A. Tau. A formal lazy replication regime for spreading conversion functions over objectbases. Lecture Notes in Computer Science, 1012:455–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sima:1995:HL

[1932] J. Sima. Hopfield languages. Lecture Notes in Computer Science, 1012:461– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stuller:1995:ICR

[1933] J. Stuller. Inconsistency conflict resolution. Lecture Notes in Computer Science, 1012:469–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tambouris:1995:MPS

[1934] E. Tambouris and P. Van Santen. A methodology for performance and scalability analysis. Lecture Notes in Computer Science, 1012:475–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuma:1995:ESV

[1935] M. Tuma and M. Rozloznik. On the efficiency of superscalar and vector computer for some problems in scientific computing. Lecture Notes in Computer Science, 1012:481–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vojtas:1995:LPR

[1936] P. Vojtas and L. Paulik. Logic programming in RPL and RQL. Lecture Notes in Computer Science, 1012:487– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zizka:1995:RHC

[1937] J. Zizka and I. Snajdarkova. Recognition of handwritten characters using instance-based learning algorithms. Lecture Notes in Computer Science, 1012:493–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ceri:1995:DAR

[1938] S. Ceri, E. Baralis, P. Fraternali, and S. Paraboschi. Design of active rule applications: Issues and approaches. Lecture Notes in Computer Science, 1013:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Motakis:1995:CTE

[1939] I. Motakis and C. Zaniolo. Composite temporal events in active database rules: A logic-oriented approach. Lecture Notes in Computer Science, 1013: 19–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baralis:1995:RTD

[1940] E. Baralis, S. Ceri, and S. Paraboschi.

Run-time detection of non-terminating active rule systems. Lecture Notes in Computer Science, 1013:38–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaniolo:1995:ADR

[1941] C. Zaniolo. Active database rules with transaction-conscious stable-model semantics. Lecture Notes in Computer Science, 1013:55–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ross:1995:EFO

[1942] K. A. Ross. Efficiently following object references for large object collections and small main memory. Lecture Notes in Computer Science, 1013:73– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemp:1995:EPE

[1943] D. B. Kemp, K. Ramamohanarao, and P. J. Stuckey. ELS-Programs and the efficient evaluation of nonstratified programs by transformation to ELS. Lecture Notes in Computer Science, 1013:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lifschitz:1995:PBP

[1944] S. Lifschitz, R. N. Melo, and E. Pacitti. Practical behavior of parallelization strategies for datalog. Lecture Notes in Computer Science, 1013:109–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robles:1995:EDD

[1945] C. Robles, J. Lobo, and T. Gaasterland. An experimental distributed deductive database system. Lecture Notes in Computer Science, 1013:128– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinn:1995:IDQ

[1946] A. Dinn, N. W. Paton, M. H. Williams, and A. A. A. Fernandes. The implementation of a deductive query language over an OODB. Lecture Notes in Computer Science, 1013:143–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakonstantinou:1995:QTS

[1947] Y. Papakonstantinou, A. Gupta, H. Garcia-Molina, and J. Ullman. A query translation scheme for rapid implementation of wrappers. Lecture Notes in Computer Science, 1013:161– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kifer:1995:DOD

[1948] M. Kifer. Deductive and object data languages: A quest for integration. Lecture Notes in Computer Science, 1013: 187–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abiteboul:1995:DER

[1949] S. Abiteboul and J. Van Den Bussche. Deep equality revisited. Lecture Notes in Computer Science, 1013:213– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calvanese:1995:SOM

[1950] D. Calvanese, G. De Giacomo, and M. Lenzerini. Structured objects: Modeling and reasoning. Lecture Notes in Computer Science, 1013:229–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsukamoto:1995:IRR

[1951] M. Tsukamoto and S. Nishio. Inheritance reasoning by regular sets in knowledge-bases with dot notation. Lecture Notes in Computer Science, 1013:247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobbie:1995:RAC

[1952] G. Dobbie and R. Topor. Resolving ambiguities caused by multiple inheritance. Lecture Notes in Computer Science, 1013:265–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaede:1995:EPQ

[1953] V. Gaede and O. Guenther. Efficient processing of queries containing userdefined predicates. Lecture Notes in Computer Science, 1013:281–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finance:1995:QPI

[1954] B. Finance, V. Smahi, and J. Fessy. Query processing in IRO-DB. Lecture Notes in Computer Science, 1013:299– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quass:1995:QSH

[1955] D. Quass, A. Rajaraman, Y. Sagiv, and J. Ullman. Querying semistructured heterogeneous information. Lecture Notes in Computer Science, 1013: 319–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yanagisawa:1995:DOO

[1956] Y. Yanagisawa, M. Tsukamoto, and S. Nishio. Deductive object-oriented programming for knowledge-base independence. Lecture Notes in Computer Science, 1013:345–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Specht:1995:MGD

[1957] G. Specht and S. Seeberger. Montague grammars as deductive databases. Lecture Notes in Computer Science, 1013: 363–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

PhanLuong:1995:MAN

[1958] V. Phan Luong. A mixed approach to negation in general datalog programs. Lecture Notes in Computer Science, 1013:378–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lawley:1995:TSD

[1959] M. Lawley. Transaction safety in deductive object-oriented databases. Lecture Notes in Computer Science, 1013: 395–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerni:1995:CRT

[1960] M. Guerni, J. Ferrie, and J.-F. Pons.

Concurrency and recovery for typed objects using a new commutativity relation. Lecture Notes in Computer Science, 1013:411–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fahrner:1995:TRD

[1961] C. Fahrner and G. Vossen. Transforming relational database schemas into object-oriented schemas according to ODMG-93. Lecture Notes in Computer Science, 1013:429–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hacid:1995:CRC

[1962] M.-S. Hacid and C. Rigotti. Combining resolution and classification for semantic query optimization in DOOD. Lecture Notes in Computer Science, 1013: 447–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:SQO

[1963] Y.-W. Lee and S. I. Yoo. Semantic query optimization for object queries. Lecture Notes in Computer Science, 1013:467–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xie:1995:NCD

[1964] Z. Xie and J. Han. Normalization and compilation of deductive and objectoriented database programs for efficient query evaluation. Lecture Notes in Computer Science, 1013:485–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraani-Dastjerdi:1995:MAO

[1965] A. Baraani-Dastjerdi, J. Pieprzyk, R. Safavi-Naini, and J. R. Getta. A model of authorization for objectoriented databases based on object views. Lecture Notes in Computer Science, 1013:503–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seki:1995:AAQ

[1966] H. Seki, Y. Ishihara, and M. Ito. Authorization analysis of queries in object-oriented databases. Lecture Notes in Computer Science, 1013:521– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ye:1995:SVD

[1967] X. Ye, C. Parent, and S. Spaccapietra. On the specification of views in DOOD systems. Lecture Notes in Computer Science, 1013:539–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blumenthal:1995:IDA

[1968] B. Blumenthal. Industrial design and activity theory: A new direction for designing computer-based artifacts. Lecture Notes in Computer Science, 1015: 1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viehstaedt:1995:GED

[1969] Gerhard Viehstaedt and Mark Minas.

Generating editors for direct manipulation of diagrams. Lecture Notes in Computer Science, 1015:17–25, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1995:AWI

[1970] B. H. Thomas and P. R. Calder.

Animating widgets in the InterViews toolkit. Lecture Notes in Computer Science, 1015:26–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moerch:1995:AUB

[1971] A. Moerch. Application units: Basic building blocks of tailorable applications. Lecture Notes in Computer Science, 1015:45–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linard:1995:DNS

[1972] M. Linard and R. Zeiliger. Designing navigational support for educational software. Lecture Notes in Computer Science, 1015:63–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bardram:1995:SDT

[1973] J. E. Bardram and O. W. Bertelsen. Supporting the development of transparent interaction. Lecture Notes in Computer Science, 1015:79–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpineto:1995:ULB

[1974] C. Carpineto and G. Romano. ULYSSES: A lattice-based multiple interaction strategy retrieval interface. Lecture Notes in Computer Science, 1015:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chelnokov:1995:HM

[1975] V. M. Chelnokov and V. L. Zephyrova. Hypertext macrodynamics. Lecture Notes in Computer Science, 1015: 105–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrews:1995:SMO

[1976] K. Andrews, H. Maurer, and N. Scherbakov. Semantic modeling of objectoriented hypermedia databases. Lecture Notes in Computer Science, 1015: 121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prokhorov:1995:PMS

[1977] V. V. Prokhorov. PYTHAGORAS:

Multienvironment software. Lecture Notes in Computer Science, 1015:135– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koshizuka:1995:SIO

[1978] N. Koshizuka and K. Sakamura. The shared interaction objects architecture: Integration of advanced GUI techniques in distributed environments. Lecture Notes in Computer Science, 1015:149–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valkovsky:1995:PRG

[1979] V. B. Valkovsky, D. L. Krechman, I. S. Nikiforov, and D. S. Chenosov. The processing of recursive generalization in the construction of hypermedia macrostructure. Lecture Notes in Computer Science, 1015:166–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nitsche-Ruhland:1995:CKB

[1980] D. Nitsche-Ruhland and G. Zimmermann. CritiGUI — knowledge-based support for the user interface design process in smalltalk. Lecture Notes in Computer Science, 1015:179–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaptelinin:1995:ATB

[1981] V. Kaptelinin, K. Kuutti, and L. Bannon. Activity theory: Basic concepts and applications. Lecture Notes in Computer Science, 1015:189–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kranakis:1995:VDG

[1982] E. Kranakis, D. Krizanc, B. Ruf, and J. Urrutia. VC-Dimensions for graphs. Lecture Notes in Computer Science, 1017:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1995:FCS

[1983] T. Kloks, D. Kratsch, and H. Mueller. Finding and counting small induced subgraphs efficiently. Lecture Notes in Computer Science, 1017:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babel:1995:IGF

[1984] L. Babel and S. Olariu. On the isomorphism of graphs with few P4s. Lecture Notes in Computer Science, 1017: 24–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degiorgi:1995:DAL

[1985] D. G. Degiorgi and K. Simon. A dynamic algorithm for line graph recognition. Lecture Notes in Computer Science, 1017:37–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAmore:1995:IHG

[1986] F. D’Amore and R. Giaccio. Incremental hive graph. Lecture Notes in Computer Science, 1017:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djidjev:1995:PGE

[1987] H. N. Djidjev and S. M. Venkatesan.

Planarization of graphs embedded on surfaces. Lecture Notes in Computer Science, 1017:62–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krumke:1995:CAC

[1988] S. O. Krumke, H. Noltemeier, S. S.

Ravi, and M. V. Marathe. Complexity and approximability of certain bicriteria location problems. Lecture Notes in Computer Science, 1017:73–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plump:1995:TGR

[1989] D. Plump. On termination of graph rewriting. Lecture Notes in Computer Science, 1017:88–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauderon:1995:UAG

[1990] M. Bauderon. A uniform approach to graph rewriting: The pullback approach. Lecture Notes in Computer Science, 1017:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fracchia:1995:VTT

[1991] F. D. Fracchia. Visualizing twoand three-dimensional models of meristematic growth. Lecture Notes in Computer Science, 1017:116–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmann:1995:GTM

[1992] S. Hartmann. Graph-theoretical methods to construct entity-relationship databases. Lecture Notes in Computer Science, 1017:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiermeyer:1995:AAC

[1993] I. Schiermeyer. An approximation algorithm for 3-Colourability. Lecture Notes in Computer Science, 1017:146– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1995:MT

[1994] S. Fischer and L. Torenvliet. The malleability of TSP2Opt. Lecture Notes in Computer Science, 1017:152–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alimonti:1995:NOL

[1995] P. Alimonti. Non-oblivious local search for graph and hypergraph coloring problems. Lecture Notes in Computer Science, 1017:167–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1995:IRS

[1996] H. L. Bodlaender, R. B. Tan, D. M. Thilikos, and J. Van Leeuwen. On interval routing schemes and treewidth. Lecture Notes in Computer Science, 1017:181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wada:1995:HFT

[1997] K. Wada, T. Ikeo, K. Kawaguchi, and W. Chen. Highly fault-tolerant routings and diameter vulnerability for generalized hypercube graphs. Lecture Notes in Computer Science, 1017:197– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MeyeraufderHeide:1995:HPR

[1998] F. Meyer auf der Heide and M. Westermann. Hot-potato routing on multidimensional tori. Lecture Notes in Computer Science, 1017:209–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flammini:1995:DBR

[1999] M. Flammini, G. Gambosi, and S. Salomone. On devising Boolean routing schemes. Lecture Notes in Computer Science, 1017:222–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birchler:1995:TGT

[2000] B. D. Birchler, A.-H. Esfahanian, and E. Torng. Toward a general theory of unicast-based multicast communication. Lecture Notes in Computer Science, 1017:237–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rolim:1995:OCB

[2001] J. Rolim, O. Sykora, and I. Vrt’o. Optimal cutwidths and bisection widths of 2- and 3-dimensional meshes. Lecture Notes in Computer Science, 1017: 252–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damaschke:1995:SFL

[2002] P. Damaschke. Searching for faulty leaves in binary trees. Lecture Notes in Computer Science, 1017:265–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:NAPb

[2003] Z.-Z. Chen and X. He. NC algorithms for partitioning planar graphs into induced forests and approximating NP-Hard problems. Lecture Notes in Computer Science, 1017:275–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahlhaus:1995:EPM

[2004] E. Dahlhaus. Efficient parallel modular decomposition. Lecture Notes in Computer Science, 1017:290–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonizzoni:1995:MDH

[2005] P. Bonizzoni and G. Della Vedova. Modular decomposition of hypergraphs. Lecture Notes in Computer Science, 1017:303–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfaltz:1995:PCA

[2006] J. L. Pfaltz. Partition coefficients of acyclic graphs. Lecture Notes in Computer Science, 1017:318–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takaoka:1995:SCC

[2007] T. Takaoka. Sub-cubic cost algorithms for the all pairs shortest path problem. Lecture Notes in Computer Science, 1017:333–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deogun:1995:DPG

[2008] J. S. Deogun and D. Kratsch. Diametral path graphs. Lecture Notes in Computer Science, 1017:344–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galinier:1995:CGT

[2009] P. Galinier, M. Habib, and C. Paul.

Chordal graphs and their clique graphs. Lecture Notes in Computer Science, 1017:358–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustedt:1995:CDS

[2010] J. Gustedt, M. Morvan, and L. Viennot. A compact data structure and parallel algorithms for permutation graphs. Lecture Notes in Computer Science, 1017:372–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandstaedt:1995:HOG

[2011] A. Brandstaedt, F. F. Dragan, and F. Nicolai. Homogeneously orderable graphs and the Steiner tree problem. Lecture Notes in Computer Science, 1017:381–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degermark:1995:ARP

[2012] M. Degermark, T. Koehler, S. Pink, and O. Schelen. Advance reservations for predicted service. Lecture Notes in Computer Science, 1018:3–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1995:DAR

[2013] D. Ferrari, A. Gupta, and G. Ventre. Distributed advance reservation of real-time connections. Lecture Notes in Computer Science, 1018:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1995:IRR

[2014] L. C. Wolf, L. Delgrossi, R. Steinmetz, and S. Schaller. Issues of reserving resources in advance. Lecture Notes in Computer Science, 1018:28–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leydekkers:1995:CEV

[2015] P. Leydekkers, V. Gay, and L. Franken. A computational and engineering view on open distributed real-time multimedia exchange. Lecture Notes in Computer Science, 1018:41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1995:SUC

[2016] M. B. Jones, P. J. Leach, R. P. Draves, and J. S. Barrera. Support for user-centric modular real-time resource management in the rialto operating system. Lecture Notes in Computer Science, 1018:53–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffay:1995:RBE

[2017] G. K. Jeffay and D. Bennett. A ratebased execution abstraction for multimedia computing. Lecture Notes in Computer Science, 1018:64–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindblad:1995:VPM

[2018] C. J. Lindblad. VuSystem performance measurements. Lecture Notes in Computer Science, 1018:76–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cranor:1995:DUC

[2019] C. D. Cranor and G. M. Parulkar.

Design of universal continuous media I/O. Lecture Notes in Computer Science, 1018:80–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitamura:1995:NAH

[2020] H. Kitamura, K. Taniguchi, H. Sakamoto, and T. Nishida. A new OS architecture for high performance communication over ATM networks-zero-copy architecture-. Lecture Notes in Computer Science, 1018:84–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eleftheriadis:1995:MAQ

[2021] A. Eleftheriadis and D. Anastassiou.

Meeting arbitrary QoS constraints using dynamic rate shaping of coded digital video. Lecture Notes in Computer Science, 1018:89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campbell:1995:DQM

[2022] A. Campbell, D. Hutchinson, and C. Aurrecoechea. Dynamic QoS management for scalable video flows. Lecture Notes in Computer Science, 1018: 101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawachiya:1995:EQC

[2023] K. Kawachiya, M. Ogata, N. Nishio, and H. Tokuda. Evaluation of QoS-Control servers on real-time mach. Lecture Notes in Computer Science, 1018: 117–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreier:1995:SLR

[2024] L. C. Schreier and M. B. Davis. System-level resource management for network-based multimedia applications. Lecture Notes in Computer Science, 1018:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaddha:1995:EES

[2025] N. Chaddha, G. A. Wall, and B. Schmidt. An end to end software only scalable video delivery system. Lecture Notes in Computer Science, 1018:130–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cen:1995:DRT

[2026] S. Cen, C. Pu, R. Staehli, and C. Cowan. A distributed real-time MPEG audio player. Lecture Notes in Computer Science, 1018:142–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolot:1995:AAP

[2027] J.-C. Bolot, H. Crepin, and A. V. Garcia. Analysis of audio packet loss in the Internet. Lecture Notes in Computer Science, 1018:154–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lauer:1995:DAV

[2028] H. C. Lauer, C. Shen, R. Osborne, and J. Howard. Digital audio and video in industrial systems. Lecture Notes in Computer Science, 1018:166–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fall:1995:WVP

[2029] K. Fall, J. Pasquale, and S. Mc-Canne. Workstation video playback performance with competitive process load. Lecture Notes in Computer Science, 1018:170–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulzrinne:1995:WCW

[2030] H. Schulzrinne. When can we unplug the radio and telephone? Lecture Notes in Computer Science, 1018:174– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rothermel:1995:ASS

[2031] K. Rothermel and T. Helbig. An adaptive stream synchronization protocol. Lecture Notes in Computer Science, 1018:178–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1995:MAM

[2032] B. K. Schmidt, J. D. Northcutt, and M. S. Lam. A method and apparatus for measuring media synchronization. Lecture Notes in Computer Science, 1018:190–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nieh:1995:IPS

[2033] J. Nieh and M. S. Lam. Integrated processor scheduling for multimedia. Lecture Notes in Computer Science, 1018:202–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mauthe:1995:SAT

[2034] A. Mauthe and G. Coulson. Scheduling and admission testing for jitter constrained periodic threads. Lecture Notes in Computer Science, 1018:206– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yavatkar:1995:CSA

[2035] R. Yavatkar and K. Lakshman. A CPU scheduling algorithm for continuous media applications. Lecture Notes in Computer Science, 1018:210–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulzrinne:1995:DCC

[2036] H. Schulzrinne. Dynamic configuration of conferencing applications using pattern-matching multicast. Lecture Notes in Computer Science, 1018: 216–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biersack:1995:WNM

[2037] E. Biersack and J. Nonnenmacher. WAVE: A new multicast routing algorithm for static and dynamic multicast groups. Lecture Notes in Computer Science, 1018:228–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettati:1995:CEM

[2038] R. Bettati, D. Ferrari, A. Gupta, and W. Heffner. Connection establishment for multi-party real-time communication. Lecture Notes in Computer Science, 1018:240–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Almeroth:1995:RMC

[2039] K. C. Almeroth and M. H. Ammar. The role of multicast communication in the provision of scalable and interactive video-on-demand service. Lecture Notes in Computer Science, 1018:251–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:RVN

[2040] H. Zhang and E. W. Knightly. RED-VBR: A new approach to support delay-sensitive VBR video in packetswitched networks. Lecture Notes in Computer Science, 1018:258–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goyal:1995:DEE

[2041] P. Goyal, S. S. Lam, and H. M. Vin.

Determining end-to-end delay bounds in heterogeneous networks. Lecture Notes in Computer Science, 1018:273– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bansal:1995:AQB

[2042] V. Bansal, R. J. Siracusa, J. P.

Hearn, and G. Ramamurthy. Adaptive QoS-Based API for ATM networking. Lecture Notes in Computer Science, 1018:285–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1995:BSN

[2043] S. S. Lam and G. G. Xie. Burst scheduling networks: Flow specification and performance guarantees. Lecture Notes in Computer Science, 1018: 289–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1995:NVD

[2044] S. W. Lau and J. C. S. Lui. A novel video-on-demand storage architecture for supporting constant frame rate with variable bit rate retrieval. Lecture Notes in Computer Science, 1018:294– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaney:1995:DIR

[2045] A. J. Chaney, I. D. Wilson, and A. Hopper. The design and implementation of a RAID-3 multimedia file server. Lecture Notes in Computer Science, 1018:306–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buddhikot:1995:EDL

[2046] M. M. Buddhikot and G. M. Parulkar. Efficient data layout, scheduling and playout control in MARS. Lecture Notes in Computer Science, 1018:318– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoller:1995:SRL

[2047] S. D. Stoller and J. D. DeTreville. Storage replication and layout in videoon-demand servers. Lecture Notes in Computer Science, 1018:330–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paek:1995:SMV

[2048] S. Paek, P. Bocheck, and S.-F. Chang.

Scalable MPEG2 video servers with heterogeneous QoS on parallel disk arrays. Lecture Notes in Computer Science, 1018:342–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neufeld:1995:DVB

[2049] G. Neufeld, D. Makaroff, and N. Hutchinson. The design of a variable bit rate continuous media server. Lecture Notes in Computer Science, 1018:354–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1995:CMC

[2050] O. Mueller and T. Nipkow. Combining model checking and deduction for I/Oautomata. Lecture Notes in Computer Science, 1019:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1995:COP

[2051] K. G. Larsen, B. Steffen, and C. Weise.

A constraint oriented proof methodology based on modal transition systems. Lecture Notes in Computer Science, 1019:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1995:UGH

[2052] T. H. Henzinger, P.-H. Ho, and H. Wong-Toi. A user guide to HyTech. Lecture Notes in Computer Science, 1019:41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mader:1995:MCM

[2053] A. Mader. Modal-calculus, model checking and Gauss elimination. Lecture Notes in Computer Science, 1019: 72–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henriksen:1995:MMS

[2054] J. G. Henriksen, J. Jensen, M. Joergensen, and N. Klarlund. MONA: Monadic second-order logic in practice. Lecture Notes in Computer Science, 1019:89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engberg:1995:ESB

[2055] U. H. Engberg and K. S. Larsen. Efficient simplification of bisimulation formulas. Lecture Notes in Computer Science, 1019:111–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roscoe:1995:HCM

[2056] A. W. Roscoe, P. H. B. Gardiner, M. H. Goldsmith, and J. R. Hulance. Hierarchical compression for model-checking CSP or how to check 100200 dining philosophers for deadlock. Lecture Notes in Computer Science, 1019:133–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1995:FEG

[2057] R. Cleaveland, E. Madelaine, and S. Sims. A front-end generator for verification tools. Lecture Notes in Computer Science, 1019:153–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tofts:1995:ALA

[2058] C. Tofts. Analytic and locally approximate solutions to properties of probabilistic processes. Lecture Notes in Computer Science, 1019:174–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFrancesco:1995:MCN

[2059] N. De Francesco, A. Fantechi, S. Gnesi, and P. Inverardi. Model checking of non-finite state processes by finite approximations. Lecture Notes in Computer Science, 1019:195–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bohn:1995:AID

[2060] J. Bohn and S. Roessig. On automatic and interactive design of communicating systems. Lecture Notes in Computer Science, 1019:216–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssen:1995:LKT

[2061] W. Janssen. Layers as knowledge transitions in the design of distributed systems. Lecture Notes in Computer Science, 1019:238–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knoop:1995:PFB

[2062] J. Knoop, B. Steffen, and J. Vollmer. Parallelism for free: bitvector analyses =¿ no state explosion! Lecture Notes in Computer Science, 1019:264–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunt:1995:ECB

[2063] J. Hunt. Evolutionary case based design. Lecture Notes in Computer Science, 1020:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griffiths:1995:FKC

[2064] A. D. Griffiths and D. G. Bridge. Formalising the knowledge content of case memory systems. Lecture Notes in Computer Science, 1020:32–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1995:BVB

[2065] M. Brown and N. Filer. Beauty vs. the beast: The case against massively parallel retrieval. Lecture Notes in Computer Science, 1020:42–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oehlmann:1995:SQE

[2066] R. Oehlmann, P. Edwards, and D. Sleeman. Self-questioning and experimentation: An index vocabulary of situated interaction. Lecture Notes in Computer Science, 1020:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dearden:1995:III

[2067] A. M. Dearden. Improving the interfaces to interactive case memories. Lecture Notes in Computer Science, 1020: 73–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Funk:1995:CMD

[2068] P. J. Funk and D. Robertson. Capturing and matching dynamic behaviour in case-based reasoning. Lecture Notes in Computer Science, 1020:85–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bichindaritz:1995:ICL

[2069] I. Bichindaritz. Incremental concept learning and case-based reasoning: For a co-operative approach. Lecture Notes in Computer Science, 1020:91– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marir:1995:CCI

[2070] F. Marir and I. Watson. Can CBR imitate human intelligence and are such systems easy to design and maintain? A critique. Lecture Notes in Computer Science, 1020:107–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Price:1995:DPV

[2071] C. J. Price and I. S. Pegler. Deciding parameter values with case-based reasoning. Lecture Notes in Computer Science, 1020:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellman:1995:ACB

[2072] J. Ellman. An application of case based reasoning to object oriented database retrieval. Lecture Notes in Computer Science, 1020:134–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dattani:1995:CBR

[2073] I. Dattani and M. Bramer. Case-based reasoning: A technique for“decision support systems” in residential valuation and the construction of residential housing. Lecture Notes in Computer Science, 1020:142–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pal:1995:HSD

[2074] K. Pal and J. A. Campbell. A hybrid system for decision-making about assets in English divorce cases. Lecture Notes in Computer Science, 1020:152– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maguire:1995:ACB

[2075] P. Maguire, V. Shankararaman, R. Szegfue, and L. Morss. Application of case-based reasoning (CBR) to software reuse. Lecture Notes in Computer Science, 1020:166–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgin:1995:CAF

[2076] E. Georgin, F. Bordin, S. Loesel, and J. R. McDonald. CBR applied to fault diagnosis on steam turbines. Lecture Notes in Computer Science, 1020:175– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perera:1995:NIC

[2077] S. Perera and I. Watson. Nirmani:

An integrated case-based system for strategic design and estimating. Lecture Notes in Computer Science, 1020: 185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khan:1995:CBC

[2078] T. Khan and Yau Jim Yip. Case-based curriculum planning for computeraided instruction. Lecture Notes in Computer Science, 1020:201–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiederhold:1995:MSM

[2079] G. Wiederhold. Modeling and system maintenance. Lecture Notes in Computer Science, 1021:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1995:ASD

[2080] L. Liu. Adaptive schema design and evaluation in an object-oriented information system. Lecture Notes in Computer Science, 1021:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thelemann:1995:ACW

[2081] S. Thelemann. Assertion of consistency within a complex object database using a relationship construct. Lecture Notes in Computer Science, 1021:32– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraani-Dastjerdi:1995:CMO

[2082] A. Baraani-Dastjerdi, J. Pieprzyk, R. Safavi-Naini, and J. Getta. A cryptographic mechanism for objectinstance-based authorization in objectoriented database systems. Lecture Notes in Computer Science, 1021:44– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liddle:1995:UMP

[2083] S. Liddle, D. Embley, and S. Woodfield. Unifying modeling and programming through an active,. objectoriented, model-equivalent programming language. Lecture Notes in Computer Science, 1021:55–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1995:DQA

[2084] M. Gogolla. A declarative query approach to object identification. Lecture Notes in Computer Science, 1021:65– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fung:1995:VQF

[2085] C. Fung and Q. Li. Versatile querying facilities for a dynamic object clustering model. Lecture Notes in Computer Science, 1021:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vermeer:1995:RER

[2086] M. Vermeer and P. Apers. Reverse engineering of relational database applications. Lecture Notes in Computer Science, 1021:89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vidal:1995:RAS

[2087] V. Vidal and M. Winslett. A rigorous approach to schema restructuring. Lecture Notes in Computer Science, 1021:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouneffa:1995:MSC

[2088] M. Bouneffa and N. Boudjlida. Managing schema changes in objectrelationship databases. Lecture Notes in Computer Science, 1021:113–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teisseire:1995:BCW

[2089] M. Teisseire. Behavioural constraints:

Why using events instead of states. Lecture Notes in Computer Science, 1021:123–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreft:1995:BCE

[2090] M. Schreft and M. Stumptner. Behavior consistent extension of object life cycles. Lecture Notes in Computer Science, 1021:133–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burg:1995:CXE

[2091] J. Burg and R. Van de Riet. COLOR-X even model: Integrated specification of the dynamics of individual objects. Lecture Notes in Computer Science, 1021:146–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SrinivasanHands:1995:DDB

[2092] P. Srinivasan Hands, G. Vignes, and A. Srinivasan. Database design with behavior and views using parameterized Petri nets. Lecture Notes in Computer Science, 1021:158–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oh:1995:SSE

[2093] Y.-C. Oh and S. Navathe. SEER: Security enhanced entity-relationship model for secure relational databases. Lecture Notes in Computer Science, 1021: 170–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellmer:1995:NNT

[2094] E. Ellmer, C. Huemer, D. Merkl, and G. Pernul. Neural network technology to support view integration. Lecture Notes in Computer Science, 1021:181– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halpin:1995:DST

[2095] T. Halpin and H. Proper. Database schema transformation and optimization. Lecture Notes in Computer Science, 1021:191–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Missaoui:1995:MEE

[2096] R. Missaoui, J.-M. Gagnon, and R. Godin. Mapping an extended entityrelationship schema into a schema of complex objects. Lecture Notes in Computer Science, 1021:204–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1995:BRT

[2097] T. Jones and I.-Y. Song. Binary representation of ternary relationships in ER conceptual modeling. Lecture Notes in Computer Science, 1021:216–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diskin:1995:VSF

[2098] Z. Diskin and B. Cadish. Variable sets and functions framework for conceptual modeling: Integrating ER and OO via sketches with dynamic markers. Lecture Notes in Computer Science, 1021:226–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeTroyer:1995:LFS

[2099] O. De Troyer and R. Meersman. A logic framework for a semantics of object oriented data modeling. Lecture Notes in Computer Science, 1021:238–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saeki:1995:OOM

[2100] M. Saeki. Object-oriented meta modeling. Lecture Notes in Computer Science, 1021:250–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarzabek:1995:CMB

[2101] S. Jarzabek and T. W. Ling. A conceptual model for business re-engineering methods and tools. Lecture Notes in Computer Science, 1021:260–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruhn:1995:DME

[2102] V. Gruhn, C. Pahl, and M. Wever. Data model evolution as a business process management. Lecture Notes in Computer Science, 1021:270–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castano:1995:REP

[2103] S. Castano and V. De Antonellis. Reengineering processess in public administrations. Lecture Notes in Computer Science, 1021:282–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knapp:1995:UCR

[2104] J. Knapp. Uniqueness conditions for ER representations. Lecture Notes in Computer Science, 1021:296–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1995:RPI

[2105] I. Mitchell, I. Ferguson, and N. Parrington. Rapid prototyping: An integrated CASE based approach. Lecture Notes in Computer Science, 1021: 308–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Venable:1995:ISE

[2106] J. Venable and J. Grundy. Integrating and supporting entity relationship and object role models. Lecture Notes in Computer Science, 1021:318–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saake:1995:OST

[2107] G. Saake, S. Conrad, and C. Tuerker.

From objects specification towards agent design. Lecture Notes in Computer Science, 1021:329–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casati:1995:CMW

[2108] F. Casati, S. Ceri, B. Pernici, and G. Pozzi. Conceptual modeling of work flows. Lecture Notes in Computer Science, 1021:341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hawryszkiewycz:1995:OOA

[2109] I. Hawryszkiewycz. An object oriented approach for CSCW system design. Lecture Notes in Computer Science, 1021:355–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1995:STV

[2110] C. Jensen and R. Snodgrass. Semantics of time-varying attributes and their use for temporal database design. Lecture Notes in Computer Science, 1021:366–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gal:1995:HCM

[2111] A. Gal and O. Etzion. Handling change management using temporal active repositories. Lecture Notes in Computer Science, 1021:378–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kouramajian:1995:GQL

[2112] V. Kouramajian and M. Gertz. A graphical query language for temporal databases. Lecture Notes in Computer Science, 1021:388–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmitt:1995:MOI

[2113] I. Schmitt and G. Saake. Managing object identity in federated database systems. Lecture Notes in Computer Science, 1021:400–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Getta:1995:MII

[2114] J. R. Getta and L. A. Maciaszek. Management of inconsistent information in federated systems. Lecture Notes in Computer Science, 1021:412–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:RSC

[2115] M.-L. Lee and T. W. Ling. Resolving structural conflicts in the integration of entity-relationship schemas. Lecture Notes in Computer Science, 1021:424– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michelsen:1995:UCD

[2116] H. E. Michelsen. Use and control of data models: A repository approach. Lecture Notes in Computer Science, 1021:434–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moody:1995:SHH

[2117] D. Moody. The seven habits of highly effective data modelers (and object modelers?). Lecture Notes in Computer Science, 1021:436–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tasker:1995:ORC

[2118] D. Tasker. Object responsibility categories in support of top-down object discovery. Lecture Notes in Computer Science, 1021:438–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tonkin:1995:DOE

[2119] D. Tonkin. Denormalisation as an OO extension. Lecture Notes in Computer Science, 1021:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Unhelkar:1995:DFM

[2120] B. Unhelkar. Development of financial markets analytical products: a MOSES case study. Lecture Notes in Computer Science, 1021:441–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ha:1995:PNI

[2121] J. Ha, K. Jeong, S. Kim, and S. Choi.

The practice of nuclear integrated design and engineering database establishment using OO DBMS. Lecture Notes in Computer Science, 1021:444– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dampney:1995:ACD

[2122] K. Dampney and M. Johnson. Application of “consistent dependency” to corporate and project information models. Lecture Notes in Computer Science, 1021:445–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkins:1995:UAC

[2123] C. Atkins. The use of agglomerative categorisation to build a conceptual view of statistical data. Lecture Notes in Computer Science, 1021:447– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turner:1995:IPE

[2124] D. A. Turner. Invited paper: Elementary strong functional programming. Lecture Notes in Computer Science, 1022:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keravnou:1995:ICS

[2125] E. T. Keravnou. Introducing computer science undergraduates to principles of programming through a functional language. Lecture Notes in Computer Science, 1022:15–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davison:1995:TCA

[2126] A. Davison. Teaching C after miranda. Lecture Notes in Computer Science, 1022:35–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquot:1995:RIF

[2127] J.-P. Jacquot and J. Guyard. Requirements for an ideal first language. Lecture Notes in Computer Science, 1022: 51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nuthez:1995:SYC

[2128] M. Nuthez, P. Palao, and R. Petha.

A second year course on data structures based on functional programming. Lecture Notes in Computer Science, 1022:65–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thompson:1995:FPT

[2129] S. Thompson and S. Hill. Functional programming through the curriculum. Lecture Notes in Computer Science, 1022:85–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarvis:1995:ULE

[2130] S. Jarvis, S. Poria, and R. Morgan.

Understanding LOLITA: Experiences in teaching large scale functional programming. Lecture Notes in Computer Science, 1022:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karczmarczuk:1995:FPM

[2131] J. Karczmarczuk. Functional programming and mathematical objects. Lecture Notes in Computer Science, 1022: 121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fokker:1995:EAT

[2132] J. Fokker. Explaining algebraic theory with functional programs. Lecture Notes in Computer Science, 1022:139– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lester:1995:ISI

[2133] D. Lester and S. Mintchev. Inducing students to induct. Lecture Notes in Computer Science, 1022:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burton:1995:CSR

[2134] C. T. P. Burton. Conceptual structures for recursion. Lecture Notes in Computer Science, 1022:179–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ODonnell:1995:TCA

[2135] J. O’Donnell. From transistors to computer architecture: Teaching functional circuit specification in hydra. Lecture Notes in Computer Science, 1022:195– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koopman:1995:FPB

[2136] P. Koopman and V. Zweije. Functional programming in a basic database course. Lecture Notes in Computer Science, 1022:215–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kluge:1995:URT

[2137] W. E. Kluge, C. Rathsack, and S.- B. Scholz. Using -RED as a teaching tool for functional programming and program execution. Lecture Notes in Computer Science, 1022:231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hilsdale:1995:CCU

[2138] E. Hilsdale, J. M. Ashley, R. K. Dybvig, and D. P. Friedman. Compiler construction using scheme. Lecture Notes in Computer Science, 1022:251– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartel:1995:BPS

[2139] P. H. Hartel, B. Van Es, and D. Tromp.

Basic proof skills of computer science students. Lecture Notes in Computer Science, 1022:269–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clack:1995:DFS

[2140] C. Clack and C. Myers. The dysfunctional student. Lecture Notes in Computer Science, 1022:289–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albacea:1995:LS

[2141] E. A. Albacea. Leapfrogging samplesort. Lecture Notes in Computer Science, 1023:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1995:ST

[2142] L. Cai. Spanning 2-Trees. Lecture Notes in Computer Science, 1023:10– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1995:MLI

[2143] M.-Y Kao. Minimal linear invariants. Lecture Notes in Computer Science, 1023:23–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baffi:1995:PMM

[2144] L. Baffi and R. Petreschi. Parallel maximal matching on minimal vertex series parallel digraphs. Lecture Notes in Computer Science, 1023:34–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1995:EET

[2145] Tak Wah Lam and Ka Hing Lee. Efficient equality-testing and updating of sets. Lecture Notes in Computer Science, 1023:48–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1995:BSP

[2146] Viet Hai Nguyen and P. Widmayer. Binary space partitions for sets of hyperrectangles. Lecture Notes in Computer Science, 1023:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Byun:1995:TOT

[2147] S. Byun, R. Kennaway, and R. Sleep. Transformation of orthogonal term rewriting systems. Lecture Notes in Computer Science, 1023:73–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Satpathy:1995:ADF

[2148] M. Satpathy, A. Sanyal, and G. Venkatesh. An automaton-driven frame disposal algorithm and its proof of correctness. Lecture Notes in Computer Science, 1023:88–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gandhe:1995:LLC

[2149] M. Gandhe, G. Venkatesh, and A. Sanyal. Labeled lambda-calculus and a generalised notion of strictness. Lecture Notes in Computer Science, 1023:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinchinat:1995:PPR

[2150] S. Pinchinat, E. Rutten, and R. Shyamasundar. Preemption primitives in reactive languages. Lecture Notes in Computer Science, 1023:111–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraczak:1995:MAP

[2151] W. Fraczak. Multi-action process algebra. Lecture Notes in Computer Science, 1023:126–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mycroft:1995:CBC

[2152] A. Mycroft, P. Degano, and C. Priami. Complexity as a basis for comparing semantic models of concurrency. Lecture Notes in Computer Science, 1023:141–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raja:1995:CFC

[2153] N. Raja and R. K. Shyamasundar.

Combinatory formulations of concurrent languages. Lecture Notes in Computer Science, 1023:156–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niehren:1995:CFC

[2154] J. Niehren and M. Mueller. Constraints for free in concurrent computation. Lecture Notes in Computer Science, 1023:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jalote:1995:FIP

[2155] P. Jalote. Formalizing inductive proofs of message recovery in distributed systems. Lecture Notes in Computer Science, 1023:187–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ye:1995:DDT

[2156] X. Ye and J. Keane. Detecting distributed termination in the presence of node failure. Lecture Notes in Computer Science, 1023:195–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1995:EGM

[2157] F. Liu, T. Peinkenkamp, and W. Damm. An extended gradient model for NUMA multiprocessor systems. Lecture Notes in Computer Science, 1023:210–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:EPP

[2158] Shao Dong Chen, H. Shen, and R. Topor. Efficient parallel permutationbased range-join algorithms on meshconnected computers. Lecture Notes in Computer Science, 1023:225–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1995:PBJ

[2159] H. Lu and K.-L. Tan. Pipelined band join in shared-nothing systems. Lecture Notes in Computer Science, 1023: 239–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamamoto:1995:ILP

[2160] K. Yamamoto, S. Naito, and M. Itoh.

Inductive logic programming based on genetic algorithm. Lecture Notes in Computer Science, 1023:254–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nienhuys-Cheng:1995:EST

[2161] S.-H. Nienhuys-Cheng and R. De Wolf. The equivalence of the subsumption theorem and the refutationcompleteness for unconstrained resolution. Lecture Notes in Computer Science, 1023:269–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ngo:1995:PLP

[2162] L. Ngo and P. Haddawy. Probabilistic logic programming and Bayesian networks. Lecture Notes in Computer Science, 1023:286–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1995:DAL

[2163] P. Anderson and D. Basin. Deriving and applying logic program transformers. Lecture Notes in Computer Science, 1023:301–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tong:1995:PDPb

[2164] B.-M. Tong and H.-F. Leung. Performance of a data-parallel concurrent constraint programming system. Lecture Notes in Computer Science, 1023: 319–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bharadwaj:1995:FIP

[2165] R. Bharadwaj, A. Felty, and F. Stomp. Formalizing inductive proofs of network algorithms. Lecture Notes in Computer Science, 1023:335–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Achuthan:1995:TOM

[2166] R. Achuthan, V. S. Alagar, and T. Radhakrishnan. TROM — an object model for reactive system development. Lecture Notes in Computer Science, 1023:350–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1995:DCS

[2167] P. Chan and D. Van Hung. Duration calculus specification of scheduling for tasks with shared resources. Lecture Notes in Computer Science, 1023:365– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iglewski:1995:TSN

[2168] M. Iglewski, M. Kubica, and J. Madey. Trace specifications of non-deterministic multi-object modules. Lecture Notes in Computer Science, 1023:381–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:EME

[2169] X. Chen, P. Inverardi, and C. Montangero. ESP-MC: An experiment in the use of verification tools. Lecture Notes in Computer Science, 1023:396– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1995:MBV

[2170] T. S. Huang and L.-A. Tang. 3D modelbased video coding: Computer vision meets computer graphics. Lecture Notes in Computer Science, 1024:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shah:1995:AMR

[2171] S. Shah and J. K. Aggarwal. Autonomous mobile robot navigation using fish-eye lenses. Lecture Notes in Computer Science, 1024:9–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1995:HPT

[2172] J. Huang and J.-Z. Wang. Highperformance tracking system. Lecture Notes in Computer Science, 1024:17– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hung:1995:PVT

[2173] Y.-P. Hung, C.-Y. Tang, Z. Chen, and S.-W. Shih. A 3D predictive visual tracker for tracking multiple moving objects with a stereo vision system. Lecture Notes in Computer Science, 1024:25–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnes:1995:VGC

[2174] N. Barnes and Z.-Q. Liu. Vision guided circumnavigating autonomous robots. Lecture Notes in Computer Science, 1024:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1995:FDO

[2175] W. H. Wong and H. H. S. Ip. Forcedriven optimization for correspondence establishment. Lecture Notes in Computer Science, 1024:43–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liang:1995:DCV

[2176] X. Liang, B. Li, and S. Liu. The deformed cube: A visualization technique for 3D velocity vector field. Lecture Notes in Computer Science, 1024:51– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ren:1995:IPT

[2177] J. Ren, G. Zeng, and S. Liu. Interactive particle tracing algorithm for unstructured grids. Lecture Notes in Computer Science, 1024:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teo:1995:FRU

[2178] P. C. Teo and C. D. Garfinkle. Fast resampling using vector quantization. Lecture Notes in Computer Science, 1024:66–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1995:BSS

[2179] M. Xu, Z. Tang, and J. Deng. A Bspline surface interpolation technique for reconstructing 3D objects from serial arbitrary shaped planar contours. Lecture Notes in Computer Science, 1024:74–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lengagne:1995:TMR

[2180] R. Lengagne, O. Monga, C. Ge, and Ma Song De. Two methods for a reliable comer detection in 2D images. Lecture Notes in Computer Science, 1024:83–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watzel:1995:DPT

[2181] R. Watzel, K. Braun, A. Hess, and H. Scheich. On the deletability of points in 3D thinning. Lecture Notes in Computer Science, 1024:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

You:1995:RTT

[2182] J. You, W. P. Zhu, H. A. Cohen, and E. Pissaloux. Real-time textured object recognition on distributed systems. Lecture Notes in Computer Science, 1024:99–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murshed:1995:LSV

[2183] N. A. Murshed, F. Bortolozzi, and R. Sabourin. Off-line signature verification without requiring random forgeries for training. Lecture Notes in Computer Science, 1024:107–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oommen:1995:NSR

[2184] B. J. Oommen and R. K. S. Loke. Noisy subsequence recognition using constrained string editing involving substitutions, insertions, deletions and generalized transpositions. Lecture Notes in Computer Science, 1024:116–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ng:1995:GGE

[2185] H. N. Ng and R. L. Grimsdale. GE-OFF — A geometrical editor for fold formation. Lecture Notes in Computer Science, 1024:124–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamae:1995:SPS

[2186] E. Nakamae, J. Chai, H. Inuyama, and F. Kato. Simplification of polygonal surface with attributes. Lecture Notes in Computer Science, 1024:132– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishioka:1995:RPD

[2187] D. Nishioka and M. Nagasawa. Reducing polygonal data by structural grouping algorithm. Lecture Notes in Computer Science, 1024:140–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1995:OOA

[2188] I. S. B. Wong and A. C. Naiman. An object-oriented architecture for Chinese character composition. Lecture Notes in Computer Science, 1024:152– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anisimov:1995:BCR

[2189] V. Anisimov, N. Gorski, D. Price, and O. Baret. Bank check reading: Recognizing the courtesy amount. Lecture Notes in Computer Science, 1024:161– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishijima:1995:AEA

[2190] M. Nishijima and T. Watanabe. An automatic extraction approach of road information on the basis of recognition of character regions. Lecture Notes in Computer Science, 1024:173–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seales:1995:IMM

[2191] W. B. Seales and A. Rajasekar. Interpreting music manuscripts: A logicbased, object-oriented approach. Lecture Notes in Computer Science, 1024: 181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1995:LCC

[2192] J. Liu, M. M. Y. Chang, and W. K.

Chain. On-line Chinese character recognition with attributed relational graph matching. Lecture Notes in Computer Science, 1024:189–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1995:LHA

[2193] X. Li and D.-Y. Yeung. On-line handwritten alphanumeric character recognition using feature sequences. Lecture Notes in Computer Science, 1024:197– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1995:ASM

[2194] R. W. H. Lau. An adaptive supersampling method. Lecture Notes in Computer Science, 1024:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ng:1995:DMM

[2195] C. M. Ng. Dynamic memory mapping for window based display system. Lecture Notes in Computer Science, 1024:215–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1995:CNC

[2196] Y. Zhou and Z. Tang. Convert non-convex meshes to convex meshes for depth sorting in volume rendering. Lecture Notes in Computer Science, 1024:223–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuen:1995:NCC

[2197] H. Yuen and L. Hanzo. A new chain coding scheme for cursive script and line drawings. Lecture Notes in Computer Science, 1024:233–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellaire:1995:AHI

[2198] G. Bellaire and M. Luebbe. Adaptive hierarchical indexing and constrained localization: Matching characteristic views. Lecture Notes in Computer Science, 1024:241–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:SEE

[2199] C.-N. Lee and R. M. Haralick. Statistical estimation for exterior orientation from line-to-line correspondences. Lecture Notes in Computer Science, 1024:250–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dance:1995:HLS

[2200] S. Dance and Z.-Q. Liu. High level scene interpretation using fuzzy belief. Lecture Notes in Computer Science, 1024:258–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tian:1995:ESR

[2201] Y. Tian and H. Tsui. Estimating shape and reflectance of surfaces by color image analysis. Lecture Notes in Computer Science, 1024:266–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pudney:1995:FDU

[2202] C. Pudney, P. Kovesi, and B. Robbins. Feature detection using oriented local energy for 3D confocal microscope images. Lecture Notes in Computer Science, 1024:274–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baciu:1995:AFM

[2203] G. Baciu and R. K. W. Tsang. Advancing front meshing for radiosity solutions. Lecture Notes in Computer Science, 1024:283–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myszkowski:1995:ECB

[2204] K. Myszkowski and T. L. Kunii. An efficient cluster-based hierarchical progressive radiosity algorithm. Lecture Notes in Computer Science, 1024:292– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamae:1995:MSC

[2205] E. Nakamae, G. Jiao, K. Tadamura, and F. Kato. A model of skylight and calculation of its illuminance. Lecture Notes in Computer Science, 1024:304– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fung:1995:RTS

[2206] G. K. P. Fung, H. H. S. lp, and K. C. K. Law. Ray tracing stereoscopic images. Lecture Notes in Computer Science, 1024:313–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lasseter:1995:CMC

[2207] J. Lasseter. Creating memorable characters with computers. Lecture Notes in Computer Science, 1024:322–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanazawa:1995:CAL

[2208] K. Kanazawa, M. Kubo, N. Niki, and H. Satoh. Computer assisted lung cancer diagnosis based on helical images. Lecture Notes in Computer Science, 1024:323–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carreira:1995:CAL

[2209] M. J. Carreira, D. Cabello, M. G.

Penedo, and J. M. Pardo. Computeraided lung nodule detection in chest radiography. Lecture Notes in Computer Science, 1024:331–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sammouda:1995:NNS

[2210] R. Sammouda, N. Niki, and H. Nishitani. Neural networks for the segmentation of magnetic resonance images. Lecture Notes in Computer Science, 1024:339–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vemuri:1995:MAK

[2211] B. C. Vemuri, S. Rahman, and J. Li.

Multiresolution adaptive K-means algorithm for segmentation of brain MRI. Lecture Notes in Computer Science, 1024:347–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawata:1995:CAA

[2212] Y. Kawata, N. Niki, and T. Kumazaki.

Computer-assisted analysis and 3D visualization of blood vessels based on cone-beam CT images. Lecture Notes in Computer Science, 1024:355–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baciu:1995:ISM

[2213] G. Baciu and B. P. W. Lee. An interface for synthesizing 3D multibody structures. Lecture Notes in Computer Science, 1024:363–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fearing:1995:IOR

[2214] P. Fearing. Importance ordering for real-tune depth of field. Lecture Notes in Computer Science, 1024:372–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ho:1995:BCF

[2215] C. S. N. Ho and R. L. Grimsdale. A behavioural control framework for computer animation using AI techniques. Lecture Notes in Computer Science, 1024:381–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hel-Or:1995:TSW

[2216] H. Z. Hel-Or. Time-space weighting for image sequence quantization. Lecture Notes in Computer Science, 1024: 389–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahuja:1995:NIC

[2217] N. Ahuja. A nonfrontal imaging camera. Lecture Notes in Computer Science, 1024:397–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ma:1995:MCS

[2218] L. Ma and Q. Peng. Modifying and controlling of smooth closed surfaces. Lecture Notes in Computer Science, 1024:398–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarfraz:1995:DRO

[2219] M. Sarfraz. Designing of 3D rectangular objects. Lecture Notes in Computer Science, 1024:411–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chun:1995:CAM

[2220] J. Chun and I. R. Greenshields. Classification algorithm for multi-echo magnetic resonance image using Gibbs distributions. Lecture Notes in Computer Science, 1024:419–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luo:1995:PES

[2221] F. Luo, L. Lu, and Z. Houkes. Parameter estimation for SAR image by a model based approach. Lecture Notes in Computer Science, 1024:427– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1995:SCR

[2222] C. Wang, H. Lu, and S. Ma. A synthesized computer recognition system for human hands. Lecture Notes in Computer Science, 1024:435–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gao:1995:IHG

[2223] W. Gao and S. Wang. Independent hand gesture recognition in HandTalker. Lecture Notes in Computer Science, 1024:443–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dehmeshki:1995:AES

[2224] J. Dehmeshki, M. F. Daemi, B. P. Atkin, and N. J. Miles. An adaptive estimation and segmentation technique for determination of major maceral groups in coal. Lecture Notes in Computer Science, 1024:451–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ando:1995:EMC

[2225] M. Ando and S. Morishima. Expression and motion control of hair using fast collision detection methods. Lecture Notes in Computer Science, 1024:463–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawakami:1995:DES

[2226] F. Kawakami, M. Ohkura, H. Yamada, and H. Harashima. 3-D emotion space for interactive communication. Lecture Notes in Computer Science, 1024:471–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peng:1995:IHF

[2227] A. Peng and M. H. Hayes. Iterative human facial expression modeling. Lecture Notes in Computer Science, 1024:479–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1995:SSU

[2228] S.-L. Kao and C.-S. Fuh. Shape from shading using near point light sources. Lecture Notes in Computer Science, 1024:487–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1995:FRL

[2229] Tak Keung Cheng, L. Kitchen, and Z.- Q. Liu. Feature replenishment for longterm visual motion tracking. Lecture Notes in Computer Science, 1024:489– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1995:IRT

[2230] A. M. K. Cheng and X. Huang. An imprecise real-tune video transmission algorithm. Lecture Notes in Computer Science, 1024:491–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adjeroh:1995:MAE

[2231] D. Adjeroh and Moon Chuen Lee.

Mechanisms for automatic extraction of primary features for video indexing. Lecture Notes in Computer Science, 1024:493–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1995:ELB

[2232] H. Li, Q. Yao, J.-H. Choi, and H.-S.

Kwak. An estimation of low bound for two-dimensional image compression coding. Lecture Notes in Computer Science, 1024:495–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiu:1995:TCB

[2233] D. K. Y. Chiu and D. A. Gadishev. Texture comparison based on selected texture primitives. Lecture Notes in Computer Science, 1024:497–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parikh:1995:TAU

[2234] J. A. Parikh, J. DaPonte, and M. Damodaran. Texture analysis of ultrasonic images using backpropagation neural networks. Lecture Notes in Computer Science, 1024:499–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cong:1995:BSB

[2235] G. Cong and S. Ma. B-spline based multiscale signal derivative filtering. Lecture Notes in Computer Science, 1024:501–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuen:1995:CDU

[2236] P. C. Yuen, S. D. Ma, J. Liu, and Y. S. Yeung. Contour decomposition using dominant points and moment difference method. Lecture Notes in Computer Science, 1024:503–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1995:IPA

[2237] Z. Hu and S. Ma. An inherent probabilistic aspect of the Hough transform. Lecture Notes in Computer Science, 1024:505–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buse:1995:FSA

[2238] R. Buse and Z.-Q. Liu. A fuzzy structural approach to handwritten word recognition. Lecture Notes in Computer Science, 1024:507–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raafat:1995:AOU

[2239] H. Raafat and G. Auda. An Arabic OCR using neural network classifiers. Lecture Notes in Computer Science, 1024:510–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zidouri:1995:ACR

[2240] A. Zidouri, S. Chinveeraphan, and M. Sato. Arabic characters recognition based on MCR. Lecture Notes in Computer Science, 1024:512–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mariani:1995:GMU

[2241] R. Mariani, M. P. Deseilligny, J. Labiche, and Y. Lecourtier. Geographic map understanding. algorithms for hydrographic network reconstruction. Lecture Notes in Computer Science, 1024:514–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohta:1995:AAW

[2242] K. Ohta, Y. Nonaka, and F. Miyawaki.

Automatic analyzing of a weaving design with the spatial frequency components. Lecture Notes in Computer Science, 1024:516–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sablatnig:1995:VIW

[2243] R. Sablatnig. Visual inspection of watermeters used for automatic calibration. Lecture Notes in Computer Science, 1024:518–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitulano:1995:CAD

[2244] S. Vitulano, C. Di Ruberto, and M. Nappi. Computer aided diagnosis in radiology. Lecture Notes in Computer Science, 1024:520–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wenger:1995:IAD

[2245] E. Wenger, V. N. Karnaukhov, A. Haidinger, and N. S. Merzlyakov. Image analysis for dating of old manuscripts. Lecture Notes in Computer Science, 1024:522–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamaoka:1995:DLA

[2246] M. Yamaoka and O. lwaki. Document layout analysis using pattern classification method. Lecture Notes in Computer Science, 1024:524–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1995:RED

[2247] Y. Xu and D. Yi. Recognition of engineering drawings based on frame structure theory. Lecture Notes in Computer Science, 1024:526–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griffin:1995:FLS

[2248] P. A. Griffin and H. Noot. FERSA:

Lip-synchronous animation. Lecture Notes in Computer Science, 1024:528– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanOorschot:1995:DCS

[2249] P. C. Van Oorschot. Design choices and security implications in implementing Diffie–Hellman key agreement. Lecture Notes in Computer Science, 1025:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korjik:1995:BKD

[2250] V. Korjik, M. Ivkov, Y. Merinovich, and A. Barg. A broadcast key distribution scheme based on block designs. Lecture Notes in Computer Science, 1025:2–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ashikhmin:1995:MSL

[2251] A. Ashikhmin and A. Barg. Minimal supports in linear codes. Lecture Notes in Computer Science, 1025:13– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shin:1995:SDS

[2252] Sooyoung Kim Shin and P. Sweeney.

Sequential decoding for a subcode of reed solomon codes. Lecture Notes in Computer Science, 1025:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caballero-Gil:1995:LSA

[2253] P. Caballero-Gil and A. Fuster-Sabater. Linear span analysis of a set of periodic sequence generators. Lecture Notes in Computer Science, 1025:22–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1995:MWK

[2254] Y. Han, D. Gollmann, and C. Mitchell. Minimal weight k-SR representations. Lecture Notes in Computer Science, 1025:34–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirschfeld:1995:MCM

[2255] J. W. P. Hirschfeld. The main conjecture for MDS codes. Lecture Notes in Computer Science, 1025:44–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norton:1995:SDA

[2256] G. Norton. Some decoding applications of minimal realization. Lecture Notes in Computer Science, 1025:53–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fan:1995:SPS

[2257] P. Z. Fan and M. Darnell. The synthesis of perfect sequences. Lecture Notes in Computer Science, 1025:63–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Penzhorn:1995:CLW

[2258] W. T. Penzhorn and G. J. Kuehn. Computation of low-weight parity checks for correlation attacks on stream ciphers. Lecture Notes in Computer Science, 1025:74–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1995:SCB

[2259] C. J. Mitchell. A storage complexity based analogue of maurer key esablishment using public channels. Lecture Notes in Computer Science, 1025:84– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ho:1995:SDD

[2260] H. P. Ho and P. Sweeney. Soft decision decoding of reed solomon codes using the dorsch algorithm. Lecture Notes in Computer Science, 1025:94–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacKay:1995:GCB

[2261] D. J. C. MacKay and R. M. Neal.

Good codes based on very sparse matrices. Lecture Notes in Computer Science, 1025:100–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phoenix:1995:QCP

[2262] S. J. D. Phoenix and P. D. Townsend.

Quantum cryptography: Protecting our future networks with quantum mechanics. Lecture Notes in Computer Science, 1025:112–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radu:1995:PEC

[2263] C. Radu, R. Govaerts, and J. Vandewalle. Prepaid electronic cheques using public-key certificates. Lecture Notes in Computer Science, 1025:132–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucks:1995:HTS

[2264] S. Lucks. How traveling salespersons prove their identity. Lecture Notes in Computer Science, 1025:142–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1995:ECA

[2265] A. Smith and C. Boyd. An elliptic curve analogue of McCurley’s key agreement scheme. Lecture Notes in Computer Science, 1025:150–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahmadian-Attari:1995:MDR

[2266] M. Ahmadian-Attari and P. G. Farrell. Multi-dimensional ring TCM codes for fading channels. Lecture Notes in Computer Science, 1025:158–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanTilborg:1995:ACA

[2267] H. C. A. Van Tilborg. Authentication codes: an area where coding and cryptology meet. Lecture Notes in Computer Science, 1025:169–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sendrier:1995:EGB

[2268] N. Sendrier. Efficient generation of binary words of given weight. Lecture Notes in Computer Science, 1025:184– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinch:1995:DRS

[2269] R. G. E. Pinch. Distribution of recurrent sequences modulo prime powers. Lecture Notes in Computer Science, 1025:188–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cachin:1995:LSS

[2270] C. Cachin. On-line secret sharing. Lecture Notes in Computer Science, 1025: 190–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oleshchuk:1995:CRC

[2271] V. A. Oleshchuk. Church–Rosser codes. Lecture Notes in Computer Science, 1025:199–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canteaut:1995:NAF

[2272] A. Canteaut. A new algorithm for finding minimum-weight words in large linear codes. Lecture Notes in Computer Science, 1025:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stansfield:1995:CCS

[2273] E. V. Stansfield and M. Walker. Coding and cryptography for speech and vision. Lecture Notes in Computer Science, 1025:213–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zyablov:1995:SCG

[2274] V. Zyablov, S. Shavgulidze, and J. Justesen. Some constructions of generalised concatenated codes based on unit memory codes. Lecture Notes in Computer Science, 1025:237–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geiselmann:1995:NHF

[2275] W. Geiselmann. A note on the hash function of Tillich and Zemor. Lecture Notes in Computer Science, 1025:257– 263, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veron:1995:CHI

[2276] P. Veron. Cryptanalysis of Harari’s identification scheme. Lecture Notes in Computer Science, 1025:264–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quirke:1995:ASS

[2277] T. M. Quirke and M. Darnell. Analysis of sequence segment keying as a method of CDMA transmission. Lecture Notes in Computer Science, 1025: 270–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buttigieg:1995:CVL

[2278] V. Buttigieg and P. G. Farrell. Constructions for variable-length errorcorrecting codes. Lecture Notes in Computer Science, 1025:282–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paterson:1995:LMD

[2279] M. Paterson, S. Tassa, and U. Zwick. Looking for MUM and DAD: text-text comparisons do help. Lecture Notes in Computer Science, 1026:1–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1995:BBN

[2280] S. Jain. Branch and bound on the network model. Lecture Notes in Computer Science, 1026:11–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dasgupta:1995:NOA

[2281] P. Dasgupta, P. P. Chakrabarti, and S. C. DeSarkar. A near optimal algorithm for the extended cow-path problem in the presence of relative errors. Lecture Notes in Computer Science, 1026:22–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roychoudhury:1995:EAV

[2282] A. Roychoudhury and S. Sur-Kolay. Efficient algorithms for vertex arboricity of planar graphs. Lecture Notes in Computer Science, 1026:37–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diwan:1995:CTC

[2283] A. A. Diwan and N. Usharani. A condition for the three colourability of planar locally path graphs. Lecture Notes in Computer Science, 1026:52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumann:1995:FSR

[2284] P. Baumann and K. Lermer. A framework for the specification of reactive and concurrent systems in Z. Lecture Notes in Computer Science, 1026:62– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zwiers:1995:SDD

[2285] J. Zwiers, U. Hannemann, Y. Lakhneche, and W.-P. De Roever. Synthesizing different development paradigms: Combining top-down with bottom-up reasoning about distributed systems. Lecture Notes in Computer Science, 1026:80–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooman:1995:VPA

[2286] J. Hooman. Verifying part of the AC-CESS.bus protocol using PVS. Lecture Notes in Computer Science, 1026:96– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petrone:1995:RBP

[2287] L. Petrone. Reusing batch parsers as incremental parsers. Lecture Notes in Computer Science, 1026:111–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panangaden:1995:EPI

[2288] P. Panangaden. The expressive power of indeterminate primitives in asynchronous computation. Lecture Notes in Computer Science, 1026:124–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garrigue:1995:TC

[2289] J. Garrigue. The transformation calculus. Lecture Notes in Computer Science, 1026:151–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiuryn:1995:EAB

[2290] J. Tiuryn. Equational axiomatization of bicoercibility for polymorphic types. Lecture Notes in Computer Science, 1026:166–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raynal:1995:CCS

[2291] M. Raynal and A. Schiper. From causal consistency to sequential consistency in shared memory systems. Lecture Notes in Computer Science, 1026:180– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomlinson:1995:OSD

[2292] A. I. Tomlinson and V. K. Garg. Observation of software for distributed systems with RCL. Lecture Notes in Computer Science, 1026:195–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dendris:1995:PAS

[2293] N. D. Dendris, L. M. Kirousis, Y. C.

Stamatiou, and D. M. Thilikos. Partiality and approximation schemes for local consistency in networks of constraints. Lecture Notes in Computer Science, 1026:210–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1995:MES

[2294] D. Kapur and G. Sivakumar. Maximal extensions of simplification orderings. Lecture Notes in Computer Science, 1026:225–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuler:1995:APT

[2295] R. Schuler. Average polynomial time is hard for exponential time under snreductions. Lecture Notes in Computer Science, 1026:240–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravikumar:1995:STG

[2296] S. Ravikumar and D. Sivakumar. On self-testing without the generator bottleneck. Lecture Notes in Computer Science, 1026:248–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1995:OBC

[2297] M. Nielsen and A. Cheng. Observing behaviour categorically. Lecture Notes in Computer Science, 1026:263– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caspi:1995:ARB

[2298] P. Caspi, J.-C. Fernandez, and A. Girault. An algorithm for reducing binary branchings. Lecture Notes in Computer Science, 1026:279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boreale:1995:CBV

[2299] M. Boreale and L. Trevisan. On the complexity of bisimilarity for valuepassing processes. Lecture Notes in Computer Science, 1026:294–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dsouza:1995:EPC

[2300] A. Dsouza and B. Bloom. On the expressive power of CCS. Lecture Notes in Computer Science, 1026:309– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Odersky:1995:PNP

[2301] M. Odersky. Polarized name passing. Lecture Notes in Computer Science, 1026:324–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balasubramanian:1995:PBH

[2302] R. Balasubramanian and V. Raman. Path balance heuristic for selfadjusting binary search trees. Lecture Notes in Computer Science, 1026:338– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RaoKosaraju:1995:PMC

[2303] S. Rao Kosaraju. Pattern matching in compressed texts. Lecture Notes in Computer Science, 1026:349–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arikati:1995:APM

[2304] S. R. Arikati, S. Chaudhuri, and C. D.

Zaroliagis. All-pairs min-cut in sparse networks. Lecture Notes in Computer Science, 1026:363–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biswas:1995:MSU

[2305] S. K. Biswas and S. Kannan. Minimizing space usage in evaluation of expression trees. Lecture Notes in Computer Science, 1026:377–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edelsbrunner:1995:SSM

[2306] H. Edelsbrunner. Smooth surfaces for multi-scale shape representation. Lecture Notes in Computer Science, 1026: 391–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levcopoulos:1995:PCP

[2307] C. Levcopoulos, A. Lingas, and C. Wang. On parallel complexity of planar triangulations. Lecture Notes in Computer Science, 1026:413–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Follert:1995:CLE

[2308] F. Follert, E. Schoemer, J. Sellen, and M. Smid. Computing a largest empty anchored cylinder, and related problems. Lecture Notes in Computer Science, 1026:428–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krznaric:1995:CHC

[2309] D. Krznaric and C. Levcopoulos. Computing hierarchies of clusters from the Euclidean minimum spanning tree in linear time. Lecture Notes in Computer Science, 1026:443–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klarlund:1995:DBA

[2310] N. Klarlund, M. Mukund, and M. Sohoni. Determinizing buechi asynchronous automata. Lecture Notes in Computer Science, 1026:456–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asarin:1995:ATC

[2311] E. Asarin and O. Maler. Achilles and the tortoise climbing up the arithmetical hierarchy. Lecture Notes in Computer Science, 1026:471–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Browne:1995:GTV

[2312] I. A. Browne, Z. Manna, and H. B.

Sipma. Generalized temporal verification diagrams. Lecture Notes in Computer Science, 1026:484–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bianco:1995:MCP

[2313] A. Bianco and L. De Alfaro. Model checking of probabilistic and nondeterministic systems. Lecture Notes in Computer Science, 1026:499–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1995:CSC

[2314] Silvia M. Muller and Wolfgang J. Paul.

The complexity of simple computer architectures, volume 995 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60580-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 270 pp. LCCN QA76.9.A73 M84 1995.

Sudan:1995:ECP

[2315] Madhu Sudan. Efficient checking of polynomials and proofs and the hardness of approximation problems, volume 1001 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60615-7 (Berlin: softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 87 pp. LCCN QA267.S83 1996.

Kistler:1995:DOD

[2316] James Jay Kistler. Disconnected operation in a distributed file system, volume 1002 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60627-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 249 pp. LCCN QA76.9.D5K57 1996.

Nayak:1995:AMP

[2317] P. Pandurang Nayak. Automated modeling of physical systems, volume 1003 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60641-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xx + 232 pp. LCCN QA76.9.C65N39 1995.

Bosselaers:1995:IPS

[2318] Antoon Bosselaers and Bart Preneel. Integrity primitives for secure information systems: final RIPE report of RACE Integrity Primitives Evaluation (R1040), volume 1007 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CO-DEN LNCSD9. ISBN 3-540-60640-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). 239 pp. LCCN QA76.9.A25 I553 1995.

delPobil:1995:SRM

[2319] Angel Pasqual del Pobil and Miguel Angel Serna. Spatial representation and motion planning, volume 1014 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60620-3. ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 242 pp. LCCN TJ211.4 .S63 1995.

Cipolla:1995:AVI

[2320] Roberto Cipolla. Active visual inference of surface shape, volume 1016 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60642-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 193 pp. LCCN TA1634 .C57 1996.

Needham:1997:TE

[2321] Roger M. Needham and David J.

Wheeler. TEA extensions. Report, Cambridge University, Cambridge, UK, October 1997. URL http://www.

movable-type.co.uk/scripts/xtea.

pdf. See also original TEA [1811] and extension XXTEA [2322].

Wheeler:1998:CX

[2322] David J. Wheeler and Roger M. Needham. Correction to XTEA. Report, Cambridge University, Cambridge, UK, October 1998. URL http:/

/www.movable-type.co.uk/scripts/

xxtea.pdf. See also original TEA [1811] and first extension XTEA [2321].